A Complete Bibliography of Publications in IEEE Spectrum (1980–1989)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

20 January 2020 Version 1.00

Title word cross-reference

$1.9 [Ada87e]. $10 [Ler84h]. 3 [Mok81g]. $5 [Zor87b]. ^{∘} [Ano82-190, Hil86].

*85 [Gut85d].

-40 [Hil86]. -D [Mok81g]. -to-75 [Ano82-190].

0 [Ano82-190]. 0.5 [Ano81-184]. 000-element [PW86].

1/30 [Ano84-237]. 1/4 [Voe87j]. 1/4-inch [Voe87j]. 10 [Ano80-193, Per87b]. 10-bit [Ano82-195]. 10-kHz [Per87b]. 10-ns [Ano81-180]. 10-ns/division [Ano81-180].

12 [Ano83-187]. 12-MHz [Ano83-187]. 128-kilobyte [Ano83-184]. 13 [Sha80f]. 14-year [And80]. 16-b [Ano80-195]. 16-bit [Ano81-185, Ano82-193, Ano82-190, Ano82-194, Ano83-184, Gut83g]. 16-kiloword [Ano82-195]. 1880s [Ano85-125]. 1930s [KW86]. 1954 [Wol85]. 1960s [MRRS84]. 1969 [Sti88a]. 1980s [Gib80]. 1984 [Fle85c]. 1k [San88a]. 1k-Bit [San88a].

2 [Ano81-180, Ano84-247]. 2-mV [Ano81-180]. 2/L [Gut89]. 20 [Ano82-195]. 20-year [Ler82a]. 200-fold [Voe87j]. 200-ns [Ano82-185]. 2000 [Jur85a]. 20th [Kel87]. 21 [Gut83b]. 23-year [Wol83]. 25-micrometer [Ano82-191]. 256-kb [Ber82h, Fis83i]. 260-km [Dup82].

260-km/h [Dup82]. 27 [Gut82f]. 28-kilobyte [Ano80-195]. 29 [Kap85e].

3-day [Sha80f]. 3-megabyte-per-second [Ano82-186]. ’30s [Ang80]. 30th [Sha81c]. 32-bit [Ano81-187, Ano83-187, Bor89, Kid81, Sha82, Zor85e]. 32-kiloword [Ano81-184].

4-inch [Voe87j]. 4-mm [Per87b]. 4-year-baccalaureate [Chr85e]. 4.8 [Kap83d]. 4.8-MW [Kap83d]. 40 [Hil86].

50 [Ano81-180]. 50-MHz [Ano81-180].

60 [MFNL85]. 60-Hz [MFNL85]. 64 [Ber81k, PW85b]. 64- [Ano83-184]. 64-kb [Ber81k]. 64-Rkb [Ber81k].

75th [Mas80a]. 765 [Sti88e]. 767 [Ber80a].

8-b [Ano80-193]. 8-kilobyte [Ano81-183, Ano83-184]. ’80 [Chr80m, Kap80d, Sha80e, Sha80f]. 80386-based [Voe87i]. ’82 [Ber82i, Chr82n, Gut82f, Gut82g]. ’83 [Ano83-239, CHG^{+}83, Ano83-162, Gut83b, Ano83-48]. ’84 [Chr84m, Gut84j, Ano84-247, BPEB84, Ler84d, Ano84-52]. ’85 [Chr85m, Ano85-45, Ano85-46]. ’86 [Jur85b, Jur86e, Ano86-49]. 86/10 [Ano81-186]. ’87 [Ano87-231]. ’88 [Ada88f, Bel88f, Cor88a, Fit88f, Fit88g, Jur88d, Jur88e, Kap88, KM88a, MW88, Voe88f, Voe88g, Wol88b, WZ88, Wol88c, Zor88e, Jur87b]. ’89 [BA89, Bel89e, Fit89d, Fit89e, Jur89c, Jur89d, Mur89d, Mur89e, San89d, San89e, Sti89d, Sti89e, Voe89c, Voe89d, WZ89, Zor89i, Zor89h].

’90 [Jur89b].

A/D [Pin80]. Abandoning [Per82d]. ability [DTS86]. able [Ber80d, Zor87e]. Aboard [Spi86, Moy80, Mur88a]. abound [TD84, VWZ86a]. absence [Mos81]. abstraction [Wal84d]. abundance [TD84]. AC [KA86, Kap87c]. academia [Wal82a, Tru83d]. academic [Ano84-86, Gra82c, Gut84f, Per81b]. academics [Kah82]. accelerate [Kap85b, Mel80]. accelerated [Bel87e]. Accelerating [KM82]. accelerators [HTW84]. acceptance [Gut84h, Mok80d]. accepts [Ano83-185]. Access [Tur87, Gra82b, OK86, Wal84c]. accessible [Bee83, Wal83a]. accessing [Mok81h]. accident [DeV81, FAEH84, LP80]. accidental [Kap82a]. accidents [Ano81-151, Fis86b, HOD82, KLCPT80, Ler81f, SS84]. accommodating [Kir80]. accomplished [Wit84, Wol85]. account [Ryd84]. accuracy [Ano82-190, Ano82-241, Fit87c, Wal87h]. accurate [Jur86a, KS86a]. achievable [Ber82j]. achieve [Ham86]. achievement [Fis83b]. Acid [Zor87e]. acoustically [MR82]. acoustooptic [GA88]. acquire [Cor87a]. acquisition [Ano81-181, Ano82-189]. across [Bal84b, Jur85a, Ler84f]. act [Jur87b]. action [AJ85, Bel86c, Chr87a, Eis84, Gib85, Kah81, Kap80a, Kat80, KLCPT80, Ler84c, Red80]. activate [Ber81d]. activist [Wol88d]. activities [Gut84i, Kap84b]. activity [Lla81, Tru83b]. actress [Kah84]. acts [FP86, Fru83]. actuators [BD83]. Ada [Voe87a, Wal85d]. adapt [LH87]. adaptability [Bel85d]. adaptive [Kap85b]. add [Voe87h]. added [PW83]. addition [Sha80c]. Additional [Bel87b]. address [Ano81-187]. addressed [Wet82b]. addresses [Cog84, Smi87]. adds [Jur82c]. adhesion [Kap87c]. Administration [Per80c, Per83f, FWVH86, Ler82a, Lom80b]. Administrative [Kir80]. admissions [Pyk87]. admit [Van84]. adroit [DJ85]. ADS [Ano88a]. advance [VB86]. Advanced [Bel81, DeL82, Jay86, Kap83b, Ler84h, Lof80b, Voe87g, Ano86-66, Bow85, DJ85, Fis87b, Kah83, Kap83c, KS86b, Riv87, Sug80a, Whi87, Zor84c]. Advances [PW85a, Ber80b, Fis86a, Sum80, Wal83a, Wal83c, Wal87f]. advantage [Aml82]. advantages [KS86b]. advent [Bel84e]. adversary [Ada87c, Zor86b]. adverse [And80]. Advertisement [Ano80l, Ano80m, Ano80n, Ano80o, Ano80p, Ano80a, Ano80b, Ano80c, Ano80d, Ano80e, Ano80f, Ano80g, Ano80h, Ano80i, Ano80j, Ano80k, Ano81a, Ano81-27, Ano81-28, Ano81-29, Ano81-30, Ano81-31, Ano81-32, Ano81-33, Ano81-34, Ano81-35, Ano81-36, Ano81b, Ano81c, Ano81d, Ano81e, Ano81f, Ano81g, Ano81h, Ano81i, Ano81j, Ano81k, Ano81l, Ano81m, Ano81n, Ano81o, Ano81p, Ano81q, Ano81r, Ano81s, Ano81t, Ano81u, Ano81v, Ano81w, Ano81x, Ano81y, Ano81z, Ano82c, Ano82-29, Ano82-30, Ano82-31, Ano82-32, Ano82-33, Ano82-34, Ano82-35, Ano82-36, Ano82-37, Ano82-38, Ano82-39, Ano82d, Ano82e, Ano82f, Ano82g, Ano82h, Ano82i, Ano82j, Ano82k, Ano82l, Ano82m, Ano82n, Ano82o, Ano82p, Ano82q, Ano82r, Ano82s, Ano82t, Ano82u, Ano82v, Ano82w, Ano82x, Ano82y, Ano82z, Ano82-27, Ano82-28, Ano83d]. Advertisement [Ano83-30, Ano83-31, Ano83-32, Ano83-33, Ano83-34, Ano83-35, Ano83-36, Ano83-37, Ano83-38, Ano83-39, Ano83-40, Ano83-41, Ano83-42, Ano83-43, Ano83-44, Ano83-45, Ano83-46, Ano83-47, Ano83e, Ano83f, Ano83g, Ano83h, Ano83i, Ano83j, Ano83k, Ano83l, Ano83m, Ano83n, Ano83o, Ano83p, Ano83q, Ano83r, Ano83s, Ano83t, Ano83u, Ano83v, Ano83w, Ano83x, Ano83y, Ano83z, Ano83-27, Ano83-28, Ano83-29, Ano84c, Ano84-29, Ano84-30, Ano84-31, Ano84-32, Ano84-33, Ano84-34, Ano84-35, Ano84-36, Ano84-37, Ano84-38, Ano84-39, Ano84-40, Ano84-41, Ano84-42, Ano84-43, Ano84-44, Ano84-45, Ano84-46, Ano84-47, Ano84-48, Ano84-49, Ano84d, Ano84e, Ano84f, Ano84g, Ano84h, Ano84i, Ano84j, Ano84k, Ano84l, Ano84m, Ano84n, Ano84o, Ano84p, Ano84q, Ano84r, Ano84s, Ano84t, Ano84u, Ano84v, Ano84w, Ano84x, Ano84y, Ano84z, Ano84-27, Ano84-28]. Advertisement [Ano85c, Ano85-29, Ano85-30, Ano85-31, Ano85-32, Ano85-33, Ano85-34, Ano85-35, Ano85-36, Ano85-37, Ano85-38, Ano85-39, Ano85-40, Ano85-41, Ano85-42, Ano85-43, Ano85-44, Ano85d, Ano85e, Ano85f, Ano85g, Ano85h, Ano85i, Ano85j, Ano85k, Ano85l, Ano85m, Ano85n, Ano85o, Ano85p, Ano85q, Ano85r, Ano85s, Ano85t, Ano85u, Ano85v, Ano85w, Ano85x, Ano85y, Ano85z, Ano85-27, Ano85-28, Ano86b, Ano86-28, Ano86-29, Ano86-30, Ano86-31, Ano86-32, Ano86-33, Ano86-34, Ano86-35, Ano86-36, Ano86-37, Ano86-38, Ano86-39, Ano86-40, Ano86-41, Ano86-42, Ano86-43, Ano86-44, Ano86-45, Ano86-46, Ano86-47, Ano86c, Ano86d, Ano86e, Ano86f, Ano86g, Ano86h, Ano86i, Ano86j, Ano86k, Ano86l, Ano86m, Ano86n, Ano86o, Ano86p, Ano86q, Ano86r, Ano86s, Ano86t, Ano86u, Ano86v, Ano86w, Ano86x, Ano86y, Ano86z, Ano86-27, Ano88b, Ano88c]. advertisement [Ano82-157, Ano82-171, Ano82-241, Ano83-143, Ano83-169]. Advertisements [Ano80q, Ano80r, Ano80s, Ano80t, Ano80u, Ano80v, Ano80w, Ano80x, Ano80y, Ano80z, Ano80-27, Ano80-28, Ano80-29, Ano80-30, Ano87b, Ano87-27, Ano87c, Ano87d, Ano87e, Ano87f, Ano87g, Ano87h, Ano87i, Ano87j, Ano87k, Ano87l, Ano87m, Ano87n, Ano87o, Ano87p, Ano87q, Ano87r, Ano87s, Ano87t, Ano87u, Ano87v, Ano87w, Ano87x, Ano87y, Ano87z]. advertisers [Ano80-160, Ano80-161]. advertising [Ano80-69, Ano80-70, Ano80-71, Ano80-72, Ano80-73, Ano80-74, Ano80-75, Ano80-76, Ano80-77, Lur84]. Advice [Chr80c, Chr81c, Ano82-220, Kap80a]. Advisor [Per83f, Ber80d]. advisor/companion [Ber80d]. Aegis [Ada88d]. aerie [Per85b]. aeronautics [KM82]. Aerospace [Ada84a, Ada85a, AM86, Ada87a, Ada88f, Den81, Gar81, HW83, KLCPT80, Ler83a, Lom80a, LM81, Moy80, BA89, Lom81a, Lom82]. affect [Kap85c, Per83a]. affected [Per87d]. affection [Fag80f]. affects [DTS86]. After [Gre80, LM81, FZ87a, Lom81b, PW85a, Ver86]. aftermath [CC85]. again [Cor87c, Tor85]. against [Mas85, SPM^{+}80, VWZ86d, Wal82a]. Age [Chr87a, Chr80a, DJ85, Fis86f, Fls83, Ful80, Hef84, Wil87]. age-old [DJ85]. aged [Per80b]. agencies [Ler83e, Wil80]. Agency [Moy80, Gib80, Per80c, Wil80]. Agenda [Gle83, FW87]. aggression [Per84a]. aggressively [Bel85c]. agile [Ada87b]. Aging [Mur89a, PW86, Voe87h]. ago [Bro81a, Hec85, Voe87j]. agreements [Per87f]. agrees [Fis87d]. ahead [Ano81-230, Bel85f, Chr84k, Cos88, Key83, Mar85, Per84a, Sha81e, Voe87g, Wet82a]. AI [WZ87]. aid [Hor84d, Sar82, Wal84f]. aided [Bel81, Ber82j, Hor84b, Jur89b, Kat87, Ler81a, Ler82e, Swe82, Tri82, Ver86, Wer87a]. Aiding [Ano80-40]. aids [Bel84b, Fag80a, Fis83e, Fis84a, HOD82, Jur80, Jur82a, Mas80b, Pic82, Sti88c, Voe87d]. ailing [KA86]. aim [Hor86b]. aimed [KW86]. Air [Cha82, Ler82a, Lom80b, SS84, Ada87c, Ada87e, Kap84e, Lom80d, Wal82c, Wal82d, Zor86a, FP86]. air-traffic [Lom80d]. airborne [Zor87e]. Aircraft [PFMP88, Wal87d, Ada87b, Ada88c, Ano87-240, Fit89b, Mas80b, Mur89a, PW86, Spi86, TBH87, Voe86a, Voe86b, Wal83c, Wal86a]. aired [Ano86-152]. airline [Fis86b, Ler83a]. airliner [Ber80a]. airlines [Dup82, FWVH86, FP86, Voe86d, Wal86a]. airport [Lom80a]. airspace [Lom80d, PW86]. akin [VWZ86e]. al [Ano83-123]. Alaska [Hil83, You84]. Alaskan [Hil83]. AlGoAs [MS84]. algorithm [Ber80f, Emm85]. algorithms [Jay86, Wal87f]. alike [Zor84c]. alive [Chi81]. all-CMOS [Ano81-184]. All-digital [Spi86]. All-electronic [Jur81a]. all-important [Ler83b]. all-optical [KK84]. all-out [Jur82a]. all-purpose [Den81]. alleged [Zor86d]. alliances [Jur87d]. Allied [Cor87a, Kah84]. allocated [Bel87b, Wal84a]. allocations [Bel84e]. allotted [Fis84d]. allow [Voe86a, Voe86b]. allowed [Lep81]. allows [Fis83e, Ros83]. ally [Wal82a]. almost [KM88b, Pan87a, Tri82]. aloft [Esc86]. alone [Ber82g]. along [Ler82j]. already [FW87, VWZ86c, Wal87f]. also [Car83, Gar84, Ham86, Hic83, Par84, Sha84]. alter [Eis84]. alteration [Fis84b]. altering [Lom80b]. alternate [Rub80]. Alternative [Jur87e, Wil80, Fis83f, Gut84f, Jur87a, KM82]. alternatives [Har83, Ril81]. alters [Sti88b]. Although [DTS86, Fis86e, Fis87b, MRRS84, VWZ86e, Ler84a, Vae86]. altitude [SD81]. Alto [Pak85a, Pak85b]. Am [BPEB84, VWZ86c, Per87b]. amass [Fis84g]. Amazon [Ada88a]. ambiguities [McC85]. Ambition [Per84a]. ambitions [de 82]. ambitious [Ber80a, Ram82, Wal84e]. America [Fis86d]. American [Bha85]. Americanizing [Tru82]. amid [Chr80f, Chr80g]. amidst [Wal84c]. ammonia [Jur86c]. ammonia-water [Jur86c]. Among [Ano88d, Car83, Nah83, Wol83, Ber82g, Cau82, DD86, Gut82g, Hef84, Kap81a, Per85c, Wal83e]. Amorphous [Mad86, Car80]. Ampex [Per88a]. amplifiers [Kap83b, KK84]. amplify [KK84]. amps [Ber82f, Wal87h]. amusement [GT82]. Anaheim [Ano84-247, Gut82g, Sha80e]. Analog [Fit88a, Ano86-153, Ano87-239, CR88, JJ87, NS80]. Analog-to-digital [Fit88a]. analysis [Ano81-181, BPEB84, Fre86, KSB87, Kap83c, Lla81, Mil87, Wal84d, Wil81]. analyst [Doy81]. analysts [Sta87a]. analyze [Kap84f, Sha81d]. analyzed [Mor81b]. analyzer [Ano82-189, Gut82h]. analyzes [Dan82]. analyzing [Fre86, Sha87]. anatomies [Fan89]. Anatomy [Ano87-28]. anchor [Fis84d]. Anderson [Ano80-245]. Andrew [Wol84a]. anew [Ano86-152]. Angeles [Cor84]. animate [BKW83]. annealing [Bro81b]. anniversary [DeV81, Mas80a]. announced [Ano87-234, Sha82]. Annoying [Dut80]. Annual [Ano85a, Ano84-72, Ano86-65, Ano87-61]. anonymity [Chr81k, Chr86d]. answer [DJ85, MFNL85, Per83a, Wei84, Wol83]. Antarctica [Beg83]. anticipate [Bel87c]. antilock [Jur85b]. antipollution [Ano86-66]. antitheft [Jur85b]. Antitrust [Per82c, Bel85b]. anxiety [Chr82f]. any [Doy81, Fis86d, Per83a]. anything [Per88e]. ape [Hor86b]. Apollo [Her83c]. appear [Voe87g, Wal87g]. Apple [Ano83-185, Gut83b, Gut84e]. Apple-compatible [Ano83-185]. appliance [Rum85]. appliances [Pan87a]. Application [Ber80b, Bro81b, Ano81-177, FG86, Fis87d, For86, Gut82f, KS86a, Mok80f]. application-specific [FG86, Fis87d]. Applications [Ano80-31, Ano80-32, Ano80-34, Ano80-33, Ano80-38, Ano80-41, Ano80-39, Ano80-40, Ano80-42, Ano80-43, Ano81-44, Ano81-43, Ano81-40, Ano81-45, Ano81-42, Ano81-41, Ano82-48, Ano82-43, Ano82-44, Ano82-46, Ano82-47, Ano82-45, Ano83-51, Ano83-55, Ano83-54, Ano83-56, Ano83-52, Ano83-53, Ano84-54, Ano84-55, Ano84-56, Ano84-57, Ano84-58, Ano84-59, Ano85-53, Ano85-54, Ano85-55, Ano85-56, Ano85-57, Ano85-58, Ano86-50, Ano86-51, Ano86-52, Ber81e, OK86, Wal83c, Ano80-186, Ano80-187, Ano80-196, Ano80-188, Ano80-189, Ano80-193, Ano80-190, Ano80-191, Ano80-197, Ano80-192, Ano80-194, Ano80-195, Ano81-178, Ano81-186, Ano81-184, Ano81-187, Ano81-185, Ano81-179, Ano81-183, Ano81-181, Ano81-182, Ano81-180, Ano82-185, Ano82-193, Ano82-195, Ano82-191, Ano82-192, Ano82-190, Ano82-189, Ano82-194, Ano82-187, Ano82-184, Ano82-186, Ano82-188, Ano83-187, Ano83-182, Ano83-185, Ano83-184, Ano83-183]. applications [Ano83-186, Ber82c, BFB87, Bro88, Cas81, Far86, FEK80, Gut82g, Gut83c, Jak85, Jur87a, Kap85f, LH87, Mal84, MS81, Pro81, TD84, VWZ86b, VWZ86c, Voe87a, Wal87f, Wal87g, Zor85e]. applied [Dou83b]. apply [Her84, Zor84b]. applying [Gev83]. approach [Gra82a, Per83d, WAO85]. approached [Voe87i]. approaches [Chr82a, Cle82, Gut84g, Kap82b, PK81, SS84, Wil80]. Approaching [Kap83f]. approval [PK81]. approved [Zor87b]. April [Gut83b]. Aquarius [Zor88f]. arcade [Per83e]. Arch [Cor88b]. archaic [Rab81]. architect [PW85c]. architects [Sti88c]. architectural [Kat87]. architecturally [AF87]. Architecture [VWZ86d, Dav83b, FHRW^{+}83, Gut83c, Her83b, Ler84b, MMB83, Mok82b, NS80, Wil87]. architectures [Bac82, VH86]. arcs [RK80]. area [Ano84-247, Bel85a, Fis86d, Gra82b, Gut84l, Jur81a, Mok82b, SD81]. areas [CA82, Cau82]. aren’t [Jur82d]. argue [LP80]. arid [Rog81]. arise [Ber83a]. arm [LM81]. armada [Ada87a]. armed [Ano83-54]. armies [Wal82d]. Arms [Chr82b, Chr82c, Chr82d, Ram82, Ada88g, Ano82-156, Ano88o, Che89c, Chr86j, Gut88, Hor86a]. Army [Ada85a, Wal82d, Cla80a, Dan82]. arouse [Flo82]. arrangements [Tru83a]. Array [Ano82-185, Ano82-186, Ber82e, Che89a]. arrays [Ber80g, Bot85, Fre88a, Lof80a, Rog81, Wol80]. arrival [Ber81k]. arrive [Voe87a]. arsenal [Ant87, Fis85c, SPM^{+}80]. arsenals [Wal82d]. arsenide [Bel85c, ELW83, NH86]. Art [Ano84-50, Fre86, Bal81, Ber82j, Che89d, Fis83a, Jur87c, Lem83a, Mcc88, Per85d, Smi87]. article [Per83a]. artifacts [Ano84-146, Ano84-148, Ano84-147]. Artificial [FHRW^{+}83, Wal85d, Wal87g, Ada85c, Ber80d, Cor87b, Fis83b, Kah83, KL85, VWZ86a]. Artificial-intelligence [Wal85d, Wal87g]. artist [Zor86c]. arts [BJ88, Cye84]. ASIC [de 89]. asking [DMW87, Wol83]. asks [Ano80-245]. assemble [VWZ86e]. Assembly [Wer87b, Gro83, Jur81b, KP86, Sha84, Sug80a, Whi87]. assess [Ada84b, Dav87a, Mor81a]. Assessing [Sta86]. assessment [Mor81a, Mor81b, MFNL85, Pak85b, Per82b, Gre80]. assist [Ada87a]. assistance [Voe89b]. Associate [Fit87a]. associates [Fag80f]. association [Fis83c]. associativity [Smi87]. assortment [CC86a]. assuring [SA85]. AT&T [Fle85c]. Atari [Ano82-171, PW83]. athlete [Fag80a]. atom [PC80]. atomic [Gib85]. attack [Ten82]. attaining [Row82]. attempted [Hec85]. attempts [Wal82a]. attendees [Ler84d]. attention [KP86]. Attitudes [Hor86a]. attraction [Ano83-162]. attractions [Gut82g]. attractive [Fag80e, Zor84c]. attributes [Sha80a]. Audio [Gut85a, Mok82a, Che86, Cor88b, Miy84, Per81a, Per84d, Per87b, TV89, Tru83e, Mok81f]. Augmenting [Rau87]. Aupperle [AMG^{+}85]. Author [Ano80-35, Ano82-40]. authors [Ano80-147, Ano81-37, Ano81-132, Ano82b, Ano82a, Ano83b, Ano83c, Ano83a, Ano84b, Ano84a, Ano85b, Ano86a, Ano87a, Chr81c, WAO85]. Auto [JT81, JCH80, Jur81b, Wal87a]. automakers [Fis87a, Riv80]. automaking [Voe87e]. automate [Ber81a, CR88, Sha87]. Automated [Che82, Kap83a, Tul83, Wal84d, BA83, Gut83i, IK84, Kap80c, Ler82e, Ler83a, Mok82b, Sta87a, Wal83d]. automates [Ano81-177]. automatic [BS81, Chr83j, Gro83, Jur87e]. automatically [Fis87b, Jur87b]. Automating [BA83, Gau85, Kat87, Kut83, Ler82a, Ler82g, Mil87, Tri82, Voe88b, Wal83b]. Automation [Gut83i, Kap87a, Sha84, Wal87a, Ano83-107, Ber82b, Gut83f, Jur83c, Kap85b, KG83, Mer83, MW88, Mur89d, Per87b, Rab81, Wal82d, Wal83d, Wer87a]. automatons [KS86a]. automobile [JV87, Jur87d, Riv87]. automobile-manufacturing [Jur87d]. Automobiles [JCH80]. Automotive [JT81, Jur81a, Jur81b, Jur82a, Jur83a, Riv80, Ano83-169, Jur86b]. autonomous [AM86]. available [Ano82-190, DD86, Gar81, Kap85f, Pic82, Wal85d, Wel85]. avalanche [For86]. avenues [Tru83d]. Aviation [FWVH86, Ler82a, Lom80b]. Avionics [Ber80a, Voe87a]. avoid [Bel87c, Chr84j, DMW87, Mar85, Tru84b]. avoidance [KLCPT80]. Avoiding [Bel86c]. await [Per84d]. awaits [WZ85]. Award [Ano85-45, Har83, Ano80-36, Ano81-38, Ano82-135, Ano83-48, Ano84-51, Ano86-48]. Award/83 [Ano83-48]. awarded [Kap84b]. Awards [Ano80-36, Ano80-37, Ano81-38, Ano81-39, Ano82-41, Ano83-49, Ano84-51, Ano86-48, Ano83-48, Ano83-144, Ano84-149, Ano85-46, Ano85-144, Ano86-133, Ano87-152, Ano87-234, Ano83-48, Ano84-52, Ano85-45, Ano85-46, Ano86-49]. Awards/83 [Ano83-49]. Awards/84 [Ano84-52]. Awards/85 [Ano85-45, Ano85-46]. Awards/86 [Ano86-48, Ano86-49]. awareness [Bel84f]. away [Con86, Zor86d]. awry [Chr87h, Lar85]. axed [AM86].

b

[Ano80-193, Ano80-197, Ano80-195, Per84a]. baby [Sum82a]. baccalaureate [Chr85e]. Back [Ano82-42, Ano83-50, Ano85-47, Ano85-48, Ano85-49, Ano85-50, Ano85-51, Ano85-52, Ano87-31, Ano87-29, Ano87-30, Ano87-32, Ano87-33, Ano87-34, Ano87-35, Ano87-36, Ano87-37, Ano87-38, Ano87-39, Ano87-40, Ano87-43, Ano87-41, Ano87-42, Ano87-44, Ano87-45, Ano87-46, Ano87-47, Ano87-48, Ano87-49, Ano87-50, Ano87-51, Ano87-52, Ano88e, Ano88f, Bel89a, Ver86, Ano82-154, Ano83-159, Bel84a, Rog81]. backbone [Fis85c]. backs [Fre86]. bad [Ano88k, Chr80f, Chr80g, Chr85d]. bake [Tus86]. baking [Mur89b]. balance [Zor86b]. balanced [Fis83d]. Ballistic [Eas86, Bel86c, NH86]. Band [Fis84a, Cla80a, PHG86]. Band-aids [Fis84a]. bands [Per86]. bandwidth [Ano80-193, EMM82, Jur83b, Pow84, Per87b]. bang [PTW82]. banged [Joh84]. bangs [Chr88a]. Banishing [Fis87a]. bank [Sho88]. banking [Per88b]. barrier [Lep81, Tan86]. base [Ano81-180, Bee83, Gut82d, Hor84c, Kah82, PD86, Wol80]. based [Ano81-186, Ano82-188, Bal84b, Bel87f, BKBC85, Fis80, Flo88, GGZ86, Gut83f, Gut83h, HOD82, Hic83, Jur82d, Kap82b, KP86, Kuw84, Ler84b, Mas80b, MNS81, Mil87, Nai83, PHG86, Rac80, Tri82, TBH87, Voe87i, Wei84, Zor84b, Zor85a]. basic [Gro83]. basin [Ada88a]. basis [Fis84a]. batch [Fag82c, MMB83]. batch-process [Fag82c]. bath [HGF85]. batteries [Kap80b, Zor84c]. battery [Ano84-85, Bel88d, Kap80c]. battle [Bel87e, FZ87a, JCH80, Per89a]. battlefield [AW85, Flo88, Ler82d, Wal82d]. Bay [Fis86d]. bazaar [Zor85f]. be [Ada84c, AH85, Ano84-237, AJ85, Bac80, Bel87c, Ber80f, Ber82h, Cal82, Chr80l, Dav83a, ELW83, Emm85, Far86, Fis84c, Fis85c, FP86, Fis80, GP81, Gro83, GT82, Gut82f, Gut82g, Hic83, Hil85, Hor85d, Kap80a, Kap83d, KL85, KS86a, KK84, Kuw84, Ler80c, Ler82c, Lof80a, McC85, MFNL85, Mos81, Par84, Pea81, Per81d, PW83, PW84, Per84e, Per85d, Per87b, Pic82, PZ86, Rab81, Rac80, Row82, Rum85, Sha80e, Sti81, Sug80a, Voe87e, Wal83d, Wal87d, Wal87h, Wet82b, Whi87, Wil80, Win84, Zor85d, Zor86d]. beachheads [FM84a]. beam [Bro81b, Eid81, Lem81c, PC80]. beams [Cal82, Fis85c, Lem81c]. beat [Bha85, Tha83]. Beating [Mok80e, Fis86e]. beautiful [Ler80c]. beauty [KM88b, Zor85a]. became [Fle85c, JI84, Wer87a]. Because [Har83, Ber80a, Fis87b]. become [Fis84e, Fre88b, Tul83, Van84]. becomes [Ber80c, Ler82e, Wal85d]. becoming [Hor86a, Per82a, Per83g]. beef [Jur85b]. been [Ada84c, Ada87f, Del80, Gut83i, Ler81b, MRRS84, Mur88b, Tor80, Voe87a, WZ85]. Before [Ano84-53, Ada87e, BFB87, DD86, FZ87a, Gut83c, Tal83, Tor80]. begin [Fis84d, Gut84l, Zor86a]. Beginning [CC86a, ELW83, Fis83b, Kah82, Ler82g]. begins [FM84a, Lom80a]. behave [Voe86c]. behavior [Nag83]. behaviors [GGZ86]. Behind [Chr82e, Chr85a, Cor84, Bel84e, Bel84d]. Behind-the-scenes [Cor84]. being [Aml82, CA82, Cas81, Cye84, Fis87c, Jak85, Ler82b, Ler84g, Per84a, Wei84]. believe [Gra82c, Per83e]. Bell [Bel84a, Bel85b, Bel88a, Dut80, Fle85c, Lep81, Wol83]. Bells [Jur84]. below [Ler84g]. belts [NK87]. Bench [Gut85b, Ler81e, VB86]. benchmark [DMW87]. benchmarking [DMW87]. benders [BJ88]. beneath [Ada85c, Fre86]. beneficiaries [Sug80a]. benefit [Ber82c]. benefits [GBDC86, Gut83h]. Bernard [Mur85]. Best [Ano80-38, Ano80-41, Ano80-39, Ano80-40, Ano80-42, Ano80-43, Ano81-44, Ano81-43, Ano81-40, Ano81-45, Ano81-42, Ano81-41, Ano82-48, Ano82-43, Ano82-44, Ano82-46, Ano82-47, Ano82-45, Ano83-51, Ano83-55, Ano83-54, Ano83-56, Ano83-52, Ano83-53, Ano84-54, Ano84-55, Ano84-56, Ano84-57, Ano84-58, Ano84-59, Ano85-53, Ano85-54, Ano85-55, Ano85-56, Ano85-57, Ano85-58, Ano86-50, Ano86-51, Ano86-52, Bel88d, Chr84g, Fag82a, For86, Fro83, Kap83g, Ler80d, PL89, San89a, San89f, Wil80, Win84, Wol83, Zor87f]. bet [Ler83d]. bets [Jur81b, Ril81]. Better [Voe87c, Wal87a, CS87, Fag81a, Fis84a, Fis84c, Fis86c, HK87, HTW84, Ler82c, Lla81]. betting [Wol84a]. between [Bel85b, Cog84, Dup82, Gau85, HM87, Kah84, Kir80, San88a]. Beware [McC85]. bewilder [BFB87]. Beyond [Bot83, Chr82a, Chr85b, KK84, Rub84, SF82, Bel85c, Bry84, Wit84]. Bhopal [Sti89a]. BiCMOS [San89a]. big [Ano87-226, Fis85b, Jur88b, Kap84e, PTW82, Per82d, You84, Gut83b]. Bigger [Chr88a]. biggest [Wol84a]. billion [Ada87e, Zor87b]. Billions [JCH80, Wil87]. biochemical [Lla81]. biochemical-activity [Lla81]. Biological [Ler80b, Ler84a, Con86, MFNL85]. Biomation [Gut82h]. Biomechanical [Fag80a]. Biomedical [Fis83b, Hor84d]. Biopolar [Mur85]. biosensors [Hor85b]. Bipolar [Ber81j, Gut89, Fis87d]. birth [Wis82]. bist [Ano83-51]. Bistable [ST81]. Bit [San88a, Ano81-187, Ano81-185, Ano82-193, Ano82-195, Ano82-190, Ano82-194, Ano83-187, Ano83-184, Bor89, Gut83g, Jay86, Kid81, Kuw84, Sha82, Whi83, Zor85a, Zor85e]. bit-based [Kuw84]. bites [Ano88g, Ano88h]. Bits [Ano88g, Ano88h, Ano80-38, Ano80-41, Ano80-39, Ano80-40, Ano80-42, Ano80-43, Ano81-44, Ano81-43, Ano81-40, Ano81-45, Ano81-42, Ano81-41, Ano82-48, Ano82-43, Ano82-44, Ano82-46, Ano82-47, Ano82-45, Ano83-55, Ano83-54, Ano83-56, Ano83-52, Ano83-53, Ano84-54, Ano84-55, Ano84-56, Ano84-57, Ano84-58, Ano84-59, Ano85-53, Ano85-54, Ano85-55, Ano85-56, Ano85-57, Ano85-58, Ano86-50, Ano86-51, Ano86-52, Zor85a]. bizarre [Bel87c]. black [Lem81b]. blame [Sta87a]. blamed [Fis86b]. blanking [Eid81]. blast [Flo88]. blatant [Ano88p]. Bleaching [Sha84]. blind [SAPM80, Win84]. block [Hor85e, Jur82c, Mok80d]. blocks [FHRW^{+}83, Sha80f]. blossoms [Fis87c]. blow [Per81d]. blower [Per81d]. blue [Sha84, Sug80a]. blue-collar [Sug80a]. board [Ano81-186, Ano82-193, HHH87, Jur83a, PP88, Tul83]. boards [Tul83]. Body [Fis85a, Bot83, Cor87b, Fis83e, Hor84d]. Boeing [Ber80a, Fit89b]. boggling [AW85]. bogs [FW87]. bold [Bel84a]. Bolstering [San88b]. bolts [PFMP88]. bomber [Ada85a]. bombs [Lem81b, Ler81f]. Bombsight [Sea89]. bona [Fis84e]. bonanza [Gut82g, WZ85]. bones [SH85]. Book [Ait87, Ano80-44, Ano80-45, Ano80-46, Ano80-47, Ano80-48, Ano80-49, Ano80-50, Ano80-51, Ano80-52, Ano80-53, Ano80-54, Ano80-55, Ano81-46, Ano81-47, Ano81-48, Ano81-49, Ano81-50, Ano81-51, Ano81-52, Ano81-53, Ano81-54, Ano81-55, Ano81-56, Ano81-57, Ano81-58, Ano82-49, Ano82-50, Ano82-51, Ano82-52, Ano82-53, Ano82-54, Ano82-55, Ano82-56, Ano82-57, Ano82-58, Ano82-59, Ano82-60, Ano83-57, Ano83-58, Ano83-59, Ano83-60, Ano83-61, Ano83-62, Ano83-63, Ano83-64, Ano83-65, Ano83-66, Ano83-67, Ano83-68, Ano84-60, Ano84-61, Ano84-62, Ano84-63, Ano84-64, Ano84-65, Ano84-66, Ano84-67, Ano84-68, Ano84-69, Ano84-70, Ano84-71, Ano85-59, Ano85-60, Ano85-61, Ano85-62, Ano85-63, Ano85-64, Ano85-65, Ano85-66, Ano85-67, Ano85-68, Ano85-69, Ano85-70, Ano86-53, Ano86-54, Ano86-55, Ano86-56, Ano86-57, Ano86-58, Ano86-59, Ano86-60, Ano86-61, Ano86-62, Ano86-63, Ano86-64, Ano87-53, Ano87-54]. Book [Ano87-55, Ano87-56, Ano87-57, Ano87-58, Ano87-59, Ano87-60, Bro88, Gol87, Oli87, Pan87b]. Book-shelf [Ano81-58]. boom [San88b]. booming [Zor88a]. boost [FZ85, Riv80]. Booth [BPEB84]. borders [Ler84f]. born [FP84, San88a]. borrowed [Fag82b]. borrows [Che89a]. bosses [Chr86g]. Boston [Gut82f, GB85, Jur85a, Ler84d, Sha80f]. Both [Fis85b, Kap80f, Lof80b, Mok80d, AH85, And80, Bot85, Cog84, CC87, Dor83, Fit87c, Fle85c, Fro83, Gau85, Har83, Hor84b, Kah83, Ler80a, LH87, Mok80e, Per82a, PW84, Rum85, San89a]. bottle [Ano82-157]. bottle-warmer [Ano82-157]. bottleneck [Gra82a]. bottlenecks [Gan82, HHH87]. bounce [SD81]. bows [Bow85]. Boys [Kid81]. brain [Ano86-145, Ano87-228, HN88, Tru83b]. brainwasher [Ler84e]. braking [Jur85b]. branches [Per83g]. brand [Wol84a]. Brazil [de 82]. Breaching [Ber82a]. break [Bel88c]. break-up [Bel88c]. breakdown [Kap84b]. breakers [RK80]. Breaking [Gra82a, Zor87a, Tan86]. breaks [Lep81]. breakthrough [Ber80f, Cor88b, Gut89]. breakthroughs [Ber80d, Row82, Tor80]. breed [Fag81a, Sha81d, Tru83d, Zor86c]. breeder [Blu80, Gut82b]. brewing [Gut83c, Ler84f]. bright [Hef84, Ler80c]. brightening [Ada84c]. brighter [Jur83b, Kap83e]. brightness [PW85a]. Brilliant [Aug82]. bring [Bel84c, FZL86, Per81d]. Bringing [Wit84, Che87, Hic83]. brings [Gut83h, WB86]. Britain [Blu80, Oak83]. British [Zor87a]. Broad [Kap82b, HMDF80, KS86a, Wal83a]. Broad-based [Kap82b]. broadcast [Bel84c, Bel84d, Jac80, Jac82, PK81, Wat87]. broadcaster [Bel84e]. broadcasting [Aml82]. broadcasts [HMDF80, Hil83]. brought [DS81, PW84]. brushfire [Kap82a]. bubble [Ber80b, Ber80g, LS81]. Bubbles [Ber80b]. budget [Ler84h]. build [PC80, Sta87b, Wal84g]. Building [Sta86, FHRW^{+}83, Fis85a, Lem81c, Zor85d]. builds [AM86, Per87a]. built [MAN80, Tal83, Wal84e]. bulb [Wis82]. bullet [ABEI84]. burden [Ler82g, Per81e]. burdened [PWZ^{+}86]. Burning [Ano86-66]. bus [DD86]. buses [Bor89, VB86]. busier [Mas80a]. Business [Fis84b, Mok82b, Bel87f, Ber81d, Ber82g, Fis85b, Per82a, Sha81c, Wal83d]. bust [San88b]. buyer [Bel86a, Per81a]. Buying [Tru84a, Gut88]. buys [Ada88b]. Bypassing [Fis85b].

C [Ano82-190, Per84b]. C3

[Ler82i, Ler82j, Din82]. C3I [Mor85].

cabinet [HHH87]. cable

[Bel84d, Bel87b, GZ85, Mok80c]. cables [Kap80f, Kec83, Mok82d, Per83d]. Cache [Smi87]. CAD [Fit87a, Gut83b, Ros83, Wal83b]. Caddy [Ano80-43]. CAE [Fit87a, Wal84c]. cage [Ano83-187]. calculators [Mok82e]. Calendar [Ano80-56, Ano80-57, Ano80-58, Ano80-59, Ano80-60, Ano80-61, Ano80-62, Ano80-63, Ano80-64, Ano80-65, Ano80-66, Ano80-67, Ano81-59, Ano81-60, Ano81-61, Ano81-62, Ano81-63, Ano81-64, Ano81-65, Ano81-66, Ano81-67, Ano81-68, Ano81-69, Ano81-70, Ano82-62, Ano82-63, Ano82-64, Ano82-65, Ano82-66, Ano82-67, Ano82-68, Ano82-69, Ano82-70, Ano82-71, Ano82-72, Ano82-73, Ano83-80, Ano83-69, Ano83-70, Ano83-71, Ano83-72, Ano83-73, Ano83-74, Ano83-75, Ano83-76, Ano83-77, Ano83-78, Ano83-79, Ano84-83, Ano84-84, Ano84-73, Ano84-74, Ano84-75, Ano84-76, Ano84-77, Ano84-78, Ano84-79, Ano84-80, Ano84-81, Ano84-82, Ano85-81, Ano85-82, Ano85-71, Ano85-72, Ano85-73, Ano85-74, Ano85-75, Ano85-76, Ano85-77, Ano85-78, Ano85-79, Ano85-80, Ano86-67, Ano86-68, Ano86-69, Ano86-70, Ano86-71, Ano86-72, Ano86-73, Ano86-74, Ano86-75, Ano86-76, Ano86-77, Ano86-78, Ano87-70, Ano87-71, Ano87-72, Ano87-73, Ano87-62, Ano87-63, Ano87-64]. Calendar [Ano87-65, Ano87-66, Ano87-67, Ano87-68, Ano87-69]. Calif. [Ano84-247]. California [Ril81, Ano83-162, Zor88a]. call [Fis86c, Wal83e]. called [Nah83, Rum85, Zor84b]. calls [Ano80-148, Ano81-134, Ano81-135, Ano81-136, Ano81-137, Ano82-136, Ano82-137, Ano82-138, Ano82-139, Ano83-145, Ano83-146, Ano83-204, Bel88a, LH87]. CAM [Wal83b]. camcorder [Per87b]. camera [Mok82f, Per89a]. Cameras [Gut82a]. Campus [Chr84a, Bry84, FM84a, Kah82]. Can [Fis86f, Hor85d, Lak82, MFNL85, PW84, PZ86, Ten82, Ano81-184, Ano84-168, Apt85, Bal84b, Ber81a, Ber81g, BA83, Cal82, Chr80l, CN83, Cun80, DD85, DMW87, Fis84h, Fle82, Fly87, Gar81, Gau85, HMDF80, HW83, HTW84, Jur87a, KSB87, Kap80a, Kap82c, Kap85f, KS86a, KK84, Kut83, Ler82c, Lom81c, Mar85, Par84, Pro81, Rab81, Rac80, Rog81, Sar82, Sha81d, Sha84, SD81, Sug80b, SW83, Tru84a, Wal82d, Wal87h, Win84]. Canada [ABEI84]. canceling [AF87, Wal86a]. cancer [Jak85, SPM^{+}80]. candidates [Mok80f]. cannot [Fis87b, Jur87a]. can’t [MRRS84, Pyk87]. capabilities [Hor86a, Mok82e, Sha80c, Voe87h]. capability [WZB86]. capable [Bal81]. capacitor [Sol88]. capacity [Bel84c, CA82, EMM82, Kec83, Voe87j]. capital [Per82a, Row82]. Capitalizing [Per87f]. capsule [Chr84f]. captain [Per87a]. capture [Per81a, Sha80c]. captures [Ada87a]. Car [Jur82a, Jur87b, Kap80b, Kap87c, KW86]. card [Ano83-187, Sho88, VB86]. cards [Ano81-44, BFB87, Wei84]. care [Bug84, Gar84]. career [Gut84c]. Careers [Ano80-68, Fag82a, Flo82, Kap81a, Ler83b, Lom80c, Per80b, Per83a, Sha80a, Car84, Per87d, Tru84b]. careful [Ber81g, Mar85]. caring [Ler84h]. Carlson [Wol89]. carriage [Ano82-157]. cars [Jur81a, Jur89b, Kap80c, Kap85d, Kel87, Wal87a]. cartel [Zor86d]. Carter [Per80c, Ano80-245]. Carter-created [Per80c]. cartooning [BKW83]. Case [Ada87b, Ano87-74, Asa87, Bel87a, KM87, NK87, Ril87, Voe87b, Voe87c, Wal87b, Wer87b, BE89, Dav83a, ELW83, FW82, Gut82c, Gut82h, Gut84e, HTW84, PW83, PW85b, Per88e, Sha81e, Sha81f, Zde83]. Cases [Mar85]. cash [Ano81-44]. cashless [Wei84]. cat [Ano84-85]. catch [JCH80]. catch-up [JCH80]. catches [Ler83a, Tan86]. catching [Ber81j, WM84]. catechism [Tri80]. cause [Doy81, Gar81, MFNL85, SW83, Zor86a]. caution [Voe88b]. cautious [Wal82a]. Caveat [Chr84b]. CCDs [Ano81-235]. CD [Per87b]. cease [Mar86]. celebrates [Sha81c]. celebration [Cog84]. Celebrities [Ano87-208]. cell [Fis84e, Kap83d, Lof80a, MS81, Rog81]. cells [Bro81b, Car80, Fis84e, FZ85, Lof80b, Wet82b, Wol80]. Cellular [CN83, Bel85a]. Centennial [Ano84-50, Ano84-86, Ano84-193, Chr83b, Cog84]. Center [Pak85a, Mok80b, Pak85b, Pak85b]. centers [Mas80c]. Centimeter [PHG86]. Centimeter-band [PHG86]. central [Ler83b]. centura [Cog84]. century [Ano88d, Ano88m, Hec85, Kel87, Tru84b]. certain [Bel85d]. Challenge [Get82, AH85, AW85, Ang80, Chr85h, CC85, CC87, DD85, Fis84h, Fls83, GBDC86, Hor85e, Ler80a, Log83, Mer83, Mok80e, Per81a, Sti89c, Tur87, Wal83b]. challenged [Gut84e]. challenger [Eid81, Ano87-75, BE87]. Challenges [VWZ86a, Ada85c, Ber81e, Fis84c, Gut84g, Jur83b, Ler82e, Tor84, Tru83c]. challenging [Gut83c, Mok80e]. chamber [PC80]. champion [Tri80]. chance [Fle85b]. change [Swe82, Voe87h]. changed [Voe87e]. changes [Ano81-133, Ano88p, Cor87b, FZ87a, Ler84a, Tal83]. Changing [CC85, Kat87, Zor89j]. channel [Ano80-192, Ano83-245]. channels [Ano80-194, Ano81-181, Ano82-189, Ano82-194]. characteristics [Sar82]. characterized [ZJ86]. characters [Ano82-192]. charge [Gar81, LP80]. charged [Gar81]. charged-up [Gar81]. Charles [Hou85]. chart [Ano83-248]. charter [Lom80b]. charts [WM84]. Chasing [Jur89a]. Cheap [And80, Bac80]. cheaper [Bar81]. chemical [Lea85]. chemistry [Bot83]. Chernobyl [Fis86c, FZ87a]. chevron [LS81]. chief [Fis83d, Ler83b]. chills [Per82c]. China [XM89]. Chip [Ano81-42, CS87, Gut83c, McC85, Wal84a, Ano81-43, Ano81-187, Ano82-194, Bel86a, Cor87c, Dut80, Fis84d, Fis84g, Fis85f, Fis87d, Gut84h, GWF84, Gut84l, Gut85b, MAN80, Mur85, Rum85, SA85, Tri82, TP83, Wal87b, Zor85e]. chip-design [Gut84l]. chip-testing [SA85]. Chips [Bor83, AF87, Ant87, Bow85, Cor87c, Gut84g, Gut85b, Hor86b, HHH87, Mel80, Rau87, Rum85, Wal82c, Wal84e]. choice [For86, Kap85f]. Choices [Bel85f, Cle82, Dav83a, PW85a, Per87d]. choke [SF82]. choose [Mor81a]. Choosing [Bel88d, Mor81a, DD86, Per87d]. chores [Fis85e, VWZ86c]. chosen [Bal85]. Christmas [Wol85]. chunks [HM87]. circa [GGPY89]. circle [Chr86e]. circuit [Ant87, Bar81, Dut80, GWF84, Har83, LH87, PP88, RK80, Tul83]. circuitry [Ber80b, Per89a, San89a]. circuits [Bar81, Bel81, Ber80c, Ber81h, Ber81j, ELW83, Fis85f, FG86, Fis87d, Kap83b, Lea85, Ler82e, Mas80c, Mok81b, Rum85, SA85, Sum80, Tul83, Tur87, Wal84e, Wal87b]. Circumventing [Har83]. cities [Wei84]. citing [Per86]. City [Fis83b, Bag84, Ver86]. civilian [Den81, Pro81]. civilization [Cog84]. Clarke [Fis83g]. clash [Bel85b, Per84a]. class [Chr81b, MAN80, Zor84b]. Classic [Ano86-152, Tor84]. classics [Ano88d]. Classified [Ano80-69, Ano80-70, Ano80-71, Ano80-72, Ano80-73, Ano80-74, Ano80-75, Ano80-76, Ano80-77]. classroom [MD84]. claws [CC86b]. Clean [Lof80a]. Cleaning [DeV81]. cleanly [Ano86-66]. cleanup [Ada84c]. clear [Ste81, Tru82]. clearer [Fis84c]. clearly [Har83]. clerks [IK84]. climate [Kap83e]. clinch [Gut82b]. clips [Bor83]. clock [Ano82-194]. clones [Voe87i]. close [KS86b]. closely [Zor86b]. closer [Ear84, Hil85, Tru83e]. cloud [Ada87c, Ler80c]. clouded [Per82c]. clumsy [Har83]. cluster [Ano82-188]. cm [Ano82-193, Ano82-187]. cm/s [Ano82-187]. Cmos [Ful80, Ano81-184, Ber81j, Dav83a, Fis83i, Zor85e]. coal [Ano86-66, Fag80c]. coast [Ber81a]. cochlear [CP86]. cockpit [Ler83a]. cockpits [Ler83a]. code [Zor87a]. coded [Jur87c]. codes [Ano82-195]. Coding [Jay86]. coffee [Ano81-42]. Cogeneration [Fag80e, FZ85]. Cogenerators [FZ87a]. Coherent [BA86, LH87, Bel87b]. collar [IK84, Sha84, Sug80a]. colleagues [Fag80f, Per81c]. collection [Ano84-148, Ano84-147, Gev83]. college [Cor87c, Hor84b, Pyk87]. colleges [Per80b, Van84]. Collision [KLCPT80, Lom80d]. Collision-proof [Lom80d]. collisions [Bel86c, HTW84, Lom80d]. color [Ano81-177, Bal85, Fis87c, WM84]. colorized [Fis87c]. colors [Ano82-189]. Colossus [Zor87a]. Columbia [LM81]. combination [Bar81, Ler80a]. combinations [Dou81, Sti81]. combined [Cor88b, Gut84k]. combustion [Ano86-66]. come [Ano81-230, Ear84, Fis83f, Joh84, Moy80, Per87e, Ros80, San89f]. comeback [Ada84c, JCH80]. comes [Ful80, Hef84, Wil87]. comet [Ada87a]. comfort [Jur87b]. comfortable [Kap85d]. comforts [Kap85a]. Coming [Ano80-78, Ano80-79, Ano80-80, Ano80-81, Ano80-82, Ano80-83, Ano80-84, Ano80-85, Ano80-86, Ano80-87, Ano80-88, Ano80-89, Ano81-71, Ano81-72, Ano81-73, Ano81-74, Ano81-75, Ano81-76, Ano81-77, Ano81-78, Ano81-79, Ano81-80, Ano81-81, Ano81-82, Ano82-74, Ano82-75, Ano82-76, Ano82-77, Ano82-78, Ano82-79, Ano82-80, Ano82-81, Ano82-82, Ano82-83, Ano82-84, Ano82-85, Ano83-81, Ano83-82, Ano83-83, Ano83-84, Ano83-85, Ano83-86, Ano83-87, Ano83-88, Ano83-89, Ano83-90, Ano83-91, Ano83-92, Ano84-87, Ano84-88, Ano84-89, Ano84-90, Ano84-91, Ano84-92, Ano84-93, Ano84-94, Ano84-95, Ano84-96, Ano84-97, Ano84-98, Ano85-83, Ano85-84, Ano85-85, Ano85-86, Ano85-87, Ano85-88, Ano85-89, Ano85-90, Ano85-91, Ano85-92, Ano85-93, Ano85-94, Ano86-79, Ano86-80, Ano86-81, Ano86-82, Ano86-83, Ano86-84, Ano86-85, Ano86-86, Ano86-87, Ano86-88, Ano86-89, Ano86-90, Ano87-76, Ano87-77, Ano87-78, Ano87-79]. Coming [Ano87-80, Ano87-81, Ano87-82, Ano87-83, Ano87-84, Ano87-85, Ano87-86, Ano87-87, Jur86a, dS83, Ano84-168, Jur81a, Kap85e, Tul83, Wei87]. Command [Ler82j, FHW85, Voe87d]. commanders [Ada87a]. commandments [Tri80]. commands [Kap80a]. comments [Fei82]. commerce [Ber81e]. commercial [ABEI84, Cas81, Dav87a, FM84b, Fis86b, Kat80, Kec83, Ler83a, MRRS84, Spi86, VH86, Wet82a]. Commercialization [Per83b, Blu80]. Commission [Lom80e, Mas80a, Eis84, Ler81f]. commitment [Cur83, Per81e]. Committee [Gra82b]. Commodore [PW85b]. Common [Gut84j, Nas83, Ham86, Pan87a, Wal87h, Wer87a]. communicate [Ler82h, Vae86]. communicates [HT83]. Communicating [Bel84b, Del80, Far86, PHG86, Cor84, Kam86]. Communication [Mok82c, Per85b, BA86, Bel88b, Ber82b, Dor83, Gau85, GZ85, LP80]. Communications [Bel84c, Bel85a, BA86, Bel87b, Bel81, Ber82b, Cla80a, CN83, Del80, Dor83, Dou81, Dut80, EMM82, Gra82b, Har83, Hil83, Jac82, Kap81b, Kec83, Kir80, Lom81a, MNS81, Mel80, Mok80a, Mok81a, Mok81b, PK81, Rab81, Sha81a, Voe87d, Wet82a, Ada84a, Ada85c, Ano84-236, Bel85c, Bel88f, Che82, Cla80b, Cor84, Fei82, Fis84b, FHW85, Gib80, Gre85, Gut84k, HMDF80, Hec85, Hil85, HM87, Jac80, Kah84, Kuw84, Lem81b, Ler82j, Mok81h, Mok82b, Pan87a, Per80c, Pow84, ST81, Sti89d, Zor87a]. communicative [Sha80d]. community [Ano84-236, GP81, Hil83, Hon80]. commuter [Per85b]. Compact [Ano87-74, AM86, Asa87, Bru87, Che86, Che87, Cor88b, Ina89, Kir83, Miy84, OK86, Sol88, Tan86]. compactness [Spi86]. companies [Cor87a, Fis84g, Fle85c, Fre84, Mac84, MRRS84, Per83a, Per87f, Tru82, Zor84c]. companion [Ano81-183, Ber80d]. company [Bel87f, DTS86, Kid81, Per84b, Per83c, Per84a, Per84b, Per84c]. Compare [Per81c]. compared [Cle82, Gra84, JT81, Jur82b]. comparing [Mac84]. compatibility [Gut83g, Jur83b, Vae86]. compatible [Ano80-196, Ano80-194, Ano81-184, Ano83-185, OK86]. compete [Cor87a, Fis84h, FP86, Per84e]. competes [Dup82]. Competing [Flo82, Lem83a]. Competition [Sti88b, Chr83g, GBDC86, San88a, Sha81e, Wal86c]. Competitive [Fis86a, Fit87a, Kid81, Lof80a, Per84a, Voe86d, Voe87e]. competitiveness [Kap87a]. compilers [Ano87-234, Wal84e, Voe87a]. complaints [Wal84c]. complement [Gut84i]. complete [Fly87, Wal83b, Wal84a]. completely [Jur83a, Luc80]. complex [Ber81g, Che82, Fis84d, Fis86d, GGZ86, Pow84, Sug80b, Tul83, Ver86, VWZ86e, de 82]. complexion [Zor89j]. complexity [AW85, Sha82, Sta87b, Wal84c, WB86]. complicate [Ler84f]. complicated [Cor84]. Component [Per83d, Fag82b, MMB83, Mok82c]. component-manufacturing [MMB83]. components [Ano87-202, Bel87b, Ber82f, Ber82j, Fis83i, JJ87, Kap82c, KP86, Mur89b, Sha87, Tul83]. composition [Kap84b]. compound [Eas86]. compounding [Zor86d]. comprehension [Pic82]. comprehensive [Lom80c]. Compressing [Fit87a]. compression [EMM82, PHG86]. comprises [Ano82-194]. compromise [Fis84c]. computation [Ano83-243]. compute [Con86]. Computer [Ano81-230, Ano84-53, Bel81, Chr82f, Dav83b, DMW87, Gar84, Gra82a, KP86, Kid81, Ler81a, Ler82e, LS81, Mal84, PW83, PW85a, Swe82, Tal83, Tri82, AM86, Ada87e, Ano81-186, Ano82-193, Ano82-192, Ano83-143, Ano83-205, Ano83-206, Ano83-207, Ano84-194, Ano84-195, Ano84-196, Ano84-197, Ano84-198, Ano84-199, Ano84-236, Ano85-191, Ano85-192, Ano85-193, Ano85-194, Ano85-195, Ano85-196, Ano86-185, Ano86-186, Ano86-187, Ano86-188, Ano86-189, Ano86-190, Bee83, Ber83b, Ber80e, Ber82j, Cye84, DM85, DD86, DS81, Dou83a, Fag80b, FHRW^{+}83, Fit89c, Gev83, Gut83f, Gut83h, Gut84e, Gut84k, GB85, HT83, HOD82, Hon80, HK87, Hor84b, Jur82d, Kap84c, Kap85d, Kap87a, Kat87, Lem81a, Ler81c, Ler82g, Ler83e, Liu84, Mas85, MMB83, Mil87, Mok80b, Per83e, PW84, Per84e, SAPM80, Sug80b, Tru83e, Vae86, VWZ86e, Voe87b]. computer [Voe88a, Wal83e, Wal84b, Wer87a, Win84, Zde83, Zor84b, Zor87a]. computer-accessible [Bee83]. Computer-aided [Bel81, Ler81a, Ler82e, Swe82, Tri82, Ber82j, Kat87, Wer87a]. Computer-based [KP86, Gut83h, HOD82, Jur82d, Mil87, Tri82, Zor84b]. computer-controlled [Ada87e]. computer-game [Tru83e]. computer-generated [Per83e]. computer-science [HK87]. computerese [Gut84j]. Computerization [Sha84]. Computerized [KSB87, Van84]. Computers [Ano83-107, Ber80b, Ber80c, Ber80d, Ber81c, Ber81d, Ber81b, Ber82c, BKW83, Chi81, Chr83e, DS81, Fag80a, Fag81a, Fag82c, FZL86, Gev83, Her83a, Hon80, Kap80a, Kap83a, Lem81a, Ler80a, Ler81a, Ler81c, Ler81b, Ler82c, Ler82b, Mas85, SF82, Ste81, Sug80b, SW83, Swe82, Wal84b, Whi83, Zor85b, Ada87a, Ano82-171, Ano87-239, AMG^{+}85, BPEB84, Bor89, Cor87c, ELW83, FHW85, Fri84, Gut83g, Gut84g, Gut84j, Gut84i, Gut85b, Gut85d, Hef84, Hou85, Jur82d, Kah83, Lem81b, Ler82h, Ler84b, Liu84, Mil87, Mok80c, Mok80f, Mok81d, Mok82e, MD84, Pan87a, Per85c, PW86, SAPM80, Sta87a, SM81, Sti81, Tor83, Vae86, VWZ86c, VWZ86f, Voe87i, Voe88f, Voe89c, Wal84f, Wal85e, Wer87a, Zad84, Zor84b, ZJ86]. computers/calculators [Mok82e]. Computers/software [Kap80a]. Computertrends [Nai83]. Computing [Ber82c, Ano86-153, Bac82, Bel85c, Bel86b, Con86, Cor87d, Dou83b, HM87, Kah83, Zor85e]. concentrate [Sha81c]. concentrates [Fis83c]. Concentrators [Bac80]. Concept [JJ87, Mok80e, Per88a]. concepts [Gut83c]. concern [Flo82, Nai83, Per83g]. Concerned [Chr83f]. concluded [San88c]. concludes [Roi86]. condition [BD83, Fis86f]. conditions [Per87a]. conduit [Sti88d]. Conference [Kir80, Mok80d]. conferences [Ano87-70, Ano87-71, Ano87-72, Ano87-73]. conferencing [Mok80d]. confidence [Blu80, Din82, Mar85]. confident [Hor85d]. configuration [LS81]. configurations [Bel81]. confinement [Ler80d]. confining [KS86b]. conflicting [Kut83]. conforming [Ano84-247]. Confronted [Gra82b]. confuse [DMW87]. congestion [Mok80a]. Congress [FP86, Gre80, Lom80b]. conjunction [Miy84]. Connecting [Wet82b]. connection [Pan87a, WZ87]. connections [Ano83-143, Fit89b]. conquest [Jur86b]. cons [Ler81b]. consequences [HOD82]. conservation [Jur85a]. conservators [Fre86]. conserving [Fre86]. consideration [VWZ86d]. considered [Per81d]. consistent [KS86a]. constants [Wal87h]. constraints [Fis86a]. constructed [Ler84g, Lom80c]. construction [Jur85a, You84, de 82]. consultants [Ber80d]. Consumer [Cle82, Jur82b, Jur83b, Jur89c, Luc80, Mok80c, Mok81g, Mok81d, Mok81e, Mok81f, Mok81c, NS80, Per81a, PTW82, Per83d, Per83e, Per84d, Per85a, PC86, Per87b, Pin80, Sch81, SD81, Tru82, CC85, Dav87a, Jur88d, Ler82f, VWZ86e, Wal87d]. consumer-electronics [Per85a]. Consumers [Tru83e]. consumes [Ano81-184]. consumption [Dav83a]. contact [Cla80a, Wal84d]. contain [VWZ86e]. contains [Ano81-186, McC85]. contemplates [Fit88d]. contenders [Ano80-245, For86, Per85c]. content [Jur85b]. Contention [Gra82b]. Contents [Ano80-90, Ano80-91, Ano80-92, Ano80-93, Ano80-94, Ano80-95, Ano80-96, Ano80-97, Ano80-98, Ano80-99, Ano80-100, Ano80-101, Ano81-83, Ano81-84, Ano81-85, Ano81-86, Ano81-87, Ano81-88, Ano81-89, Ano81-90, Ano81-91, Ano81-92, Ano81-93, Ano81-94, Ano82-86, Ano82-87, Ano82-88, Ano82-89, Ano82-90, Ano82-91, Ano82-92, Ano82-93, Ano82-94, Ano82-95, Ano82-96, Ano82-97, Ano83-93, Ano83-94, Ano83-95, Ano83-96, Ano83-97, Ano83-98, Ano83-99, Ano83-100, Ano83-101, Ano83-102, Ano83-103, Ano83-104, Ano84-99, Ano84-110, Ano84-100, Ano84-101, Ano84-102, Ano84-103, Ano84-104, Ano84-105, Ano84-106, Ano84-107, Ano84-108, Ano84-109, Ano85-95, Ano85-105, Ano85-106, Ano85-96, Ano85-97, Ano85-98, Ano85-99, Ano85-100, Ano85-101, Ano85-102, Ano85-103, Ano85-104, Ano86-91, Ano86-92, Ano86-93, Ano86-94, Ano86-95, Ano86-96, Ano86-97, Ano86-98, Ano86-99, Ano86-100, Ano86-101, Ano86-102, Ano87-88, Ano87-89, Ano87-90, Ano87-91, Ano87-92, Ano87-93, Ano87-94]. Contents [Ano87-95, Ano87-96, Ano87-97, Ano87-98, Ano87-99, Ano88i, Ano80-153, Ano80-154, Ano80-155, Ano80-156, Ano80-157, Ano80-158, Ano80-159, Ano80-149, Ano80-150, Ano80-151, Ano80-152, Ano81-138, Ano81-139, Ano81-140, Ano81-141, Ano81-145, Ano81-146, Ano81-147, Ano81-148, Ano81-149, Ano81-142, Ano81-143, Ano81-144, Ano82-140, Ano82-141, Ano82-149, Ano82-150, Ano82-142, Ano82-143, Ano82-144, Ano82-145, Ano82-146, Ano82-151, Ano82-147, Ano82-148, Ano83-147, Ano83-148, Ano83-149, Ano83-150, Ano83-151, Ano83-152, Ano83-153, Ano83-154, Ano83-158, Ano83-155, Ano83-156, Ano83-157, Ano84-150, Ano84-151, Ano84-152, Ano84-153, Ano84-154, Ano84-155, Ano84-156, Ano84-157, Ano84-158, Ano84-159, Ano84-160, Ano84-161, Ano85-145, Ano85-146, Ano85-147, Ano85-148, Ano85-149, Ano85-150, Ano85-151, Ano85-152, Ano85-153, Ano85-154, Ano85-155, Ano85-156, Ano86-134, Ano86-135, Ano86-136, Ano86-137, Ano86-138, Ano86-139, Ano86-140, Ano86-141]. contents [Ano86-142, Ano86-143, Ano86-144, Ano87-153, Ano87-155, Ano87-156, Ano87-157, Ano87-154, Ano87-158, Ano87-159, Ano87-160, Ano87-161, Ano87-162, Ano87-163]. Contest [Ano84-50, Ano84-193, Bel85b]. Contiguous [LS81]. Contiguous-element [LS81]. continue [GS80, Gut84l, Her84, KP86, Tru83c]. continues [BA86, Mok82c]. Continuing [Her84, Lem81c, Mok81g]. continuous [Kat80, Red80]. continuous-speech [Kat80, Red80]. continuously [Jur87e]. Continuum [Ano83-105, Ano83-106, Ano84-115, Ano84-111, Ano84-112, Ano84-113, Ano84-114, Ano85-112, Ano85-107, Ano85-108, Ano85-109, Ano85-110, Ano85-111, Ano86-103]. contrasts [BMoF^{+}83]. contribute [Hor84c, Kap81a]. contributors [Ano84-86]. Control [HOD82, Ano80-195, Ano83-185, Bel89d, Ber81a, BD83, Chr82b, Chr82c, Chr82d, Chr86j, Fag81b, Fag82c, Fit87c, Gra82c, GT82, Hor86a, Jur86a, Jur86b, Jur87e, KA86, Ler81h, Ler82a, Ler82j, Lom80d, Lut88, MNS81, Mic81, Mur88a, Pea81, RK80, Ros83, STK82, Spi86, TBH87, Voe87d, Voe87g, Wal82b, Wal82a]. Control-room [HOD82]. Controlled [Ler80d, Ada87e, Ano81-177, Gro83, Jur83a]. Controller [Ano82-186, Ano80-196, Ano82-194, Jur82c, Kap87a]. Controllers [Per86, Jur81b, KP86, PW86]. Controlling [Fag80c, Mih81]. controllvi [Ano82-188]. controls [Bug84, Fag80c, Fag82c, Jur85b, Kap80c, Kap85b, Kap85d, Lom81c, Riv80, Tru84a]. controversy [Gre80, Ler80b]. convention [Gut83b, Ler84d, Tus86]. conventional [KM82, Lem83a, Mar86]. conventions [Ano87-70, Ano87-71, Ano87-72, Ano87-73]. conversations [Dut80]. conversion [Jur86c, MNS81]. conversions [Pin80]. convert [Per84e]. converted [Kap80a]. converter [Ano82-190]. Converters [Ber82f, Fit88a, FEK80]. converts [Wal85d]. Conway [Cor87c]. Coolidge [Wol84b]. cooling [Fis85d]. Cooperating [Cor87a, Wal86c]. cooperation [Ano87-225, Row82, Wal82b]. coordinating [Mas85]. cope [Esc86, Sha81c]. copies [WM84]. Coping [Per81b, Tru83a, Tul83, Jur88c]. copy [Voe87i]. Cordless [Ril87]. core [San88a]. corporate [Car83, Fis85b, Fle85b]. correct [BE87, Mos81, Voe87h]. Correcting [Sch81]. Corrections [Ano87-100]. corrective [Kah81]. corresponding [Mal84]. corridor [Zor89d]. corrosion [GBDC86]. cost [Ano80-41, Ano82-152, Bel84c, BD83, Car80, Cun80, Gel81, Gut84e, HMDF80, Jac82, Jur86a, Jur87a, Kap87a, Ler84g, Mel80, Mid80, Mok82e, Rau87, Red80, Voe87g, Wal87d, Wer87a]. cost-effective [Jur86a, Wer87a]. Costly [Aml82]. costs [Ano86-66, Bel81, Bug84, CC86a, Fag80e, Fly87, Gau85, Her83b, Kap82d, Kap82c, Kap83e, Kap85d, Mas85, Pow84, SA85, Voe86d, Wal85e, Whi83]. could [Ano84-236, Bug84, GP81, Lem81b, Ler81f, ST81, TD84]. couldn’t [Chr80l]. count [Mok82c]. counterbalancing [Nai83]. counterinvasions [Kap85c]. countermeasures [Ber82d, Ler82j]. counterparts [Tri82]. Counting [Ada86a]. country [de 82]. course [Ada84b]. courses [Bal84a, Pyk87]. Court [Per82c]. cousin [KS86b]. cover [Ada87c, Ano82-42, Ano82-134, Ano82-154, Ano82-155, Ano83-50, Ano83-142, Ano83-159, Ano83-160, Ano87-31, Ano87-29, Ano87-30, Ano87-32, Ano87-33, Ano87-34, Ano87-35, Ano87-36, Ano87-37, Ano87-38, Ano87-39, Ano87-40, Ano87-43, Ano87-41, Ano87-42, Ano87-44, Ano87-45, Ano87-46, Ano87-47, Ano87-48, Ano87-49, Ano87-50, Ano87-51, Ano87-52, Ano87-127, Ano87-128, Ano87-129, Ano87-130, Ano87-131, Ano87-132, Ano87-133, Ano87-134, Ano87-135, Ano87-136, Ano87-137, Ano87-138, Ano87-139, Ano87-140, Ano87-141, Ano87-142, Ano87-143, Ano87-144, Ano87-145, Ano87-146, Ano87-147, Ano87-148, Ano87-149, Ano87-150, Ano87-234, Ano88e, Ano88f, Ano88l, SFV^{+}88]. coverage [SD81]. covers [Chr83d]. CPU [Ano81-186, Ano82-194]. Crack [Mas80b]. Crack-detectives [Mas80b]. cracks [Kah82]. craft [Ano81-45]. crammed [DJ85]. crashes [Zor86a]. Cray [WZ87]. crazy [Jur88c]. create [Bel84b, Bel88b, Cor87d, JI84, Per81e]. created [AJ85, Per80c, Tri82]. creates [Ano83-162, Bot83]. Creating [Rei80, Sha87, Tru84b, Wil80]. creation [GZ85]. creative [Bel87f, PW83]. Creativity [Ber82h]. creators [Sha81f]. credibility [Per82b]. credit [Wei84]. credited [Wol83]. creeps [Gol81]. crew [Ler83a]. crime [PW84]. crippler [Lem81b]. crises [Hor85d]. crisis [Kah81, Per81b]. critical [Ver86]. crossed [Fit89b]. Crossroads [Ler83c, Gib80]. crowd [Per85b]. crowded [Zor85f]. CRT [Bal85, Fle82, Jur81a, Tan89b]. CRTs [Apt85, Bal85, Tan89a]. crucial [Fit87a, Hor86a, Ler81g]. crunching [Gut83d]. Cryptology [Kah84]. crystal [Dou81, Jur81a, Per85c]. crystallizing [Con86]. crystals [Tha83]. cues [Sar82]. Culham [Fis84h]. culture [Ler84e]. curator [Ano84-148]. curmudgeon [Nah83]. current [Ada84b, Ano80-153, Ano80-154, Ano80-155, Ano80-156, Ano80-157, Ano80-158, Ano80-159, Ano81-145, Ano81-146, Ano81-147, Ano81-148, Ano81-149, Ano82-149, Ano82-150, Ano82-151, Ano83-158, Ano87-155, Ano87-156, Ano87-157, Ano87-158, Ano87-159, Ano87-160, Ano87-161, Ano87-162, Ano87-163, Che87, Fis83f, Fis86e, Wal83a, Zor84a, Lom81c]. currents [HW83]. curricula [Cye84, Tru83a, Zor84b]. curricular [Kap84b]. curriculum [Jur87c]. curtail [Ler84f]. curve [Cun80]. customer [Bel87c, Dav87a, STK82]. customers [Bel84c, Gau85, Gel81, Zor84c]. cut [Gau85, Jak85, Kap81c, Kap82d, Kap85d, Mas85]. cutting [CC86a, Kap87a, Zor89a]. cybernetics [Ano82-249]. cycle [Fit87a, Jur86c, Ler81g, Whi87]. cylinder [Ano80-43].

D [Mus85, Wol84b, Ano81-177, Ano82-190, Mok81g, Per84c, Pin80]. D-size [Ano81-177]. D/A [Ano82-190, Pin80]. daily [Sha80f]. Dallas [Per85d]. Damaging [Ano81-151]. darkness [Ada87c]. Darpa [Lem83b]. dashboards [Jur81a]. Data [Ano83-107, Cas81, Hon80, Jur83c, KG83, Ler84b, Mer83, Mok81h, Ano82-189, Ano84-168, AJ85, BCUY82, Bee83, Chr80i, Chr82h, Cye84, Fis84g, Jur82c, Kut83, Ler84f, Pan87a, Par84, PHG86, SF82, Sha80c, Sti89d, TV89, Kid81]. data-capture [Sha80c]. Data-driven [Ano83-107, Jur83c, KG83, Mer83]. Data-flow [Ler84b]. database [Che86, Mil87]. daunting [Per87a]. day [Dou83a, Fis84d, KM82, Ros83, Sha80f]. days [Chr87l, GWF84, Ros83, Wal84a]. Dazzling [LH87, Per87e]. Dead [Chr82g]. deaf [Fle82, Sar82]. deal [Ber83a]. dearth [Gut84f]. death [Chr84h]. debate [Gut82i, Zor87f]. debates [Ano86-152, Flo82, Ler82k, Tru84b]. Debating [AH85, Ber82b]. debit [Wei84]. debris [KLCPT80]. debug [Swe82]. Debut [Per87b]. debuts [Jur83a]. DEC [Ano80-196]. decade [Dav83a, Del80, Fis83c, FZ87b]. decades [Bal84a, Con86, Ryd84, Wei87]. deceived [BJ88]. deceiving [Mok81g]. decentralized [Ler84b]. decided [Jur87a, Wol85]. deciphered [Zor87a]. decision [AJ85, Bel85b, HOD82]. decisions [Ler83b]. decisive [Aml82]. declared [FW87]. decline [FZ87a, Per81b]. declining [Ros80]. decontamination [DeV81]. decrease [Mok82c]. deemed [Lem83a]. deep [PHG86]. deeply [Bel85b]. defeat [Box89]. defect [Chr86b, Ham86]. defend [Hor86d]. Defense [Gan82, WB86, Ada87a, Ada87e, Ada88d, Bel89a, DeL82, Fei82, Per89b, Sti89b, Wal87g, Ada87e, Wal85d]. defenses [Ada87f]. Defensive [Bel87c, Fit87c]. defibrillators [CP86]. defined [Fis84c, Vae86]. defines [MMB83]. defining [Cla80b, Dav87a]. definition [Jur83b, Jur88a, Mok81i, Roi86]. defueling [Ada84c]. degree [Kap84b, Ros84, Smi87]. degrees [Bal84a, Bal84b, Kap84b]. delayed [Per87b]. delays [Per86]. deliver [Bag84, STK82]. delivered [Bal84a, GWF84]. delivering [Kap83d]. Demand [Gel81, Fis83f, Fis84f, Fis86a, Kap80f, Kap84a, WM84]. Demand-side [Gel81]. demise [Chi81]. democratization [Wit84]. demonstrate [Per81e]. demonstrated [And80]. demonstrations [Ada84a, Fis87b]. denied [Har83]. denser [Ber82j]. densities [Chi81, Whi83]. density [Ber80b, Chr81l, Mel80, Miy84, MR82]. Department [Ada87e, Wal85d, Car83]. departments [Van84]. depend [Bel86a, Ber80b, HT83]. dependability [VWZ86d]. Dependent [HK87, CA82]. Depending [For86]. depict [BJ88]. deployment [War82]. depth [EG82]. derailed [KW86]. Deregulation [Mun88, Voe86d, BA86, Fis86a, FWVH86, Ler84f, O’C86]. describe [GWF84]. described [Ano84-146]. Design [Fis84b, FW82, Gut82c, Gut82h, Gut84e, Ler81i, PW83, PW85b, Sha81e, Sha81f, Wal83b, Wal84a, Wer87a, Wil81, AW85, Ada88c, Ano82-188, Ano83-186, Bee83, BE87, Bel81, Ber82h, Ber82j, BKBC85, Chr80j, Chr86b, Dav87a, Dav83a, DD86, Fag80a, Fag80b, Fis85f, Fit87a, Gra82a, Gut84h, Gut84l, HT83, HOD82, Jur84, Kap87b, Kat87, Kid81, Ler83b, Mel80, Mil87, MW88, Mur89d, Nag83, PP88, Ros83, Sha80a, Sha80d, Smi87, SD88, Tri82, TD84, Tru83d, Tul83, Tus86, Voe86a, Voe86b, VWZ86d, VWZ86e, WB86, Wal87d, WSC89, Whi87, Zde83, Zor84b, de 89, Ano82-152]. design-to-manufacturing [Wal83b]. Design/manufacturing [Wal83b]. designed [Per81a, Per83d, Per85b]. designer [Ber81g, Kap85f, Per87c, Voe87a, Zor85a]. Designers [Dav87a, Lem81a, Smi87, Tan86, Ant87, Ber82b, Blu80, CS87, CC87, DD86, Fag81a, Gut83c, Gut84e, Hor86d, KSB87, PW83, PW85a, Rau87, Sha80a, Sha84, Sta87b, Swe82, TBH87, VWZ86a, VWZ86c, VWZ86d, Wal84a, Wal85e]. Designing [FHW85, JJ87, SA85, TBH87, Wal83a, Whi87, Fis84h, Gut85b, Kid81, PV89, Sha87]. designs [Bel87c, CR88, Cor87c, Fag82b, Fis87b, Lla81, McC85, Swe82, Tri82, Voe87j]. Desktop [Jur87a, Mil87, Voe87i]. desperately [GP81]. Despite [Chi81, Lem81c, SW83, Wal86c, FP84, Hor86d, ST81]. destroys [Ada85a, AW85]. details [Voe87e]. detect [Ada85a]. Detecting [Ada87c, Mas80b, GBDC86]. detection [BFB87, LH87]. detectives [Mas80b]. detector [Dou81]. detectors [BCUY82, For86, Jur81b, Tha83]. deteriorating [Kah82]. determine [Bel85f, Zor86b]. deterrent [Hor86d]. Detroit [JCH80, Jur81b, Jur83a, Jur86a, Jur84, Jur85b, Jur86b, Jur87b, Jur88b]. Deutsches [Ano84-146]. developed [Aml82, Bel87b, Cas81, Gut83i, Har83, Jur80, Kah84, Kir80, Wal82b]. developers [Ano86-152, HM87, Mac88, Sha81e, Sti89c]. developing [Ada87e, Ber81k, Mas80c, SF82, Wal82b]. development [Ano80-197, Ano81-183, Ano83-187, Bac82, Bal81, BKBC85, CC87, Cun80, DTS86, Fag80c, FZL86, HOD82, Her83b, MNS81, Pak85a, Pro81, WB86, Zor84c]. developmental [Spi86]. developments [Bel87e, CP86, Jur86d, Kap80e, Per84d, PC86, Roi86, Wil80, ZJ86]. device [MS84]. devices [Ano80-249, Bal81, CS87, Con86, CP86, DGZ88, Fag80a, Fag81c, Fag82b, GA88, Kap81c, Ler82b, MAN80, PC80, Pic82, Rum85, Sha82, ST81, Sta86, Wal87d]. devices/systems [Kap81c]. devise [GP81]. devising [Gut85b]. diagnose [Sha81d]. diagnoses [Fis83e]. diagnosis [Jur82c, Wel84]. diagnostic [Bot83, Cor87b, Jur80, Lla81]. diagnostics [Gar84, Jur83a, Jur85b, Jur86a, KA86, Riv87]. diagram [Jur82c]. Diana [Cla80a]. Diapers [Ano82-157]. Dichotomous [Lom81c]. did [Per82a, Sta87a]. didn’t [Chr82m, Nah83]. diesel [JT81]. Diesels [JT81]. differ [BS81]. difference [Ber83a, Per84d]. different [Bac82, PK81, SS84]. difficult [Ham86, Jur86d, Ler80a, Ler82b, Ste81, Wil87]. difficulties [ELW83, FP84]. dig [Ada89a]. Digging [GB85]. Digital [Ano80-193, Ano80-192, Ano81-181, Ano81-182, Fit88a, GA88, Gut84k, Kuw84, Ler83d, Miy84, Pan87a, Per87b, Sha80c, TV89, Wer87b, AF87, AJ85, Ber80a, Che86, Dor83, EMM82, Kap80a, Ler83a, Luc80, Lut88, Mel80, Mok82c, NS80, Pin80, Riv80, Sha80d, Spi86, Fis87d]. Digitized [Sti88c]. digitizing [Chu88]. dikes [Fis86d]. dilemma [Wal82b]. dilemmas [Per81c]. dimensional [Bro81b, Cor87b, PHG86]. Diminishing [Luc87a]. dinosaur [Ada89a]. dinosaurs [Her83a, Luc87a]. diode [Ano84-253, For86]. diodes [Bot85, KK84]. DIP [Bow85]. direct [Bac80, Bel84c, Bel84d, Fis83f, Hil85, PK81, Wat87]. direct-broadcast [Bel84c, Bel84d, PK81, Wat87]. direct-current [Fis83f]. directed [Ada87e, Car83]. Directions [VWZ86b, Tru84b, Wal85c]. director [Ler84h, Pak85b, Wol83, Zor89e]. dirty [Fis85e]. disabled [FP84]. disadvantaged [Ano80-40, FP84]. disappear [WZB86]. disappointed [Ada84c]. disarmament [Ler82k]. disaster [Ada87f, BE87, Esc86]. disc [Cor88b]. discern [Cas81]. discharge [Box89]. discipline [Kap84a]. disciplines [Kap84f, Wal83e, Wal84b]. disclosures [Sha80e]. discourage [GP81]. discouraging [Pro81]. discover [Fis87a, IK84]. discovered [Ano86-152]. discoveries [Ano88d, Ano88m]. discovery [Kap81a]. Discrete [Fag82b, GGZ86, Fag81c]. Discrete-event [GGZ86]. discrimination [Chr80a]. discuss [Gut83f, Her84]. discussed [Bel84e, Dan82, SS84]. discussion [Wal83e]. discussions [Ano80-68]. dishes [Bag84]. disk [Asa87, Che86, Chi81, Mok82f, OK86, Per83e]. disks [Ber80b, Bru87, Cle82, Fre88b, Miy84, VW86, Voe87j, Whi83]. Disney [GT82]. Disneyland [Sha80e]. displace [GZ85, LS81, Tan89a]. displaced [Luc80]. display [Fle82, Sil89]. displays [Ano82-189, Apt85, Bal85, Jur81a, PW85a, Per85c, Tan89a, Tan89b]. disposal [Lom80f]. dispute [Mai83]. dissipation [ST81]. distance [Bel84c, Bel85a, LH87, LP80, Sha84, de 82]. distances [Voe87d]. distortions [Sch81]. distracting [Per84c]. distribute [Cor87a, Fis84f]. distributed [Ano83-243, Mok82c, Sha81f]. distributed-processing [Sha81f]. distribution [Ano87-234, Ber81a, Che82, Kap81c]. divergent [Cla80b]. Diverse [Kap84f, Che82]. divest [Bel85b]. divestiture [Bel85a, Chr85b]. divide [HM87]. division [Ano81-180, Ler84h]. dizzying [WZ87]. DMA [Ano82-194]. Do [Car84, Ano82-220, Ano82-248, DD85, Fis85e, Fis86e, Fre84, Jur87a, Kat87, Mai83, Mok82e, Swe82, VWZ86c, Zor87e, Per83a]. ’Doc’ [Bel89b]. doctor [Bug84, Wal86b]. DOD [Fis84b, Lem83b, Lom80a, Voe87a, Cor87d, Fei82]. DOE [Fis83d, Ler84h]. does [KP86, Per83a, Per84b]. dog [SW88]. dollar [Fis85b]. dollars [JCH80]. dolphins [Hor86b]. dominant [Fle85c]. dominate [Fis84e, FG86]. dominating [Bel85c]. domino [FAEH84]. Donald [Fis83d]. done [Chr80l, Fis87c, Gut83i, Mih81, Rab81, Wal84e]. Don’t [Chr85c]. doped [Wol80]. Dougherty [Dou85]. Dounreay [Blu80]. down [Ano88p, Fis87b, FW87, Fly87, Liu84, Tal83, Wal85e, Whi83, Zor85b]. downtime [Ber80e, Kap80e]. dozens [Ler82b]. drafting [Ano82-187]. DRAM [San88a]. drama [Fru83]. Dramatic [Hil85, FZL86]. dramatically [Pea81]. drastic [CC86b]. draw [KP86]. drawing [Wet82a]. draws [Ano82-191]. dream [Moy80, Per85b]. dreams [Ler81c]. drill [Ril87]. drip [Ano81-42]. drive [Ano80-196, Ano82-191, Ano85-125, CC86b, Jur82a, Ler84c]. Driven [Kap81a, Ano83-107, Bel87e, Her83c, Jur83c, KG83, Mer83]. Driver [Jur87b, Ano80-196, Bel87e, Her83c, Jur81a, Jur82a]. Driver-friendly [Jur87b]. Drivers [Jur82a]. driving [Bha85, Gut84h, Whi83]. drudgery [BKW83]. drug [Hor85b]. dry [Wol89]. dual [Ano81-180, Bow85, Gut89, Log83]. dual-in-line [Bow85]. Dubrovnik [Roi86]. due [Fre84, Sti81, Tul83]. duress [Ler82i, Wal87c]. During [Bha85, And80, Bel86c, Gut84l, Jur85a, Kah84, Kap81c, PW86, Voe86a, Voe86b, Wol83]. duties [Ler83b]. duty [Chr86f]. Dynamic [Kat80, Ber81k, Mer83]. dynamically [RK80].

each [Ano82-186]. eagerly [Per85d]. Eaglecrest [Per85b]. ear [Bel88e]. earlier [Voe87h]. early [Ano85-125, Jak85, Mok80f, Sug80a, Tha83, Whi87, Zor87a, Ear84]. earn [Bal84b]. earned [Bal84a]. Earth [Ano83-248, HMDF80, Hos83]. Earthquake [Bel88e, Ver86]. ease [Gut83i]. easier [Ler82e]. easily [KSB87]. Eastern [Lom81c]. easy [Gut84e]. easy-to-use [Gut84e]. eat [SW88]. eating [Zor86d]. echo [AF87, Dut80]. echo-killer [Dut80]. echoes [Dut80]. Eckert [Ste81]. economic [Tor80, Wet82b]. economical [Bro81b, Dor83]. economically [Fis86f, MS81]. Economics [FG86, Wal82a, Fis87b, Ler80c, Mok80d]. economy [Jur82a, Kap85d, Riv80, Sha81c]. edge [Cor87d, Per84a, PHG86, Wal82d, Wal86c, Zor89a]. Edgerton [Bel89b]. Edison [JI84, Jur85a, Ril81, Wis82]. Editor [Fit87a]. editors [Bag84]. educated [Mac84]. Education [Bag84, Ber83a, Che89b, Chr85d, Cye84, Gut84a, Hor84a, Kah81, Kah82, MD84, Per81b, Tor84, Wit84, Bal84a, Ber83b, Chr80d, Chr80e, Gra84, Gut83f, Gut84i, Her84, Jur87c, Kap84b, PZ86, Ros84, Ryd84, Wol88b, Zor84a]. educational [Bel85d, Gra84]. Educator [Bry84, Ler84e, Ryd84]. Educators [Gra82c, Fri84]. Edutainment [MD84]. EE [Ano87-111, Bel89b, Car84, Gra84, Gut84b, Kap84b, Per80b, Ros84, Ryd84, Tru84b, Wal84b, Zor84a, de 83]. EES [Tru84b, Ada84b, Ano80-68, Ano80-102, Ano80-103, Ano80-104, Ano80-105, Ano80-106, Ano80-107, Ano81-95, Ano81-96, Ano81-97, Ano81-98, Ano81-99, Ano81-100, Ano82-98, Ano82-99, Ano82-100, Ano82-101, Ano82-102, Ano83-108, Ano83-109, Ano83-110, Ano83-111, Ano83-112, Ano83-113, Ano83-114, Ano83-115, Ano83-116, Ano84-116, Ano84-117, Ano84-118, Ano84-119, Ano84-120, Ano84-121, Ano84-122, Ano84-123, Ano84-124, Ano84-125, Ano84-126, Ano84-127, Ano85-113, Ano85-114, Ano85-115, Ano85-116, Ano85-117, Ano85-118, Ano85-119, Ano85-120, Ano85-121, Ano85-122, Ano85-123, Ano85-124, Ano86-104, Ano86-105, Ano86-106, Ano86-107, Ano86-108, Ano86-109, Ano86-110, Ano86-111, Ano86-112, Ano86-113, Ano86-114, Ano87-101, Ano87-102, Ano87-103, Ano87-104, Ano87-105, Ano87-106, Ano87-107, Ano87-108, Ano87-109, Ano87-110, Ano87-112, Kat87, Nah83, Van84]. effect [Che87, Gra82a, Gut83i, Ler84a, Voe89a]. effective [Jur86a, Wer87a]. effectiveness [Fit87a]. effects [And80, BKW83, Esc86, Gar81, Ler80b, LP80, Mor81b, WAO85]. efficiencies [Jur86c, Kap80e, Lof80b]. efficiency [EMM82, FZ85, Jur87d, Wal87a]. efficient [Ber82b, Fag80b, HM87, JCH80, Jur84, Kir83, SA85, Wal83c, Wol80, Zor85d, Kap81c]. efficiently [Ano86-66, Fis83f, Fis84f, Kel87, STK82, Whi87]. efficiently-produced [Kel87]. effort [Wal84a]. efforts [Gra82c, Kap82b, Mai83, Mal84]. eight [Ano82-187, Ano82-186]. either [Nah83]. EJAO [Ano82-187]. EJAO-size [Ano82-187]. Elaborate [Mai83]. elasticity [MR82]. election [Per84e]. Electric [Ano84-193, Fis86a, Ano85-125, Ano86-66, Ber81a, CA82, Cog84, Hor84d, HHH87, Kap80b, KW86, Lur84, STK82, TD84, Ver86, Wis82, Zor86d, CC86b]. Electrical [Gar81, Ano84-86, Ano84-148, Ant87, Gol81, JI84, Jur86c, Jur87c, Jur88e, Kah81, Per80b, Per89b, Tru83b, Van84]. electricity [And80, Fis84f, Fis84h, Fis86d, Gau85, Lur84, MS81, Mun88, You84]. electrics [Dup82]. electrification [You84]. Electro [Gut82f, Gut83b, Ler84d, Sha80f]. electrohydraulic [Fis83b]. Electroluminescence [Tan86]. electroluminescent [Jur81a, Per85c, Tan86]. electrolytes [MO89]. Electromagnetic [Chr87j, Lem81b, Ler80b, Tel82, KM82, Ler81f, Ler84a, Ler84c, SH85, Wel85, WAO85]. Electromagnetics [Kap84a, Ano87-234, KZG^{+}84]. electron [Bel86c, Bro81b, Eid81, Gib85]. electron-beam [Bro81b, Eid81]. Electronic [Ber82d, BS81, Hef84, Hor85a, JT81, Kap80c, Per88b, Ano84-148, Bal84b, Ber80d, Ber81d, Ber82b, Chr87i, Fag80c, Gar84, Jur81a, Jur83a, Jur85b, Jur86b, KSB87, Kat87, Lem83a, Luc80, Mok80a, Mok80e, Mok82f, PTW82, PFMP88, Riv80, Sha80b, Sha81e, Tru83e, Wal83e]. Electronically [Lem83a]. Electronics [Fag80a, Jur80, Ler82d, Sha80e, SAPM80, SPM^{+}80, Voe88c, Wal83c, Ano81-236, Ano84-53, Cle82, CP86, Cor87b, Cor88a, DeL82, Fit89e, Fle82, GA88, Gut84i, Hor84d, Hor85b, Hor85e, JT81, Jur81a, Jur81b, Jur82a, Jur82b, Jur82c, Jur82d, Jur83a, Jur83b, Jur84, Jur88d, Jur89b, Jur89c, Kap84c, Kap85b, KP86, Kap87a, Kap88, Lem81b, Ler81f, Ler82f, Ler84d, Lla81, Luc80, Mok80c, Mok80f, Mok81g, Mok81d, Mok81e, Mok81f, Mok81c, Per81a, PTW82, Per82e, Per83d, Per83e, Per84d, Per85a, PC86, Per87b, Per87f, Pic82, Riv80, San89d, Sar82, Sch81, SK89, SD81, Spi86, Tru82, Tru83b, Voe87d]. electronics-aided [Jur89b]. electrons [Ano86-184, Eas86]. electrostatic [Box89]. Electrotechnical [Mas80a, Ada87f, Bel87e, Kap84f]. electrotechnological [Ano84-147]. Electrotechnology [Bel87e, Fre86, Kap80c, Ano84-50, Cog84, de 83]. elegant [Emm85, Har83]. element [LS81, PW86]. elements [GGZ86, HM87, Voe87j]. ELF [Ano81-233]. elicits [Wal83e]. Elihu [Car83]. eliminated [Dut80]. eliminating [Ler81g]. else [Ler84h]. elsewhere [AH85, Hor84c, Par84]. elude [Sti81]. elusive [Gut83g, Ham86]. elusiveness [Mos81]. emanates [Mok82a]. embedded [Voe87a]. embrace [Jur82a]. embraces [CC86a]. emerge [Her83b]. emerged [Wis82]. emergency [Fis86c, Gar84, Zor87b]. emerges [Ano80-68, Per80b]. emerging [Bel84d, Gev83, Hor84b, Voe87j]. EMI [WAO85]. emission [JT81]. emphasis [Ber80c, Cor87d, Wal84b]. employ [Van84]. employees [Bel87f, Fle85c]. employer [Bel87f, Per81d]. employing [Cal82]. employment [CC86b]. EMPs [Ler81f]. emulate [WZ87]. emulation [Ano80-197]. emulator [Ano81-183]. enabled [Sha81f]. encoding [EMM82]. encountered [Voe86a, Voe86b]. encounters [Gol81]. encourage [Gel81, Her83b, Ler84a, Zor84c]. Encouraged [SH85]. encourages [Bug84, Voe86d]. end [Ano83-184, Chr82g, Dav83a, Fre84, Mil87]. Endangered [Mur88b]. Ending [Sim83]. endowed [VWZ86c]. enemy [Ada87c, Fit87c, Zor87a]. energize [Zor88a]. energized [PZ86]. Energizing [XM89]. Energy [Ano80-108, Ano80-109, Ano80-110, Ano80-111, Ano80-112, Ano80-113, Ano80-114, Ano80-115, Ano80-116, Ano80-117, Ano80-118, Ano80-119, Ano81-101, Ano81-102, Ano81-103, Ano81-104, Ano81-105, Ano81-106, Ano81-107, Ano81-108, Ano81-109, Ano81-110, Ano81-111, Ano81-112, Ano82-103, Ano82-104, Ano82-105, Ano82-106, Ano82-107, Ano82-108, Ano82-109, Ano82-110, Ano82-111, Ano82-112, Ano82-113, Ano82-114, Ano83-117, Ano83-118, Ano83-119, Ano83-120, Ano83-121, Ano83-122, Fag80b, Per80a, Ang80, Ano83-244, Bel84a, Ber81a, Bro81a, CA82, Che82, Chr80j, Dou85, Fag80e, Fis83d, Fis83c, Fis83f, Fis84f, FZ85, FZL86, FZ87a, Gel81, Hic83, HTW84, Jur85a, Jur86c, Kap80e, Kap81c, Kap82d, Kap82c, Kap83d, Kap83e, Kir83, Ler80c, Ler80d, Ler84h, Lom80f, Lom81b, MS81, Pro81, RK80, Ril81, Rog81, Row82, STK82, Wal83c, Wet82b, WZ88, WZ89, de 82, Fis83d]. energy-conservation [Jur85a]. Energy-efficient [Fag80b, Wal83c]. energy-independent [Ang80]. energy-management [FZ85]. enforcement [Fis80, Ler83e]. engagement [Fle85b]. engine [Ano83-169, Jur85b, Kap80c]. Engineer [Fag80f, Per87a, Ada87d, AFG87, Ano80-120, Ano80-121, Ano80-122, Ano80-123, Ano80-124, Ano80-125, Ano81-113, Ano81-114, Ano81-115, Ano81-116, Ano81-117, Ano81-118, Ano82-115, Ano82-116, Ano82-117, Ano82-118, Ano82-119, Ano82-120, Ano83-123, Ano83-124, Ano83-125, Ano83-126, Ano83-127, Ano83-128, Ano84-128, Ano84-129, Ano84-130, Ano84-131, Ano84-132, Ano84-133, Ano85-126, Ano85-127, Ano85-128, Ano85-129, Ano85-130, Ano85-131, Ano86-115, Ano86-116, Ano86-117, Ano86-118, Ano86-119, Ano86-120, Ano87-113, Ano87-114, Ano87-115, Ano87-116, BJ88, Bha85, Fag81a, Fag82a, Fre84, Gut84i, JI84, Ler83b, Sha80a, Tru83c, Tru83d, Wal84d, Wal84f, Wol84b]. Engineer-manager [Per87a]. Engineering [Chr80d, Chr80e, Chr86d, Per81b, Tru83a, VWZ86c, Wal84c, Wei80, Ano80-245, Bel84b, Bel84f, Bel89c, BE89, Car83, Fag80a, Fag80d, FHRW^{+}83, Fis83b, Flo82, Gol81, HT83, Hil86, JI84, Jur87c, Jur88e, Kah81, KZG^{+}84, Kat87, Ler83b, Mur88b, Per80b, PZ86, Pyk87, Riv87, Sti88a, Swe82, Tru83c, Van84, Wal84f, Wer87a, Yeh83, Zor89a, Ano81-133]. Engineers [Bal84b, Ber83a, FP84, Hor86b, Kah84, Per82d, Per84b, Tri80, Ano80-68, Ano84-86, Ano87-208, Ano88j, Bal84a, Bel87c, Chr83f, DTS86, Del80, Fis84h, Fre84, Get82, GT82, Her84, Hor84d, Lep81, Mac88, Per81a, Per83g, Per84c, Per87d, Per89b, Ros88, Tru84a, Ver86, Voe88a, Wal83e, Wal84c, Fis87d, Zor87e]. Engines [Jur83a, JT81, Jur87e]. England [ABEI84, Fis84h]. enhance [Ano81-185, Kap81a]. Enhanced [Gut83d, Lla81]. enhances [KSB87, KG83]. Eniac [Ste81]. enjoy [Ber80g]. enjoys [Cau82]. Enlarging [Chr86e]. enormous [Wel85]. enough [Nag83, Pea81]. enrolled [Bal84a]. enrolling [Bal84b]. ensure [HM87, Wal82d, Wol84a]. ensured [Hil83]. ensuring [Mih81]. enter [Hef84, Tru82, Tru84b, Wet82a]. entering [Per87e]. enterprise [Kut83]. enters [Fis87d]. Entertain [Luc84]. Entertainment [GT82, Lem83a, Jur82a]. Entertainment/control [GT82]. enthusiastic [IK84]. entice [Gut83b]. entire [Fis85f]. entrenched [Bel84d, Gut83c]. Entrepreneur [Wol84a, Per82a, Pin85]. entrepreneurial [WZ85]. entrepreneurs [Per82d, Ste81]. environment [Fri84, Kah82, Per81b]. Environmental [Tus86, Bel84f, Fis86f]. environments [Ada85c, Ber81e, KS86b]. envisaged [Lom81b]. envision [Mok81g]. Epcot [GT82]. Epicurean [Fle85a]. epitaxy [PC80]. EPRI [Fis83c]. EPROM [Ano82-193]. equation [Sti88b]. equations [Cen89, Emm85]. equipment [Bel88d, DD86, DS81, Fag80a, Fag82c, Gut83h, Hou85, Mok82b, Mok82f, Sha81b, Sha81d, VB86, Wal85e, Zor85f, Zor86d]. era [Hef84, Kap83b, San88a, Voe88a, Win84]. erasable [Fre88b]. Ergonomics [Jur87c, Fis83a]. Eric [AMG^{+}85]. error [Ada87f, Ano84-53, Ber81g, Mil87]. errors [BE87, Ham86]. ESA [Gib80]. escalating [Kap82a]. eschews [Mcc88]. ESS [Tus86]. essay [Dan82]. essays [Dan82]. establish [Fag81a, Jac82]. established [Cla80a, Kap84a, VWZ86d]. Establishing [Bee83, Mok81h, FM84a, Mor81b]. establishment [Bug84, Nah83]. esteem [Per81d]. etching [Lea85]. Ethernet [Ano83-184]. ethical [Per81c]. Ethics [Chr87a, Per81c, Per81d, Chr81d, Flo82]. Europe [Ang80, BMoF^{+}83, FW87, Kap80d, Lom81c, Moy80]. European [FZ87a, Hon80, Ler84f, Moy80, Nas83, Zor86d]. Evacuation [Zor87b]. Evaluating [Chr83g]. evaluation [DS81, Tal83]. eve [Mas80a]. Even [Per87e, SS84, Tru82, Zor87f, Ber81g, Fag80e, Fly87, GT82, Hic83, HHH87, Ler84b, Van84, Wel84, Win84, Wit84]. event [Fit88g, GGZ86, Zor89i]. Events [BE87, Ano88d, Ano88m]. eventually [Fis83b]. ever [Ano80-247, Ano80-248, Ano80-249, Ano81-231, Ano81-232, Ano81-233, Ano81-234, Ano82-171, CC87, Dav87a, Ear84, Gut83g, Mas80a]. every [Hil83, Kap80b, Rum85]. everyone [Dou83b, HMDF80]. everything [Per85b, Zor85f]. evidence [Ler83e]. evident [DM85]. Evoked [Tru83b]. evolution [Mok82c]. evolutionary [Jur87a]. evolve [Sha80c]. evolving [Smi87]. examination [MR82]. examined [Bel84d, Kli84, PK81]. examines [Chr81f]. exceeding [Voe86a, Voe86b]. excelled [Sta86]. exceptionally [Gut84e]. excess [Wal84c]. exchange [Tru84a]. exchanges [Kap82a]. exciting [Ano88d, Ano88m, Per84e, Wol83]. exclusive [Per83f]. executing [Wil87]. executive [Fis87a, Ler83b]. exercise [Chr81d]. exercises [Cye84]. exhibits [Gut83b, Ler84d]. exist [Wal83d]. existing [Eis84, Her83b, Tal83, Wet82b, Wil80]. exists [Doy81, Mok81g]. Exotic [Fis85c]. expand [CN83]. expandable [Ano81-186]. expanded [Hor84c, Mok82b]. expands [Mok80c]. Expansion [GZ85]. expansive [Chr86i, PZ89]. Expectations [WZ85, Bel85d, FZ87b]. expected [ABEI84, Fis85c, Wol80, You84, ZJ86]. expense [Fis83i]. expensive [Wal84a]. experience [Blu80]. experiences [Ada84b]. experiment [AF85, Bro81a, Jur85a]. experimental [LH87, MS84, Red80, SS84]. experimenter [Zor86c]. experimenters [Cla80a]. Experiments [HT83, Hec85, Wit84]. Expert [Gev83, Ano84-168, AMG^{+}85, Fei82, Kap84c, KP86, Per82b, Wal84g]. expertise [Cor87a]. Experts [Jur86d, Tru83d, BPEB84, Cla80b, Kap82a, PZ89, Voe87e]. explains [Fis83d]. explanation [Doy81]. explode [Fis87b]. exploded [Lem81b]. exploit [Lem81a, Per82e, STK82, Wol80]. exploitation [Bel84a]. exploited [Bac80, Lur84]. Exploiting [Jur80, Hor84d]. exploits [Red80]. explore [Kap81c]. explored [Gut82f]. explores [Ada87a]. explosion [FZ87a]. Explosions [Tel82, Zor86b]. explosive [Chr86i]. export [Tru84a]. exports [Ano82-156, Lom81c]. exposure [Chr83a, MFNL85]. Express [Voe88e]. extend [Mok82e, Wal87f]. extended [Bry84, FZ87a, Voe87d]. extends [Wel84]. Extensive [Esc86, ST81]. Extracting [Ada88a]. extrasensitive [Ant87]. extremely [Voe86a, Voe86b]. eye [Ano80-243, Ano88j, Mok81g, Sha80b, Tan86]. Eyes [KS86a, Ada86a].

F [Hil86]. FAA [FWVH86, Lom80b, Lom80d]. Fabrication [Fag82b, Fis84d, Fag80b, Fis84e, Fis85f, Gut84l]. fabricators [Fis84d]. face [Ano81-152, Mor81a, Tru83c]. face-lift [Ano81-152]. Faced [Ano86-66]. faces [Ano81-151, Ler82a, Par84]. facets [Car83]. facilitate [Mok81h]. facilitated [Bag84]. facing [Kah81]. fact [BCUY82, Fit88e, Win84]. factor [Jay86]. factories [Tru83d]. factors [Fag80a, HT83, Jur87c, Mus85, Per81a, Voe86a, Voe86b]. factory [Asa87, Ber82h, Chr83j, Fag80b, Fis87a, Gro83, Kam86, Mal84]. facts [Ano81-151, Bee83, Gev83]. fail [Doy81, SM81, TBH87, Zor88d]. fail-safe [TBH87]. failed [KW86, Mac84, Mar85]. failure [Doy81, Esc86, Gar81, Mas80b, Wil81, Zor87f]. failures [Ano87-117, Ano87-118, Sti81, Voe87f, VDT87]. faire [Fis83d]. Falcon [Fis87c]. falsifying [Chr80i]. fame [Ano84-86, Chr83b]. familiar [Per83a]. family [Bel85c, Bel87a, WZ87]. Famous [Per87d, Wol84b]. fancy [Fit88e, Per81a]. far [Bry84, FW87, Joh84, Kap80b, Ler80c, Per85b, Voe86a, Voe86b, Wal83a]. Fast [Che87, Eid81, Lem81a, Wal84e, Blu80, Kah83]. Faster [KS86a, Ber82c, Ber82j, Ler82e, Tri82, Wal83a, Wal85e, Wer87a]. fastest [Lep81]. fatal [BE87]. fated [BE87]. Fault [Ber81h, Chr82g, Zor85b, Zor85e, Zor87f]. fault-tolerant [Zor85b, Zor85e]. Faults [Ano87-117, Ano87-118, Voe87f, VDT87, Kap87a]. favorably [Dup82]. favored [Wil81]. favorites [Fis87c]. FCC [BA86, Jac82, PK81]. Fear [Hor85e, Per87b]. feasibility [Lem81c]. feature [Ber82f]. features [Ano81-181, NS80, Riv87, Sha80f, VB86, Voe87h]. fed [Ano81-177]. Federal [FWVH86, Hor84a, Ler82a, Ler84c, Lom80b, McC85]. feeling [Per82a, Pic82]. fellow [Nah83]. female [FP84]. Ferroelectrics [BG89]. fester [ABEI84]. few [Ano88j, Gol81, Gut83i]. Fewer [ZJ86, Jur86a, Kec83]. Feynman [Ano87-100]. Fiber [Bel88b, GZ85, Kap83b, Kap85f, KS86b, MNS81, Mok82d, Ano83-141, Bel84c, Bel85a, GBDC86, Gre85, Gut82f, Hec85, KK84, Mid80, Voe87g, Bel87b, Mok81a]. Fiber-optic [Kap85f, KS86b, MNS81, Ano83-141, Bel84c, Bel85a, GZ85, Hec85, KK84, Bel87b]. fibers [Bel88b, Kap83b, Kec83, Shu89]. fide [Fis84e]. fidelity [Mok82a]. Field [Ano80-36, Ano81-38, Ano82-135, Ano84-51, Ano85-45, Ano86-48, Bel86b, Ber80d, Che87, Kec83, Kir83, Mos81, Wal84c, Ler81d]. field-effect [Che87]. Fielding [Fit88b]. fields [Ano87-234, Bot83, Jur86d, Ler84a, Ler84c, MFNL85, SH85]. fifth [Dan82, Mo83]. fight [Wal82a]. fighter [Ada87b]. fighting [Zor88c]. figuring [Voe87h]. film [BKW83, MO89]. films [BKW83]. filtering [AF87]. filters [Sol88]. final [Ano87-75, Bow85, Ler81g, Mok80e]. Finally [Per87e]. find [BFB87, Doy81, Zor85f, Zor87f]. Finding [Her82, DM85, Wol83]. findings [Ler84a]. finds [AW85]. fine [LM81]. fines [Jur81b]. finesse [Cal82]. finish [Fis84d]. fire [Chr80d, Chr80e]. fired [Ler84g, Zor85d]. fires [Far86]. firing [Chr80a]. firm [Rac80]. firms [FM84b]. Firmware [Rac80, Ano80-42]. Firmware-based [Rac80]. first [Ano85-125, Ano86-152, Apt85, Bal85, Bel84e, BFB87, Cla80a, DD85, Fis83c, Fis84g, FZ87b, Kap84b, Ler82k, LM81, Lur84, MRRS84, Pak85b, Per87d, Sum82a, Voe88a, Wet82b, Wol85, XM89]. firsthand [Ryd84]. Firsts [Ano88k, LM81]. fishing [HW83]. fit [Mas80a, Miy89]. fits [BFB87]. fitted [Wal83d]. Five [Ano84-147, Per81c, Bel88c]. fix [KSB87]. Fixed [Wol80]. Fixed-base [Wol80]. Flamboyant [JI84]. Flash [PL89]. Flat [MO89, Tan89a, Wol80, Apt85, Sil89]. Flat-panel [Tan89a, Apt85, Sil89]. flatness [PW85a]. flaw [BE87]. flawed [Ano87-151]. flaws [Her82, Mas80b, Tus86, Voe87h]. flee [Kah82]. fleet [Dup82]. fleeting [Ada87a]. Flex [Voe88d]. flexibility [PW83, Spi86]. flexible [Dor83, Kel87]. flight [BE87, Chr87k, LM81, Wal86a]. flights [Voe86a, Voe86b, Wal86a, Zor86b]. floating [Ano82-185]. floating-point [Ano82-185]. flooded [Fis86d]. floor [Fis87a, HW83, Kap84c]. flourished [Voe87i]. flow [Ler84b, Ler84f, Wal82a]. fluid [Ano82-250]. flux [Bel86b]. fly [MRRS84, Mur89a]. Flying [Ano82-46, HW83, Wal82c, Wal86a]. FM [Voe87d]. Focus [Ano84-247, WB86]. focused [Miy84]. Focusing [Bac80]. foil [Mas80b]. fold [Voe87j]. follow [ABEI84, Fis87a]. foot [Voe88c]. foothills [Per85b]. footsteps [Luc87b]. Force [Cha82, DD85, DeL82, Gut84h, KM82, Mac84, Wal82c, Wal82d, War82]. forced [Fis84f]. forces [Aml82, Cor88b, Den81, Fis86a, Voe86a, Voe86b]. forecasting [Ler82c, Mar85]. forecasts [Cun80, Mar85]. foreign [FP84]. foreign-born [FP84]. foremost [Car83]. Forensic [Ler83e]. foreseeable [SW83]. forgeries [Fre86]. form [Bar81, Lem83a, Pic82]. Formal [Kap84b, Zor84a, Her84]. former [Dan82, Fei82]. formidable [CC87, Fag80d, FP84]. forming [Gro83]. fortress [Per84c]. Forum [Act87, Ano80-127, Ano80-128, Ano80-129, Ano80-130, Ano80-131, Ano80-132, Ano80-133, Ano80-134, Ano80-135, Ano80-136, Ano80-137, Ano80-138, Ano81-120, Ano81-121, Ano81-122, Ano81-123, Ano81-124, Ano81-125, Ano81-126, Ano81-127, Ano81-128, Ano81-129, Ano81-130, Ano81-131, Ano82-122, Ano82-123, Ano82-124, Ano82-125, Ano82-126, Ano82-127, Ano82-128, Ano82-129, Ano82-130, Ano82-131, Ano82-132, Ano82-133, Ano83-129, Ano83-130, Ano83-131, Ano83-132, Ano83-133, Ano83-134, Ano83-135, Ano83-136, Ano83-137, Ano83-138, Ano83-139, Ano83-140, Ano84-134, Ano84-135, Ano84-136, Ano84-137, Ano84-138, Ano84-139, Ano84-140, Ano84-141, Ano84-142, Ano84-143, Ano84-144, Ano84-145, Ano85-132, Ano85-133, Ano85-134, Ano85-135, Ano85-136, Ano85-137, Ano85-138, Ano85-139, Ano85-140, Ano85-141, Ano85-142, Ano85-143, Ano86-121, Ano86-122, Ano86-123, Ano86-124, Ano86-125, Ano86-126, Ano86-127, Ano86-128, Ano86-129, Ano86-130, Ano86-131, Ano86-132, Ano87-119, Ano87-120, Ano87-121, Ano87-122, Ano87-123, Ano87-124, Ano87-125]. Forum [Ano87-126, Hur87, San87, Ten87]. forward [Bor89]. forward-looking [Bor89]. fossil [Fis84h]. fossil-fuel [Fis84h]. foster [Lom81c]. found [Ber81e, MFNL85]. Foundation [Ros80, Wei80]. founder [Pak85b]. foundries [Fis84b]. Four [Mok80f, Ano80-43, Ano83-141, Ano83-245, Cor87d, Gut84i, Jay86, Kli84, Per83a, de 83]. four-channel [Ano83-245]. four-cylinder [Ano80-43]. four-year [Gut84i]. four-year-old [Cor87d]. Fractals [Zor88b]. fractured [Bel85b, SH85]. fragmented [Hec85, Hor85d, Kel87]. fragments [Ano84-236]. frailty [Chr81l]. frames [BKBC85]. framework [DD86, FM84a, Mor81b]. France [Cur83, Dup82, Gra84, Pin80, Wei84]. Francisco [Sha81c]. fraud [Zor88c]. Frederick [Bry84]. free [Per84c, Wal82a]. French [And80, Dup82, Gut88, Kap83a]. frequencies [Che87, EMM82]. frequency [Bel84e, BA86, Hic83, Sar82]. friendliness [Mo83, VWZ86a]. friendly [And80, Jur87b, Per83c, PV89, SH85]. frigate [Ada87f]. Front [Ano80-139, Ano80-140, Ano80-141, Ano80-142, Ano80-143, Ano80-144, Ano80-145, Ano80-146, Ano82-134, Ano83-142, Ano87-127, Ano87-128, Ano87-129, Ano87-130, Ano87-131, Ano87-132, Ano87-133, Ano87-134, Ano87-135, Ano87-136, Ano87-137, Ano87-138, Ano87-139, Ano87-140, Ano87-141, Ano87-142, Ano87-143, Ano87-144, Ano87-145, Ano87-146, Ano87-147, Ano87-148, Ano87-149, Ano87-150, Ano88l, SFV^{+}88, Ano82-155, Ano83-160, Ano83-184, Mil87, Wal83a]. front-end [Ano83-184, Mil87]. frontiers [Gut82f, Jur88b]. fronts [PW85a]. frozen [Hil86]. fruition [Tor80]. fuel [Fis84h, FZ85, JT81, JCH80, Jur82a, Kap80c, Kap83d, Kap85d, Mil81, Riv80, Wet82b]. fuel-cell [Kap83d]. fuel-efficient [JCH80]. fuels [Fag80e]. fulfilled [Jak85]. full [Ler84h, Zor87b]. full-time [Ler84h]. fully [Fis83b]. fun [Ada85b, Bel85d, Chr85f, Zor86c]. Function [Bac82, Ber80d, FHW85]. Function-level [Bac82]. functional [Ano83-186, Jur84, Wal87d]. functions [Fag80b, Gro83, GWF84, Hor85d, Jur85b, MMB83, Vae86, Wal85e, Zor85e]. Fundamental [Gro83, Doy81, Fis84b, HTW84]. funded [Kap80c]. funding [Tru83a]. further [Ano80-126, Ano81-119, Ano82-121, Ano82-247, Ano83-107, Ano83-240, Ano83-241, Ano83-239, Ano84-240, Ano84-238, Ano84-239, Ano85-233, Ano85-234, Ano85-235, Ano86-231, Ano86-232, Ano86-233, Ano87-232, Ano87-233, IK84, Mok82e]. furthers [Mok82c]. fusing [SH85]. fusion [Fis84h, Kap84f, Ler80d, Wel85]. Future [Fei82, Kap82d, Ano80-153, Ano80-154, Ano80-155, Ano80-156, Ano80-157, Ano80-158, Ano80-159, Ano81-145, Ano81-146, Ano81-147, Ano81-148, Ano81-149, Ano82-149, Ano82-150, Ano82-151, Ano83-158, Ano87-155, Ano87-156, Ano87-157, Ano87-158, Ano87-159, Ano87-160, Ano87-161, Ano87-162, Ano87-163, AMG^{+}85, Bel85f, Bel89a, Ber82b, Bot85, Chr86i, FM84a, FAEH84, Flo88, Gol81, Jur81a, Jur83b, Key83, Ler84b, Roi86, Tor85, Tru83d, Wal84e]. futuristic [GT82]. fuzzy [Zad84].

G. [Dou81]. GaAs [MS84]. gain

[FZ87a, Wal86c, Wal87g]. gained [Bel87f]. gaining [AH85]. gains [Fis83i, Gut84j, Zor89f]. Galileo [DJ85]. gallium [Bel85c, ELW83, NH86]. game [Jur87d, Mok80e, PW83, Tru83e]. Games [Cor84, Ada85b, Mok81e, PTW82, Per83e]. gamut [Ler83b]. gap [Chr80k, Mey86]. garage [Kap80b]. Garbage [Wol88a]. Gas [Lea85, Voe88c]. Gas-plasma [Lea85]. gases [Ano83-56]. gasification [Fag80c]. gasoline [JT81]. gate [Ano83-186, Che87, Fre88a]. gathered [Gar81]. gathering [Bag84]. gave [PW83]. gazers [Ano80-243]. GE [Wol84b]. Gems [Ano84-146, Ano84-148, Ano84-147]. General [Ano84-240, Gut83d, Per87e, CC86b, Kid81]. general-purpose [Gut83d]. generate [Eis84, Fis84f]. generated [Per83e, You84]. generates [MS84, Ros83]. generating [FZ87a, Mas85]. generation [Bel89d, Dou83a, Dou83b, FW87, Gut83f, Kah83, Kap83a, Kir83, Mok82d, Mo83, Pea81, Wal83a]. generations [Ano83-141, Bow85]. generator [Ano80-192, Ano82-194]. generators [Jur87g, Kir83]. generic [PK81]. genesis [Fit88c]. geniponics [MS87]. genius [Nah83]. geometries [Lea85]. geophysics [Che89a]. George [Per83f]. Germany [ABEI84, Gra84, Zor87a]. gestation [Sum82a]. get [ABEI84, Ber80c, CC87, Jur82a, MAN80, Pyk87, Sha80d]. gets [Bel84e, Din82, Jur82c, Moy80, VB86]. Getting [Chr81e, PK81, Per84a, Wal86c]. GF [Ano81-186]. giant [Chr86c, CC86b, Hil83, HTW84, Sti88a, Wet82a]. Giants [Ber82e, Ano88j, Car83, Fle85b, Luc87b, Nah83, Wol83]. gigabyte [Voe87j]. gigawatts [Flo88]. Give [CC87, Aml82, Ano80-245, Bel84b, Cog84, Con86, Cor87d, Fle85b, Lur84]. give-and-take [Cog84]. given [Hic83]. gives [Ryd84]. giving [Ant87, Rau87]. glass [Chr84l]. glasses [RD81]. gleaned [Sha80a]. Glitch [BFB87]. Global [Get82, Jur89b, Ano84-236, JV87, Jur87d, Kap82b, Mok80d, Per87f]. Go [Jur87d, AF87, BA86, Dor83, Fis85e, Luc80, Mel80, Per87f]. goal [Bel86c, CC86a, Gut83g, Ham86, Kap80b, Ler82j, Lom80d, Row82]. goals [Cun80, FW87]. goes [Del80, Fis85a, Kap83d, Lar85, Mas80c, Per88b, Pin80]. Going [Chr83h, JT81, Kap85e]. gold [GB85, Voe86a, Voe86b]. Gold-plated [Voe86a, Voe86b]. gone [Chr87h]. Good [Chr80f, Chr80g, Chr85d, Ano88k, Chi81, Chr86g, Fag82a, Kap87b, Kat87, Ler82f, Pea81]. goods [VWZ86e]. Gordon [Bel88a]. got [Bro81a]. Govermnent [Per83f]. govern [Mos81]. Government [Chr81j, FM84a, Hor85d, Per81b, Wal82b, Zor88c, Hor84a, Ler84c, Per83g, Wal82a, Wal86a, Zor87e]. grabbed [Per85a]. grades [Ano82-190]. graduate [Hor84b]. grand [AF85]. grande [Fis86d]. graphics [Ano82-191, Ano82-188, Ano84-53, Ano84-237, Bal85, Ber82g, Gut84j, Hef84, Jac80, Lem81a, Ler81c, Per83e, Per84e, Voe87c, WM84, Wer87a]. graphs [WM84]. grapple [CS87, Tru83a]. grapples [Hon80]. grappling [RD81]. grasp [Mac84]. grass [MD84]. grass-roots [MD84]. Great [FZ87b, Bee83, Bel85d, Cal82, CN83, ELW83, KS86b, Ler82c, Ler84e, Oak83]. Greater [EMM82]. greatest [AH85, Kap84f, Zor86a]. greatly [Bac82, CN83]. greet [Ler84d]. grid [CA82, Gau85, Ver86]. grids [Wet82b]. grip [Per88d]. gripping [DD85]. ground [Bel88e, Gut84j, Hil86]. grounds [HW83]. group [Ano82-157, Kap80c]. growing [Ant87, Cau82, Fri84, Ler82g, WM84, Zor84b]. grows [WZ87]. growth [AMG^{+}85, Fis86a, PZ89, ZJ86, Zor86d]. GS [Zor84b]. GS/E [Zor84b]. guarantee [Kah84]. guessing [Ano82-241]. guidance [Fit87c]. Guide [Ano80-147, Ano81-132, Ano82a, Ano83a, Ano84a, Ano85b, Ano86a, Ano87a, Apt85, Hor84b]. guided [Wol84b]. guides [Hor84b]. guiding [Win84]. gun [AM86, Ada87e, Zor89e]. Gunn [Voe89a]. Gutting [SA85]. gyro [KS86b]. gyroscope [KS86b]. Gyroscopes [Mar86, KS86b].

H [Hou85, Dup82, Ler81f]. H-bombs [Ler81f]. hack [Jur84]. Haggerty [Fag80f, Lom80e]. half [Bot85, Dav83a]. hall [Ano84-86, Chr83b]. halt [Wei80]. halting [Ada87e]. hand [Ano83-143, Chr86a]. hand-held [Ano83-143]. handicapped [Hor84d]. handicapping [Apt85]. handle [Fit89a, Lom80f, SW83]. handles [Kat80]. handling [Bal81, BA83, Jur82a, Jur87b, Rum85, Sha81f]. hands [Gut84i, Nag83]. hands-on [Gut84i]. handsomely [Zor86c]. hang [Chr85c]. happened [Ada87f, Ano80-247, Ano80-248, Ano80-249, Ano81-231, Ano81-232, Ano81-233, Ano81-234, Ano81-235, Ano81-236, Ano82-249, Ano82-250, Ano82-251, Ano82-252, Ano82-253, Ano83-243, Ano83-244, Ano83-245, Ano83-247, Ano83-246, Ano84-252, Ano84-253, Ano84-248, Ano84-249, Ano84-250, Ano84-251, Ano85-242, Ano85-243, Ano85-244, Ano85-245, Ano85-246, Ano85-247, Ano86-240, Ano86-241, Ano86-242, Ano86-243, Ano86-244, Ano86-245, Ano87-239, Ano87-240, Jur87e, Jur87f, Jur87g, MS87, Wat87]. happenings [de 83]. hard [Bha85, Cla80b, Mos81, Tru83a, Voe87j]. hard-driving [Bha85]. harder [Ada85a, Tul83]. hardly [FHW85, Sug80b]. Hardware [Fis84a, Ano81-133, Ano82-61, Ano84-237, Ber81f, Ber82b, DD86, Gut84h, Her83b, Jay86, Kap83c, Ler80a, Ler81c, Ler81b, PW83, Sta87b, Sti81, Wal84d]. hardwire [HHH87]. Hardy [Kid81]. Harmony [Gut83a]. harness [Zor88a]. harnessing [TBH87]. Harpy [Red80]. Harris [Gut84d, Gut84b, Gut84c, Gut84a]. harsh [Ada85c, KS86b]. hasn’t [Tor80]. hasten [Ber81k]. haven’t [Tru84b]. having [Wal84c]. hazardous [Fis85e, Lom80f]. hazards [Bel87c, Chr84k, Chr87j, Ler80b]. HDTV [Jur89a, Tan89b]. head [Fre84, GBDC86, Ler80c]. headway [Kap87c]. healing [SH85]. Health [Bug84, MFNL85]. Health-care [Bug84]. Hear [Chr82h]. Hearing [Pic82, Jur80]. heart [Ano82-48, Bee83, Fis83b]. hearts [Cor87b]. heat [Ler84g, MS84]. heatedly [Ber82b]. Heating [Fis85d, SPM^{+}80, FHW85, Ler84g]. Heaviside [Nah83]. heavy [Miy89, Tan89a]. heavy-fit [Miy89]. Hedy [Kah84]. held [Ano83-143, Bel85b, Bel85d, Kap83c, Per85a]. Hello [Tor85]. helm [Gut88]. Help [Asa87, Ano81-150, Bel81, BFB87, Cun80, DTS86, Dav87a, Fis86e, HOD82, Hor84a, Hor84c, Hor86d, KSB87, Kap83a, KA86, Ler83e, Mar85, Per82a, Per83d, Per83a, PW86, Pro81, Tri80, VB86, Wei80]. helped [Bel87e, Eis84, Fit87c, JI84, Lur84, SS84, Sta87a, Zor88a]. Helping [Chr80h, DTS86, Vae86, Chr86a, Ler82h, Per81e]. helps [Mel80]. HEMT [MS84]. Here [Lom80c, Per84c, Fis87c, Gar84, Per84b, Per87e, PK81, Tru84a]. heritage [Chr81f]. Hero [Fis86b]. heroes [Ano88j, Chr85k, de 83]. hiatus [Sim83]. hidden [FW87]. High [Ada89a, Ano82-152, Bor89, CA82, Chr84c, Eid81, HW83, Jur88a, Kap80e, Kap83c, Miy84, Mok82a, Per83b, SD81, Zor84c, Bal81, Bel85c, BA86, Bel87f, Ber81h, ELW83, Fis84f, Fis85a, Fis85c, FHW85, Fle85a, GS80, Jur83b, JP85, Jur87d, Kap80b, Kap87a, Kec83, Kel87, Lar85, Lea85, Lom80a, Lur84, Mel80, Mok81i, Nai83, Per85b, Per85d, RK80, Roi86, SA85, Sha80e, ST81, Spi86, Sug80b, Sum80, Tru83e, Voe86a, Voe86b, Win84, Zor85c]. High-altitude [SD81]. High-capacity [CA82]. High-definition [Jur88a, Jur83b, Mok81i, Roi86]. High-density [Miy84, Mel80]. High-fidelity [Mok82a]. High-flying [HW83]. high-frequency [BA86]. high-performance [Kap80b, Lea85]. high-power [RK80]. high-precision [Kap87a]. high-quality [Kel87]. High-speed [Bor89, Bel85c, ELW83, Fis85c, GS80, Lom80a, Sug80b]. high-stakes [Jur87d]. High-tech [Ada89a, Chr84c, Zor84c, Fis85a, FHW85, Fle85a, JP85, Lar85, Nai83, Per85b, Per85d, Tru83e, Win84, Zor85c]. high-volume [Ber81h]. Higher [Chi81, Ber83b, Ler84b, Lof80b]. highlight [LM81]. highlighted [Ada84a, Wal83c]. Highlights [Bel85a]. highly [Car80, Gut83i, Lem81c, VWZ86e]. highways [Jur82c]. him [Fag80f, Nah83]. himself [JI84, Wol84b]. hinges [Kap80b]. hints [Jak85]. Hiring [Chr80a]. historical [Ano84-146, Ano84-148, Ano84-147]. history [And80, Ano83-169, Ano87-151, Chr81l, FW82, Gut82c, Gut82h, Gut84e, PW83, PW85b, Per88e, Sha81e, Sha81f, Wal86c]. hit [Jay86, Riv80]. hitherto [Cye84]. hobbyhorse [Zor85c]. Hodel [Fis83d]. hold [DD85, Roi86, Wal85e]. holds [ELW83]. Hologram [Ina89]. Holography [Cau82]. Home [Cla80b, Jac80, Ano82-171, Bel85d, Dor83, Han89, Jur85a, JP85, Kap85a, MD84, Mor85, Per84e, Per85b, Per87b, Tru83e, Wit84]. Homegrown [Miy89]. Homeostatic [STK82]. homeowner [Mok80b]. homes [Shu89]. Homo [Hor86b]. homogeneous [Car80]. honed [Bel87f]. honing [Fit87a]. honor [Wal86a]. Hope [Kap82c, KS86a]. hopes [Ada88d, Kap83e, Kap83c]. hordes [Voe87i]. Horizon [Aug82, Ber80a, Wal84e]. horizons [Ano84-236]. hospital [Bug84, Gar84]. host [Bel84e, Liu84]. hot [Cor87c, Fit89a, KSB87, Sti88e]. hour [Hil86]. hours [Ano87-75, PW86, Ros83, Ver86]. house [FHW85, Hou85]. household [Pan87a, Rum85, Tru83e]. Howard [Per87a]. however [McC85]. howling [Hil86]. huge [Kap85c, de 82]. Human [HT83, Ano81-150, Ano84-168, Ano84-236, AJ85, Fag80a, Fis83e, HN88, Jur87c, MFNL85, Mus85, Per81a, Sha80b, Tri82, Ada87f]. Human-factors [HT83, Per81a]. humanistic [Nai83, Wol84a]. humanities [Kli84]. Humanization [Ler83c]. humans [Ada85c, DD85, KS86a, MFNL85, Sta87b]. hungry [Per84b, Tan89a]. hurdles [FP84, Sti89c, TD84]. hurting [Per87b]. husband [Per85d]. husband-and-wife [Per85d]. hut [TD84, Wal84f, ZJ86]. HVDC [FZL86, FEK80, Zor85d]. Hybrid [Hic83]. Hybrids [Fro83]. hydraulics [Ano83-55]. hydro [Fis86d, Ler80c]. hydroelectric [Fis86d, Ler80c, You84, de 82]. hydropower [Zor88a]. Hypercube [Wil87]. Hypermedia [Per87e]. hypersonic [Voe88e]. hypertext [Per87e]. hyperthermia [SPM^{+}80]. Hz [MFNL85].

I/O [Ano80-196, Ano83-185]. I/sup [Gut89]. IBM [Ano80-194]. IBM-compatible [Ano80-194]. IC [Ano83-186, Bow85, CR88, HHH87, Lea85, Per88c, Rum85, Sum80]. ICBM [Ada85a]. ICs [Ber81j, Ber82f, Ber82j, Bro81b, Dav83a, DS81, Eid81, Fis85f, Fis87d, Kap83b, Lom80a, Mok82c, Pea81, Sha81c]. idea [Ros80]. ideal [Zde83]. ideals [Ear84]. ideas [Lom80b, Wal82a]. identify [HW83, Jur86d, Ler83e, Tus86]. identifying [KSB87]. IEC [Mas80a]. IEEE [Ano85-45, Ano80-36, Ano80-37, Ano80-148, Ano80-153, Ano80-154, Ano80-155, Ano80-156, Ano80-157, Ano80-158, Ano80-159, Ano80-149, Ano80-150, Ano80-151, Ano80-152, Ano80-174, Ano80-175, Ano80-176, Ano80-177, Ano80-178, Ano80-179, Ano80-180, Ano80-181, Ano80-182, Ano80-183, Ano80-184, Ano80-185, Ano81-38, Ano81-39, Ano81-134, Ano81-135, Ano81-136, Ano81-137, Ano81-138, Ano81-139, Ano81-140, Ano81-141, Ano81-145, Ano81-146, Ano81-147, Ano81-148, Ano81-149, Ano81-142, Ano81-143, Ano81-144, Ano81-165, Ano81-166, Ano81-167, Ano81-168, Ano81-169, Ano81-170, Ano81-171, Ano81-172, Ano81-173, Ano81-174, Ano81-175, Ano81-176, Ano82-41, Ano82-135, Ano82a, Ano82-136, Ano82-137, Ano82-138, Ano82-139, Ano82-140, Ano82-141, Ano82-149, Ano82-150, Ano82-142, Ano82-143, Ano82-144, Ano82-145, Ano82-146, Ano82-151, Ano82-147, Ano82-148, Ano82-172, Ano82-173, Ano82-174, Ano82-175, Ano82-176, Ano82-177, Ano82-178, Ano82-179]. IEEE [Ano82-180, Ano82-181, Ano82-182, Ano82-183, Ano83-144, Ano83a, Ano83-145, Ano83-146, Ano83-147, Ano83-148, Ano83-149, Ano83-150, Ano83-151, Ano83-152, Ano83-153, Ano83-154, Ano83-158, Ano83-155, Ano83-156, Ano83-157, Ano83-170, Ano83-171, Ano83-172, Ano83-173, Ano83-174, Ano83-175, Ano83-176, Ano83-177, Ano83-178, Ano83-179, Ano83-180, Ano83-181, Ano84-50, Ano84-51, Ano84-149, Ano84a, Ano84-150, Ano84-151, Ano84-152, Ano84-153, Ano84-154, Ano84-155, Ano84-156, Ano84-157, Ano84-158, Ano84-159, Ano84-160, Ano84-161, Ano84-169, Ano84-170, Ano84-171, Ano84-172, Ano84-173, Ano84-174, Ano84-175, Ano84-176, Ano84-177, Ano84-178, Ano84-179, Ano84-180, Ano85-144, Ano85b, Ano85-145, Ano85-146, Ano85-147, Ano85-148, Ano85-149, Ano85-150, Ano85-151, Ano85-152, Ano85-153, Ano85-154, Ano85-155, Ano85-156, Ano85-167, Ano85-168, Ano85-169, Ano85-170, Ano85-171, Ano85-172, Ano85-173, Ano85-174, Ano85-175, Ano85-176]. IEEE [Ano85-177, Ano85-178, Ano86-48, Ano86-133, Ano86a, Ano86-134, Ano86-135, Ano86-136, Ano86-137, Ano86-138, Ano86-139, Ano86-140, Ano86-141, Ano86-142, Ano86-143, Ano86-144, Ano86-160, Ano86-161, Ano86-162, Ano86-163, Ano86-164, Ano86-165, Ano86-166, Ano86-167, Ano86-168, Ano86-169, Ano86-170, Ano86-171, Ano87-152, Ano87a, Ano87-153, Ano87-155, Ano87-156, Ano87-157, Ano87-154, Ano87-158, Ano87-159, Ano87-160, Ano87-161, Ano87-162, Ano87-163, Ano87-170, Ano87-171, Ano87-172, Ano87-173, Ano87-174, Ano87-175, Ano87-176, Ano87-177, Ano87-178, Ano87-179, Ano87-180, Ano87-181, Ano87-234, Chr81f, Cog84, Gra82b, Lom80c, Per85d, Ano84-72, Ano84-193, Ano85a, Ano86-65, Ano87-61]. IEEE-member [Per85d]. If [Kap83d, Whi87, Cal82, Fis87b, Tor80]. iffy [Voe88e]. ignore [Wal86a]. II [Ano81-150, Ano82-152, Ber81c, Ber81j, Chr80e, Chr80g, Chr82c, Fle82, HMDF80, Kah84, Ler81e, Ler82g, Mok81d, Mok81b, Per83f, Pin80, Red80, Rei80, Wal82a, Wal87d, Wol80]. III [Ano81-151, Ano82-153, Bac80, Ber81d, Chr82d, Cla80b, Fag81c, Kap81b, Kat80, Ler82h, Mok81e, NS80, Sar82, Sha81b]. Ill [Ano84-85, BE87]. ill-fated [BE87]. Ill-starred [Ano84-85]. illegal [Hor85e, Sti81]. illiteracy [Chr83i]. illustrations [WM84]. image [AJ85, BJ88, Fre86, Mok81g]. Images [Lur84, AJ85, Kap81a, Lla81, Per84e]. imaginatively [Cye84]. Imaging [Fis83e, Bot83, Cor87b, EG82, Jur80, Lla81]. imitation [Lem83a]. imminent [Kec83]. immunity [GBDC86]. impact [Bel84e, Jur85a]. impacts [Gut83f]. impasse [Roi86]. impeding [Bug84]. impending [Chi81, de 83]. imperiling [Voe86d]. Implantable [CP86, Fis83b, Hor85b, Wal87d]. implementable [Fis83c]. implementation [Bug84]. importance [Ber80f]. important [Ber82h, Hic83, Ler83b, Mih81, Mos81, Smi87, Voe87h, Wal84f, Wer87a, ZJ86]. Imports [Ano82-156, JCH80]. imposing [Ler84c]. impossible [GGZ86]. impromptu [Cla80a]. improve [Fit87c, Jur82a, Jur87b, Jur87c, Kap82d, Kap85d, Kap87c, Liu84, Ros88, Tus86]. Improved [HOD82, Ano81-133, Ler82c, Pea81, Wal87a].

Improvements [Nag83, Fag82c, FZL86, Hil85, Hor86d, Jel81]. improving [Fis84a, Kap83e, Kap83c, Lom80b]. inability [BE87]. incentives [Mac84]. inch [Voe87j]. Incidents [Hor84c]. incites [Cor87c]. include [Ano88j, Bel85a, Fis86e, Gib80, Kap82a, Lom80b]. Includes [Ano83-184, Ano80-195]. incompatible [Cle82]. incomplete [Ler84a]. incorporated [Ler82b]. incorporating [GS80]. increase [Bel86c, Bel81, CC86b, EMM82, FZL86, LS81, Mid80, Sha84, VH86, Voe87j, Wal85e, You84]. increased [Ano81-152, Bel84c, Ber80b, Kap87a]. increases [Cor87b]. Increasing [DD85, Jur85b, FP86]. Increasingly [Che82]. Incredible [Bel85b]. incubation [Sum82a]. incumbent [Dan82]. independent [Ang80, Fly87]. independently [FHW85]. Index [Ano80-160, Ano80-161, Ano80-35, Ano82-40, Ano84-72, Ano85a, Ano86-65, Ano87-61]. India [Gra84]. indicate [Ada84b]. individual [Bot85, DTS86, Hon80, Rub84]. individuals [Fag82a]. induced [Mor81a, Mor81b, SD81]. induction [Kap84e]. Industrial [Fag82c, Jur82c, Jur82d, Kap84a, Kap84c, Kap85b, KP86, Kap87a, Ano84-86, Cas81, CC86a, FZ85, FEK80, Gut83f, Gut83i, Kap88, Kli84, San89d, Zor87e, Fag81b]. industrialist [Car83]. Industry [Ano80-202, Ano80-204, Ano80-205, Ano80-206, Ano80-208, Ano80-209, Ano81-188, Ano81-189, Ano81-190, Ano81-192, Ano81-194, Ano81-195, Ano81-196, Ano81-198, Ano81-199, Ano82-201, Ano82-207, Ano83-188, Ano83-193, FM84a, Gan82, Gut84f, Jur82c, MMB83, San89b, Wal82b, Ano80-198, Ano80-199, Ano80-200, Ano80-201, Ano80-203, Ano80-207, Ano81-191, Ano81-193, Ano81-197, Ano82-196, Ano82-197, Ano82-198, Ano82-199, Ano82-200, Ano82-202, Ano82-203, Ano82-204, Ano82-205, Ano82-206, Ano83-189, Ano83-190, Ano83-191, Ano83-192, Ano86-152, Bel86a, Ber81e, Bry84, Cor87d, Dan82, Fag81a, Fis83c, Fis84b, FM84b, Fis84e, FAEH84, Fis86e, FG86, FZ87a, Gut84h, Hor84a, Hor85e, Kah81, Kah82, Kap83e, KA86, KW86, Kel87, Ler84c, Lom81b, Mai83, Mun88, Per81b, Per81e, Per88a, SW88, Sha84, Sti88b, Tru83d, Voe87e, Wal82a, Wal86c, Zor89j]. inertia [Fis87b, Rab81]. inertial [Fit87c]. inertial-guidance [Fit87c]. inevitable [Bel85b]. inexpensive [Dut80, Sol88, Tus86]. infants [Ler84h]. inferences [Gev83]. infinite [Gut82c]. inflated [WZ85]. inflationary [Ano86-66]. influence [Bry84, Fre84]. influences [Bag84, Per80c]. inform [DMW87]. Information [Gra82c, Par84, Per80c, Sta87b, Wal82b, Wal82a, Ber82g, Chr80k, Hor85e, Jur81a, Kap84d, PW84, PW85c, Tru84a, Kap81b]. informed [Zor85a]. infrared [Ada84a]. infrastructure [San88b]. ingenious [Per88a]. Inherently [Fis87b]. initial [Voe87j]. Initiative [WB86, Wal87g]. initiatives [Hor84a]. injection [JT81]. injuries [Bel87c]. innovate [Voe88d]. Innovation [Fag80b, Jur82d, Wis82, Ber82h, Gra82c, VWZ86d, WZ85]. Innovations [Ano84-162, Ano84-163, Ano84-164, Ano84-165, Ano84-166, Ano84-167, Ano85-157, Ano85-158, Ano85-159, Ano85-160, Ano85-161, Ano85-162, Ano86-146, Ano86-147, Ano86-148, Ano86-149, Ano86-150, Ano86-151, Ano87-164, Ano87-165, Ano87-166, Ano87-167, Bel87d, BBJ87, Jur87b, Mok82a]. Innovative [Sha80d, Bel81, Kuw84, Pyk87, Van84, Wol84b]. INPO [Per81e]. input [Ano82-195, Jur82c, Jur87b, MNS81, SAPM80]. input/output [Jur82c]. inputs [BD83]. inroads [Jur86b]. insider [Lom80e]. Insidious [Ano88p]. inspection [Gro83, Wal87f]. inspiration [Fag80f]. installations [MS81]. instantaneous [Wel85]. Instead [DD86, Ros83]. instilled [Blu80]. Institute [Ano80-231, Ano80-232, Ano80-233, Ano80-234, Ano80-235, Ano80-236, Ano80-237, Ano80-238, Ano80-239, Ano80-240, Ano80-241, Ano80-242, Ano81-212, Ano81-213, Ano81-214, Ano81-215, Ano81-216, Ano81-217, Ano81-218, Ano81-219, Ano81-220, Ano81-221, Ano81-222, Ano81-223, Ano82-227, Ano82-228, Ano82-229, Ano82-230, Ano82-231, Ano82-232, Ano82-233, Ano82-234, Ano82-235, Ano82-236, Ano82-237, Ano82-238, Ano83-213, Ano83-214, Ano83-215, Ano83-216, Ano83-217, Ano83-218, Ano83-219, Ano83-220, Ano83-221, Ano83-222, Ano83-223, Ano83-224, Ano84-206, Ano84-207, Ano84-208, Ano84-209, Ano84-210, Ano84-211, Ano84-212, Ano84-213, Ano84-214, Ano84-215, Ano84-216, Ano84-217, Ano85-203, Ano85-204, Ano85-205, Ano85-206, Ano85-207, Ano85-208, Ano85-209, Ano85-210, Ano85-211, Ano85-212, Ano85-213, Ano85-214, Ano86-199, Ano86-200, Ano86-201, Ano86-202, Ano86-203, Ano86-204, Ano86-205, Ano86-206, Ano86-207, Ano86-208, Ano86-209, Ano86-210, Ano87-209, Ano87-210, Ano87-211, Ano87-212]. Institute [Ano87-213, Ano87-214, Ano87-215, Ano87-216, Ano87-217, Ano87-218, Ano87-219, Ano87-220, Per80a, Per81e]. institution [Bal84b]. institutional [Bel85f, Fis87b, Tor80]. institutions [Fri84]. instruction [Wal85e]. Instructional [Bal84a, Wit84]. instrument [Fit88a, Voe87g, Lom80e]. Instrumentation [Fag80c, Gut84f, Gut84g, Gut85b, Ler81d, Ler81e, Ler82e, MR82, Sha80c, Sha81b, Tul83, VB86, Voe87g, Ano80-230, DJ85, Fis84a, Fit88f, Fit89b, Fit89d, HOD82, Hou85, Jur84, KZG^{+}84]. Instruments [Sha80d, Ano84-168, BFB87, Gut83h, Gut85b, Lem83a, Wol85]. insufficient [Wal84c]. insulated [Per84c]. insurance [Ano82-157]. integrate [BA83, Rab81]. Integrated [BCUY82, ELW83, Bar81, Ber81j, CA82, Fis85f, Fis87d, Gro83, GWF84, Kap83b, Kap85b, Lea85, Mas80c, MAN80, Per89a, Riv87, Ros83, Rum85, Sum80, VWZ86a, Wer87a, Mok81b, Pan87a]. Integrated-Services [Pan87a]. Integrating [Wal83b, Fri84]. integration [Ano82-253, Ano83-162, Bee83, FHRW^{+}83, Jur84, MRRS84, Rau87]. Intel [Per89c]. Intelligence [Gut82d, Ada85c, Ber80d, FHRW^{+}83, Hor85e, Kah83, Ler82j, Nag83, PW84, VWZ86a, Wal85d, Wal87g, Zor87a]. Intelligent [Mok80c, AW85, Cor87d, Cor87c, FW87, Han89, Jur82c, Mok80e, Sha80d]. intelligibility [Jay86]. intelligible [KL85]. intense [Wal86c]. intensify [Kap84c]. intensity [Sar82]. intensive [Gar84]. Intent [Bel85c, Ham86]. interacting [GGZ86]. interaction [Bel84f, Ber83a]. Interactive [Ano82-188, Per83e, Bru87, Gut84j, Lem81a, Lut88]. intercept [Ada85a]. intercepted [Par84]. interception [Hor85e]. intercity [Mid80]. Interconnection [Vae86]. interconnections [Bel86a, HHH87]. interconnects [HHH87]. interest [Ano80-229, Den81, Fis84f, Mai83, Rub80]. interests [BMoF^{+}83, Ler81b, Per84c]. interface [Cla80b, Ham86, Jur82d, PV89, Wer87a]. interfaces [MMB83, Mok82e, Voe87g]. interfacing [Fag82a, NS80]. interference [GBDC86, WAO85]. internal [Bot83]. International [Ano82-156, Che89c, Cla80b, Jur87d, Kap82b, Ler84f, Mas80a, GP81, Kap85c, Mas80c, Row82, Wal86c, Ada87a].

Internationalization [Gut83a]. internship [Bha85]. interpretive [Mas80b]. interrelationships [Sha87]. interrupt [Ano82-194]. intertie [Zor88a]. Interviews [Wal84c, Lom80c, Per80b, Sha80a]. intra [Mok81h]. intra-office [Mok81h]. intriguing [GT82]. introduced [Bel87b, Ber80g, Cye84, WZB86]. Introducing [Pin85, Tru83a]. Introduction [Chr81a]. introspective [Sha80a]. intruder [Fit89c]. intruder-proof [Fit89c]. intuitive [VWZ86b]. invader [Mas85]. invades [Ber83b]. Invasions [Kap85c]. invented [Hec85]. inventing [Wis82]. inventions [Ano88d, Ano88m, Rac80, Wol84b]. inventor [Car83, JI84, Wol88d]. inventories [CC86a]. inventors [Ste81]. inventory [Ano88p, Wal87f]. inverter [Fag82c]. inverters [FEK80, Kap84e]. investigated [Rub80]. investigative [Joh84]. investment [Row82]. investments [Fle85b]. invisible [Ada88c]. invited [Ano83-204, Ano83-200, Ano83-201, Ano83-202, Ano83-203, Ano84-181, Ano84-182, Ano84-183, Ano84-184, Ano84-185, Ano84-186, Ano84-187, Ano84-188, Ano84-189, Ano84-190, Ano84-191, Ano84-192, Ano85-179, Ano85-180, Ano85-181, Ano85-182, Ano85-183, Ano85-184, Ano85-185, Ano85-186, Ano85-187, Ano85-188, Ano85-189, Ano85-190, Ano86-172, Ano86-173, Ano86-174, Ano86-175, Ano86-176, Ano86-177, Ano86-178, Ano86-179, Ano86-180, Ano86-181, Ano86-182, Ano86-183, Ano87-190, Ano87-191, Ano87-192, Ano87-193, Ano87-194, Ano87-195, Ano87-196, Ano87-197, Ano87-198, Ano87-199, Ano87-200, Ano87-201]. involve [Fis84c]. involved [Per83g]. involvement [Kap82b]. ionized [SD81]. ions [SD81]. irradiation [Ler84a]. irrigation [MS81]. ISDN [Bel87b]. Island [DeV81, FAEH84, Lom81b, Zor87b, dS83]. islands [BA83]. isolate [IK84]. isolated [BA83]. Israel [Bro81a]. Israeli [Ros88]. issuance [de 83]. issue [Chr82e, Chr84i, Gra82c, JP85, Ler82k, Wal82a]. issues [AH85, Ano80-245, Bel84f, Bel85f, Ber83a, Chr84j, CN83, Fis84c, Her84, Kli84, Per80c, Per83g, Tru84b, WB86, Wet82b, Zor84a]. Itaipu [de 82, dS83]. itself [Ang80, BE87, KA86, Ler82g, Lur84, Mih81, Mur88b, XM89]. IV [Ano81-152, Lof80b, Lom81a, Mok81f, Whi87]. Ivory [Kah82].

J [Dou85, Wol83]. James [Ear84, Fis86d]. Jan. [Fle85c]. Japan [ABEI84, BMoF^{+}83, Bel85c, Bel87b, Che89d, Fis86e, Fis87d, FW87, Gra84, Jur81b, Jur89a, Kap83a, Kap84d, Kuw84, Ler81h, Miy89, Mo83, Sta86, Tan89b]. Japanese [Bel85c, Bha85, Fre84, Zor86d]. jaw [Fle82]. Jennys [Chr86k]. jeopardy [Ros80]. jets [Ber80a, Chr86k, Spi86]. JIT [CC86a]. job [Gut84d, Per82d, Wal84d]. jobs [Kah82, Kat87, Per87d, Swe82]. John [Dou85, Mus85]. joined [Bel88b]. joining [Gro83]. joins [FM84a]. joint [Ano87-151]. jolts [PFMP88]. Joseph [Wis82]. journalism [Bag84]. joust [Fis84e]. joy [Tru84b]. joys [Chr84k]. Jr. [Gre85]. juggle [PW85a]. jump [Mo83]. Jumping [Bha85, Per82d]. jungle [Gre80]. Jupiter [DJ85]. Just [Ada86a, Chr85f, Zor88b, CC86a, DMW87, Gut85c, MS84, San88c]. just-concluded [San88c]. just-in-time [CC86a]. JVC [Ano87-74, Per88a].

Kalina [Jur86c]. Karen [Fit87a]. Karmarkar [Emm85]. Kay [Wol84a]. Kaypro [Wol84a]. kb [Ber82h, Fis83i, Ber81k]. keep [Ada84b, Ada86a, Fly87, Voe87h]. Keeping [Wal86a, Zor89b, Bel88e, Den81, Gar84]. Kelly [Wol83]. key [Ano84-237, AMG^{+}85, Gut82h, Row82]. keyboard [SAPM80]. keys [Per84a]. Keyworth [Per83f]. kHz [Per87b]. Kilby [Per88c]. Killer [Lem81c, Dut80]. kilobyte [Ano80-195, Ano81-183, Ano83-184]. kilobytes [Ano81-186, Ano81-183, Ano82-193, Ano82-192]. kilometers [SD81]. kiloword [Ano81-184, Ano82-195]. kinds [Gra82b]. kinematic [Nag83]. Kingdom [Gra84]. kits [BKBC85]. km/h [Dup82]. knack [Fag82a]. knot [Mok81h]. know [Nah83, WZB86]. Knowing [Per81d, Per82a, Voe87h]. knowledge [Bel87f, Chr83c, Fis87a, Gra82c, Gut83f, Hec85, Hor84c, Nai83, Sar82]. knowledge-based [Gut83f, Nai83]. known [GGZ86, Gut84e, Mar85]. Korea [Per87f]. Korean [Per87f]. kV [Sti88e]. Kwajalein [Zor89c].

L [BPEB84, Gut89, Pow84]. lab [Per85c]. labeled [Ber80f]. laboratories [Bel85c, Gar84, VH86, Wol83]. Laboratory [Ler81e, Wol84b]. Laboratory/bench [Ler81e]. Labs [Jur88c, Lep81]. Lack [Zor87b]. ladder [Jur82c]. laid [Gut83i]. laissez [Fis83d]. laissez-faire [Fis83d]. Lake [Fis83b]. Lamarr [Kah84]. Lamm [Wol88d]. land [AM86]. Lander [Sti88a]. lanes [Ler82a]. language [Ano82-171, Ano83-185, Voe87a]. languages [Ler82h]. LANs [Ano84-247, Gre85]. lap [Per85c]. laptops [Voe87i]. large [Ada87d, AFG87, Ano80-120, Ano80-121, Ano80-122, Ano80-123, Ano80-124, Ano80-125, Ano81-113, Ano81-114, Ano81-115, Ano81-116, Ano81-117, Ano81-118, Ano82-115, Ano82-116, Ano82-117, Ano82-118, Ano82-119, Ano82-120, Ano83-123, Ano83-124, Ano83-125, Ano83-126, Ano83-127, Ano83-128, Ano83-162, Ano84-128, Ano84-129, Ano84-130, Ano84-131, Ano84-132, Ano84-133, Ano85-126, Ano85-127, Ano85-128, Ano85-129, Ano85-130, Ano85-131, Ano86-115, Ano86-116, Ano86-117, Ano86-118, Ano86-119, Ano86-120, Ano87-113, Ano87-114, Ano87-115, Ano87-116, Ber81g, BS89, FHRW^{+}83, GGZ86, GT82, Jur81a, Mas80c, Per80c, Sha87, Sta86, Sti89c, Sug80b, Tan89a, Wal87g, Wei84]. large-area [Jur81a]. large-LCD [Sti89c]. large-scale [Ano83-162, FHRW^{+}83, Wal87g, Wei84]. largely [Ber80f, DTS86, Mai83]. larger [Che82]. Laser [Bot85, Cal82, Jel81, Bro81b, Fag82c, KS86b, KK84, Miy84, PHG86]. laser-based [PHG86]. Lasers [Fis85c, Jak85, Wel85, BCUY82, Bel88b, Jur81b]. Last [Bal84a, Del80, Ver86, Wal87g, Wil87]. late [Ano85-125, Kel87]. latest [Jur86b]. laugh [MD84]. launch [BE87, Esc86, PHG86]. launched [Cor87d]. launcher [Miy89]. launchers [Wel85]. launches [Gib80]. launching [KM82, Pak85b]. Law [Fis80, Ler83e, Hos83, McC85, Bel89c]. law-enforcement [Ler83e]. laws [Ano86-66, Bel85b]. lawsuits [Bel87c]. layers [SD81]. layout [Kat87, Tri82]. LCD [Sti89c]. lead [ABEI84, CP86, CC87, Fag82c, Gra82c, Gut84g, GZ85, Ina89, Sha80d, ST81]. leaders [Dan82, Kah84, Per81b, Voe87e]. leadership [HK87, Mas80c, Ros80]. leading [BE87, Per87d]. leads [Ber81j]. lean [Per84b]. leap [Per82d]. learn [Bel84b, Ber83a, MD84, Sar82]. learned [Ano86-152, VWZ86a, VWZ86f]. Learning [Her83a, Hor84c, Voe87h, Ada84b, Cun80, Fri84, Her84, Wit84]. learns [Bha85, FW82]. least [Jay86, Tor80]. leave [Bel87f]. lector [Chr84b]. led [Jur82d]. leg [Fis84d]. legacy [Fis83g]. legal [Wet82b]. legislation [Wil80]. leisure [Lom80c]. leisure-time [Lom80c]. lends [Per83e]. lengths [Kat80]. lenses [Ina89]. Less [Jur85b, Ano81-184, Ber81f, CS87, Chr87a, Fis83c, GWF84, Jur86a, Kap80e, Mok82e, Wal87a, ZJ86]. less-than-expected [ZJ86]. lessen [Ber81k]. lesson [Gut82h]. Lessons [Ano86-152, Ber81e, HOD82, VWZ86a, VWZ86f, Wil81, Mai83, Ric88]. let [Fis87b, PW85a]. letting [Bel86a]. level [Bac82, Ber82f, Gra82c]. leveling [Zor84c]. levels [Ano83-186, Che87]. leverage [DeL82, Sum82b]. liability [Per82c]. liberal [Cye84]. liberal-arts [Cye84]. license [Voe88d]. licensing [Per87f]. lies [Mok81g, Per83e]. life [Ano82-157, Ano83-247, Ano84-236, Bal85, Chr84h, Dou83a, FM84a, FZ87a, Fit88d, Ros80, Sti88e, Wil81]. lifelong [Ada84b]. lifespan [Sum82a]. Lifestyle [Bug84, Gar84, Luc84, Win84]. lift [Ano81-152]. Light [HHH87, KK84, MNS81, Mok80a, Wis82]. lighter [Fis83b, Wal87d]. Lighting [Wal85a, FHW85]. lightly [Wol80]. lightness [PW85a]. lightning [PFMP88]. lightweight [Ano80-230, Ano87-74]. Like [Fis87c, Bro81a, Jur85b, Pan87a, Per82d, Sha80c, Voe86c, Voe87a, Zad84]. liked [Chr82l]. likely [Bel84e, Gra82c]. Lilco [Fit88d, Zor87b]. Lille [Kap83a]. limit [Ber82c, Gra82c, GZ85]. limitation [BPEB84]. limitations [Fis80, Jur87a]. Limited [Gev83, Bel85d, Fis87b, Voe89b, Wal87f]. limiting [WAO85]. limits [Per86, SW83]. line [Ano81-177, Bow85, Chr86f, Fis83f, Mai83, MFNL85, Smi87, Sug80a, Ver86]. line-width [Ano81-177]. Linear [Ber82f, Kap84e, Ber80f, Ber81j, Emm85, GS80, Ber81j]. Linear-induction [Kap84e]. linearize [BD83]. Lines [Chr85d, Chr86c, ABEI84, Ano83-51, Ano86-218, Chr80c, Chr80d, Chr80e, Chr80i, Chr80f, Chr80g, Chr80h, Chr80j, Chr80b, Chr80k, Chr80l, Chr81k, Chr81c, Chr81l, Chr81d, Chr81e, Chr81f, Chr81b, Chr81g, Chr81h, Chr81m, Chr81i, Chr81j, Chr82b, Chr82c, Chr82d, Chr82e, Chr82f, Chr82g, Chr82h, Chr82i, Chr82j, Chr82l, Chr82k, Chr82m, Chr83j, Chr83b, Chr83f, Chr83e, Chr83g, Chr83h, Chr83c, Chr83d, Chr83i, Chr83k, Chr83l, Chr84b, Chr84a, Chr84c, Chr84j, Chr84k, Chr84d, Chr84e, Chr84f, Chr84g, Chr84h, Chr84l, Chr84i, Chr85b, Chr85a, Chr85h, Chr85c, Chr85f, Chr85g, Chr85j, Chr85k, Chr85i, Chr85e, Chr85l, Chr86e, Chr86d]. lines [Chr86g, Chr86a, Chr86k, Chr86f, Chr86b, Chr86i, Chr86h, Chr86j, Chr87j, Chr87i, Chr87k, Chr87g, Chr87l, Chr87h, Chr87b, Chr87c, Chr87d, Chr87e, Chr87f, CC86b, Kap83a, Zor89b, de 82]. link [Ano80-195, Wal87h]. linked [Bac82, Gut85b]. linking [Ler84a, SD88]. links [Dor83, Ler81g, Mid80, WSC89]. lip [Sar82]. liquefaction [Fag80c]. Liquid [Per85c, Jur81a]. Liquid-crystal [Per85c]. List [Ano82-158, Ano83-161, Ano86-211, CP86, GBDC86, Mok82f]. literate [Chr83k]. literature [Ano80-31, Ano80-32, Ano80-33]. literatures [Ano80-34]. lithography [Eid81, Fis84g, Lep81, San89f]. lithotripters [CP86]. Little [Gut83i, Chr83c, Fre84, Gut84e, MS84, Zor87e]. little-known [Gut84e]. live [Hor86d, Mor81a]. lived [Ano84-85, Bac80]. lively [Flo82]. living [Gut85a, Per83d, SK89]. Load [Zor84c, Gel81, Hic83]. loads [Ber81a]. Local [Gra82b, Mok81h, Mok82b, Ano84-247, Bel85a, Gut84g, LH87]. local-area [Bel85a]. localities [Ler84c]. locally [Zor85d]. locomotive [KA86]. log [Ano87-182, Ano87-183, Ano87-184, Ano87-185, Ano87-186, Ano87-187, Ano87-188, Ano87-189]. Logic [Ano82-189, de 89, Ano82-250, Ano83-186, Ber80g, Fis85f, FG86, Gut82h, Ham86, Jur82c, Rum85, Swe82]. logic-design [Ano83-186]. London [Ano84-147]. Long [Bel84c, Mid80, Bac80, Bel85a, Fis83c, Kap85f, LH87, Mok82d, de 82, Zor87b]. Long-distance [Bel84c, Bel85a, LH87, de 82]. long-lived [Bac80]. long-successful [LH87]. long-term [Fis83c]. Long-wavelength [Mid80, Kap85f, Mok82d]. longer [Tru83c]. longest [Per88d]. look [Ano80-68, Ano81-230, Ber80f, Cos88, Fag82a, Fis83e, Kap81a, Kid81, Ler83b, Lom80c, Per80b, Per82d, Sha80a, SH85, Voe87i, de 83]. Looking [Bel84e, Bel84d, Cog84, Key83, Mar85, Bor89, Chr84k, Chr84l, Smi87]. looks [Ber80g, Jur83b]. loop [STK82]. lose [Ber81g]. loses [Per88d]. losing [Bow85, Jay86]. loss [Bel88b, Mid80, Wei80]. losses [Kap81c, Kap82d]. Lost [Chr84d]. lots [Bal81, PW83, Per87e, Zor85d]. love [Per84b]. Low [Ano80-41, Bel88b, BD83, HMDF80, Jac82, Ler80c, Ler84g, Mel80, Mid80, Pin80, Ano81-182, Ano82-152, Bel84c, Ber82f, Car80, Dav83a, Gut84e, Jay86, Jur87a, Kap87a, Red80, Voe87g]. Low-cost [Ano80-41, BD83, HMDF80, Mel80, Bel84c, Car80, Gut84e, Jac82, Jur87a, Kap87a, Red80, Voe87g]. Low-head [Ler80c]. Low-loss [Bel88b, Mid80]. Low-power [Jac82]. Low-temperature [Ler84g]. LSI [Ber81h, Ber81i, Ber82j, Fag82b, Fis83i, Mel80, San88a]. lure [Ano86-153, Con86]. luxuries [Riv87]. Luxury [Voe87b]. Lynn [Cor87c]. Lyons [Dup82].

M [AMG^{+}85, Ear84]. M.D. [Sha81d]. machine [Ano84-168, Ano85-125, BA83, Chr87g, KS86a, Ler83c, Ler84e, Sha81e, Sug80a, SW83, Zde83, WZ87]. Machines [Wal83a, Ber81d, Ber82g, CA82, Cor87d, Fis84g, FW87, Kir83, Per87a, SF82, Sta87a, Wil87]. Machining [Gro83]. Macintosh [Gut84e]. madding [Per85b]. made [CC86a, Dav83a, Hor85b, KK84, Sti88a]. magic [Cen89]. maglev [ABEI84, Ano82-251, GS80]. Magnetic [Ler80d, Whi83, Bot83]. magnetism [GBDC86]. Magnetohydrodynamic [Fag80d]. magnetooptical [Whi83]. maiden [LM81]. mail [Mok80a]. main [Ano80-194, Fit88g, Gut82g, Zor89i]. mainframe [Ano81-187, Ano86-152, Gut83c, Liu84, Wal85b, WZB86]. mainframes [Ano83-239, Ber81c, Ber82i, DM85, Her83b, Wal85b, Wal85d, WZB86, WZ87, Zor88e, Zor89h]. mains [Wel85]. maintain [Hor86d, Jur81b, Tru83d, Tru84b]. Maintenance [Liu84, Kap83e, Kap87c, Wal86a, WSC89]. Major [Gut82e, Ler82k, Ano80-37, Ano81-39, Ano82-41, Ano84-86, Ano84-149, Ano85-144, Ano86-133, Ano87-152, Bee83, Ber82g, Fag82b, FHRW^{+}83, Fit88c, Jur86b, MRRS84, Per85c, Per88a, Voe87j, Wal83b, Wal83c]. make [Ang80, Ber81h, Eid81, Fis87d, Kap85d, Lem81b, Lep81, MRRS84, Nah83, Per82d, Per84b, Per84d, Pow84, Wal86a, Wal87d]. maker [JCH80, Jur81b]. makers [Ano86-152, FZ87a, Fis87d, Jur82a, Kel87, Ler83d, WZ87]. makes [Ano81-42, Per82d, Wal82a]. makeup [Aml82]. Making [Ano82-47, Bru87, HGF85, Mur89b, Voe86c, Wal87c, Zad84, AJ85, Bar81, Buc82, Gev83, Ler81c, Moy80, Pak85b, Sta86, VWZ86e]. malfunction [Kap83c]. Malfunctions [Sti81, Gar81]. malignancy [SPM^{+}80]. malignant [SPM^{+}80]. Malo [And80]. Maltese [Fis87c]. Man [Car83, Ler83c, Ano88d, Ano88m, Ros83, Sti88a, Cla80a]. man-days [Ros83]. man-hours [Ros83]. manage [Ber81a, Gel81, Mor81a, Tru84b]. manageable [HM87]. Management [Cun80, Luc87c, Per82a, Ber82g, Car84, CC87, FZ85, FZL86, Gel81, Her83d, Mac84, Pak85b, Pin85, Wol84a, Ano81-151]. management-oriented [Ber82g]. Manager [Wol83, Fag80f, Per87a, Wol84b]. managers [IK84]. manages [Flo82]. Managing [Ano85-163, Ano85-164, Ano85-165, Ano85-166, Ano86-154, Ano86-155, Ano86-156, Ano86-157, Ano86-158, Ano86-159, Ano87-168, Ano87-169, Bel89c, RK80, Wol87a, Wol87b, Wol87c, WC87, BS89, Mac84, Mor81a]. Manhattan [Kap83d]. manipulate [DD85]. manipulated [Ano84-237]. Manipulating [AJ85]. manipulators [Nag83]. manpower [CC86b, Kah81]. manual [Joh84]. manually [Tri82]. manufacturability [Whi87]. manufacture [Per88d]. manufacturer [Ler81b, Zor86d]. manufacturers [CS87, FZ87a, Har83, Per87f, Wal86c, Zor87f]. Manufacturing [Fag81b, Mal84, Pea81, Sti89c, VWZ86e, Whi87, Ano83-107, Bee83, CC87, ELW83, Fag82a, Gut84i, Jur82d, Jur87d, Kap85b, KP86, Lea85, Ler81a, MMB83, Mil87, RD81, Ros83, SD88, Tal83, Wal83b]. Many [Ada86a, Mur88b, Sta87a, Car83, Chr87k, Dav83a, DD86, Fre84, Gau85, JI84, Jur87a, Ler83a, Par84, Per81d, Per83d, Tru83c, Wal87f, Wet82b]. map [Fis86c, HW83]. mapping [Ada84a]. mare [Wal84e]. margins [FP86, PW86]. mark [Jur87d, Lom82, ZJ86]. market [CC86b, FP86, KM87, Per87e, Per88b, Rei80, Sta87a, Voe87i, Wil87, Wol85, Zor86d, Zor89f, Nas83]. marketable [Con86]. marketplace [Bal84a, Fis84e, Fis87d, Gra82b, Hor85b, Kel87, Per84d, Per87b, Per87f, Tru82, Wal86c, Wet82a]. markets [Cor87a, Tus86, Wei80, Zor85f]. marriage [BCUY82, HT83, Hil85, Hou85, Ros83]. Massive [Row82, Ada87e]. Master [Bal84a, Bal84b]. match [Apt85, DM85, GA88, Tri82]. matching [BFB87]. material [Car80]. Materials [BA83, RD81, Bel85c, Dou81, Fls83, Kut83, Lof80b, MR82, ST81, Whi83]. Materials-handling [BA83]. Math [Pyk87, Rau87, Chr82f]. mathematicians [Nah83]. Matrix [NK87, Gut82c]. matter [Ano80-139, Ano80-140, Ano80-141, Ano80-142, Ano80-143, Ano80-144, Ano80-145, Ano80-146, Ano82-153, Ano85-47, Ano85-48, Ano85-49, Ano85-50, Ano85-51, Ano85-52, Ano88d, Ano88m, DM85, Hic83, HTW84, PW86, Ros84]. matures [Cor87b]. maturing [Ano83-162]. Mauchly [Ste81]. maverick [Kid81]. mavericks [Ano88j]. maximum [Ano82-186, Fit87a]. Maxwell [Cen89]. may [Ada84c, AH85, Ada87f, Aml82, Ano84-168, Bac82, Bel86a, Ber82h, Bow85, BFB87, Dav83a, Eid81, Fis84c, FP86, Fis80, FW87, For86, GT82, GZ85, Ham86, Hil85, Hor84a, Hor84c, Jur86c, Kah83, KL85, KA86, KM82, Lea85, Ler80a, Ler80b, Ler80c, Ler83b, Ler84f, LS81, Mar86, Mok81h, MFNL85, Per81d, Per82a, Per82d, Per84d, Per85d, Pic82, PHG86, Pyk87, Rac80, Rum85, Sha80c, Sha84, Sti81, WZB86, Wet82b, Zor85d, Zor86a, Zor86d, Zor87b, Zor87f, Gut82f, Ler84d, Sha80f]. maze [Per83d]. MC68000 [Ano81-185]. MCC [Fis86e]. me [Luc84]. mean [Che82, Dav83a]. means [Ada86b, Ber81f, CC86b, Gel81, Kap81c]. Meanwhile [Ber80g, Lom82, VB86]. measure [Chr81b, HW83, Kap84f]. measured [BD83]. Measurement [Ano80-195, Ano83-185, Fis80, Wal87h]. measurements [Ano87-234, Chu88]. measures [Ano82-193]. measuring [DMW87, Mus89]. mechanical [Jur86c, Kap87c, Mar86, Mil87, Nag83]. Mechanics [Ano81-43]. medalists [Ano80-37, Ano81-39, Ano82-41, Ano84-149, Ano85-144, Ano86-133, Ano87-152]. media [BJ88]. mediate [Mai83]. Medical [CP86, Cor87b, Fle82, Hor84d, Hor85b, Lla81, Mok80f, Pic82, Sar82, Tru83b, Wel84, Bug84, Cor88a, Fit89e, Jak85, Sha81d, Sha81b]. medicine [Fit89f, Jur80, SAPM80, SPM^{+}80]. medium [KM82, Wit84]. meet [Bar81, JT81, Per81b, WM84]. Meeting [Ano80-162, Ano80-163, Ano80-164, Ano80-165, Ano80-166, Ano80-167, Ano80-168, Ano80-169, Ano80-170, Ano80-171, Ano81-153, Ano81-154, Ano81-155, Ano81-156, Ano81-157, Ano81-158, Ano81-159, Ano81-160, Ano81-161, Ano81-162, Ano81-163, Ano82-159, Ano82-160, Ano82-161, Ano82-162, Ano82-163, Ano82-164, Ano82-165, Ano82-166, Ano82-167, Ano82-168, Ano82-169, Ano82-170, Ano83-162, Ano83-163, Ano83-164, Ano83-165, Ano83-166, Ano83-167, Ano83-168, Gut82f, Gut82g, Gut83b, Sha80e, Sha81c, Ano84-247, Sha80f, Wal87f]. Meetings [Ano80-172, Ano80-173, Ano87-71, Ano87-72, Ano87-73, Ano80-229, Ano80-230, Ano87-70, Hef84]. meets [GA88]. megabyte [Ano82-186]. megabytes [Ano80-194, Ano83-187, Ano84-237]. megawatts [dS83]. melt [Fis87b]. member [Lom80c, Lom80e, Per85d]. members [Bha85, Ler83a]. Memories [Fis84g, Ler84d, Ber80b, Ber80g, Chi81, Fit88a, LS81, PL89, San88a]. memory [Ano80-194, Ano81-184, Ano81-187, Ano81-183, Ano82-195, Ano83-187, Ber81k, Par84, Smi87, Zor85e]. men [Chr81k]. menacing [Zor86a]. mentor [Bry84]. menu [HMDF80]. menus [PV89]. mere [SF82]. merge [Fag80b, Tru83e, WZ87]. mergers [San89b]. Mervin [Wol83]. message [EMM82, Lur84]. metal [PD86]. metal-base [PD86]. Metallic [RD81]. meters [Ber81a]. method [Ada87c, Bac82, Emm85, Ler80d]. methods [Gev83, Nag83, Nah83, Pea81, Voe87e, Whi87]. Meting [Ano81-164]. metric [Gol81]. metrication [Gol81]. metro [Kap83a]. Mexico [Ver86]. MHz [Ano80-193, Ano80-197, Ano80-192, Ano81-181, Ano81-182, Ano81-180, Ano82-195, Ano83-187]. Miami [Kap87c]. mice [Chr81k, PV89]. micro [Gut82g, TBH87]. micro-based [TBH87]. microburst [Zor86a]. Microcomputer [Ler81b, Ano81-184]. Microcomputers [Cye84, Riv80, Sha81d, WZB86, Ano83-185, Gut83g, PW84]. microcontroller [Gut83d]. Microelectronic [Pic82]. Microelectronics [Kah83]. microgrip [Ano82-191]. microkids [Kid81]. micrometer [Ano82-191]. Microprocessor [Ano81-177, Ano81-183, Ano81-185, Ano82-194, Ano82-188, Ber80c, Che82, DD86, Fis83e, Gut84h, Sha81e, Sha81c, Sug80b, TBH87, Wei84, WAO85, Zor85a]. microprocessor-based [Wei84, Zor85a]. Microprocessor-controlled [Ano81-177]. Microprocessors [GGPY89, Gut82h, Gut82g, Gut83c, Gut83d, Gut84h, Kap87c, MAN80, Mok80e, PW83, Sha80f, Sha81e, Sha81f, Sha81d, Sha82, VH86, Zor85e, AF87, Ano80-38, Ano80-41, Ano80-39, Ano80-40, Ano80-42, Ano80-43, Ano81-44, Ano81-43, Ano81-40, Ano81-45, Ano81-42, Ano81-41, Ano82-48, Ano82-43, Ano82-44, Ano82-46, Ano82-47, Ano82-45, Ano83-51, Ano83-55, Ano83-54, Ano83-56, Ano83-52, Ano83-53, Ano84-54, Ano84-55, Ano84-56, Ano84-57, Ano84-58, Ano84-59, Ano85-53, Ano85-54, Ano85-55, Ano85-56, Ano85-57, Ano85-58, Ano86-50, Ano86-51, Ano86-52, Fag82c, Jur80, Kap84e, KA86, Per87a, Rau87, STK82, Sha81a, VB86, VWZ86b, WZB86]. Micros [Ano83-55, Gut82f, Ler82f, Ano83-54, Gut83c, Gut83b, Ber81b]. microscopes [Gib85]. Microwave [Ano87-240, Bag84, Bel81, Del80, Flo88, Jur81b, Mok80a, Rog81]. microwave-based [Flo88]. mid [Per83f]. mid-term [Per83f]. midair [Lom80d]. middle [Tan86]. middle-range [Tan86]. midst [Hon80]. might [Ada88e, Bel87c, Cal82, WZ85, Wil80, Zor86b]. Mile [DeV81, FAEH84, Lom81b]. miles [Hil86]. milestone [Gut84e, Kuw84]. Milestones [Ano88m]. Militarization [GT83]. Military [Aml82, Cal82, Lem81b, Lem81c, Ler81f, Lom82, Zor88d, Ada84a, Ada85a, AM86, Ada87a, Ada88e, Ada88f, Ano88o, BA89, Ber81e, Cas81, Che89d, Cor87c, Dan82, Dav87a, DeL82, Den81, Fei82, FT85, Hor85d, Lem83b, Ler82j, Lom80a, Pro81, Spi86, Tus86, Voe87d, Zor87a, Lom81a]. Military/aerospace [Lom82, Lom81a]. millimeter [Bel88b]. millimeter-wave [Bel88b]. million [Ano82-185, Bal81, Ler84h, PW86, TP83, Wal85b]. million-transistor [TP83]. millionaires [Ano88j]. milliwatts [Bot85]. Mind [AW85, Ano82-153, Ano84-168]. Mind-boggling [AW85]. mine [Del80]. Miniature [Asa87]. Minicameras [Bag84]. minicomputer [Ano86-152, Red80, Voe88a, WZ87]. minimize [Bel87c, Bug84, HOD82, Wal84d]. Minimizing [Buc83, Gar81]. minimum [Bel89c]. minimum-risk [Bel89c]. Minis [Ano83-239, Her83b, Wal85b, WZB86, WZ87, Wal85d, Zor88e, Zor89h, Ber81c, Ber82i]. minisupers [WZ87]. Minnesota [Mai83]. Minorities [Tru83c]. mirror [San88c]. misdirected [Mar85]. misfire [Ada87e]. Misreading [KM87]. misses [Per86]. Missile [Zor86b, Ada84a, Ada85a]. missiles [Aug82]. Missing [Ano87-226]. mission [BM81, EMM82]. missions [PHG86]. misspelling [Ano87-100]. misunderstanding [DMW87]. misuse [Buc83]. misused [Bel87c, Fis80, Win84]. Mitaka [Kuw84]. mitigate [Esc86]. Mitigating [LP80]. Mitsubishi [Bha85]. mix [DM85]. mix-and-match [DM85]. Mixed [Chr85g, Ano88k, Bel88c, Fis85f, Wal84c]. Mixing [NS80]. mixture [Jur86c]. MLS [Lom80a]. mm [Per87b]. mobile [Bel87b, CN83, Hil85]. mock [Ada85a]. mode [Bot85, Kec83, Mok82d, Syk89]. Model [FM84b, Beg83, Bro81a, Kuw84, Vae86]. Modeled [Voe87b]. Modeling [Ler81g, Sin87, Mil87, Sug80b, Wal84d]. models [Ber81h, Jur84, Jur85b, Jur86a, Jur86b, Ler82c, Mos81, Nag83, Voe87j, Wal87f]. Modern [Ant87, DeL82, Doy81, JI84, Wal82d, Wis82]. modernize [Ler82a]. modest [Cau82]. modified [AF87, Ler80b]. modifying [Wil80]. modular [Mok82a, NS80, VB86]. modulated [MNS81]. modulation [Jur87c]. module [Ano80-196]. modules [Ano83-185, Jur82c, MMB83]. Molecular [PC80, Ano81-236, Con86]. molecules [PC80]. money [Fis83c, FZ85, HK87, VWZ86c]. money-saving [Fis83c]. monitor [Kap87c, Pro81, Sha81d]. Monitoring [Hor85c, Zor86b, GP81, Hor86a, Kap87a]. monochrome [Bal85]. Monolithic [Ano82-190, BCUY82, Bar81, Mok82c]. month [GWF84, Kap83d, Wol85]. Moon [Sti88a, Cla80a]. morass [Tri80]. MOS [Che87, Fis87d, Mur85]. MOSFETs [Lep81]. Most [Fis86b, Wal83d, You84, Ano83-169, Ber80a, Ber81g, Ber82b, DMW87, Gut84l, Ham86, Hor84a, Jur86d, Lom80f, Pyk87, SA85, Ver86, Wol84b]. motion [Cor87b]. motivate [Mac84]. motivator [Wol83]. motor [Fag82c, GS80, Kap85d]. motors [Ano82-191, Ano85-125, FEK80, Kap84e]. Motto [CC87]. mounted [KP86]. mounting [Per86, Tur87]. move [Bha85, FEK80, Gut83g, MD84, Per85d, Tru83e]. movement [MD84]. movements [Fle82]. mover [Kap80c, Kap87c]. moves [Hil85, Voe87g, Wal84d]. Moving [Zor85d, Zor89d, Mar86, Tus86, VH86, KG83]. much [Bel85b, Dou85, Fis87a, Fit89f, Hic83, Mor81a, PW85a, Sta87b, Wal83a]. much-sought-after [PW85a]. Multibus [Ano81-184, Ano83-187]. Multibus-compatible [Ano81-184]. multicolored [Ano82-191]. multidisciplined [Fag82a]. multifunction [Jur81a]. multimedia [OK86]. multimegabyte [Che86]. multimeter [Sha80c]. multimode [Kec83]. Multiple [Chr83a]. Multiplexed [Jur86a]. multiplier [CN83, DeL82]. multiprocessing [Zor85e]. Munich [Ano84-146]. Murphy [Bel89c, Mur85]. Musa [Mus85]. museum [Fre86, Ano84-146, Ano84-147, GB85]. music [Lem83a]. must [Ano84-237, Bel87c, DD86, Gel81, Gut84l, Hic83, HM87, Mok81g, Per81d, Voe87e, Wal84a, Wal87d, Wet82b, Whi87]. mV [Ano81-180]. MW [Kap83d]. MX [Ada84a, Chr85a]. mysterious [DJ85]. Mythinformation [Win84].

N [Fru83, For86]. name [Ano87-100, Per87e, Whi85]. names [Zor84b]. NASA [BE87, Chr86b, Esc86, Log83, Lom82, SF82, Wil81]. nation [Chr87i, Gut83e, Kuw84, Lem81b]. National [Dup82, Per80c, Ros80, Wei80, BMoF^{+}83, Cur83, HK87, Hor85d, Wal83e, Wal85c]. nations [Gut83i]. NATO [Ler82k, Per82b]. natural [Wal87h]. nature [Hor86b, HTW84]. navies [Gut82i]. navigation [Voe87d]. navigational [Voe87d]. Navstar [Den81]. Navy [Ano87-228, Lak82, Wal82d, Ada87a]. near [Fis86d, LM81, Per86, Zor85a]. near-misses [Per86]. near-optimum [Zor85a]. near-perfect [LM81]. necessary [Fis83d, Tru83d]. necktie [Fis87a]. need [HK87, Hor84a, Hor86d, Jur82d, Kec83, MS84]. needed [Fag81a, Kap80f, Lem81a, Ler83a, Nag83, Rog81, Tru83d, Zor85d]. Needs [Dou83b, Dav87a, Gra82b, HK87, Jur87c, LM81, WM84]. neglected [Gut84l]. negotiation [Fis84c]. nets [Dor83, Gra82b, Mok81h]. network [Bel88a, Cor84, Fis84b, Hil83, Kap84d, Ler82c, Wal83e, Pan87a]. Networking [AMG^{+}85, Han89, Sti88d]. Networks [Gut83e, Jur86a, Ano84-236, Ano84-247, Bel84c, Bel85a, BS89, Cor87c, Fri84, Gra82b, Gut84g, GZ85, LH87, Mel80, Mok81h, Mok82b, Per84e]. Neurocomputing [HN88]. neurologists [Tru83b]. neutralized [Cal82]. Neutralizing [Gar81]. neutrons [Fis84h]. Nevada [Per85b]. never [BFB87, Gut83c, JI84, Zor85b]. nevertheless [ZJ86]. new-generation [Dou83a, Dou83b]. newcomers [Bow85]. newly [Hor85d]. News [Ano80-198, Ano80-199, Ano80-200, Ano80-201, Ano80-202, Ano80-203, Ano80-204, Ano80-205, Ano80-206, Ano80-207, Ano80-208, Ano80-209, Ano80-210, Ano80-211, Ano80-212, Ano80-213, Ano80-214, Ano80-215, Ano80-216, Ano80-217, Ano80-218, Ano80-219, Ano80-220, Ano80-221, Ano80-222, Ano80-223, Ano80-224, Ano80-225, Ano80-226, Ano80-227, Ano80-228, Ano81-188, Ano81-189, Ano81-190, Ano81-191, Ano81-192, Ano81-193, Ano81-194, Ano81-195, Ano81-196, Ano81-197, Ano81-198, Ano81-199, Ano81-200, Ano81-201, Ano81-202, Ano81-203, Ano81-204, Ano81-205, Ano81-206, Ano81-207, Ano81-208, Ano81-209, Ano81-210, Ano81-211, Ano82-196, Ano82-197, Ano82-198, Ano82-199, Ano82-200, Ano82-201, Ano82-202, Ano82-203, Ano82-204, Ano82-205, Ano82-206, Ano82-207, Ano82-208, Ano82-209, Ano82-210, Ano82-211, Ano82-212, Ano82-213, Ano82-214, Ano82-215, Ano82-216, Ano82-217, Ano82-218, Ano82-219, Ano83-188, Ano83-189, Ano83-190, Ano83-191, Ano83-192, Ano83-193, Ano83-194, Ano83-195, Ano83-196]. News [Ano83-197, Ano83-198, Ano83-199, Ano87-182, Ano87-183, Ano87-184, Ano87-185, Ano87-186, Ano87-187, Ano87-188, Ano87-189, Ano87-203, Ano88k, Bag84, Chr80f, Chr80g, Chr85d, WBK87]. newsman [Joh84]. Next [Bel89d, FW87, Gut83f, Gle83, JT81, Mok82d, Pea81, Per83e, Wal83a, XM89]. Next-generation [Bel89d, FW87, Gut83f, Mok82d]. niche [Cor87a]. nine [Her84]. NMCS [Ber81j]. NMOS [Fis83i]. NMR [Hor85b]. No [NK87, Wal84f, Ada87c, Ano82-191, Ber80e, Chr80b, Chr82h, Hec85, Ler82k, Ler84h, SW83, Tru83c]. no-downtime [Ber80e]. noise [Ber82f, Jur87c, Pin80]. nonconformist [JI84]. nondegreed [FP84]. nondestructive [Cau82, Mas80b, MR82]. none [Apt85]. nonetheless [FG86]. nonexistent [GWF84]. nonlinearity [SW83]. nonoptical [Fis84g]. nonprogrammers [Ler82h]. nonsurgical [Jak85]. nonvolatile [BG89]. Nor [Wal87h]. Norden [Sea89]. normal [SD81]. north [Zor87b, Fis86d]. northeast [Zor89d]. notch [DTS86]. notes [Ano83-205, Ano83-206, Ano83-207, Ano84-194, Ano84-195, Ano84-196, Ano84-197, Ano84-198, Ano84-199, Ano85-191, Ano85-192, Ano85-193, Ano85-194, Ano85-195, Ano85-196, Ano86-185, Ano86-186, Ano86-187, Ano86-188, Ano86-189, Ano86-190, Ano87-203, Ano87-204, Ano87-205, Ano87-206, Wal87e, WBK87]. Noteworthy [Ano84-146]. notification [Fis86c]. Nov [Ano84-247]. novel [RD81, Sha81f]. November [Far86]. nows [Ler80d]. NPF [Ros80]. NRC [Ano81-152, Eis84]. ns [Ano80-192, Ano81-181, Ano82-185]. ns/division [Ano81-180]. NTIA [Per80c]. Nuclear [Bot83, Chr84e, DeV81, FAEH84, Fit89a, GP81, HOD82, Kap83c, Ler82k, Lla81, LP80, Lom80e, Lom80f, Per81e, Rub80, Tel82, TD84, AW85, Ano80-41, Chr81m, Eis84, Fis85e, Fis86f, FZ87a, Her82, Hor84c, Hor86d, Kap82c, Kap82a, Kap83e, Ler81f, Ler82d, Ler84g, Lom81b, Mil81, Pro81, Zor85d, Zor87b, Zor89j, Eis84]. Nuclear-electric [TD84]. nuclear-plant [Kap83c]. nuclear-power [Fis85e]. Nuclear-power-plant [Kap83c]. Nuclear-reactor [TD84]. Number [Kap84b, FWVH86, Fri84, Gut83d, Per86]. number-crunching [Gut83d]. numbers [BKBC85]. numerical [Ros83]. numerical-control [Ros83]. numerous [Sha87]. nursery [Hor85c]. Nurturing [Ler84h].

O [Ano80-196, Ano83-185]. object [Wal84g]. object-oriented [Wal84g]. objective [GWF84, Zor84b]. objectives [Fis86e, Kut83, Per80a]. objects [Ano83-248, DD85]. obsolescence [Mur88b, Tru84b]. obsolete [Chr85e, Fis86f]. obstacles [Fag80d]. obvious [BE87, Ber81g, CC86a]. occur [Ham86]. occurred [Ano88p]. ocean [Ano84-252, HW83]. oceans [HW83]. Oct. [Ano84-247]. OEM [Kap87a]. off [DD86, Kap80b, Spi86, VWZ86d, Zor86c]. offensive [Fit87c]. offer [Gut84i, Ler82j]. offered [Apt85]. offering [Cye84, Sha80c]. offerings [Ber82g, FM84b]. offers [Ano82-193, Ano82-189, Ano83-187, Gut84f, KS86b]. Office [IK84, Wal83d, Ber82b, Dor83, Far86, Gut83f, Mok81h, Mok82b, Per83a, Gre80]. officials [Fei82, Zor87e]. offing [CS87, Fis83b, Hil85, Jac82, Kir83, LH87, VWZ86b]. offs [Ano83-229, Ano83-230, Ano83-231, Ano84-224, Ano84-225, Ano84-226, Ano84-227, Ano84-228, Ano84-229, Ano85-221, Ano85-222, Ano85-223, Ano85-224, Ano85-225, Ano85-226, Ano86-219, Ano87-227, Ano87-228, Dav87b, Dav87c, Dav87d]. Offshoot [Bel87a, Che86]. often [Fis84c, Fis86b, IK84, KL85, Per87a]. often-daunting [Per87a]. ohm [Tay89]. oil [CA82, Kap80e, Ler84g, You84, Zor85d]. oil-dependent [CA82]. oil-fired [Ler84g, Zor85d]. oil-generated [You84]. old [Bal85, Cor87d, DJ85, Fag80e, Fis86f, FZ87a, Fis87c, Gut82h, Hor84c, Ler83a, Mur89a, Mur89b, NS80, Tru84b]. old-time [Mur89b]. Oliver [Nah83]. Olympics [Cor84]. ominous [Per81b]. omitting [PW83, Wal85e]. omnipresence [Luc84]. omnipresent [Sug80b]. on-board [Jur83a]. on-chip [Bel86a]. once [Per83g]. One [Chr86h, Chr87k, Ler84h, Ros83, Sti88a, BMR88, BFB87, FHW85, GWF84, Hec85, Her84, KM82, LH87, Lom80f, MAN80, Mok81g, Per82a, Tru83e, Wal87a]. one-month [GWF84]. ones [Fag82a, Wal83a, ZJ86]. Only [Kel87, VWZ86c, Bug84, Cal82, Del80, Fis83b, Gar84, Sta87a, Wal87a]. op [Ber82f]. open [Zor89b, Vae86]. Opening [Chr84f, Sha80f]. openings [WZ85]. operate [Ler84b, Wal85b]. operates [Ano82-187]. operating [FZ87a, Fle85c, Gut84h, Kap82c, OK86, Voe86d]. operation [Ano82-185, Che82, Kap83e, Kap85d, Liu84, Smi87, Ver86, Zde83, Zor87b]. operational [Fag80d]. operations [Ano80-192, Ano82-185, Fis84a, Gro83, Rau87, Sti81, Wal85b, Wil87, Per81e]. operator [Ano81-150]. opinions [Ano82-248]. Oppenheimer [Cog84]. opponent [Aml82, Mok80e]. opportunist [WZ85]. opportunities [Zor85f]. optic [Ano83-141, Bel84c, Bel85a, Bel87b, GZ85, Hec85, Kap85f, KS86b, KK84, MNS81]. Optical [Bel86b, Cas81, For86, Fre88b, GBDC86, HHH87, Mar86, Mid80, Shu89, Ano86-192, Ano86-145, Bel85c, BA86, Bel87b, Bel88b, Hec85, Kap83b, KK84, LH87, ST81]. Optical-fiber [GBDC86]. Optics [SK89, Bac80, Bel88b, Gre85, Gut82f, GZ85, Kap83b, Mok81a, Mok82d, Voe87g]. optimistic [Kap80b]. optimists [Ada84c]. optimizes [Kap87c, OK86]. Optimizing [Wal87d]. optimum [Zor85a]. option [Ano81-182]. options [Bel84e, BFB87, CC85, Gau85, Jur82a, Rau87]. optoelectronics [BCUY82]. oral [Bel84b]. Orbit [Mok80a, Ano83-248]. ordeal [Lom81b]. order [CC86a]. organisms [Con86, Ler84a]. organization [Fis86e, Per84b]. organize [Ler83e]. organizers [Fag82a]. Organizing [Kli87]. Orient [Voe88e]. oriented [Ber82g, Red80, Wal84g]. original [Fis86f]. Origins [Kli84, Gut89, Kah84, Ros84]. orphans [Ler84h]. oscilloscope [Ano80-193]. Oscilloscopes [BFB87, Ano81-180, Sha80c, Voe87g]. OTA [Gre80]. Other [Ano80-229, Ano80-230, Fis83i, Fis86d, Fle82, Hef84, Hor84c, Hor85d, Tus86, Voe87a, Wal87h, WM84, Whi87]. otherwise [BA83]. Our [Chr86i, Jur82d, PWZ^{+}86, Ano88p, Chr81l, Chr85k]. ourselves [Ano80-68, Fag82a, Kap81a, Ler83b, Lom80c, Per80b, Per82d, Sha80a]. oust [Bor83]. Out-sourcing [Fly87]. outages [BS81, Mas85]. outer [Cla80a]. outgrowth [Fis86a, Jur87a]. Outguessing [Ano83-56]. outlet [Pan87a]. outlook [Ada84c, AMG^{+}85, Bel84d, Fis84b, Lom80e, Zor87e]. outperform [Kec83]. output [Jur82c, Sha84]. outstanding [Ano84-147, Jur86d, Ler84d]. overcome [ELW83, MRRS84]. Overcoming [FP84]. overcrowded [Per86]. overlapping [LS81]. overlapping-triangle [LS81]. overload [Sta87b]. Overlooking [Ber81g]. overseeing [Ler84h]. oversight [Zor88c]. overstress [Ant87]. overwhelm [Sta87b]. overwhelmed [FWVH86]. overwhelming [CC85]. Ovonic [Ano80-249]. own [Bel87f, Gre80, Lem83a, Per81c, Per81e, Per82a, Rab81, Sug80c]. owner [Mor85]. owns [Hos83].

p [For86]. p-i-n [For86]. pace [WZ87]. Pacemaker [Cor87b]. pacemakers [Hor85b]. Pacific [Bel87b, Zor88a]. package [BCUY82, Ber82g, Bow85]. packages [Ber82e]. packaging [Bow85]. packed [Bot85]. packs [Wel85]. pact [Ada88g]. pacts [Hor86a]. pad [Esc86, Per85d]. padlocked [VW86]. pair [Gra82b]. Palo [Pak85a, Pak85b]. panacea [Wal84f, Win84]. panel [Apt85, Sil89, Tan89a, Wal83e]. panels [Wol80]. Paper [Ano80-36, Ano81-38, Ano82-135, Ano83-48, Ano80-148, Ano81-134, Ano81-135, Ano81-136, Ano81-137, Ano82-136, Ano82-137, Ano82-138, Ano82-139, Ano83-145, Ano83-146, KG83, Par84]. papermill [Luc83]. Papers [Ano83-204, Ano83-200, Ano83-201, Ano83-202, Ano83-203, Ano84-52, Ano84-181, Ano84-182, Ano84-183, Ano84-184, Ano84-185, Ano84-186, Ano84-187, Ano84-188, Ano84-189, Ano84-190, Ano84-191, Ano84-192, Ano85-46, Ano85-179, Ano85-180, Ano85-181, Ano85-182, Ano85-183, Ano85-184, Ano85-185, Ano85-186, Ano85-187, Ano85-188, Ano85-189, Ano85-190, Ano86-49, Ano86-172, Ano86-173, Ano86-174, Ano86-175, Ano86-176, Ano86-177, Ano86-178, Ano86-179, Ano86-180, Ano86-181, Ano86-182, Ano86-183, Ano87-190, Ano87-191, Ano87-192, Ano87-193, Ano87-194, Ano87-195, Ano87-196, Ano87-197, Ano87-198, Ano87-199, Ano87-200, Ano87-201]. parachute [Per82d]. Parallel [CC87, Her83b, HM87, Lem81a, Wil87]. parallel-architecture [Her83b]. parameters [BD83, Dav87a, Smi87]. PARC [Pak85a, Pak85b, PW85c]. paring [Jay86]. Paris [Dup82]. park [GT82]. Parkinson [Zor89e]. Part [Voe87h, Wal87d, Whi87, Bel84e, Bel84d, Chr80g, Doy81, Hor86a, Kap87a, Ros83, Rum85, Wal84a, Dav87a, Kat87]. participant [Bel85b]. participation [Kli84]. particle [Fis85c, HTW84, Lem81c]. particle-beam [Lem81c]. particularly [DeL82, Jak85]. partitioned [Ano80-195]. partner [Ano84-168]. parts [Ano83-244, Doy81, Fly87, Kut83, Mar86, Mas80a, VWZ86e]. pass [Zor88d]. passengers [ABEI84]. passing [Gra82b]. passive [FHW85]. past [Chr86i, Joh84]. pastures [Per89b]. patent [Har83]. patentable [Rac80]. Patents [Ano87-165, Ano87-167, Rac80, Ano87-164, Ano87-166, Bel87d, BBJ87]. path [BA86]. Paths [DMW87]. patients [Fis83e]. Patrick [Fag80f, Lom80e]. Pattern [Cas81]. Paul [Gre85]. paved [Fag80d]. paycheck [Per82d]. payload [LM81]. payloads [KLCPT80, KM82]. pays [Zor86c]. PBXs [Bel84c, Gre85]. PC [Ano86-152, BFB87, Voe86c, VWZ86b, VWZ86c, VWZ86d, VWZ86e, Voe87c, Wol84a]. PCC [KW86]. PCjr [KM87]. PCs [VWZ86e, Voe87g, Voe87i]. PDP [Voe88a]. PDP-8 [Voe88a]. Peace [GT83, Kap82b, Kap82a, Hor86a, Rub82]. Peaceful [Ada89b, Per89b]. Peacekeeping [Ada86b, Chr82a, Kap82b]. peak [Gel81, Zor84c]. peak-demand [Gel81]. Peer [Chr81g]. peers [Kap81a]. pejorative [Chr87a]. Penetrating [Ada87c, BPEB84]. pens [Ano82-187]. Pentagon [Ada88b, Sum80]. people [Flo82, Kap87c, Wol83, Zad84]. perceived [Mor81b]. percent [Jur86c]. perceptions [Bel84a, San88c]. perfect [LM81, Mok81i, Per85d]. perfected [Wel84]. Perfecting [Apt85]. perfection [Kap83f]. perform [Zor85e]. performance [Ano81-185, Apt85, Bel81, Che87, DD85, DMW87, Jur82a, Kap80b, Lea85, Lem81a, Rau87, Riv87, Wal85b]. performs [Ano82-185]. perhaps [KK84]. period [Wol83]. peripheral [Ber80c, Mok82e]. Peripherals [Ano81-185, Sti89e]. permalloy [LS81]. permalloy-chevron [LS81]. permanent [WM84]. permanently [Hil86]. permit [BKW83, Jur86c, Ler84b, MR82, Per85b, Tal83]. permits [Ano81-182, Ano82-187]. permitting [FZL86]. perplexed [Nah83]. persistent [Fag80f]. Personal [Gut83g, Mok82e, VWZ86f, Voe87i, Wer87a, ZJ86, Ber83b, Ber80d, Fri84, Gut84e, Gut84g, Gut84j, Gut85b, Gut85d, Hef84, Mok80c, Nai83, Sha80a, VWZ86c, Voe88a, Voe88f, Voe89c, Win84, Mok81d]. personnel [Zde83]. perspective [Bel89a]. persuasive [Fag80f]. pervades [Rub80]. pervasive [Chr87g, Wal84b]. pet [Cye84]. Phase [Chu88, Gut84l]. phenomena [ST81, Sug80b]. phenomenon [Tru83c]. philatelic [de 83]. philatelist [de 83]. Philips [Cor88b]. philosophies [Bel85b]. philosophy [Pak85b, Voe87e]. Phone [Wal87b, Mok80a, Zor89b]. phones [Hil85]. phonographs [Ano85-125]. Photo [Ano84-193]. Photo-Electric [Ano84-193]. photodiode [For86]. photoelectric [Bro81b]. photoeonductor [For86]. photography [Ano84-193]. Photons [Ano86-184]. Photovoltaic [Fis84e, MS81]. Photovoltaic-cell [Fis84e]. Photovoltaics [Bac80, Car80, Lof80b, Lof80a, Wol80, Zor89f, Wet82b]. physical [Ano83-187, Bel81, Bot83]. physicians [Fis83e]. physics [Ant87, RK80]. Pickard [Dou81]. picking [HN88]. pickup [Miy84]. picoseconds [MS84]. PICs [Rum85]. Picture [Zor88b, Apt85, Chr83l, Per80b]. Picturephone [Ano82-252]. pictures [Bot83]. pie [Kap85c]. piecemeal [Fis84a, Gra82a]. pieces [Kap85c, Wal83d]. pilots [Fis86b, FWVH86]. pink [IK84]. Pinning [Ada88d]. pinpointing [Ada87c, Sti81]. pioneer [Kat80]. Pioneered [Wei84, Cor87c, Per88a]. pirates [McC85]. pitfalls [Ano86-191, DMW87, Per81d]. pitted [Miy84]. place [Apt85, Esc86, Gra82a, Jur85a, Per83c]. places [Zor87f]. plan [Eis84, Ler82a]. planar [Mok81g]. plane [Per82d]. planes [FWVH86]. planet [DJ85]. planned [de 82]. planners [Dav87a, Fag82a]. Planning [DJ85, Hil85, Fle85b, Mas85, Zor87b]. plans [Esc86, Gut83i, JCH80, Wet82a, You84]. plant [And80, Bro81a, Fag80d, FAEH84, Fis86f, Kap83d, Kap83e, Kap83c, Kuw84, Ler81f, Zor87b]. plants [Ano81-230, Fis84h, Fis85e, Gau85, Her82, Ler84g, Sin87, Wal87a, Zor85d]. Plasma [RK80, Gar81, Lea85, Per85c]. plastic [Sho88]. plated [Voe86a, Voe86b]. platform [Gib80]. platforms [Wer87a]. play [Fag82b, FHRW^{+}83, Kap84e, Per87a]. playback [Cle82]. played [Kah84]. players [Asa87, BMoF^{+}83]. plentiful [Zor85d]. plethora [Gut83h]. Plodding [Wal82d]. plot [HW83]. plots [Kir80]. Plotter [Ano82-191, Ano81-177, Ano82-187]. plowshares [Sti89b]. Plugging [Luc87d, Pan87a]. plume [Fis86c]. plunge [Sta87a]. plus [Bel88c, FAEH84, HOD82, Lom81b, Rog81, de 82]. Plying [Gut82i]. Pocket [Ano82-192, Sho88]. Pocket-size [Ano82-192]. point [Ano82-185, KSB87]. Poised [Fis86a, GZ85, Voe87j]. policeman [Per81e]. policies [Chr81h]. policy [Ano80-245, Bel84f, Fis83d, Fru83, Gra82a, HK87, Ler83b, Lom81c, Per83f]. Politburo [Fre84]. political [AH85, Gre80, Kap82b, Kap83e, Tri80]. Politics [Gre80, Ros80, Wei80, Wil80, FG86, Fru83, Hor86a, Kap85c, Tri80, Zor87c]. poll [Gut84d, Gut84b, Gut84c, Gut84a, San88c]. pollution [Riv80, Zor87e]. polycrystalline [Lof80b]. polymer [MO89]. pond [Bro81a]. poor [Car84]. poorly [Pyk87]. popular [Bry84, DMW87, Gut84k]. portable [Bel88d, Per85c, Wal87h]. portend [MAN80]. portends [Bro81b]. portfolio [Ano87-202]. Portia [Per85d]. portrait [Lom80c]. pose [Ber81e, Fis86c, Per81a]. posed [Per83a]. poses [Ada85c, Ler82e]. posit [Ler84a]. position [Aml82, BD83, Bow85, Cau82, Fis84e]. positioned [Fis86e]. possible [Ano81-151, Dou81, Ham86, Per81d, Pow84, Sti88a, Tor80]. Post [Ric88]. Post-quake [Ric88]. Postal [Rab81]. PostScript [Per88e]. potent [AJ85]. Potential [Lem81b, Den81, HOD82, KSB87, Ler80c, OK86, Zor85a]. potentially [Bee83]. potentials [Tru83b]. Power [Ang80, Bal81, Ber81a, Blu80, Bro81a, CA82, Che82, Che87, DM85, Fag80e, Fis83d, Fis83c, Fis84f, FZ85, FZL86, FZ87a, Gel81, Hic83, Kap80e, Kap81c, Kap82d, Kap82c, Kap83d, Kap83e, Kap83c, Kir83, Ler80c, Ler80d, Lom80f, Lom81b, MFNL85, RK80, Rog81, Row82, Rum85, STK82, Wet82b, Zor87c, de 82, Ada85c, Ada88a, Ano83-246, Ano87-234, Bel84a, Bot85, CS87, Chr81m, Con86, Dav83a, DeV81, DS81, Dou85, Fag80d, Fis84h, FAEH84, Fis85e, Fis85f, FG86, Fis86f, FEK80, Gau85, HOD82, Her82, Jac82, Kap80f, Kap83f, KZG^{+}84, Ler81f, Ler82k, LP80, Lom80e, Luc87c, Mai83, Mas85, MS81, MO89, Mun88, Per81e, PZ86, Ric88, Rub80, Sin87, ST81, Syk89, Tan89a, TD84, TBH87, VWZ86c, Voe87i, Wal84c, Wel85, Wol88a, WZ88, WZ89, Zor84c, Zor85d]. power [Zor87b, Zor89a, Zor89d]. power-handling [Rum85]. power-hungry [Tan89a]. power-industry [Mai83]. Power-line [MFNL85, Mai83]. power-logic [FG86]. power-packed [Bot85]. power-supply [Ada85c, CS87]. power-system [Kap82c]. Power/energy [Ang80, Ber81a, Bro81a, CA82, Che82, Fag80e, Fis83d, Fis83c, Gel81, Hic83, Kap80e, Kap81c, Kap82d, Kap83d, Kap83e, Kir83, Ler80c, Ler80d, Lom80f, Lom81b, RK80, Rog81, Row82, STK82, Wet82b, de 82]. powered [Ano87-240]. Powerful [VB86, Ano83-143, Bot83, Chr88b, Fri84, Gev83, Her83b, Nah83, Sta87b, VWZ86b, Wal84f, WZB86]. powerhouses [Fis86d]. Powers [Pow84]. practical [Ber81h, Dou81, Fis84h, Fis80, Ler80d]. practice [DS81, Fag80e, Kli84, Mad86, Voe87a]. practices [CC86a, Fis86b, Per81d, Voe86d]. practitioner [Wal84d]. pragmatic [Gut84f]. praise [Wal84c]. precise [BD83, Sol88]. precisely [Voe87e]. Precision [Kap84c, Ano87-234, Kap87a, Miy84]. precision-tracking [Miy84]. Precursors [Hec85]. predict [PZ89]. prediction [Bel88e]. predictions [Wil81]. prematurely [KW86]. premier [Kap84f]. prepare [Tru82]. prepared [Esc86]. Preparedness [Dan82]. prepares [Fis83i, Jac82]. Preparing [Tay89, Cun80]. preprocess [BD83]. prerequisite [Kap82b]. presage [Fis84b, Kap83b]. prescribes [Wal86b]. present [Fei82, Fis80, Gut84g, Sha80c]. presentations [Bel84b]. presents [SPM^{+}80]. Presidency [Ano80-245]. president [Car83, Lom80e]. Presidential [Per83f]. press [Blu80]. presses [Lom80a, Wal87g]. pressing [Jur82d]. pressure [Fag81a, GBDC86, MD84]. pressures [Bel87e, Per84c]. Prestel [Rei80]. prestigious [Ano87-234]. Pretty [Zor88b]. prevent [Ant87, Bel87c, KLCPT80, Zor87b]. preview [Ano83-162, Gut82f, Gut82g, Gut83b, Sha80e, Sha80f, Sha81c]. previews [Ano80-162, Ano80-163, Ano80-164, Ano80-165, Ano80-166, Ano80-167, Ano80-168, Ano80-169, Ano80-170, Ano80-171, Ano81-164, Ano81-153, Ano81-154, Ano81-155, Ano81-156, Ano81-157, Ano81-158, Ano81-159, Ano81-160, Ano81-161, Ano81-162, Ano81-163, Ano82-159, Ano82-160, Ano82-161, Ano82-162, Ano82-163, Ano82-164, Ano82-165, Ano82-166, Ano82-167, Ano82-168, Ano82-169, Ano82-170, Ano83-163, Ano83-164, Ano83-165, Ano83-166, Ano83-167, Ano83-168]. previous [Bel87f]. Previously [Gut84l]. price [Ada88e, Apt85, Chr88a, FP86, PW85a, Tan86]. pricing [Cun80]. pride [Cur83]. primary [Fag80e]. prime [Ano83-162, Ano84-247, Kap80c]. primer [Syk89]. primitive [Per87e]. Princeton [Fis84h]. principal [VWZ86d]. principles [Fis80]. print [Chr85j]. printed [Tul83]. printer [Gut82c, NK87, Wol89]. printers [Hef84]. printing [WM84]. prints [Per88e]. prior [Mih81]. priorities [Wal82b]. privacy [Gut83f, Hon80, Ler84f]. Prize [Ano80-36, Ano81-38, Ano82-135, Ano83-48, Ano84-52, Ano85-46, Ano86-49, Joh84]. prize-winning [Joh84]. probe [Ano82-247, Ano83-107, Ano83-240, Ano83-241, Ano83-239, Ano84-240, Ano84-238, Ano84-239, Ano85-233, Ano85-234, Ano85-235, Ano86-231, Ano86-232, Ano86-233, Ano87-232, Ano87-233, EG82, Kap84f]. probes [Lom80a]. Probing [Ada85c, Ano83-51, Fis83e, Fit89b, Mor81b, DJ85, Fre86, Gib85]. problem [KL85, Lom80f]. problems [Ber83a, Ber81k, Fis84c, Gol81, Hor84c, HM87, Jur83b, Jur86d, Kah81, KSB87, Ler82b, MRRS84, MFNL85, Per81a, PW84, SW83, Wal86a, Wil87, Zor87f]. procedures [Ber81h, Kap83e, Pea81]. proceed [Voe88b]. Process [Zde83, Ano84-168, Fag80c, Fag82c, Her84, KP86, Mih81, MFNL85, SF82, Sha84, Wal83b]. processes [Ano87-164, Ano87-165, Ano87-166, Ano87-167, Bel87d, BBJ87, CC87, Fag80b, Tri80]. processing [AF87, Ber82j, Cas81, Che89a, Fag82b, Fis87d, Jur87a, Kuw84, Lem81a, MAN80, Sha80c, Sha81f, Wal87f]. processor [Ano80-194, Ano80-195, Ano82-185, Ano83-184, Mok82c, Sha81f]. processors [Ano82-186, Ano83-187, Ber82e, Gut83d, HM87, VH86, Wal84f]. proctors [Liu84]. procurement [Ada88e, Mcc88]. produce [Cor88b, Fis86d, Gut84e, Kah83, Tri82, TD84, Wal85d]. Produced [Voe87i, Kel87]. producers [Bag84, FZ85, JT81]. producibility [Whi87]. producing [Eid81, Fis84d, KW86, MS81]. product [Ano80-186, Ano80-187, Ano80-196, Ano80-188, Ano80-189, Ano80-193, Ano80-190, Ano80-191, Ano80-197, Ano80-192, Ano80-194, Ano80-195, Ano81-178, Ano81-177, Ano81-186, Ano81-184, Ano81-187, Ano81-185, Ano81-179, Ano81-183, Ano81-181, Ano81-182, Ano81-180, Ano82-185, Ano82-193, Ano82-195, Ano82-191, Ano82-192, Ano82-190, Ano82-189, Ano82-194, Ano82-187, Ano82-184, Ano82-186, Ano82-188, Ano83-187, Ano83-182, Ano83-185, Ano83-184, Ano83-183, Ano83-186, Bag84, Ber82h, CC86b, Cun80, Dav87a, Gut82h, Mok80e, Pin85, Ril87, Sha84, Tus86]. production [CC86a, Jur81b, Kap84c, Ler81g, Mih81, RD81, Wal87a, Mer83]. Productions [GT82]. Productivity [Mus85, Fis87a, KG83, WB86, Wei80]. Products [Ler82b, Ano87-164, Ano87-165, Ano87-166, Ano87-167, Bel87c, Bel87d, BBJ87, Con86, Cor87a, CC87, KSB87, Lur84, Mih81, Mok81f, Mok81c, Per81a, Per87b, Per87e, Tan86, VH86, VWZ86e, Wal87d, Whi87]. profession [Ano84-86, Gol81, Joh84]. professional [Ano84-168, DTS86, FM84b, Gut83b, JI84, Ler84d, Per85b, Per87d, Sha80a]. professionally [Ler82h]. professionals [DTS86, Tru83d]. Professions [Ros80]. professor [Bry84]. profile [Ano80-68, Ano88o, BM81, Lom80c, Lur84]. profit [MMB83, Zor86c]. Profitability [Voe86d]. profits [CC86b, Per83b]. Program [Ano83-205, Ano83-206, Ano83-207, Ano84-194, Ano84-195, Ano84-196, Ano84-197, Ano84-198, Ano84-199, Ano85-191, Ano85-192, Ano85-193, Ano85-194, Ano85-195, Ano85-196, Ano86-185, Ano86-186, Ano86-187, Ano86-188, Ano86-189, Ano86-190, Ano87-203, Ano87-204, Ano87-205, Ano87-206, Sha80f, Wal87e, WBK87, Ada87e, Cor87d, Gut83b, Hor84b, HM87, Ler84d, Mos81, Roi86, Sta87a, Sum82a]. Programmable [Jur81b, Ano82-194, Fre88a, Jur82c]. programmed [Ano82-187]. programmer [Jur87c]. programmers [PW83, SAPM80]. Programming [Ler82h, Bac82, Ber80f, Jur82c, Kat80, Ler82g, Nag83, Sha82, Voe87a, Wal84g]. Programs [Ano87-236, Ano87-238, Bel84f, KSB87, Ano83-205, Ano83-206, Ano83-207, Ano84-194, Ano84-195, Ano84-196, Ano84-197, Ano84-198, Ano84-199, Ano85-191, Ano85-192, Ano85-193, Ano85-194, Ano85-195, Ano85-196, Ano86-185, Ano86-186, Ano86-187, Ano86-188, Ano86-189, Ano86-190, Ano87-235, Ano87-237, Cor87c, Cun80, Fag81a, FM84b, Gut84f, Gut84k, Gut84i, Hor84b, KZG^{+}84, Kap87d, KC87, Lom82, Pyk87, Tru83a, Zor84a, Zor84b]. Progress [Gre85, PD86, Sha82, Wal84g, Ada84c, DeV81, Fis86e, Hon80, Hor85b, Kap80b, Kap80f, Lur84, Wil80]. progresses [Ada87a]. progressing [DeV81]. project [Sum80, Wol85, Jur87f]. projectiles [Fis85c]. Projection [Sch81, Tan89b]. projectors [Hef84]. Projects [Gib80, Ler84h, Tor80, Wal87g, Ler84h]. proliferate [Jur83a, Mok80c]. proliferated [Wer87a]. Proliferating [KLCPT80]. proliferation [GP81, PW84]. PROM [Ano83-184]. promise [BA86, BKBC85, Dor83, ELW83, GS80, Hor85b, Jak85, Jur81a, Jur86c, Lof80b, Mad86, Mid80, MO89, ST81, SW83, Sum82b, Swe82, Voe87a, Wil87, Zor85e]. Promises [Bel85d, BCUY82, Car80, HHH87, Jur83b, Jur86a, Ler82k, Riv87, Vae86]. promising [Bot85, KW86, KK84, Wit84]. promote [Lur84]. prompt [KLCPT80]. prompts [de 83]. proof [Fit89c, Lom80d].

Propelling [ABEI84]. proper [Liu84].

propitious [Ros83]. Proposal [Wei80, Ram82]. propose [WAO85]. proposed [Lom80d, Tal83]. proposing [Gra82b]. propulsion [KA86, Kap87c]. pros [Ler81b]. prospective [HW83]. Prospects [Ano86-192, Ano86-191, Kap83e]. prosper [Kel87]. prosthetic [Fag80a]. protect [McC85]. protecting [Hon80, PW84]. protection [PFMP88, Voe87i, Wal82b]. protectionism [Kap85c, Ler84f]. protective [Ant87]. protocol [Kap85b]. Protocols [Kam86, Far86]. prototyping [GGZ86]. prove [Bro81a, For86, KM82]. proved [BE87, Wol84b]. proven [Fag80e]. proves [AJ85]. provide [BMoF^{+}83, BD83, Cor87a, Lem81a, Ler84g, Lla81, Miy84, Mok82b, SAPM80, Wal87h]. provided [Voe87g]. provides [Ano80-197, Ano80-194, DeL82, Per84b]. providing [Ano84-168]. ps [Ano81-182]. psychologists [Tru83b]. Psychology [Ano82-153]. public [Ano88j, BJ88, Chr83k, GT82, Jur85a, Lom80f, Mai83, Per81d, Rub80, Tan86, Wit84]. public-interest [Mai83]. public-television [Jur85a]. Publication [Chr81h]. publications [Ano80-148, Ano80-153, Ano80-154, Ano80-155, Ano80-156, Ano80-157, Ano80-158, Ano80-159, Ano80-174, Ano80-175, Ano80-176, Ano80-177, Ano80-178, Ano80-179, Ano80-180, Ano80-181, Ano80-182, Ano80-183, Ano80-184, Ano80-185, Ano81-134, Ano81-135, Ano81-136, Ano81-137, Ano81-145, Ano81-146, Ano81-147, Ano81-148, Ano81-149, Ano81-165, Ano81-166, Ano81-167, Ano81-168, Ano81-169, Ano81-170, Ano81-171, Ano81-172, Ano81-173, Ano81-174, Ano81-175, Ano81-176, Ano82-136, Ano82-137, Ano82-138, Ano82-139, Ano82-149, Ano82-150, Ano82-151, Ano82-172, Ano82-173, Ano82-174, Ano82-175, Ano82-176, Ano82-177, Ano82-178, Ano82-179, Ano82-180, Ano82-181, Ano82-182, Ano82-183, Ano83-145, Ano83-146, Ano83-158, Ano83-170, Ano83-171, Ano83-172, Ano83-173, Ano83-174, Ano83-175, Ano83-176, Ano83-177, Ano83-178, Ano83-179, Ano83-180, Ano83-181, Ano84-169, Ano84-170, Ano84-171, Ano84-172, Ano84-173, Ano84-174, Ano84-175, Ano84-176, Ano84-177]. publications [Ano84-178, Ano84-179, Ano84-180, Ano85-167, Ano85-168, Ano85-169, Ano85-170, Ano85-171, Ano85-172, Ano85-173, Ano85-174, Ano85-175, Ano85-176, Ano85-177, Ano85-178, Ano86-160, Ano86-161, Ano86-162, Ano86-163, Ano86-164, Ano86-165, Ano86-166, Ano86-167, Ano86-168, Ano86-169, Ano86-170, Ano86-171, Ano87-155, Ano87-156, Ano87-157, Ano87-158, Ano87-159, Ano87-160, Ano87-161, Ano87-162, Ano87-163, Ano87-170, Ano87-171, Ano87-172, Ano87-173, Ano87-174, Ano87-175, Ano87-176, Ano87-177, Ano87-178, Ano87-179, Ano87-180, Ano87-181]. publishing [Ber81d, Chr84c, Jur87a, Voe87i]. pull [Kap80f]. pulleys [NK87]. Pulsed [Wel85]. Pulses [Tel82, Lem81b, Ler81f, Wel85]. pumped [FZ85]. pumps [Hor85b, MS81]. punch [Wel85]. Purchasing [Fly87]. purists [Per83e]. purpose [Den81, Gut83d, Liu84]. purposes [Bel85d]. pursue [FZ85, Per84c]. pursuit [GWF84]. push [Kap80f, Wal85b]. pushed [Mur88b]. pushes [Wal85d]. put [LM81, Rum85, Sha81e, Voe87a, Wol85]. puts [Gut88, Roi86, Voe88c]. Putting [Sha81e, Wal83e, Wal87f, MD84, STK82]. puzzle [Fis86c, Wal83d]. quake [Ric88]. Quality [Ler81h, Mic81, Pea81, Bag84, CC85, Fly87, Jur81b, Jur82a, Jur87d, Kel87, Mih81, Miy84, Mus85, Ros80, Wal87a]. quandary [Ber82b]. Quantum [DGZ88, Mo83]. Quantum-tailored [DGZ88]. quarter [Ano88d, Ano88m]. Que. [Fis86d]. Quebec [Fis86d]. quest [Ano83-241, Bel86c, Fis83b, Fit89c, Mok81g]. question [Gut83i, Mor81b, Per83a]. questionable [CC85]. Questionnaire [Ano82-220]. questions [Ada87f, Ano80-245, DJ85, Ler83a, Wol83]. quickly [Fis83d, Fit88b, Wal84e]. Quiescent [Per84b]. Quitting [Bel87f].

R&D

[Fei82, Fis83c, Kli87, Mar85, San89c, Sta86]. race [Apt85, FW87, Jur89a]. races [Per87a]. Radar [Aml82, Fit87c, Ada87c, Ada87e, Ano81-45, Bro88, Fis80, Jel81]. radar-directed [Ada87e]. radars [EG82, PW86, Zor89c]. radiation [Fis86c, Ler80b, SPM^{+}80]. radically [Bac82, Swe82]. Radio [Chr87l, Del80, Kir80, Ano80-243, Bel85a, Bot83, Cla80a, Fit87c, Hil83, LH87, Per86, Per87b, Tha83, Wol85]. radios [PW86, Voe87d]. Rail [Kap84e, Kap87c]. railroad [Chr80c, KA86, Dup82]. railroads [TBH87]. railways [ABEI84, Kap83a]. rain [Zor87e]. raised [Ada87f, Bag84]. RAM [Ano80-195, Ano81-186, Ano81-183, Ano82-192, Ano83-184, Ber81k, Ber82h, Fis83i]. rampant [Hon80]. RAMs [Ber81k, BG89]. range [Ano82-190, DD85, Jac82, Tan86]. ranging [Den81, GBDC86, Gut84i, Wol84b, Zor85e]. rapid [Kap80c, War82]. rapid-deployment [War82]. Rapidly [Ano83-162, WM84]. rare [Tru83c]. rate [Ada84b, Ada84c, Ano82-186]. rates [Ano81-181, Ano82-195, CC85, Fis84f, Jay86]. rather [Gut82h, Jur87b, Ler80a, Wal86a, Wil80, Zde83]. rating [DMW87]. ratings [Che87]. ratios [Jur87c]. ravages [Zor87e]. raw [AJ85]. ray [Lep81, San89f]. rays [Bac80]. RCA [Har83]. re [PZ86]. re-energized [PZ86]. reach [Fis86f, Kir83, Shu89, Voe87a, Voe87j, Wal85b, WZB86, Wil87]. reached [Bry84]. reaches [Bel85c, Hor85b]. Reaching [TP83]. react [Fis83d]. reaction [FAEH84]. reactor [DeV81, Kap84f, TD84, Blu80]. reactors [Eis84, Fis84h, Fis86f, Fis87b, Fit89a, Ler84g, Wel85]. read [Ber81a]. readable [Tan86]. readers [Ada84b, Chr82l, Sar82]. readiness [Dan82, Ler82k]. reading [Ano80-126, Ano81-119, Ano82-121]. ready [Dan82, Fis84g]. Reagan [Ano80-245, Per83f]. Real [Gau85, Ano80-197, Fle82, Mor81b, Per83e, Per84c, Per86, Wal84d]. Real-time [Gau85, Ano80-197, Per83e]. Realism [KL85]. realistic [Far86, Mok81g]. reality [Hil85, Ler81c, Per83e]. really [Ada87f]. reap [KW86]. reason [Chi81]. reasons [Per87d, Zor84c]. reassessment [Cau82]. Rebirth [Mil81, Chr81m]. receiver [Ano80-243, Jac80, Per88d]. receivers [Mok80c]. recently [VWZ86c]. reception [SD81]. recession [Flo82, Tru83a]. Recipe [Tus86]. recognition [Cas81, DS81, Red80]. recognized [Kap80a]. recognizer [Kat80]. recognizers [Wal87f]. recommended [Gra82a, Ros80]. Reconfigured [Ada87a]. reconnect [Ver86]. record [Bel85a, Gar84, Kid81, Per88c, Wer87b]. record-keeping [Gar84]. recorder [Ano82-195]. recorders [Ano80-230, Mok82a]. recording [Cle82, Whi83]. recovery [BS81]. redefining [Jur87c]. redesign [Ler81g]. Redesigning [Voe87h]. rediscovered [Fly87]. reduce [Bel81, BS81, Kap82c, Liu84, Pow84, Riv80, SA85, Voe86d]. reduced [CS87, Her83b, Wal85e]. reduced-instruction-set [Wal85e].

Reducing [Lom80d, CC86a, Kap83e]. reduction [Cun80, Pin80, Ram82]. Redundancy [Voe86a, Voe86b, BS81, Jur86a, Wil81]. reed [Luc80]. reels [CC85]. Reemployment [Voe89b]. referee [Eis84]. references [Hor84b]. refined [Jur80]. Reflections [Ano82-221, Ano82-222, Ano82-223, Ano82-224, Ano82-225, Ano82-226, Ano83-208, Ano83-209, Ano83-210, Ano83-211, Ano83-212, Ano84-200, Ano84-201, Ano84-202, Ano84-203, Ano84-204, Ano84-205, Ano85-197, Ano85-198, Ano85-199, Ano85-200, Ano85-201, Ano85-202, Ano86-193, Ano86-194, Ano86-195, Ano86-196, Ano86-197, Ano86-198, Ano87-207, Ano87-208, Luc83, Luc87a, Luc87b, Luc87c, Luc87d]. Reflector [Rog81]. reflects [Joh84]. region [Ano80-210, Ano80-211, Ano80-212, Ano80-213, Ano80-214, Ano80-215, Ano80-216]. regional [Kap82b]. regions [Bac80]. regular [Per82d]. regulate [Ler84c]. Regulation [Fit87b, Ano81-152]. regulations [Ler84c, Ler84f]. Regulators [LP80, FZ87a]. Regulatory [Eis84, Ler81f, CN83, Fis86a]. rehabilitate [Sha81d]. rehabilitation [Hor84d]. reigning [Bow85]. Rejuvenating [Ril87]. rekindled [Kap82c]. relate [Sar82]. relationships [Ano84-236]. relative [SA85]. relatively [VWZ86e]. relays [Kap87c, Luc80]. relearned [Gut82h]. released [Far86, Per87b]. Reliability [Ber81h, Chr81i, Ber81e, KSB87, Kap82d, Liu84, Mus89, SA85, Spi86, Tus86]. Reliable [Kap83f, Ler81i, Ano82-47, BCUY82, Bel88b, Fit88b, Kap83d, Mar86, TBH87, Wal87d]. relieve [HHH87]. reluctantly [Gol81]. rely [Dav87a, Wal84a]. remain [CN83, Fis84c, Jur83b, Lem81c]. remains [Gut83g, Ler80d]. remarkable [Fit88a]. remedies [Wal86b]. remedy [Ros80]. Remember [Ano88k, Ano88a, Fag80f].

Remote [Kap83g, Ada89b, Bro88, LM81, Per84c, Ver86, Voe87g, Wal83c]. renaissance [Kap84a]. rendering [Fis83e]. Renewable [Row82, Wet82b]. renowned [Ryd84, Wol84b]. repairs [Ver86]. repeater [Mid80]. repeaters [Kec83, KK84]. repetitive [Ano81-182, Ler81g]. replace [Ano83-55, HHH87, Kap87c]. replaces [Gut84j, Mil87]. report [Ada86b, Ada88e, Ano80-108, Ano80-109, Ano80-110, Ano80-111, Ano80-112, Ano80-113, Ano80-114, Ano80-115, Ano80-116, Ano80-117, Ano80-118, Ano80-119, Ano81-101, Ano81-102, Ano81-103, Ano81-104, Ano81-105, Ano81-106, Ano81-107, Ano81-108, Ano81-109, Ano81-110, Ano81-111, Ano81-112, Ano82-103, Ano82-104, Ano82-105, Ano82-106, Ano82-107, Ano82-108, Ano82-109, Ano82-110, Ano82-111, Ano82-112, Ano82-113, Ano82-114, Ano83-117, Ano83-118, Ano83-119, Ano83-120, Ano83-121, Ano83-122, Cor87d, DeV81, Kap87b, Per83f, Roi86, Sum82a, Tor80, Wal84c]. report/verification [Ada86b]. reported [Lep81, Sha82, Wal84g]. reporter [Joh84]. reporting [Ler82c]. reports [Ano83-205, Ano83-206, Ano83-207, Ano84-194, Ano84-195, Ano84-196, Ano84-197, Ano84-198, Ano84-199, Ano85-191, Ano85-192, Ano85-193, Ano85-194, Ano85-195, Ano85-196, Ano86-185, Ano86-186, Ano86-187, Ano86-188, Ano86-189, Ano86-190]. representative [Ano84-147]. represents [Jur85b]. reproduction [Miy84]. require [Ber80d, Wal83a, Wel85]. required [GWF84, Hil86]. requirements [Ano86-211, Jac82, Sha84]. requires [Per82a, Sum80]. requiring [Tor80]. requisite [DS81, Mok82d]. rescue [Ano84-168, Bug84, SS84, Tul83]. rescued [Kid81]. Research [Fei82, Kap82d, Pak85a, Pak85b, Per80a, Wol84b, Ada87a, Bel88b, Bel88a, Car83, Dou83b, Fis83c, Gut83i, Gut84i, HK87, Jak85, KW86, Lem81c, Ler84a, Mas80c, Moy80, Per84c, RK80, ST81, Wal86c, Wel84]. researcher [Kap81a]. Researchers [Ler82g, RD81, Con86, Kap81a, KL85, Ler84a, SH85]. reside [Ler80a]. resident [Par84, PW84]. residential [CC85]. Resistance [GBDC86]. resolution [Ano80-193, Ano80-192, Ano81-181, Ano82-191, Mok81i]. resolutions [Ano81-182]. resonance [Bot83]. Resonant [Syk89]. Resonant-mode [Syk89]. resource [Wal85c]. respect [Fag80f, Fis87a]. respects [JI84]. respond [Hor85d]. responded [Per81c, SAPM80]. responds [Cor87b]. responses [Per81c, Wal84c]. responsibility [DTS86]. responsible [Ler84h]. restored [Fis86f]. restoring [Fre86]. restructured [Lom81b]. restructuring [CC86b, Lom80b]. result [Kah81, Pin80]. resulted [Mai83]. results [Per84e]. retain [DTS86]. Rethinking [Ber82h, Zor89g]. retrack [KA86]. retrained [FP84]. retrenchment [Fit87b]. retrospective [de 83]. Return [Fle85b]. reveal [FP84]. revealed [BE87]. reveals [Ano88p, San88c]. reverse [Ano84-236]. Review [Bro88, Ano82-61, Cas81, Chr81g, Fis86e, Jur87c, Kah81, Ler82a, Wal84f, de 83]. reviewed [Chr81g, Fei82]. reviews [Ait87, Ano80-44, Ano80-45, Ano80-46, Ano80-47, Ano80-48, Ano80-49, Ano80-50, Ano80-51, Ano80-52, Ano80-53, Ano80-54, Ano80-55, Ano81-46, Ano81-47, Ano81-48, Ano81-49, Ano81-50, Ano81-51, Ano81-52, Ano81-53, Ano81-54, Ano81-55, Ano81-56, Ano81-57, Ano82-49, Ano82-50, Ano82-51, Ano82-52, Ano82-53, Ano82-54, Ano82-55, Ano82-56, Ano82-57, Ano82-58, Ano82-59, Ano82-60, Ano83-57, Ano83-58, Ano83-59, Ano83-60, Ano83-61, Ano83-62, Ano83-63, Ano83-64, Ano83-65, Ano83-66, Ano83-67, Ano83-68, Ano84-60, Ano84-61, Ano84-62, Ano84-63, Ano84-64, Ano84-65, Ano84-66, Ano84-67, Ano84-68, Ano84-69, Ano84-70, Ano84-71, Ano85-59, Ano85-60, Ano85-61, Ano85-62, Ano85-63, Ano85-64, Ano85-65, Ano85-66, Ano85-67, Ano85-68, Ano85-69, Ano85-70, Ano86-53, Ano86-54, Ano86-55, Ano86-56, Ano86-57, Ano86-58, Ano86-59, Ano86-60, Ano86-61, Ano86-62, Ano86-63, Ano86-64, Ano87-53, Ano87-54, Ano87-55]. reviews [Ano87-56, Ano87-57, Ano87-58, Ano87-59, Ano87-60, Ano87-221, Ano87-222, Bel88c, Bro81a, Gol87, Lom80e, Oli87, Pan87b, Wil81]. Revitalization [Bel86a]. revitalized [Lom82]. revival [Ber80g]. revolution [Gut83c, Ler81c, Ler83a, WZ85]. revolutionary [KW86, Sum80]. revolutionize [RD81, Sha81f]. revolutionized [Ano84-53]. revolutionizing [Che86, Gar84]. revolutions [Cor87c]. rewards [Bee83, KW86]. RF [Ler80b, SH85, SPM^{+}80]. Rhythmic [Sar82]. rich [CA82]. Richard [Ano87-100]. Riding [Wol84a, Jur82a, Jur87b]. right [Chr88a, FT85, Hor84b, Voe87e, Wol83]. rights [Hon80]. ring [KS86b]. RISC [VH86]. rise [Ano81-230, Whi83]. rising [CC85, Gel81, Mur89c, Per86]. Risk [Mor81a, Mor81b, Bel84a, Bel89c, Cor87a, Lom80d, MFNL85, Sta87b]. risks [Mor81a]. rivals [Cor88b]. river [Gut82b]. Rkb [Ber81k]. RM02 [Ano80-196]. road [Fag80d, Mok80e, Riv80]. roads [Wal87a]. Robert [Bel84a]. robot [Ano83-54, Bel85d, Sug80a]. robotic [Sug80a]. Roboticists [Hor86b]. robotics [Gut82f]. robotry [DD85, KP86]. Robots [Bel85d, Fis85e, Kap85b, Nag83, Zor86c, Hor86b, Jur81b, Kap84c, Whi87]. rockets [KM82]. role [Ano81-230, Che82, Dan82, GZ85, Kah84, Kap84c, Kap84e, Kli84, KK84, Mas80a, Sta86, Zor89c]. roles [Ano84-168, Fag82b, FHRW^{+}83, Wal84e]. roll [Fis85e, Kap84e]. ROM [Ano82-192, Che86, OK86]. romanticization [Win84]. Room [Ano84-53, Gut85a, HOD82, PW83]. rooms [Per83d]. roots [MD84]. ropes [FAEH84]. rosy [Bot85]. rough [Ler82a]. round [Ano80-68]. round-table [Ano80-68]. route [Par84, PW86]. routine [IK84, Voe86a, Voe86b]. rubber [Kap84e]. Rugged [Tan86, Mar86, Miy84]. rules [CC85, Gev83, Mos81]. ruling [Per82c]. rumors [Chi81]. run [Bel87f, Chr80c, Fis84d, Kap83a, Ler83b, Tal83]. run-throughs [Tal83]. rural [Hil85]. Russian [Ber80f].

s [Ano82-187]. S33 [Ano80-196]. S33/A [Ano80-196]. Sachs [Per87a]. safe [Fis86b, Fis87b, TBH87]. Safeguarding [Hor85d]. safely [Fis86f, Lom80f]. safer [Fis83e, Lom81b]. safety [Ano80-41, Fag80c, FP86, Jur87b, Kap83e, Kap83c, Lom80b, Per81e, Voe86a, Voe86b, Voe86d]. sag [Che89c]. saga [Zor87d]. said [Chr80l, Chr82m, Chr85l]. sailboats [Per87a]. Saint [And80]. sale [Mas85]. sales [Che89c, Per87b]. Salient [Ano88d, Ano88m]. Salt [Fis83b]. salvaged [Ada85a]. salvation [San89c]. same [Fis85f, Fis87d, Mur85, Rum85]. sample [Ano81-181]. samples [Ano82-195]. sampling [Ano81-182]. Samsung [Per87b]. San [Sha81c]. Sandia [AM86]. sapiens [Hor86b]. Sarnoff [Jur88c]. Satellite [Ber81a, Gib80, HMDF80, Ada84a, Bel84d, BA86, Den81, Dut80, Hil83, Hil85, SS84, SF82]. Satellites [HW83, Hil85, Ada85a, Ano83-246, Bel84c, EMM82, GZ85, Kap83b, PK81, Rog81, Wat87]. satisfy [Kut83]. save [FZ85, Kap80c, Wal87b]. saver [Fag80e]. Saving [Chr80j, Chr84g, Fis83c, VWZ86c]. say [Voe87e]. scale [Ano82-253, Ano83-162, FHRW^{+}83, Ler80c, Mas80c, MRRS84, Wal87g, Wei84]. scanners [Ina89]. Scanning [Ano80-231, Ano80-232, Ano80-233, Ano80-234, Ano80-235, Ano80-236, Ano80-237, Ano80-238, Ano80-239, Ano80-240, Ano80-241, Ano80-242, Ano81-212, Ano81-213, Ano81-214, Ano81-215, Ano81-216, Ano81-217, Ano81-218, Ano81-219, Ano81-220, Ano81-221, Ano81-222, Ano81-223, Ano82-227, Ano82-228, Ano82-229, Ano82-230, Ano82-231, Ano82-232, Ano82-233, Ano82-234, Ano82-235, Ano82-236, Ano82-237, Ano82-238, Ano83-213, Ano83-214, Ano83-215, Ano83-216, Ano83-217, Ano83-218, Ano83-219, Ano83-220, Ano83-221, Ano83-222, Ano83-223, Ano83-224, Ano84-206, Ano84-207, Ano84-208, Ano84-209, Ano84-210, Ano84-211, Ano84-212, Ano84-213, Ano84-214, Ano84-215, Ano84-216, Ano84-217, Ano85-203, Ano85-204, Ano85-205, Ano85-206, Ano85-207, Ano85-208, Ano85-209, Ano85-210, Ano85-211, Ano85-212, Ano85-213, Ano85-214, Ano86-199, Ano86-200, Ano86-201, Ano86-202, Ano86-203, Ano86-204, Ano86-205, Ano86-206, Ano86-207, Ano86-208, Ano86-209, Ano86-210, Ano87-209, Ano87-210, Ano87-211, Ano87-212]. Scanning [Ano87-213, Ano87-214, Ano87-215, Ano87-216, Ano87-217, Ano87-218, Ano87-219, Ano87-220]. scapegoat [Fis86b]. Scarce [KZG^{+}84]. scare [Wal84a]. scenario [Roi86]. Scenarios [Kap82a, Fis86c]. scene [Sha80e, Wal85a]. scenes [Cor84]. schedules [Wal86a]. scheduling [Mas85]. scheme [NS80, STK82]. schemes [GP81, Jur87c]. school [Hor84b, Mas80c, MD84, Pyk87]. schools [Hor84a, Kap84b, Tru83a, Wit84]. Science [Bal84a, Chr81j, Chr84h, Per83f, Fis83a, Gut84k, HK87, Pyk87, Ano84-147]. scientific [Ber82c, Fis84h, Mok80f]. scientist [Zor86c]. scientists [HK87, Voe88a, Wal83e, Zor87e]. scopes [BFB87, Fit88a, Sha80c, Wer87b]. score [Lom80a, Spi86]. scramble [Fis87d]. scrambler [Kah84]. screen [Fis87c]. screening [Pea81, Tus86]. screw [NK87]. scrutinized [Whi87]. scrutinizing [Gut83c]. scrutiny [Lom80b]. SDI [AH85, AF85, Fis85c, Zor89c]. SE3000 [Ano80-230]. sea [Ada85c, Ada87c, And80, Gut82i, HW83, SS84]. search [Ano83-247, Ano86-145, Dou81, Per89b]. seasons [Bel89b]. seat [Jur84]. second [Ano82-185, Ano82-186, Ano84-237, Bel84d, Ber80f, DeV81, Fle85b, Kap83a, Tru84b, Wal85b, Wil87]. second-generation [Kap83a]. seconds [Hic83]. secret [Ada86a, Chr80b, Per89c, Wol85, Zor87a]. Secretaries [IK84]. Secretary [Fis83d, Wal84f]. section [Ano84-148]. secure [Cau82, Kah84, Per82d, Per84b, Wal85d]. security [Ber82a, FHW85, Hor85d, Kap82b, Lom81c, Per84b]. Sedan [Voe87b]. see [Chr82h, Ler80a, Per87e]. ’Seeing [MR82, Fle82, Jur80]. seek [Kap81a, Tru83d]. Seeking [Gut83c, Ler84c, San89b]. seem [Con86, Per82d, Zor87e]. seemed [Cor84]. seems [Ber80f, Kap82b, Ler80d]. seen [FEK80, Jur85a, Mok82a, Mok82d, WZ85]. sees [Den81]. select [Ano84-86, DD86, HMDF80, Sil89, Zde83]. selected [Ano84-148, Ano84-147]. Selecting [Hor84b]. selection [Ano81-177, Ano82-187, Jur82b]. Self [Mac88, Jur82c, San88c]. Self-developers [Mac88]. self-diagnosis [Jur82c]. self-perceptions [San88c]. sell [Lur84]. selling [Tru84a]. Semiconductor [KK84, Bel86a, Car80, ELW83, Fan89, Fis84b, Fis87d, McC85, San88a, SW88, Wal86c]. Semiconductors [Per87f, Bal81, Che87, Eas86, Rum85]. seminal [Ano84-148]. seminars [Gut83b, Ler84d]. send [Rog81]. Sending [Ada85c]. sense [Gut84j]. Sensing [MNS81, MAN80, Ada89b, Bar81, Bro88, Hor86b, Kap83g, Wal83c]. sensitivity [Ano81-180]. sensor [Ano87-228]. Sensors [BD83, Bar81, DD85, GBDC86, HW83, Kap85b, MAN80, Nag83]. September [Gut82g, Ver86]. sequences [Zde83]. Sergeant [AM86, Ada87e]. SERI [Per80a]. series [Bel84e, Bel84d, Ano80-230]. serious [Eid81, Jur85b]. seriously [Del80]. serve [Wet82b]. served [Wil80]. serves [Mil87]. Service [Ano80-36, Ano81-38, Ano82-135, Ano83-48, Ano85-46, Ano86-49, Ano86-184, Hil85, Ler81d]. servicemen [Ber81e]. Services [Ano84-52, Bel87b, Cla80b, Dor83, GZ85, Hil83, Jur87a, Pow84, Zor85f, Pan87a]. Serving [Log83]. sessions [Sha80f, Sha81c]. set [Bee83, Bel87f, Dav87a, PK81, Wal85e, Zad84]. sets [Wal85a]. setting [Per82c, Roi86]. settings [Dou83b]. setup [Ber81d]. Seven [Per87d, Tri80, Gre80]. several [Bow85, Fis84g, PW85a, Sha80a, Wal87a, Wal87h]. sewing [Sha81e]. Shake [Tus86]. shape [Bel85f, HT83]. shaped [Ryd84]. share [Cor87a, CC86b, FP86, Wei80, Zor86d]. Shared [Wal87b, Lom80f]. shares [Sha81a]. Sharp [Ler80b, Ada86a]. sharpens [Chu88, CC86b]. Sharper [Ano80-243, Lla81]. sharply [Bel87e, Miy84]. shaving [Zor84c]. shear [Zor86a]. shed [Hic83]. shedding [Hic83]. sheet [Ano81-177]. sheet-fed [Ano81-177]. shelf [Ano81-58]. Shield [Aml82]. shift [Ler82g, Per80a]. shifting [Wal84b]. shifts [Gut84h]. shine [Zor87e]. Shinkansen [Kap83a]. ships [Lak82]. Shirt [Wol84b]. Shirt-sleeves [Wol84b]. shop [Sug80a, Zde83]. shopping [Ano86-211, Mok82f, Wei84]. shore [Zor87b]. Shoreham [Fit88d, Zor87b, Zor87d]. Short [Jac82, Ano84-85, KZG^{+}84, Kap85f, Per80a, Tru83d]. short-lived [Ano84-85]. Short-range [Jac82]. short-term [Per80a]. short-wavelength [Kap85f]. Shortage [Chr82i]. Shortcomings [Fis80]. shorter [CC87]. shortsighted [IK84]. should [DJ85, HOD82, Kap83d, Mor81a, War82]. show [Hor85b, Jur81a, Zor85e]. showing [Den81]. showplace [Ano84-146, Per85d]. shows [BA86, DS81]. Shrewd [JI84]. Shrinking [Per81a, SH85, Zor86d]. shuffle [Chr82j]. shunning [Gre80]. shut [Fis87b]. shutting [Tal83]. Shuttle [LM81, Ano82-47, BE89, Esc86, Lom82, Moy80, Wal83c]. side [Ano81-150, Gel81]. Sierra [Per85b]. sight [Ber81g, Pic82]. Sightless [SAPM80]. sights [Per80a]. signal [AF87, Ano80-192, Fis87d, Jur87c, MAN80, Mok82c, SD81, Wal87f, de 82]. signal-to-noise [Jur87c]. signaling [Kap87c]. signals [Ano82-195, Chr85g, Hor84d, MNS81]. significant [Wal84e]. silent [Lof80a]. Silicon [Bar81, Bel85c, Car80, Dut80, Fis84b, Lof80b, Mad86, MAN80, SW88, Wal84e, Wol80]. silver [Fis87c]. simple [AF87, Fis84c, Hor85e, KK84, VWZ86e]. simpler [Wal85e]. simplex [Emm85]. simplify [Bac82, BKBC85, Kap87c]. Simply [Ler84g, Sti81]. simulate [Swe82]. Simulating [FT85, Tal83]. simulation [GGZ86, Hor86d, SW83, Wil87]. simulations [Tal83]. simulator [Ano83-186]. simulators [FT85]. Since [FWVH86, MRRS84]. Singer [Sha81e]. Single [Ano81-186, Ano82-193, Ano82-194, Bot85, Kec83, Mok82d, BCUY82, Ber81a, Zor86a]. Single-board [Ano81-186, Ano82-193]. Single-chip [Ano82-194]. Single-mode [Bot85, Kec83, Mok82d]. site [Fis86f]. sites [Hor84c, Ler80c]. situation [Ros80]. Six [Per82a, Voe87j, Wol85]. six-month [Wol85]. size [Ano81-177, Ano82-192, Ano82-187, HT83, Smi87]. sizzle [Jur85b]. skeptic [Flo82]. skeptics [Sha81e]. skies [Ada86a, PWZ^{+}86]. skill [Chr84d]. skills [Bel87f, Gut84i]. slackened [Tru84b]. slackening [Fis86a]. slave [Gut83d]. sleeping [Gut85c]. sleeves [Wol84b]. Sleuthing [Ler83e, Doy81]. SLICs [Mok82c]. slide [Hef84, Wei80]. slow [Ada84c, Fit87a, Voe87a]. slowdown [Zor86d]. Slowed [Fis84f, Ler81b]. Slowly [Ler81c, DeV81]. slugs [Hor86b]. Small [Ada87b, Ber82g, Ler80c, Per81a, Ano81-45, Ber81d, Ber82e, Gut84e, JCH80, MMB83, Per80c, Sta86, Sti88a, Voe87j, Ber81d]. small-batch [MMB83]. Small-business [Ber82g]. small-craft [Ano81-45]. small-scale [Ler80c]. small-system [Voe87j]. smaller [Fis83b, Mok82a]. smallest [Lep81]. Smart [Ano81-44, Ano81-45, Ano82-48, Wei84, Ber80c, Kah83, Kap85b, MAN80, Nag83, Sha81d, Sho88, Ada87a]. smarter [Fis85f, Jur82c, Sha80d, Wal83a]. Smithsonian [Ano84-148]. smorgasbord [FM84b]. soar [Fag80e]. Sobering [Bel85f]. social [Per80c, Win84]. Societal [Bel84f]. societies [FM84b]. society [Ano84-86, Ber83a, Chr83e, Gut83i, Kap81a, Mai83, Nai83, PW84, Rub80, Win84]. Sociotechnology [GP81, Tri80]. Software [Ano87-204, Ano87-205, Ano87-206, Ano87-221, Ano87-222, Bac82, Ber80f, BPEB84, Gut83h, HM87, Ler82g, Ler82h, Mil87, PP88, Sha87, Wal84g, Wal85d, Wal86b, WB86, Wal87g, Yeh83, Ano80-196, Ano83-186, Ano87-203, Ber80c, Ber81b, Ber81f, Ber82g, DD86, Fag81a, FHRW^{+}83, Gra82a, Gut83g, Gut84h, Gut85b, Ham86, HK87, Kap80a, Kap83c, Kap87c, Ler80a, Ler81c, Ler81b, Mos81, Mus85, Mus89, OK86, Per84d, Sta87b, Sti81, VWZ86a, VWZ86b, Voe88b, Voe88g, Voe89d, Wal84a, Wal87e, WBK87, Wer87a]. Solar [Lof80a, MS81, Per80a, Sim83, Ano83-246, Bro81a, Car80, Fis84e, PHG86, Rog81, Per80a]. Solar-cell [Lof80a, MS81, Rog81]. solar-pond [Bro81a]. solar-power [Ano83-246]. sole [Per83g]. solicits [Ano82-220, Ano82-248]. Solid [BCUY82, Bar81, Ber80g, Ber81j, Ber81k, Ber81i, Ber82h, Bro81b, Dav83a, ELW83, Eid81, Fag81c, Fis84g, Fis85f, FG86, Fis87d, FEK80, Lep81, Mas80c, MS81, Mok80f, PC80, ST81, Sum80, Sum82a, Wal84b, Wol88c, Bal81, Bel86a, DGZ88, Ear84, Kap83b, Mur85, Mur88a, San89e]. Solid-state [Wal84b, Bal81, DGZ88, Kap83b]. solution [Hor86a, Mok81g]. Solutions [Ber81k, Ler80a, Fis84c, Hor84c, Kir80, Ler82j, Per81b, Rub80, Zor86a]. solve [Wil87]. solved [Ler82b]. solving [Emm85, KL85]. Some [Ano88j, Ano88k, Fis85e, Ano88p, CN83, Cor87c, CC87, DJ85, Ler82j, Per85d, San88c, Ste81, Sum82a, Wal84e, Wol84b]. sometimes [Mil87, Pyk87, Wal86a, Zor86c]. Sonar [Che89a]. Sony [Cor88b, Per88a]. soon [Ano84-168, Fit89f, Pic82, Rum85, Sta87b, Zor86a, Per87b]. Sophisticated [Fag80c, Per84e, Ber82f, Hor85e, Jur83a, WZB86]. sophistication [Gut82h]. sorties [Sim83]. sought [Fag80b, HMDF80, Kap80e, Kir80, PW85a, Wal82b]. sound [Bee83, CC86a, Fis80, Jur82b, Miy84, Mok82a, Per83a, Per85a, Voe87c]. sounds [KL85]. sources [Bot85, Dou85, Wet82b]. sourcing [Fly87]. Southern [Ril81, Zor88a]. Soviet [Ada86a, Ada88g, Fre84, Gra84, Lom81c, Mcc88, Mey86]. Space [Beg83, Her83c, KM82, Lom80a, Moy80, Rub83, Ten82, Ada84a, Ada85a, Ano81-187, Ano83-247, Bel87b, Buc83, Chr82j, Chr83d, Cla80a, Fei82, Fls83, Fru83, Gar81, Gib80, Hos83, KLCPT80, Key83, Lem81b, Ler83c, Ler84g, PHG86, Rog81, TD84, TBH87, Wal83c, Wal87b, Ada87a, BE89, Wal83c]. space-age [Fls83]. Spaceborne [Bro88, EG82]. spacecraft [Ada87a, Bel89d, DJ85, Gar81, Wil81]. Spacelab [Moy80]. spacing [Mid80]. span [Bel87b]. spans [Wal83e]. spark [FG86]. spawn [Nai83]. spawned [Hor85e]. spawns [Kap84a]. speak [Chr82h, FW82, Voe87e]. Speaking [Per81d, Ano80-244, Ano80-246, Ano81-224, Ano81-225, Ano81-226, Ano81-227, Ano81-228, Ano81-229, Ano82-171, Ano82-242, Ano82-243, Ano82-244, Ano82-245, Ano82-246, Ano83-233, Ano83-234, Ano83-235, Ano83-236, Ano83-237, Ano83-238, Ano84-230, Ano84-231, Ano84-232, Ano84-233, Ano84-234, Ano84-235, Ano85-227, Ano85-228, Ano85-229, Ano85-230, Ano85-231, Ano85-232, Ano86-225, Ano86-226, Ano86-227, Ano86-228, Ano86-229, Ano86-230, Ano87-229, Ano87-230, Cor87e, Cor87f, Cor87g, Cor87h, Jur82d]. Speakout [Ano82-239, Ano83-225, Ano83-226, Ano83-227, Ano83-228, Ano84-218, Ano84-219, Ano84-220, Ano84-221, Ano84-222, Ano84-223, Ano85-215, Ano85-216, Ano85-217, Ano85-218, Ano85-219, Ano85-220, Ano86-212, Ano86-213, Ano86-214, Ano86-215, Ano86-216, Ano86-217, Ano87-226, Ano87-225, Ano87-223, Ano87-224, Flo87, Gar82, Kay87]. Special [Ada86b, Ada88e, JP85, Kap87b, Liu84, Tor80]. Special-purpose [Liu84]. specialist [Ano87-234, Ler81c]. specialists [CC87]. specialities [Jur88e, Jur89d]. specialized [Gut83d, KZG^{+}84]. specialties [Gut84b, Jur86d, Jur87c, Wal84b]. species [Mur88b]. specific [FG86, Fis87d, Kap85f]. specification [Far86, Kat87]. specifications [Ler81g]. specify [Bel86a]. specs [Voe88d]. spectacular [BKW83]. Spectral [Ano86-218, Chr80c, Chr80d, Chr80e, Chr80i, Chr80f, Chr80g, Chr80h, Chr80j, Chr80b, Chr80k, Chr80l, Chr81k, Chr81c, Chr81l, Chr81d, Chr81e, Chr81f, Chr81b, Chr81g, Chr81h, Chr81m, Chr81i, Chr81j, Chr82b, Chr82c, Chr82d, Chr82e, Chr82f, Chr82g, Chr82h, Chr82i, Chr82j, Chr82l, Chr82k, Chr82m, Chr83j, Chr83b, Chr83f, Chr83e, Chr83g, Chr83h, Chr83c, Chr83d, Chr83i, Chr83k, Chr83l, Chr84b, Chr84a, Chr84c, Chr84j, Chr84k, Chr84d, Chr84e, Chr84f, Chr84g, Chr84h, Chr84l, Chr84i, Chr85b, Chr85a, Chr85h, Chr85c, Chr85d, Chr85f, Chr85g, Chr85j, Chr85k, Chr85i, Chr85e, Chr85l, Chr86e, Chr86d, Chr86g, Chr86a, Chr86k]. Spectral [Chr86f, Chr86b, Chr86i, Chr86h, Chr86c, Chr86j, Chr87j, Chr87i, Chr87k, Chr87g, Chr87l, Chr87h, Chr87b, Chr87c, Chr87d, Chr87e, Chr87f]. spectrum [Bel87b, Kah84, Kir80, Mok80a, Ano80-78, Ano80-79, Ano80-80, Ano80-81, Ano80-82, Ano80-83, Ano80-84, Ano80-85, Ano80-86, Ano80-87, Ano80-88, Ano80-89, Ano80-245, Ano81-71, Ano81-72, Ano81-73, Ano81-74, Ano81-75, Ano81-76, Ano81-77, Ano81-78, Ano81-79, Ano81-80, Ano81-81, Ano81-82, Ano82-74, Ano82-75, Ano82-76, Ano82-77, Ano82-78, Ano82-79, Ano82-80, Ano82-81, Ano82-82, Ano82-83, Ano82-84, Ano82-85, Ano82-220, Ano82-61, Ano82-248, Ano83-81, Ano83-82, Ano83-83, Ano83-84, Ano83-85, Ano83-86, Ano83-87, Ano83-88, Ano83-89, Ano83-90, Ano83-91, Ano83-92, Ano84-87, Ano84-88, Ano84-89, Ano84-90, Ano84-91, Ano84-92, Ano84-93, Ano84-94, Ano84-95, Ano84-96, Ano84-97, Ano84-98, Ano84-72, Ano84-193, Ano85a, Ano85-83, Ano85-84, Ano85-85, Ano85-86, Ano85-87, Ano85-88, Ano85-89, Ano85-90, Ano85-91, Ano85-92, Ano85-93, Ano85-94, Ano86-79, Ano86-80, Ano86-81, Ano86-82, Ano86-83]. Spectrum [Ano86-84, Ano86-85, Ano86-86, Ano86-87, Ano86-88, Ano86-89, Ano86-90, Ano86-65, Ano87-76, Ano87-77, Ano87-78, Ano87-79, Ano87-80, Ano87-81, Ano87-82, Ano87-83, Ano87-84, Ano87-85, Ano87-86, Ano87-87, Ano87-61, Ano88g, Ano88h, Gut84d, Gut84b, Gut84c, Gut84a, Lom80c, Per81c, Per83a]. Spectrum/Harris [Gut84d, Gut84b, Gut84c, Gut84a]. Speech [DS81, Ler82b, Fle82, Jay86, KL85, Kat80, Pic82, Red80, Sar82, SAPM80, Wal87f]. Speech-synthesis [Ler82b]. Speed [Ano84-237, Bel85c, Ber80b, Ber82c, Bor89, ELW83, Fis85c, Fis80, GS80, Lem81a, Lom80a, PP88, Rau87, Sha82, Sug80b, Sum80, VH86, Wal85e]. speedier [Ler82c]. speedily [Per84e]. speeding [ABEI84, Smi87]. speeds [Ano82-187, Bal81, Bel86c, Eid81, Ler84b, Wal85b, de 89]. spell [FW82]. spending [Bel89a, Sti89b]. spent [Ada87e]. Sperry [Sea89]. Spin [Ano83-229, Ano83-230, Ano83-231, Ano84-224, Ano84-225, Ano84-226, Ano84-227, Ano84-228, Ano84-229, Ano85-221, Ano85-222, Ano85-223, Ano85-224, Ano85-225, Ano85-226, Ano86-219, Ano87-227, Ano87-228, Dav87b, Dav87c, Dav87d, AM86]. Spin-offs [Ano83-229, Ano83-230, Ano83-231, Ano84-224, Ano84-225, Ano84-226, Ano84-227, Ano84-228, Ano84-229, Ano85-221, Ano85-222, Ano85-223, Ano85-224, Ano85-225, Ano85-226, Ano86-219, Ano87-227, Ano87-228, Dav87b, Dav87c, Dav87d]. spinning [Mar86]. Spinoffs [Ano86-220, Ano86-221, Ano86-222, Ano86-223, Ano86-224, DA87]. splinter [Ano84-236]. Spoken [Kap80a]. Sponsored [Jur85a]. sports [Fag80a]. spot [KSB87]. spotlight [Ano83-162, Gre80, Per85a, Sha81a]. spots [KSB87]. spread [Kah84]. spur [Kap80f, Mal84]. Spurred [Fis84h]. spurring [FZL86]. spy [Mey86]. spying [Hor85e]. squeeze [Kap80e]. stabilization [Hic83]. Staff [Ano82-240, Ano83-232, Ano88n, Ano82-158, Ano83-161]. stake [JCH80]. stakes [Jur87d]. stampede [Sta87a]. stamps [de 83]. stand [Ber82g]. stand-alone [Ber82g]. standard [CS87, Fis86b, Jur82b, Kap80a, KM88b, Per87b, VB86]. Standardization [Ler81b, Ber80b, Gut84j]. standardizes [MMB83]. Standards [Gol81, Mas80a, Per82c, Ano84-247, Cla80b, DD86, Gra82b, JT81, Jur82d, Jur83b, Ler80b, Ler81b, Mok81h, Per87b, Roi86, Sti88d, VWZ86d, Wal86a, Wal87h]. standby [Bal85]. stands [DS81]. Stanford [Bry84]. Star [Ada89c, Ano80-243, Sha80f, AF85]. Stark [Ada87f]. starred [Ano84-85]. start [Bel87f]. started [MD84]. Starting [Per82a, DD86]. State [Bal81, BCUY82, Bar81, Bel86a, Ber80g, Ber81j, Ber81k, Ber81i, Ber82h, Ber82j, Bro81b, Che89d, Dav83a, DGZ88, Ear84, ELW83, Eid81, Fag81c, Fis84g, Fis85f, FG86, Fis87d, FEK80, Hil83, Kap83b, Lep81, Mai83, Mas80c, MS81, Mcc88, Mok80f, Mur85, Mur88a, PC80, Per85d, San89e, ST81, Sum80, Sum82a, Wal84b, Wol88c, You84]. State-of-the-art [Bal81, Ber82j, Per85d]. state/energy [MS81]. states [Hor84a, Kir80, Ler84c, LP80, Bal84b, BMoF^{+}83, Fis83h, FW87, Gra84, HK87, Jur85a, Ler84c, Sta86, Zor86a]. static [Ano81-183]. station [Ano82-188]. stations [Fis84b, HMDF80, Jac82, You84, Zde83]. status [Cor87d, Gol81, Sum82a]. statute [McC85]. stay [Voe86d]. steadily [Cau82]. steam [Jur86c]. steel [Kap84e]. steering [Kap85d]. stem [Kec83, Wal82a]. step [DD85, Fis83f, Mok81i, Sti88a, Tru83e]. stepper [Ano82-191]. Steps [O’C86, Gut84j, Gut85b, Per82a]. Stereo [Per85a, Ano83-245]. Stereophonic [Jur82b]. still [Bal85, Ber81j, FW87, Flo82, Jur88c, Mok82f, Per80a, Wal84e]. stimuli [GBDC86]. stirring [Chr86c]. stirs [Ler83a]. stock [Sta87a]. Stop [Ano82-241, Zor87e]. stopped [PW84]. Storage [Whi83, Ano80-196, Ano80-193, Ano81-182, Che86, FZ85, LS81, Sha80c, TV89]. storage-module [Ano80-196]. stored [Whi83]. stores [Ano82-195, Ano82-192]. storie [Joh84]. storm [Fle85c]. story [Chr81e]. straight [Chr81e]. Strategic [Cor87d, Ler82j, Per82e, WB86, Ada85a, Wal87g, Zor86b]. Strategies [BMoF^{+}83, Ler82k]. strategy [Jur87d, Kah81, Zor87b, Zor89g]. Streams [PC80]. streetcar [KW86]. strength [San89b]. strengths [Wet82a]. stress [Bel84f, Tus86]. stretched [Per86]. striking [BMoF^{+}83]. strive [Sta87b]. striving [Log83]. structure [Ler84b]. struggle [GP81, Kap85c, Mas80c, Mun88]. student [FP84, Per80b]. Students [Bel84b, Ber83a, Cor87c, Per80b, Pyk87, Van84]. studied [Cas81, Kuw84]. Studies [Mac84, Eis84, Fis84a]. studios [Ano84-53]. study [Ada87b, Ano87-74, Asa87, Bel87a, Hor86b, Kap83c, KM87, NK87, Ril87, Tor80, Voe87b, Voe87c, Wal87b, Wer87b, Wis82]. studying [Bel85c, Ler81f, Tru83b]. stuff [FT85]. sturdy [Bel87a]. style [Wis82]. styles [Wis82]. stymie [Jur81b]. subelements [Sha87]. subjective [BE89]. submarine [Bel87b]. submicrometer [CS87, Eid81, Lep81]. submillimeter [Ano80-243]. submillimeter-wavelength [Ano80-243]. subpicosecond [ST81]. subscriber [CC85]. subside [Ler80b]. substances [Lom80f]. substantially [Ler84g]. substations [Ver86]. subsurface [MR82]. subsystems [Jur86b]. suburbs [Jur85a]. Subzero [Hil86]. succeed [FP84]. succeeded [FW87]. success [Gut82h, Kap83d, Pak85b, Wil81]. successes [Ano88d, Ano88m, Lom80a, Lom82, SH85, Sta86, Ste81]. Successful [Kec83, Ano80-68, Ano83-169, HT83, Her84, LH87, Per82a]. such [Cau82, EMM82, Lur84, Pan87a, Wal84g, Whi83]. suffer [Car84]. suggest [Mac84]. suggests [Dou83b]. suite [Fis87a]. Summer [Cor84]. sun [Jur85a, Bac80]. sunlight [Bac80]. sunshine [Sti88c]. sup [Gut89]. Supercomputer [PZ89, Ano86-192, Ano87-234, SF82, Sug80c, Wal85b, WZ87]. supercomputers [Ber82c, MS84, Per87a, Wal85b, WZB86]. Supercomputing [BMR88]. Superconducting [Kap80f, Jur87g, Kir83]. Superconductivity [Fit88e, Fit88g]. Supercool [Kir83]. supercooled [Kap80f]. superelectrification [de 82]. Superfast [Bro81b, Dup82, ELW83, MS84, SW83]. Superfine [Lea85]. superhet [Har83]. Superhigh [Per85d]. Superhigh-tech [Per85d]. superior [Har83]. superminicomputer [Kid81]. Superminicomputers [Wal85b]. Superpower [Sug80b]. superpowers [Hor86a, Ler82k]. Superproject [Her83d]. Supersmart [Gut83h]. superstructure [FHW85]. supervise [Liu84]. Supervisory [Mas85]. supplant [Lea85, Sug80b]. supplemental [Ano82-157]. supplier [Gau85]. Suppliers [Zor85f, Fis85b, Fly87]. supplies [Syk89]. supply [Ada85c, CS87, Fis83d, KZG^{+}84, Mun88, Tru83d]. Supplying [Fis83f, Bel86a]. support [AH85, Ber80b, Per84d]. supported [Gut83d]. supports [Ano82-186]. Supreme [Per82c]. surely [Ler81c]. surface [KP86, Tur87]. surface-mounted [KP86]. surface-mounting [Tur87]. surge [Bel86a]. surgeon [Jak85]. surgery [Wel84]. surgical [Jak85]. surprise [BE87, Chr80b, Fis86c]. surprised [Sha81e]. surprises [Ano88p]. survey [Ada84b, Kap84b, Lom80c, Per81c, WM84]. Survival [Jur87d, Ler82i, DM85, Get82, Tri80, Wol84a]. survive [Gre80, Lak82, WZ87]. survivor [Per88d]. survivors [SS84]. suspects [Ler83e]. suspension [Jur83a]. Suspensions [Jur86b]. Swan [Wis82]. swath [Jak85]. Swedish [Zor88a]. sweep [Ano81-180]. sweeping [FZ87a]. switch [Luc80, You84]. Switched [Sol88]. Switched-capacitor [Sol88]. switches [Kap87c, MS84, Zor85f]. Switching [Ano80-42, Bal81, Bel86c, BS81, Che87, Luc80, Mel80, Mok82c, Pin80, RK80, ST81]. swords [Sti89b]. SX [Per89a]. SX-70 [Per89a]. syllables [Sar82]. symbiosis [Whi87]. sync [Bha85]. synfuel [Fag80c]. synthesis [Ler82b, de 89]. synthesize [Hec85]. Synthesized [KL85, Lem83a]. synthetic [Jur83a, KL85, SAPM80]. system [AW85, Ada87e, Ano80-197, Ano81-183, Ano81-181, Ano82-188, Ano83-187, Ano83-185, Bel89c, Ber82a, Ber82f, BKBC85, BS81, Chr87h, Eid81, Gev83, Gol81, Gut84h, Hor85d, Kap82c, Kap83a, Kap84d, Kap87a, Kap85f, Kat80, Kuw84, Lep81, Ler81g, Lla81, Lom80d, MMB83, Mok82f, Mur88a, PW86, Per86, PHG86, PK81, Rab81, Red80, Ros83, Sim83, Voe87j, Wal85e, Wet82b, Bel85b, PW83]. system-level [Ber82f]. Systematic [Chr82a, Fis84a, Per83d, WAO85]. Systems [Ber82b, Riv87, Vae86, Ada88e, Ano81-133, Ano83-141, Ano84-168, Bel85f, Bel87b, Ber81d, Ber81g, Ber82g, BKBC85, BA83, Bug84, BS81, Che82, Cla80b, DM85, Dav87a, DeL82, Dou83b, Fag80a, Fag80c, Fag82c, Fis83b, Fis83f, FZ85, FHW85, FZL86, Fis80, Fit87c, Fit88c, Fit89c, GP81, GGZ86, Gev83, Gra84, Gut82f, Gut82e, Gut83f, Gut83g, Her83b, JJ87, Jur81a, Jur82b, Jur87b, Kap81c, Kap82c, Kap84c, Kap84e, Kap85b, Kap85d, KP86, Kap87c, KS86a, KK84, Lem81a, Ler81i, Ler83b, LH87, Lom82, Luc80, Mar86, Mas85, MS81, MNS81, Mel80, Mid80, Mok82a, Mok82d, MD84, NS80, OK86, PW84, PFMP88, PHG86, Rau87, SS84, Sha81f, Sha87, Sta86, Sta87b, Sug80a, Sug80b, Sum80, Tal83, Tan89b, Ten82, Tri82, TD84, Tul83, TBH87, Voe87h, Wal84g, Wal85d]. systems [WAO85, Zor85a].

T [Bel85a, Cor84, Mur85, Wet82a]. table [Ano80-68]. tables [Ano80-153, Ano80-154, Ano80-155, Ano80-156, Ano80-157, Ano80-158, Ano80-159, Ano80-149, Ano80-150, Ano80-151, Ano80-152, Ano81-138, Ano81-139, Ano81-140, Ano81-141, Ano81-145, Ano81-146, Ano81-147, Ano81-148, Ano81-149, Ano81-142, Ano81-143, Ano81-144, Ano82-140, Ano82-141, Ano82-149, Ano82-150, Ano82-142, Ano82-143, Ano82-144, Ano82-145, Ano82-146, Ano82-151, Ano82-147, Ano82-148, Ano83-147, Ano83-148, Ano83-149, Ano83-150, Ano83-151, Ano83-152, Ano83-153, Ano83-154, Ano83-158, Ano83-155, Ano83-156, Ano83-157, Ano84-150, Ano84-151, Ano84-152, Ano84-153, Ano84-154, Ano84-155, Ano84-156, Ano84-157, Ano84-158, Ano84-159, Ano84-160, Ano84-161, Ano85-145, Ano85-146, Ano85-147, Ano85-148, Ano85-149, Ano85-150, Ano85-151, Ano85-152, Ano85-153, Ano85-154, Ano85-155, Ano85-156, Ano86-134, Ano86-135, Ano86-136, Ano86-137, Ano86-138, Ano86-139, Ano86-140, Ano86-141, Ano86-142, Ano86-143, Ano86-144, Ano87-153, Ano87-155]. tables [Ano87-156, Ano87-157, Ano87-154, Ano87-158, Ano87-159, Ano87-160, Ano87-161, Ano87-162, Ano87-163]. tabulated [Kap84b]. tackle [VWZ86a]. tackles [Per80c]. Tactical [Ler82i]. Tactile [DD85, Pic82]. tailor [Gut83c, Kap85f]. tailored [DGZ88]. take [Ber80b, BKW83, Bow85, Cog84, Fis84h, Jur84]. Taken [Fis86c, Del80, Gut84j]. takes [AM86]. taking [Spi86]. talents [Chr85i, Per84a]. talk [Chr87a, FHW85, FW82, Ler82b, SAPM80]. talked [Ber80a]. talkies [Voe87d]. talking [Ano85-125]. Taming [Kap84f, WAO85]. tandem [Lof80b]. tank [Gre80]. tape [Chr80k, Mok82a, Per87b, TV89]. tapes [Ros83]. tapped [Par84]. Target [Row82, Aml82, Den81, Per81b]. targeted [Lof80a]. Targeting [Fit87c]. targets [AW85, Cal82, Cas81, Mok80f]. tariff [Kuw84]. Tariffs [Kap85c]. task [Cor84, Rab81, Red80]. task-oriented [Red80]. tasks [Bel85d]. tastes [Per85d]. taught [Pyk87]. Taylor [BPEB84]. teacher [Cye84]. teaches [Ber81k]. teaching [HK87, Van84, Wit84]. team [Bha85, Gut84e, Kid81, Per85d, Wal84a]. teams [BMoF^{+}83]. tech [Ada89a, Chr84c, Fis85a, FHW85, Fle85a, JP85, Lar85, Nai83, Per85b, Per85d, Tru83e, Win84, Zor84c, Zor85c]. Technical [Ano84-86, Chr83i, Hor86a, Sha81c, Wil80, Ada86b, Bel85f, CN83, Far86, Fis86a, Gut84e, Kap82b, Ler82b, MRRS84, Per80c, Per84a, PK81, Row82, San88a, Sha80f, Tor80, VWZ86d, Wal82b, Wal84b, Wet82a, Wet82b, Zor89f, Far86]. Technically [Ano80-244, Ano81-224, Ano81-225, Ano81-226, Ano81-227, Ano81-228, Ano81-229, Ano82-242, Ano82-243, Ano82-244, Ano82-245, Ano82-246, Ano83-233, Ano83-234, Ano83-235, Ano83-236, Ano83-237, Ano83-238, Ano84-230, Ano84-231, Ano84-232, Ano84-233, Ano84-234, Ano84-235, Ano85-227, Ano85-228, Ano85-229, Ano85-230, Ano85-231, Ano85-232, Ano86-225, Ano86-226, Ano86-227, Ano86-228, Ano86-229, Ano86-230, Ano87-229, Ano87-230, Cor87e, Cor87f, Cor87g, Cor87h, Chr83k]. technicians [Gut84i, Per86]. technique [Che86, Dav83a, GGZ86]. Techniques [EMM82, Ant87, BMoF^{+}83, Bro88, Cas81, Fag82b, Gar81, Hil86, Hor86d, Lem81a, Ler81g, Ler83e, LH87, Mar85, MR82, Mur88b, RD81, Sha80d, Tul83, Van84, Wel84, Whi83]. technocrats [Gut88]. technological [Chr87a, Gut82d, Kah82, Kap80f, Lur84, Mar85, Mas80c, Wal83a, WZ85, Wil80, Zor86a]. technologies [Ano84-236, Ano86-66, BPEB84, Che82, Fis83c, Fis83i, Fis84e, IK84, Kap85f, Per85a, Tru83e, Tru84a, Wal82d, Wal84b]. technologist [Sug80c]. Technologists [Gut84i, Fis86c, Lom80f, Per84e]. Technology [Ada88f, Ano80-245, Ano80-246, Ano83-239, Ano84-236, Ano87-231, Bel88f, BA89, Bel89e, Ber82i, Bug84, Cau82, Chr80m, Chr82n, CHG^{+}83, Chr84i, Chr84m, Chr85m, Cor88a, Dou83b, Fit88f, Fit88g, Fit89d, Fit89e, Fit89f, Gra82c, Gut83f, Gut84j, Gut85d, Hor86c, Jur86e, Jur88d, Jur88e, Jur89c, Jur89d, Kap88, KM88a, Lem83b, Ler82j, Lom81c, Mai83, MW88, Mur89d, Mur89e, Per83g, Pro81, Rub80, Rub82, Rub84, San89d, San89e, Sti89d, Sti89e, Van84, Voe88f, Voe88g, Voe89c, Voe89d, Wal82a, Wol88b, WZ88, Wol88c, WZ89, Zor88e, Zor89i, Zor89h, Ada85c, Ano82-152, Ano83-162, Ano85-163, Ano85-164, Ano85-165, Ano85-166, Ano86-154, Ano86-155, Ano86-156, Ano86-157, Ano86-158, Ano86-159, Ano87-168, Ano87-169, Ano87-235, Ano87-236, Ano87-237, Ano87-238, Bag84, Bal84a, Bar81, Bel84d, Bel84a, Ber83a, Ber81j, Ber81k]. technology [Che86, Chr81j, Chr82k, CN83, Cor87a, Cor87d, Cor87c, DeL82, DS81, Fag82b, FM84a, Fle85b, Gau85, Gre80, GT82, Gut82f, Her83c, Joh84, Kah84, Kap87d, KC87, Lem83a, Miy89, Mor81a, Mor81b, Mur88b, Per83e, Per83f, Rab81, Ros80, Sha80e, Sum80, Tan86, Tri80, Tru84a, Wal82b, Wal87h, WM84, Wei80, Wil80, Wol87a, Wol87b, Wol87c, WC87, Zor88a, BPEB84]. technology-induced [Mor81a, Mor81b]. technology-transfer [Chr82k, Tru84a]. Technology’81 [Chr81n]. tedious [Fis85e]. Telecommunications [Fis83g, Gut84k, Per80c, Wei87, Bel85f, Bel89e, Gut83b, Hor85d, Jay86, KZG^{+}84, Kap85c, Ler84f, Mid80, Mok80f, Sta86, Tor85, Wal85c, Wet82a, Zor85f]. Teleconferencing [Hef84]. Telephone [BS81, Dor83, Hil83, Sti88d, HT83, Jac80, Luc80, NS80, Pin80]. telephones [Mel80, Pan87a, Zor85f]. telephony [Kap83b, Mok82c]. teleterminal [HT83]. Teletext [Jac80]. Television [Ano84-237, HMDF80, Ano84-53, Bal84a, Bal84b, Bel84e, Bel84d, Hil83, Jac80, Jac82, Jur82b, Jur83b, Jur85a, Jur88a, Mok80c, Mok81g, Per83d, Per84e, Per85a, Sch81, Wit84]. television-based [Bal84b]. tell [Jur83a, Per87d]. tells [Pak85b, RK80]. Telvision [Luc84]. temperature [Ler84g]. tempered [Far86]. tended [Mos81]. tenfold [LS81]. tension [Gut83a]. Tensions [Bel87f]. terabyte [Ano81-187]. term [Fis83c, Per80a, Per83f]. Terman [Bry84]. terminal [Ano82-192, HT83, Kap87a]. Terminals [Swe82, Ber81d, Mel80, Mok80f]. Terminations [Fle85c]. terms [Whi87]. terrorism [Hor86c]. terrorists [Mur89c]. Test [Blu80, Ber81h, Hou85, Kap84f, LM81, Per84d, Swe82, Tul83, VB86, WSC89, Zor86b]. testability [SA85]. tested [Dou81, Emm85, Kap80c, Kap85d, McC85]. Testing [Gut84l, Ber81h, Cau82, Gut84g, Hor86d, Ler81g, Ler82e, Mas80b, SA85, Tul83, Wil81, Zor88d]. Tests [GS80, Ano83-186, Gut85b, Kuw84, Ler81i, Lom80a, Mok80a, Wei84, Zor86b]. Texas [Lom80e, Wol85]. Text [Jac80, Bal85, Sar82, Sha81f]. text-handling [Sha81f]. thankless [Cor84]. thanks [Hef84, VWZ86b]. Their [Jak85, Ano80-68, Ano80-245, Bel85d, Bel87f, Chi81, CC87, Fis85e, FW87, Gut82g, Hor86d, Jur86d, Kap81a, Kat87, KS86a, KK84, Liu84, MMB83, Mok82e, Per81e, PW84, Per84c, Per84d, PW86, Per87d, Per87f, ST81, Swe82, Tri82, Tru84b, Voe87i, Wal84a, Wal84e, Wal86c, Wal87f]. them [Ano84-247, Cal82, DM85, Fag81a, Fis87b, Fit87a, Har83, Hor84a, IK84, KSB87, Lem81c, Nah83, Per80b, Per83a, Per87d, Sha80a, Sta87b, Van84, VWZ86e, Wal86a, WZB86, Wal87h, Wol83]. theme [Cog84, Fle85c]. themselves [DTS86]. theoretical [Ber80f, Ear84]. theoretically [Ham86]. theories [Con86, Ler84a]. Theory [Kli84, BS89, DS81, Zad84]. therapy [Jak85, SPM^{+}80, Wel84]. There [Ang80, FM84b, Ano83-248, Lut88, Rac80, Wei84]. thereafter [Mih81]. thereby [Cye84]. thermal [Ano84-252, Jur86c, KSB87]. thermal-mechanical-electrical [Jur86c]. thermonuclear [Ler80d]. These [Wal84f, AF87, Ano84-148, Ano88k]. They’d [Fis87b]. they’ve [Wal84e]. thieves [Hor85e]. thin [MO89]. thin-film [MO89]. think [Ano82-220, Ano82-248, Gre80, Gut82a, Mor81b, Zad84]. thinking [Hor86b]. third [Kir80, Ler84f]. thirds [Hic83]. Thomas [JI84, Wis82]. Thomson [Car83]. thoroughly [Emm85]. those [Bug84, Gev83, Ler82h, Mur89b, Per81c, Per87d, Tri82, Voe86a, Voe86b, Voe87a]. Though [Wal84e]. thought [Ber82j]. thousand [LH87]. threat [Emm85, Per82c]. threaten [Lom81c]. threatening [Kah82]. threatens [FAEH84, Win84]. Threats [Kap82a]. Three [Ada85a, Ano81-187, Cle82, Fis86d, For86, Jur82b, Lem81b, PK81, Ano82-189, Bro81b, Cor87b, Dan82, Fis84e, Jur88b, PHG86, Sha80f, Wei84, DeV81, FAEH84, Lom81b]. Three-chip [Ano81-187]. three-dimensional [Bro81b, Cor87b, PHG86]. thrill [Kap81a]. throughout [You84]. throughs [Tal83]. thrusts [Fei82]. thumb [Gev83, Mos81]. thwart [SD81]. Thwarting [Hor85e]. Thyristors [FEK80, Kap82d, Che87, Kap85d]. tides [And80]. tightened [Lom81c]. tightrope [Chr82k]. tim [Wer87b]. Time [Hic83, Ano80-197, Ano81-182, Ano81-180, Ano82-185, Ano87-225, Chr84f, Chr85h, Chr88b, CC86a, Fis84d, Fle82, Gau85, Hec85, HK87, Kid81, Ler84h, Liu84, Lom80c, LP80, Mur89b, OK86, Per83e, Per83b, Ros80, VWZ86c]. time-base [Ano81-180]. times [Ber82c, CC87, FWVH86, Ste81, Tru83a]. timing [Chu88]. Tinkertoy [Jur87f]. tissues [SPM^{+}80]. TMI [DeV81, Eis84, FAEH84, HOD82, Hor84c, Lom80e, Lom81b]. Today [Ano84-86, Ano84-168, Gut82i, Wal82c, Wal82d, Ber82c, DS81, Fis83c, Gar84, Pea81, Per80b, PW84, Sha80a, Sha81c, VWZ86b]. together [SK89, Wal83d, Wal83e, Wal86c, Zor87e]. tokamak [Kap84f]. token [Gra82b]. Tokyo [Bha85]. told [Sta87a]. tolerant [Zor85b, Zor85e]. toll [Zor85f]. Tom [Pow84]. Tomorrow [Gar84, Tor83, Beg83, Bel85f]. tongue [Fle82]. Too [Sta87b, Ada85c, Bry84, Chr87k, Den81, Fit89a, Fit89f, Mur89a, Per84c, Per85d, Sha87]. took [Cor84]. tool [AJ85, Bot83, Cun80, Fit88a, Lom81c, SPM^{+}80, Tru83b, Van84, Win84]. Tools [Gib85, Mih81, Mus89, Ano80-102, Ano80-103, Ano80-104, Ano80-105, Ano80-106, Ano80-107, Ano81-95, Ano81-96, Ano81-97, Ano81-98, Ano81-99, Ano81-100, Ano81-150, Ano82-98, Ano82-99, Ano82-100, Ano82-101, Ano82-102, Ano83-108, Ano83-109, Ano83-110, Ano83-111, Ano83-112, Ano83-113, Ano83-114, Ano83-115, Ano83-116, Ano84-116, Ano84-117, Ano84-118, Ano84-119, Ano84-120, Ano84-121, Ano84-122, Ano84-123, Ano84-124, Ano84-125, Ano84-126, Ano84-127, Ano85-113, Ano85-114, Ano85-115, Ano85-116, Ano85-117, Ano85-118, Ano85-119, Ano85-120, Ano85-121, Ano85-122, Ano85-123, Ano85-124, Ano86-104, Ano86-105, Ano86-106, Ano86-107, Ano86-108, Ano86-109, Ano86-110, Ano86-111, Ano86-112, Ano86-113, Ano86-114, Ano87-101, Ano87-102, Ano87-103, Ano87-104, Ano87-105, Ano87-106, Ano87-107, Ano87-108, Ano87-109, Ano87-110, Ano87-111, Ano87-112, BA83, Bug84, Fit87a, Hef84, IK84, Kap83c, Mur88b, PP88]. tools [Voe87a, Wal84a, Wal87g, Wer87a, Kat87]. top [Ada86a, DTS86, Her84, Tan89b, Zor87a]. top-notch [DTS86]. top-secret [Zor87a]. topic [Ano84-247]. topics [Ada84b, Ano87-204, Ano87-205, Ano87-206, Wal87e]. tortuous [Mok80e]. total [Jur86b, Jur87c]. totally [Mos81]. touched [Cla80a]. tough [AM86, Che87]. tougher [Ano86-66, JT81]. toughest [Rab81]. tour [Fis86d]. Tower [Kah82]. town [Kap83g]. toys [Ano80-102, Ano80-103, Ano80-104, Ano80-105, Ano80-106, Ano80-107, Ano81-95, Ano81-96, Ano81-97, Ano81-98, Ano81-99, Ano81-100, Ano82-98, Ano82-99, Ano82-100, Ano82-101, Ano82-102, Ano83-108, Ano83-109, Ano83-110, Ano83-111, Ano83-112, Ano83-113, Ano83-114, Ano83-115, Ano83-116, Ano84-116, Ano84-117, Ano84-118, Ano84-119, Ano84-120, Ano84-121, Ano84-122, Ano84-123, Ano84-124, Ano84-125, Ano84-126, Ano84-127, Ano85-113, Ano85-114, Ano85-115, Ano85-116, Ano85-117, Ano85-118, Ano85-119, Ano85-120, Ano85-121, Ano85-122, Ano85-123, Ano85-124, Ano86-104, Ano86-105, Ano86-106, Ano86-107, Ano86-108, Ano86-109, Ano86-110, Ano86-111, Ano86-112, Ano86-113, Ano86-114, Ano87-101, Ano87-102, Ano87-103, Ano87-104, Ano87-105, Ano87-106, Ano87-107, Ano87-108, Ano87-109, Ano87-110, Ano87-111, Ano87-112, Mok81e]. track [GS80, Sha87]. tracking [ABEI84, Miy84]. tract [Fle82]. trade [Ano88o, Lom81c]. tradeoffs [CS87, Kap85f, PW85a, Wal87d]. Trading [VWZ86d, Sta87a, Tru84a]. traditional [Bel84e, Fis85b, Hor84b, Mac84, Mur88b, Voe86d, Wet82a]. traffic [Jur87e, Ler82a, Lom80d]. tragedy [Ano87-28, Sti89a]. train [Dup82, Fag81a, Mur88a]. trained [Ler82h]. Training [Ano81-150, Per80a, Tru83d]. trains [FEK80]. transducers [Bar81, MNS81]. transfer [Chr82k, Chr84i, Gra82c, Kap81b, Lom81c, Sha80c, Tru84a, Wal82b, Wal82a]. transferring [Ano82-186]. transfers [Ano82-156, Fle85c, Kut83]. transform [FM84a]. transformed [Per88a]. transformer [RD81]. transforms [Kut83]. transients [Kap83c]. transistor [Bel86c, MS84, NH86, PD86, TP83, Wol85]. transistors [BCUY82, Bel86a, Che87]. transit [Kap80c]. transition [Ada89c, Gut83i, Lom82]. transmission [Ano83-51, Ano87-234, Bag84, BCUY82, Fis83f, Jur82b, Jur85b, Kap81c, Kap82d, Kuw84, Mok82d, Rog81, de 82]. transmissions [Jur87e, PHG86]. transmit [Fis84f]. transmitted [Jac80]. transmitters [KK84]. transplant [Ano83-169]. transport [Bel86c, Voe88e]. Transportation [Ano84-236, Dup82, GS80, Kap85d, KA86, Kap87c, Kap84e, KM88a, Mur89e]. traps [Rac80]. travel [Kap85d, TBH87]. travelers [FP86]. treatment [Gar84]. Treaty [Buc82]. tree [Bel87a]. trellis [Jur87c]. trellis-coded [Jur87c]. trend [Ano84-236, DM85, Jur86b, Kah82]. Trends [Fei82, FG86]. trial [Mil87]. trials [Chr84e, Kec83, MRRS84]. triangle [LS81]. tribulations [Chr84e]. tribute [Chr83d, Fag80f]. tried [MRRS84]. trigger [Ler81f]. triggered [Ada87f, FAEH84]. trip [LM81]. triumphs [Ano88d, Ano88m, Fis84h]. trolley [KW86]. trouble [Che89b]. troubleshooting [Ler82e]. true [Moy80, Per84d]. truly [Per85c]. trying [GP81]. tube [Bel84e, Bel84d, Zor87a]. tubes [Tha83]. tumors [SH85]. tuning [Har83, LM81]. tunnel [Ano84-253]. tunnels [Del80]. turbines [CA82]. turbulence [SW83]. turbulent [Wol84a, Wol84b]. turmoil [Roi86]. turn [Ano86-66, Bel87e, JT81, Wal82d]. turned [Fit88a]. Turning [DS81, Fis84h, Riv80]. turnoff [Che87]. turns [Gut84e, Sta86]. turntables [Mok82a]. tutorial [Gut83b, Ler84d]. TV [Per88d, Bag84, Bal84a, Jac82, Ler83d, Mok80c, Mok81i, Per84e, Roi86, SD81, Wit84]. Twelve [Ano87-234]. Two [Gut84i, Lom81b, Ano82-190, Ano82-194, Ano83-187, Bal84a, Bel84e, Bel84d, Dut80, Gra82b, Hic83, PL89, Sum82a, Wis82, Wol84b, SS84]. Two- [Gut84i]. two-part [Bel84e, Bel84d]. two-way [Dut80]. type [Gev83, Kap84b]. types [Bel85d, LS81]. typesetting [Jur87a]. typewriter [Joh84]. typical [Ros83].

U.S [Dan82]. U.S. [AH85, Ada87a, Ada87e, Aml82, Ano80-245, Bel86a, Cha82, Cor87d, Dan82, Fei82, Fis83d, FAEH84, Fis86e, FP86, FZ87a, Fis87d, Fle85b, Fre84, Fru83, Gan82, Gle83, Gol81, Gra82c, Hor86d, JCH80, Jur81b, Jur82b, Kah81, Kah82, KW86, Lak82, Lem81b, Ler81f, Ler82a, Ler82k, Ler84f, Lom80b, Lom81c, Mac84, Mas80c, Per80b, Per82c, Per83d, Per83a, PW86, Per87f, Pyk87, Riv80, Ros88, Ros80, Tri80, Tru82, Tru84a, Wal82b, Wal82a, Wal85d, Wal86c, Wei80, Wil80, Zor86d]. ultimate [Bel86c, Fis83b, Per83e, ST81]. ultimately [Bel85b, BFB87]. ultrafast [SF82]. ultrasonics [Wel84]. ultrasound [Wel84]. unaffordable [FP86]. unanswered [Gut83i]. unavailable [Cye84, Pyk87]. unaware [FP86]. unbreakable [Kah84]. uncertain [Lem81c, Ler80d, Zor89f]. Uncertainties [Bel84a]. Uncertainty [Fle85c]. Uncol [Ano81-234]. uncommitted [Ber80g]. uncommon [Ano88j]. unconnected [Bel86a]. unconventional [Fis85b, Hil86]. Undaunted [Fle85b]. undergoes [FZ87a]. Underground [Hor86d, Del80]. underlie [Wal84b]. underlying [Ant87]. undersell [Fis87d]. understanding [HTW84, Ler80b]. understood [Ant87]. undertaking [Bee83]. underutilization [Chr82i]. underwater [Zor88f]. undisturbed [Per84c]. unemployment [Flo82]. unending [GP81]. unethical [Per81d]. unexpected [Bel87c, San88c, Zor87f]. unified [BS89]. Union [Gra84, Lom81c]. unit [HT83, SAPM80]. United [Bal84b, BMoF^{+}83, Fis83h, FW87, Gra84, HK87, Jur85a, Ler84c, Sta86, Zor86a]. units [FZ87a, Ler81e, Mas85, Mok82e, Per81a, Wal87h]. Univac [Ste81]. Universal [Ano80-197, Ano83-186, Pan87a]. universities [Fag81a, FM84b, Mas80c]. University [Ber83a, Bal84b, FM84a, Fri84, Gra82c, Mal84, VH86, Bry84]. UNIX [Gut84h]. unless [FP86]. Unlocking [Cen89]. unmanned [PHG86]. unmasking [Fre86]. unnecessary [Wal85e]. unprecedented [Voe87d, Wal87h, Zor85a]. unreached [Ler82j]. unsolved [CN83]. unstable [Fis83d]. unsuccessful [Fle85b]. untangle [Per83d]. until [IK84]. Untying [Mok81h]. unveiled [Ada85a]. unveils [GT82, Jur83a]. unwary [Rac80]. unwieldy [Rab81]. update [Jur88a, Tru84a, Voe87h]. upgrade [Voe87c]. upon [Ber80b]. upset [Zor86b]. Uranus [Ada87a]. urban [ABEI84, Kap83a]. urgent [Lom80f]. usage [Gel81, STK82]. use [Ber80a, Cor87b, Dou83b, EMM82, Gar84, Gut84e, IK84, KA86, Kec83, Kir80, Lem81a, Ler82k, Mar85, Mok80d, Mok82b, OK86, Per85c, ST81, Sta87b]. used [Ano85-125, Fag80b, Kap84f, Ler84g, Voe87e, Wal87h, Wei84, Whi87]. useful [Bel85d, Fag80a, Fis87a, GGZ86, Mos81, VWZ86b]. usefulness [Jak85, Wal87f, Wel84]. useless [Lem81b]. User [Fre88a, WM84, Ano80-195, Dav87a, Dou83a, Gau85, HT83, Ham86, Mok80d, PV89, VWZ86a, Voe87g]. user-friendliness [VWZ86a]. user-friendly [PV89]. user-partitioned [Ano80-195]. User-programmable [Fre88a]. users [BFB87, Den81, Gut82h, Gut83h, PW85a]. Uses [AF87, Ano82-191, Ano82-188, Ano83-187, Ano83-185, Bot83, Dou83b, Pro81]. ushered [Voe88a]. Using [Cun80, Gev83, Jur86c, Bac80, Dut80, Kap83c, Lem81c, Ler80d, Tan86, Wal84e]. USS [Ada87f]. USSR [Ler82k, Ler84g]. usurp [Ano84-168]. utilitie [Wel85]. Utilities [Fis83f, FZ85, Kap81c, Per81e, Ano86-66, DM85, Fis84f, Fis86a, FZ87a, Gel81, Kap80e, LP80, Ric88, Zor84c]. Utility [Gut82h, Fis83c, Mas85]. utilization [CC86b]. utilized [Cal82].

V [Ano81-230, Car80, Kat87]. vacuum [PC80, Zor87a]. vacuum-tube [Zor87a]. validates [Ano83-186]. values [Nai83]. vanished [Voe87i]. variable [Jur87e]. variables [MNS81]. Variations [Cog84]. variety [Cal82, Rau87, Vae86]. various [Ano83-204, OK86, Zor84b]. varying [Kat80]. vast [Fis86d]. VAX [Bel87a, WZ87]. VAXes [WZ87]. VCR [Per88a]. vehicle [AM86, GS80]. vehicles [Fly87, TBH87]. vehicular [Voe87d]. venerable [CC86b, Fit88a]. vengeance [Sum80]. ventures [Fle85b]. verge [Kap83d]. Verification [Chr86j, Ada86b, Buc82, Kat87, San89c]. verify [Ada88g]. verifying [Sha87]. versatile [Bar81, Ber80c, Sha80d]. versatility [AF87, Pin80, Sha82]. versions [Ano81-180, Fis84e, Her83b, VWZ86b]. versus [Har83, Kli84]. vertical [Ano80-193, Whi83]. very [Ano83-162, FHRW^{+}83, Mas80c, Ros80, Sho88, Sum80, Tru83e]. very-high-speed [Sum80]. very-large-scale [Mas80c]. vessels [Ada85c]. vested [Ler81b]. VHDL [WSC89]. VHSIC [Sum82a, Sum82b]. VI [Voe87h]. via [Bal84a, Dut80, SS84]. victim [Rab81]. Victorian [Hec85, Lur84]. Video [Ano83-242, Ano84-241, Ano84-242, Ano84-243, Ano84-244, Ano84-245, Ano84-246, Ano85-236, Ano85-237, Ano85-238, Ano85-239, Ano85-240, Ano85-241, Ano86-234, Ano86-235, Ano86-236, Ano86-237, Ano86-238, Ano86-239, Ano87-235, Ano87-236, Ano87-237, Ano87-238, Cle82, Kap87d, KC87, Mok82f, PTW82, Per83e, PW83, Bal84a, BKW83, Hef84, Lut88, Mok80d, Per81a, Per84d, Per85a, Tru83e, Mok81c]. video-disk [Mok82f, Per83e]. video-game [PW83]. videomovie [Ano87-74]. videotex [Tru82, Fle85b]. Videotext [Cla80b]. vie [Fis85b, Per84e, PK81]. view [Ano80-68, Ano87-226, Fle82, Her84, Joh84, Kap83g]. viewdata [Jac80]. viewed [Sum82a]. viewer [Bel84e, Per84e]. viewers [Jur85a, Mok81g]. viewpoint [Gra82c, Lom80e, Wal82a, de 83]. views [Ano80-245, Flo82]. vigil [Ada86a]. vignettes [Chr81i, Chr84a]. village [Ano84-236, MS81]. virtual [Ano81-187, Smi87, Zor85e]. virtual-memory [Zor85e]. virtually [Dou83b, Hil83]. vision [Kap84c, Kap85b, Kap87a, KS86a]. visionaries [Per87e]. visionary [Fag80f, Gra82a]. visit [Zor88f]. visited [Per83a]. visitors [Gut83b]. visual [Bel84b, BKW83, Pic82]. vital [Ber82c]. VLSI [Ano81-185, Ano83-162, Ano83-187, Ber80g, Ber81k, Ber81i, Ber82j, BKBC85, Fis83i, Gut84g, KM88b, Ler82e, Ler83d, Mas80c, Mok80f, Pow84, SA85, SD88, Sum80, Tul83]. VLSI-based [BKBC85]. VLSI/LSI [Ber81i, Ber82j, Fis83i]. vocal [Fle82]. vocal-tract [Fle82]. voice [Jay86]. voices [Jur83a]. volt [Tay89]. voltage [Che87, CS87]. Volts [Wal87h]. volume [Ber81h]. Voluminous [Eis84]. vote [Din82]. votes [Per84e]. voyage [DJ85]. Vs [Sea89, Fit88e, Hor86c]. vulnerability [Par84, Sti81]. vulnerable [Cha82]. vying [Fis84e, Jur82b].

W [Dou81, Ano81-184]. wafer

[Fis85f, MRRS84]. wafer-scale [MRRS84]. wafers [Fan89]. wait [Jur87b]. waiting [Che89b, Sti89a]. waits [Per85d]. wake [Ler80b]. walk [Fis85e, Sum82a]. walkie [Voe87d]. walkie-talkies [Voe87d]. Walking [Chr82k]. Walt [GT82]. Wang [Sha81f]. wanted [Tor80]. wants [Ler84h, Tru82]. War [Dan82, DeL82, Fei82, Kah84, Ler82k, Ler82j, Sea89, Wal82d, Bel87e, Chr87a, GT83, Kap82a, Kli87, Mur89c, Rub82]. WARC [Kir80]. warhead [Ada85a]. warmer [Ano82-157]. warn [Per86]. Wars [AF85, Ada89c, Kap82a, Ler84f, Mai83]. wartime [Ano87-202]. was [Ada87c, Ano88k, Bel85b, Bel87e, Ber80f, Ber82j, Car83, Fag80f, Har83, Jur85a, KW86, Lur84, MRRS84, Ryd84, San88a, Sti88a, Wol85]. wash [Bel85e]. Washington [Ano80-217, Ano80-218, Ano80-219, Ano80-220, Ano80-221, Ano80-222, Ano80-223, Ano80-224, Ano80-225, Ano80-226, Ano80-227, Ano80-228, Ano81-200, Ano81-201, Ano81-202, Ano81-203, Ano81-204, Ano81-205, Ano81-206, Ano81-207, Ano81-208, Ano81-209, Ano81-210, Ano81-211, Ano82-208, Ano82-209, Ano82-210, Ano82-211, Ano82-212, Ano82-213, Ano82-214, Ano82-215, Ano82-216, Ano82-217, Ano82-218, Ano82-219, Ano83-194, Ano83-195, Ano83-196, Ano83-197, Ano83-198, Ano83-199]. waste [Lom80f, Zor88c]. Wasted [Chr85i, Per84a]. Watchdog [Pro81]. watchdogs [Ada89b, Hor85a]. watched [Zor86b]. watchers [Gut83f]. Watching [Tru83b]. water [Ano82-253, Jur86c]. water-scale [Ano82-253]. waters [Wol84a]. watt [Bot85]. watts [Bal81]. wave [Bel88b, Per83e, Wol84a]. Waveform [Ano82-195, Ano81-182]. wavelength [Ano80-243, Kap85f, Mid80, Mok82d]. waves [Bot83, HGF85]. way [AF87, Ang80, Ano80-68, Bel85e, Bro81a, Den81, DMW87, Dut80, Fis83i, Fis84a, Kap82d, Kat87, KS86a, Mac84, Mas80b, Per87f, Ryd84, SA85, Swe82, Tus86, Wis82]. Ways [Ada88g, Ano84-247, Bel87c, Cog84, Hor86b]. weak [Ler81g, Ler84a]. weaken [Gra82c]. weapon [Fit88c]. weaponry [Ada86a, Fei82, Fis85c, Fit87c, Ros88]. weapons [Ada86a, Cal82, Fit88b, Fit89a, Flo88, GP81, Gut82e, Hor86d, Lem81c, Per82e, Per84e, SPM^{+}80, Zor86b, Zor89g]. Weather [Zor87e, Ler82a, Ler82c]. Weathering [Fle85c]. we’d [Fis87b]. Weighing [Ler81b]. welcome [Mas85]. well [Ano88p, Ant87, Chi81, DMW87, HT83, Jur86a, Kap83d, Kut83, Mar85, Mih81, Per81a, Per84c, PW85a, Zor85a]. well-designed [Per81a]. well-informed [Zor85a]. went [Zor87f]. were [Ano85-125, Bal84a, Esc86, Hec85, Hil86, Lur84, Sta87a, Ste81]. Wescon [Ano83-162, Ano84-247, Gut82g, Sha80e, Sha81c]. West [ABEI84, Gra84]. Western [Ang80, BMoF^{+}83]. Westinghouse [Bha85]. Whatever [Ano81-235, Ano81-236, Ano82-249, Ano82-250, Ano82-251, Ano82-252, Ano82-253, Ano83-243, Ano83-244, Ano83-245, Ano83-247, Ano83-246, Ano84-252, Ano84-253, Ano84-248, Ano84-249, Ano84-250, Ano84-251, Ano85-242, Ano85-243, Ano85-244, Ano85-245, Ano85-246, Ano85-247, Ano86-240, Ano86-241, Ano86-242, Ano86-243, Ano86-244, Ano86-245, Ano87-239, Ano87-240, Jur87e, Jur87f, Jur87g, MS87, Wat87, Voe87e]. Wheeling [Zor85d, FZL86]. wheels [Jur86a]. Where [Chr85k, Lom80f, DS81, KSB87, Mok81g, Rog81, Zor85d]. Whether [Lem83a, Esc86, Ler81f, Zor86b]. which [Bel87c, Bro81a, Mos81, Sta86]. While [Bel85d, Con86, Wal87f, Ber80c, Bug84, Fis83e, Fis84g, FZ85, Fre84, Gut84g, Jak85, Ler82k, Pro81, Rau87, Rub80, Voe86d]. whistle [Per81d]. whistles [Jur84]. white [Ler82f]. Who [Hos83, Tru83d, Ano88j, Fag80f, FP84, Ler82h, LP80, Per81c, Sta87b]. whose [Chr82g, Hor85e, Ros80]. wide [DD85, Den81, Wol84b]. wide-ranging [Den81, Wol84b]. widely [Bal85]. widen [Ano84-236]. widespread [Gut84h, Mok80d]. width [Ano81-177, Zor85a]. wife [Per85d]. Will [Bro81a, Ang80, Bac80, Bel85f, Ber81k, BFB87, Che82, Con86, Cor87d, Dou83b, EMM82, Far86, FHRW^{+}83, Fis84c, GS80, Gut82f, Gut82g, GZ85, Hor85d, Kel87, Kuw84, Ler84b, Ler84d, LM81, Luc80, Mac84, MMB83, Nai83, Pan87a, Pea81, Per84e, Pow84, Row82, Sha80e, Sha81c, Sta87b, Sug80a, Sum82a, Tru83d, Van84, Wet82a, Wol84a, Zde83]. William [Wol84b]. win [Bel87e, Pyk87]. Winchester [Voe87j]. Wind [CA82, SD81, Zor86a, Hil86]. Wind-induced [SD81]. wind-rich [CA82]. Windmills [Wet82b]. winds [HW83]. wings [Che89b]. Winners [Ano84-193, Ano87-234, Ano80-36, Ano81-38, Ano82-135, Ano84-50, Ano84-51, Ano86-48, Jur87d, Ano85-45]. winning [Joh84]. wipe [Lem81b]. wired [Fri84]. Wireless [Ber81d]. wires [Jac80, Jur86a, Per83d]. Wiring [Gut83e, Jur86a, Sti88d]. wisdom [Lem81c, Win84]. within [Ano84-237, CC87, HHH87, Voe87a]. without [Fit88d, Jay86, Mar86, Per82d, Tal83, Win84]. withstand [Ten82, Voe86a, Voe86b]. woman [Tru83c]. Women [Tru83c]. wonderful [GT82]. woodenheaded [Nah83]. word [Jur87a, Kat80, Sha81f, Wal84f]. wordlength [Ano82-193]. Words [Kap80a, Kat80, Red80]. Work [Joh84, Ano80-68, CA82, Cla80b, Fis84b, Fis85e, Fis87b, Fit87a, Mac84, Mac88, MS81, Per83a, Per85b, Per87a, Rau87, Tru83b, Wal87f, War82, Zor86c]. worked [Pak85b]. worker [Sha84]. workhorse [Bal85, Jur82c]. workhorses [Kap82d]. Working [Sti88e, Per87f, Wal84a]. workplace [Tru83c]. works [Che86, Fre86]. workstations [Voe87i, Wal84c, Wal84f, Wer87a]. World [Kap84b, Kir80, Ano88o, Chr80h, Chr86h, Fis83d, Gra84, GT82, Har83, Kap84f, Per84c, Per86, SW88, Wei80, XM89, Zor86d, Kah84, Ler84f]. worlds [Fro83, PL89, San89a]. worldwide [Bel87b, Cor87b, Sta87a]. worry [Kap82a]. worsens [Kap80e]. worsts [Ano88k]. worthy [Wal86a]. Would [Wei80, Ber82c, RD81, SF82, STK82, Wal87f]. wounds [SH85]. writing [Bel84b]. written [McC85]. wrong [Chr83l, Zor87f]. WSI [MRRS84].

X [Kap85e, Lep81, San89f]. X-29 [Kap85e]. X-ray [Lep81, San89f]. Xerox [Pak85a, Pak85b].

year [And80, Ano80-245, Bal84a, Bha85, Bro81a, Chr85e, Cor87d, Gut84i, Ler82a, Lom82, PW86, Wol83, ZJ86]. years [Ano88p, Ber80a, Gre80, Kli84, Lom81b, Sum82a, Voe87j, Wol84b]. yield [HTW84, Wal84c, Zor86a]. yields [Bal81]. yore [Tru83c]. York [AM86, Ada87e, Kap83d]. Young [Ano80-68, Per80a]. younger [Chr80h]. zapping [Ant87]. zeal [Ano81-152]. Zero [Ham86]. Zero-defect [Ham86]. zeroed [Ver86]. zones [Rog81].

References

Alscher:1984:PPF

[ABEI84] H. Alscher, I. F. Boldea, A. R. Eastham, and M. Iguchi. Propelling passengers fester than a speeding bullet: As maglev urban railways get tracking, lines in West Germany, Japan, and Canada are expected to follow England’s commercial lead. IEEE Spectrum, 21(8):57–64, August 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Acton:1987:F

[Act87] K. Acton. Forum. IEEE Spectrum, 24(3):8–11, March 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1984:AMN

[Ada84a] J. A. Adam. Aerospace and military: a new space communications satellite, demonstrations of the MX missile, and infrared mapping are highlighted. IEEE Spectrum, 21 (1):86–89, January 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1984:ELL

[Ada84b] J. A. Adam. EEs on lifelong learning: a survey of readers indicate how they keep current, assess recent learning experiences, and rate course topics. IEEE Spectrum, 21(11):92–98, November 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1984:SCO

[Ada84c] J. A. Adam. A slow comeback:

The optimists have been disappointed in the rate of progress in cleanup, but the outlook for its defueling may be brightening. IEEE Spectrum, 21 (4):27–33, April 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1985:AMT

[Ada85a] J. A. Adam. Aerospace and military: Three satellites are salvaged in space, an army intercept missile destroys a mock ICBM warhead, and a new strategic bomber that’s harder to detect is unveiled. IEEE Spectrum, 22 (1):85–88, January 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1985:MTF

[Ada85b] J. A. Adam. More than fun and games. IEEE Spectrum, 22(5):86–89, May 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1985:PBS

[Ada85c] J. A. Adam. Probing beneath the sea: Sending vessels into environments too harsh for humans poses challenges in communications, artificial intelligence, and power-supply technology. IEEE Spectrum, 22 (4):55–64, April 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1986:CWM

[Ada86a] J. A. Adam. Counting the weapons: Many top secret eyes in the skies keep vigil over Soviet weaponry: Just how sharp are they? IEEE Spectrum, 23(7):46–56, July 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1986:SRV

[Ada86b] J. A. Adam. Special report/ verification: Peacekeeping by technical means. IEEE Spectrum, 23(7):42–45, July 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1987:AMR

[Ada87a] J. A. Adam. Aerospace and military: Reconfigured spacecraft explores Uranus. International armada captures a fleeting comet. Space defense research progresses. Smart computers assist U.S. Navy commanders. IEEE Spectrum, 24 (1):62–65, January 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1987:CSF

[Ada87b] J. A. Adam. Case study fighter aircraft: Small and agile. IEEE Spectrum, 24(5):64–66, May 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1987:DEP

[Ada87c] J. A. Adam. Detecting the enemy: Penetrating darkness and cloud cover, radar was the no. 1 method for pinpointing the adversary in the air or at sea. IEEE Spectrum, 24 (8):37–41, August 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1987:ELa

[Ada87d] J. A. Adam. The engineer at large. IEEE Spectrum, 24 (6):22–23, June 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1987:SYG

[Ada87e] J. A. Adam. The Sergeant York gun: a massive misfire: The U.S. Department of Defense spent $1.9 billion on developing the computer-controlled, radardirected air defense system before halting the program. IEEE Spectrum, 24(2):28–35, February 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1987:USW

[Ada87f] J. A. Adam. USS Stark: What really happened? Human error may have triggered the disaster, but questions have been raised about the frigate’s electrotechnical defenses. IEEE Spectrum, 24 (9):26–29, September 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1988:EPA

[Ada88a] J. A. Adam. Extracting power from the Amazon basin. IEEE Spectrum, 25(8):34–38, August 1988. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1988:HPB

[Ada88b] J. A. Adam. How the Pentagon buys. IEEE Spectrum, 25 (12):26–33, November 1988. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1988:HDI

[Ada88c] J. A. Adam. How to design an ‘invisible’ aircraft. IEEE Spectrum, 25(4):26–31, April 1988. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1988:PDH

[Ada88d] J. A. Adam. Pinning defense hopes on Aegis. IEEE Spectrum, 25(6):24–27, June 1988. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1988:SRM

[Ada88e] J. A. Adam. Special report military systems procurement — the price for might. IEEE Spectrum, 25(12):24–25, November 1988. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1988:TAM

[Ada88f] J. A. Adam. Technology ’88:

Aerospace and military. IEEE Spectrum, 25(1):61–64, January 1988. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1988:WVU

[Ada88g] J. A. Adam. Ways to verify the US–Soviet arms pact. IEEE Spectrum, 25(2):30–34, February 1988. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1989:HTD

[Ada89a] J. A. Adam. High-tech dig for a dinosaur. IEEE Spectrum, 26(4):48–49, April 1989. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1989:PWR

[Ada89b] J. A. Adam. Peaceful watchdogs (remote sensing). IEEE Spectrum, 26(11):28–33, November 1989. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1989:SWT

[Ada89c] J. A. Adam. Star wars in transition. IEEE Spectrum, 26 (3):32–38, March 1989. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1985:SWS

[AF85] J. A. Adam and M. A. Fischetti.

Star Wars SDI: The grand experiment. IEEE Spectrum, 22 (9):34–35, September 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Aliphas:1987:VDS

[AF87] A. Aliphas and J. A. Feldman. The versatility of digital signal processing chips: Uses for these architecturally modified microprocessors go all the way from simple digital filtering to echo canceling. IEEE Spectrum, 24 (6):40–45, June 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1987:ELb

[AFG87] J. A. Adam, A. H. Flax, and M. Golladay. The engineer at large. IEEE Spectrum, 24 (8):18–19, August 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1985:DIG

[AH85] J. A. Adam and J. Horgan. Debating the issues: The greatest challenge of all may be gaining political support for the SDI both in the U.S. and elsewhere. IEEE Spectrum, 22(9): 55–64, September 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Aitken:1987:BR

[Ait87] H. G. J. Aitken. Book reviews. IEEE Spectrum, 24 (8):12–15, August 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Aronoff:1985:DIA

[AJ85] S. Aronoff and G. F. Jones. From data to image to action: Manipulating digital images created from raw data proves to be a potent tool in human decision making. IEEE Spectrum, 22(12): 45–53, December 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1986:AMS

[AM86] J. A. Adam and E. W. Martin. Aerospace and military: Sandia builds a tough compact computer, the autonomous land vehicle takes a spin, and the Sergeant York gun is axed. IEEE Spectrum, 23(1):76–81, January 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Aupperle:1985:EOE

[AMG^{+}85] E. M. Aupperle, J. D. Musa, P. E. Green, B. T. Murphy, C. H. House, and J. J. Dougherty. The experts’ outlook: Eric M. Aupperle on computers: Networking is key to future growth. IEEE Spectrum, 22(1):36, January 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Amlie:1982:MRS

[Aml82] T. S. Amlie. Military: Radar:

Shield or target? Costly radar being developed by the U.S. may give an opponent a decisive advantage by broadcasting the position and makeup of forces. IEEE Spectrum, 19 (4):61–65, April 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Andre:1980:CEF

[And80] H. Andre. Cheap electricity from French tides: The sea has demonstrated both friendly and adverse effects during the 14year history of the Saint Malo plant. IEEE Spectrum, 17(2): 54–57, February 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Angelini:1980:PEP

[Ang80] A. M. Angelini. Power/energy: Power for the ’30s: a challenge for Western Europe: There is a way for Europe to make itself energy-independent. But is there a will? IEEE Spectrum, 17 (10):42–45, October 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Ap

[Ano80a] Anonymous. [advertisement]. IEEE Spectrum, 17(8):1–2, August 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Aq

[Ano80b] Anonymous. [advertisement]. IEEE Spectrum, 17(8):2–5, August 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Ar

[Ano80c] Anonymous. [advertisement].

IEEE Spectrum, 17(8):1–4, August 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:As

[Ano80d] Anonymous. [advertisement].

IEEE Spectrum, 17(9):1–2, September 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:At

[Ano80e] Anonymous. [advertisement].

IEEE Spectrum, 17(9):2–7, September 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Au

[Ano80f] Anonymous. [advertisement].

IEEE Spectrum, 17(9):1–4, September 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Av

[Ano80g] Anonymous. [advertisement].

IEEE Spectrum, 17(10):1–2, October 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Aw

[Ano80h] Anonymous. [advertisement].

IEEE Spectrum, 17(10):2–5, October 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Ax

[Ano80i] Anonymous. [advertisement]. IEEE Spectrum, 17(10):1–4, October 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Ay

[Ano80j] Anonymous. [advertisement]. IEEE Spectrum, 17(11):1–2, November 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Az

[Ano80k] Anonymous. [advertisement]. IEEE Spectrum, 17(11):2–7, November 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Aaa

[Ano80l] Anonymous. [advertisement]. IEEE Spectrum, 17(11):1–4, November 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Aab

[Ano80m] Anonymous. [advertisement]. IEEE Spectrum, 17(12):1–2, December 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Aac

[Ano80n] Anonymous. [advertisement]. IEEE Spectrum, 17(12):2–3, December 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Aad

[Ano80o] Anonymous. [advertisement].

IEEE Spectrum, 17(12):110– 112, December 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Aae

[Ano80p] Anonymous. [advertisement].

IEEE Spectrum, 17(12):1–4, December 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Aa

[Ano80q] Anonymous. [advertisements].

IEEE Spectrum, 17(1):1–4, January 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Ab

[Ano80r] Anonymous. [advertisements].

IEEE Spectrum, 17(2):2–7, February 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Ac

[Ano80s] Anonymous. [advertisements].

IEEE Spectrum, 17(2):1–4, February 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Ad

[Ano80t] Anonymous. [advertisements].

IEEE Spectrum, 17(3):2–7, March 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Ae

[Ano80u] Anonymous. [advertisements]. IEEE Spectrum, 17(3):1–3, March 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Af

[Ano80v] Anonymous. [advertisements]. IEEE Spectrum, 17(4):2–5, April 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Ag

[Ano80w] Anonymous. [advertisements]. IEEE Spectrum, 17(4):1–4, April 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Ah

[Ano80x] Anonymous. [advertisements]. IEEE Spectrum, 17(5):2–7, May 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Ai

[Ano80y] Anonymous. [advertisements]. IEEE Spectrum, 17(5):88–90, May 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Aj

[Ano80z] Anonymous. [advertisements]. IEEE Spectrum, 17(5):1–4, May 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Ak

[Ano80-27] Anonymous. [advertisements].

IEEE Spectrum, 17(6):2–5, June 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Am

[Ano80-28] Anonymous. [advertisements].

IEEE Spectrum, 17(6):1–4, June 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:An

[Ano80-29] Anonymous. [advertisements].

IEEE Spectrum, 17(7):2–5, July 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Ao

[Ano80-30] Anonymous. [advertisements].

IEEE Spectrum, 17(7):1–4, July 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:ALa

[Ano80-31] Anonymous. Applications literature. IEEE Spectrum, 17(1): 113–117, January 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:ALb

[Ano80-32] Anonymous. Applications literature. IEEE Spectrum, 17 (6):77–78, June 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:ALd

[Ano80-33] Anonymous. Applications literature. IEEE Spectrum, 17(8):73, August 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:ALc

[Ano80-34] Anonymous. Applications literatures. IEEE Spectrum, 17 (7):67–68, July 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:AIa

[Ano80-35] Anonymous. Author index. IEEE Spectrum, 17(12):99–109, December 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:AIF

[Ano80-36] Anonymous. Awards: IEEE Field, Service, and Prize Paper Award winners. IEEE Spectrum, 17(8):46–55, August 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:AIM

[Ano80-37] Anonymous. Awards: IEEE major medalists for 1980. IEEE Spectrum, 17(4):48–49, April 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:BBAa

[Ano80-38] Anonymous. Best bits: Applications of microprocessors. IEEE Spectrum, 17(1):24–25, January 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:BBAc

[Ano80-39] Anonymous. Best bits: Applications of microprocessors. IEEE Spectrum, 17(5):23–28, May 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:BBAd

[Ano80-40] Anonymous. Best bits: Applications of microprocessors: Aiding the disadvantaged. IEEE Spectrum, 17(7):16–20, July 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:BBAb

[Ano80-41] Anonymous. Best bits: Applications of microprocessors: Lowcost nuclear safety. IEEE Spectrum, 17(3):18–19, March 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:BBAe

[Ano80-42] Anonymous. Best bits: Applications of microprocessors: Switching to firmware. IEEE Spectrum, 17(9):24–25, September 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:BBAf

[Ano80-43] Anonymous. Best bits: Applications of microprocessors: The ‘four-cylinder caddy. IEEE Spectrum, 17(11):22, November 1980.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:BRa

[Ano80-44] Anonymous. Book reviews. IEEE Spectrum, 17(1):102– 103, January 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:BRb

[Ano80-45] Anonymous. Book reviews. IEEE Spectrum, 17(2):66–71, February 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:BRc

[Ano80-46] Anonymous. Book reviews. IEEE Spectrum, 17(3):78–79, March 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:BRd

[Ano80-47] Anonymous. Book reviews. IEEE Spectrum, 17(4):56–57, April 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:BRe

[Ano80-48] Anonymous. Book reviews. IEEE Spectrum, 17(5):79–83, May 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:BRf

[Ano80-49] Anonymous. Book reviews. IEEE Spectrum, 17(6):64–66, June 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:BRg

[Ano80-50] Anonymous. Book reviews.

IEEE Spectrum, 17(7):60–63, July 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:BRh

[Ano80-51] Anonymous. Book reviews.

IEEE Spectrum, 17(8):64–65, August 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:BRi

[Ano80-52] Anonymous. Book reviews.

IEEE Spectrum, 17(9):108–111, September 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:BRj

[Ano80-53] Anonymous. Book reviews.

IEEE Spectrum, 17(10):66, October 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:BRk

[Ano80-54] Anonymous. Book reviews.

IEEE Spectrum, 17(11):63–65, November 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:BRl

[Ano80-55] Anonymous. Book reviews.

IEEE Spectrum, 17(12):73–79, December 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Cb

[Ano80-56] Anonymous. Calendar. IEEE Spectrum, 17(1):14–15, January 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Cd

[Ano80-57] Anonymous. Calendar. IEEE Spectrum, 17(2):8–13, February 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Cf

[Ano80-58] Anonymous. Calendar. IEEE Spectrum, 17(3):8–9, March 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Ch

[Ano80-59] Anonymous. Calendar. IEEE Spectrum, 17(4):6–7, April 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Cj

[Ano80-60] Anonymous. Calendar. IEEE Spectrum, 17(5):8–11, May 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Cl

[Ano80-61] Anonymous. Calendar. IEEE Spectrum, 17(6):6–9, June 1980.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Cn

[Ano80-62] Anonymous. Calendar. IEEE Spectrum, 17(7):6–7, July 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Cp

[Ano80-63] Anonymous. Calendar. IEEE Spectrum, 17(8):6–7, August 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Cr

[Ano80-64] Anonymous. Calendar. IEEE Spectrum, 17(9):8–9, September 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Cu

[Ano80-65] Anonymous. Calendar. IEEE Spectrum, 17(10):6–7, October 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Cw

[Ano80-66] Anonymous. Calendar. IEEE Spectrum, 17(11):8–11, November 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Cy

[Ano80-67] Anonymous. Calendar. IEEE Spectrum, 17(12):4–5, December 1980. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:CWL

[Ano80-68] Anonymous. Careers: We look at ourselves: Young EEs on their way up: From round-table discussions, a profile emerges of successful engineers and how they view their work. IEEE Spectrum, 17(7):40–43, July 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:CAa

[Ano80-69] Anonymous. Classified advertising. IEEE Spectrum, 17(1): 118–127, January 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:CAb

[Ano80-70] Anonymous. Classified advertising. IEEE Spectrum, 17(2): 88–95, February 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:CAc

[Ano80-71] Anonymous. Classified advertising. IEEE Spectrum, 17 (3):97–106, March 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:CAd

[Ano80-72] Anonymous. Classified advertising. IEEE Spectrum, 17 (4):78–86, April 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:CAe

[Ano80-73] Anonymous. Classified advertising. IEEE Spectrum, 17 (5):84–87, May 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:CAf

[Ano80-74] Anonymous. Classified advertising. IEEE Spectrum, 17 (6):79–84, June 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:CAg

[Ano80-75] Anonymous. Classified advertising. IEEE Spectrum, 17 (7):69–74, July 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:CAh

[Ano80-76] Anonymous. Classified advertising. IEEE Spectrum, 17 (8):74–77, August 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:CAi

[Ano80-77] Anonymous. Classified advertising. IEEE Spectrum, 17(9): 123–128, September 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:CSa

[Ano80-78] Anonymous. Coming in Spectrum. IEEE Spectrum, 17(1): 26–28, January 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:CSb

[Ano80-79] Anonymous. Coming in Spectrum. IEEE Spectrum, 17(2): 22–24, February 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:CSc

[Ano80-80] Anonymous. Coming in Spectrum. IEEE Spectrum, 17 (3):20–22, March 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:CSd

[Ano80-81] Anonymous. Coming in Spectrum. IEEE Spectrum, 17 (4):14–16, April 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:CSe

[Ano80-82] Anonymous. Coming in Spectrum. IEEE Spectrum, 17 (5):23–25, May 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:CSf

[Ano80-83] Anonymous. Coming in Spectrum. IEEE Spectrum, 17 (6):18–20, June 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:CSg

[Ano80-84] Anonymous. Coming in Spectrum. IEEE Spectrum, 17 (7):14–15, July 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:CSh

[Ano80-85] Anonymous. Coming in Spectrum. IEEE Spectrum, 17 (8):14–15, August 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:CSi

[Ano80-86] Anonymous. Coming in Spectrum. IEEE Spectrum, 17(9): 22–23, September 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:CSj

[Ano80-87] Anonymous. Coming in Spectrum. IEEE Spectrum, 17(10): 22–23, October 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:CSk

[Ano80-88] Anonymous. Coming in Spectrum. IEEE Spectrum, 17(11): 18–21, November 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:CSl

[Ano80-89] Anonymous. Coming in Spectrum. IEEE Spectrum, 17 (12):15, December 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Ca

[Ano80-90] Anonymous. Contents. IEEE Spectrum, 17(1):1, January 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Cc

[Ano80-91] Anonymous. Contents. IEEE Spectrum, 17(2):1–2, February 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Ce

[Ano80-92] Anonymous. Contents. IEEE Spectrum, 17(3):1, March 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Cg

[Ano80-93] Anonymous. Contents. IEEE Spectrum, 17(4):1, April 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Ci

[Ano80-94] Anonymous. Contents. IEEE Spectrum, 17(5):1, May 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Ck

[Ano80-95] Anonymous. Contents. IEEE Spectrum, 17(6):1, June 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Cm

[Ano80-96] Anonymous. Contents. IEEE Spectrum, 17(7):1, July 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Co

[Ano80-97] Anonymous. Contents. IEEE Spectrum, 17(8):1, August 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Cq

[Ano80-98] Anonymous. Contents. IEEE Spectrum, 17(9):1, September 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Ct

[Ano80-99] Anonymous. Contents. IEEE Spectrum, 17(10):1, October 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Cv

[Ano80-100] Anonymous. Contents. IEEE Spectrum, 17(11):1, November 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Cx

[Ano80-101] Anonymous. Contents. IEEE Spectrum, 17(12):1, December 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:ETTa

[Ano80-102] Anonymous. EEs’ tools toys.

IEEE Spectrum, 17(2):21, February 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:ETTb

[Ano80-103] Anonymous. EEs’ tools toys. IEEE Spectrum, 17(4):10–11, April 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:ETTc

[Ano80-104] Anonymous. EEs’ tools toys. IEEE Spectrum, 17(6):14–15, June 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:ETTd

[Ano80-105] Anonymous. EEs’ tools toys. IEEE Spectrum, 17(8):10–11, August 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:ETTe

[Ano80-106] Anonymous. EEs’ tools toys. IEEE Spectrum, 17(10):18– 19, October 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:ETTf

[Ano80-107] Anonymous. EEs’ tools toys. IEEE Spectrum, 17(12):10–11, December 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:ERa

[Ano80-108] Anonymous. Energy report. IEEE Spectrum, 17(1):20, January 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:ERb

[Ano80-109] Anonymous. Energy report.

IEEE Spectrum, 17(2):20, February 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:ERc

[Ano80-110] Anonymous. Energy report.

IEEE Spectrum, 17(3):16–17, March 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:ERd

[Ano80-111] Anonymous. Energy report.

IEEE Spectrum, 17(4):12–13, April 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:ERe

[Ano80-112] Anonymous. Energy report.

IEEE Spectrum, 17(5):16–17, May 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:ERf

[Ano80-113] Anonymous. Energy report.

IEEE Spectrum, 17(6):16–17, June 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:ERg

[Ano80-114] Anonymous. Energy report.

IEEE Spectrum, 17(7):8–9, July 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:ERh

[Ano80-115] Anonymous. Energy report. IEEE Spectrum, 17(8):12–13, August 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:ERi

[Ano80-116] Anonymous. Energy report. IEEE Spectrum, 17(9):18–19, September 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:ERj

[Ano80-117] Anonymous. Energy report. IEEE Spectrum, 17(10):20– 21, October 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:ERk

[Ano80-118] Anonymous. Energy report. IEEE Spectrum, 17(11): 16, November 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:ERl

[Ano80-119] Anonymous. Energy report. IEEE Spectrum, 17(12):12–14, December 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:ELa

[Ano80-120] Anonymous. The engineer at large. IEEE Spectrum, 17(1): 21–23, January 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:ELb

[Ano80-121] Anonymous. The engineer at large. IEEE Spectrum, 17 (3):14–15, March 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:ELc

[Ano80-122] Anonymous. The engineer at large. IEEE Spectrum, 17 (5):20–22, May 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:ELd

[Ano80-123] Anonymous. The engineer at large. IEEE Spectrum, 17 (7):12–13, July 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:ELe

[Ano80-124] Anonymous. The engineer at large. IEEE Spectrum, 17(9): 20–21, September 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:ELf

[Ano80-125] Anonymous. The engineer at large. IEEE Spectrum, 17 (11):17, November 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:FR

[Ano80-126] Anonymous. For further reading. IEEE Spectrum, 17(9): 92–93, September 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Fa

[Ano80-127] Anonymous. Forum. IEEE Spectrum, 17(1):16–18, January 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Fb

[Ano80-128] Anonymous. Forum. IEEE Spectrum, 17(2):14–18, February 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Fc

[Ano80-129] Anonymous. Forum. IEEE Spectrum, 17(3):10–13, March 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Fd

[Ano80-130] Anonymous. Forum. IEEE Spectrum, 17(4):8–9, April 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Fe

[Ano80-131] Anonymous. Forum. IEEE Spectrum, 17(5):18–19, May 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Ff

[Ano80-132] Anonymous. Forum. IEEE Spectrum, 17(6):10–13, June 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Fg

[Ano80-133] Anonymous. Forum. IEEE Spectrum, 17(7):10–11, July 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Fh

[Ano80-134] Anonymous. Forum. IEEE Spectrum, 17(8):8–9, August 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Fi

[Ano80-135] Anonymous. Forum. IEEE Spectrum, 17(9):14–17, September 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Fj

[Ano80-136] Anonymous. Forum. IEEE Spectrum, 17(10):8–17, October 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Fk

[Ano80-137] Anonymous. Forum. IEEE Spectrum, 17(11):12–15, November 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Fl

[Ano80-138] Anonymous. Forum. IEEE Spectrum, 17(12):6–9, December 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:FMa

[Ano80-139] Anonymous. [front matter]. IEEE Spectrum, 17(1):1–2, January 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:FMb

[Ano80-140] Anonymous. [front matter]. IEEE Spectrum, 17(1):2–13, January 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:FMc

[Ano80-141] Anonymous. [front matter]. IEEE Spectrum, 17(2): 1–2, February 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:FMd

[Ano80-142] Anonymous. [front matter]. IEEE Spectrum, 17(3):1–2, March 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:FMe

[Ano80-143] Anonymous. [front matter]. IEEE Spectrum, 17(4):1–2, April 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:FMf

[Ano80-144] Anonymous. [front matter]. IEEE Spectrum, 17(5):1–2, May 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:FMg

[Ano80-145] Anonymous. [front matter].

IEEE Spectrum, 17(6):1–2, June 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:FMh

[Ano80-146] Anonymous. [front matter].

IEEE Spectrum, 17(7):1–2, July 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:GA

[Ano80-147] Anonymous. Guide for authors. IEEE Spectrum, 17(1): 91–93, January 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:IPP

[Ano80-148] Anonymous. IEEE publications paper calls. IEEE Spectrum, 17 (10):72–73, October 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:ITCh

[Ano80-149] Anonymous. IEEE tables of contents. IEEE Spectrum, 17(9): 96–107, September 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:ITCi

[Ano80-150] Anonymous. IEEE tables of contents. IEEE Spectrum, 17 (10):63–65, October 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:ITCj

[Ano80-151] Anonymous. IEEE tables of contents. IEEE Spectrum, 17(11): 59–62, November 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:ITCk

[Ano80-152] Anonymous. IEEE tables of contents. IEEE Spectrum, 17(12): 70–72, December 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:ITCa

[Ano80-153] Anonymous. IEEE tables of contents for current and future publications. IEEE Spectrum, 17 (1):96–101, January 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:ITCb

[Ano80-154] Anonymous. IEEE tables of contents for current and future publications. IEEE Spectrum, 17 (3):74–77, March 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:ITCc

[Ano80-155] Anonymous. IEEE tables of contents for current and future publications. IEEE Spectrum, 17(4):52–55, April 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:ITCd

[Ano80-156] Anonymous. IEEE tables of contents for current and future publications. IEEE Spectrum, 17(5):67, May 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:ITCe

[Ano80-157] Anonymous. IEEE tables of contents for current and future publications. IEEE Spectrum, 17(6):54–63, June 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:ITCf

[Ano80-158] Anonymous. IEEE tables of contents for current and future publications. IEEE Spectrum, 17(7):50–55, July 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:ITCg

[Ano80-159] Anonymous. IEEE tables of contents for current and future publications. IEEE Spectrum, 17 (8):58–63, August 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:IAa

[Ano80-160] Anonymous. Index to advertisers. IEEE Spectrum, 17 (1):128, January 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:IAb

[Ano80-161] Anonymous. Index to advertisers. IEEE Spectrum, 17 (2):96, February 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:MPa

[Ano80-162] Anonymous. Meeting previews. IEEE Spectrum, 17(3):88–96, March 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:MPb

[Ano80-163] Anonymous. Meeting previews. IEEE Spectrum, 17(4):58–63, April 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:MPc

[Ano80-164] Anonymous. Meeting previews. IEEE Spectrum, 17(5):68–71, May 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:MPd

[Ano80-165] Anonymous. Meeting previews. IEEE Spectrum, 17(6):67–73, June 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:MPe

[Ano80-166] Anonymous. Meeting previews. IEEE Spectrum, 17(7):56–59, July 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:MPf

[Ano80-167] Anonymous. Meeting previews. IEEE Spectrum, 17(8):66–71, August 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:MPg

[Ano80-168] Anonymous. Meeting previews. IEEE Spectrum, 17(9): 112–113, September 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:MPh

[Ano80-169] Anonymous. Meeting previews. IEEE Spectrum, 17(10): 74–79, October 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:MPi

[Ano80-170] Anonymous. Meeting previews. IEEE Spectrum, 17(11): 66–71, November 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:MPj

[Ano80-171] Anonymous. Meeting previews. IEEE Spectrum, 17(12): 86–97, December 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Ma

[Ano80-172] Anonymous. Meetings. IEEE Spectrum, 17(1):104–108, January 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:Mb

[Ano80-173] Anonymous. Meetings. IEEE Spectrum, 17(2):72–78, February 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:NRIa

[Ano80-174] Anonymous. New and recent IEEE publications. IEEE Spectrum, 17(1):94–95, January 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:NRIb

[Ano80-175] Anonymous. New and recent IEEE publications. IEEE Spectrum, 17(2):64–65, February 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:NRIc

[Ano80-176] Anonymous. New and recent IEEE publications. IEEE Spectrum, 17(3):72–73, March 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:NRId

[Ano80-177] Anonymous. New and recent IEEE publications. IEEE Spectrum, 17(4):50–51, April 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:NRIe

[Ano80-178] Anonymous. New and recent IEEE publications. IEEE Spectrum, 17(5):58–59, May 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:NRIf

[Ano80-179] Anonymous. New and recent IEEE publications. IEEE Spectrum, 17(6):52–53, June 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:NRIg

[Ano80-180] Anonymous. New and recent IEEE publications. IEEE Spectrum, 17(7):48–49, July 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:NRIh

[Ano80-181] Anonymous. New and recent IEEE publications. IEEE Spectrum, 17(8):56–57, August 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:NRIi

[Ano80-182] Anonymous. New and recent IEEE publications. IEEE Spectrum, 17(9):94–95, September 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:NRIj

[Ano80-183] Anonymous. New and recent IEEE publications. IEEE Spectrum, 17(10):60–62, October 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:NRIk

[Ano80-184] Anonymous. New and recent IEEE publications. IEEE Spectrum, 17(11):57–58, November 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:NRIl

[Ano80-185] Anonymous. New and recent IEEE publications. IEEE Spectrum, 17(12):69, December 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:NPAa

[Ano80-186] Anonymous. New product applications. IEEE Spectrum, 17 (1):111–112, January 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:NPAb

[Ano80-187] Anonymous. New product applications. IEEE Spectrum, 17 (2):81–87, February 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:NPAd

[Ano80-188] Anonymous. New product applications. IEEE Spectrum, 17 (4):64–64BB, April 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:NPAe

[Ano80-189] Anonymous. New product applications. IEEE Spectrum, 17 (5):72–72T, May 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:NPAg

[Ano80-190] Anonymous. New product applications. IEEE Spectrum, 17 (7):64–64P, July 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:NPAh

[Ano80-191] Anonymous. New product applications. IEEE Spectrum, 17 (8):72–77P, August 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:NPAj

[Ano80-192] Anonymous. New product applications: Digital signal generator has 10 ns resolution up to 100 MHz, and 16 channel operations. IEEE Spectrum, 17 (10):80–86, October 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:NPAf

[Ano80-193] Anonymous. New product applications: Digital storage oscilloscope has 10 MHz storage bandwidth and 8-b vertical resolution. IEEE Spectrum, 17(6):74–74T, June 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:NPAk

[Ano80-194] Anonymous. New product applications: IBM-compatible processor provides up to 16 channels and 16 megabytes of main memory. IEEE Spectrum, 17 (11):72–83, November 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:NPAl

[Ano80-195] Anonymous. New product applications: Measurement and control link includes 16-b processor and 28-kilobyte user-partitioned RAM. IEEE Spectrum, 17(12): 80–85, December 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:NPAc

[Ano80-196] Anonymous. New product applications: S33/A storage-module drive controller is software compatible with DEC RM02 I/O driver. IEEE Spectrum, 17(3):80, March 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:NPAi

[Ano80-197] Anonymous. New product applications: Universal development system provides real-time emulation from 4 to 32 b at up to 30 MHz. IEEE Spectrum, 17(9): 118–122, September 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:NIa

[Ano80-198] Anonymous. News from industry. IEEE Spectrum, 17 (1):109, January 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:NIb

[Ano80-199] Anonymous. News from industry. IEEE Spectrum, 17 (2):79, February 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:NIc

[Ano80-200] Anonymous. News from industry. IEEE Spectrum, 17(3):81, March 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:NId

[Ano80-201] Anonymous. News from industry. IEEE Spectrum, 17(4):65, April 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:NIe

[Ano80-202] Anonymous. News from industry. IEEE Spectrum, 17(5):73, May 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:NIf

[Ano80-203] Anonymous. News from industry. IEEE Spectrum, 17(6):75, June 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:NIg

[Ano80-204] Anonymous. News from industry. IEEE Spectrum, 17(7):65, July 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:NIh

[Ano80-205] Anonymous. News from industry. IEEE Spectrum, 17 (8):78–80, August 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:NIi

[Ano80-206] Anonymous. News from industry. IEEE Spectrum, 17(9): 114–117, September 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:NIj

[Ano80-207] Anonymous. News from industry. IEEE Spectrum, 17 (10):86–88, October 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:NIk

[Ano80-208] Anonymous. News from industry. IEEE Spectrum, 17(11): 84–86, November 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:NIl

[Ano80-209] Anonymous. News from industry. IEEE Spectrum, 17 (12):98, December 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:NRa

[Ano80-210] Anonymous. News from region 1. IEEE Spectrum, 17(1):110, January 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:NRb

[Ano80-211] Anonymous. News from region 1.

IEEE Spectrum, 17(2):80, February 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:NRc

[Ano80-212] Anonymous. News from region 1. IEEE Spectrum, 17(3):82–87, March 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:NRd

[Ano80-213] Anonymous. News from region 1. IEEE Spectrum, 17(4):66–77, April 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:NRe

[Ano80-214] Anonymous. News from region 1. IEEE Spectrum, 17(5):74–78, May 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:NRf

[Ano80-215] Anonymous. News from region 1. IEEE Spectrum, 17(6):76, June 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:NRg

[Ano80-216] Anonymous. News from region 1. IEEE Spectrum, 17(7):66, July 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:NWa

[Ano80-217] Anonymous. News from Washington. IEEE Spectrum, 17 (1):19, January 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:NWb

[Ano80-218] Anonymous. News from Washington. IEEE Spectrum, 17 (2):19, February 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:NWc

[Ano80-219] Anonymous. News from Washington. IEEE Spectrum, 17 (3):16, March 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:NWd

[Ano80-220] Anonymous. News from Washington. IEEE Spectrum, 17 (4):12, April 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:NWe

[Ano80-221] Anonymous. News from Washington. IEEE Spectrum, 17 (5):16, May 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:NWf

[Ano80-222] Anonymous. News from Washington. IEEE Spectrum, 17 (6):16, June 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:NWg

[Ano80-223] Anonymous. News from Washington. IEEE Spectrum, 17(7): 8, July 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:NWh

[Ano80-224] Anonymous. News from Washington. IEEE Spectrum, 17 (8):12, August 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:NWi

[Ano80-225] Anonymous. News from Washington. IEEE Spectrum, 17 (9):18, September 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:NWj

[Ano80-226] Anonymous. News from Washington. IEEE Spectrum, 17 (10):20, October 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:NWk

[Ano80-227] Anonymous. News from Washington. IEEE Spectrum, 17 (11):16, November 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:NWl

[Ano80-228] Anonymous. News from Washington. IEEE Spectrum, 17 (12):12, December 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:OMI

[Ano80-229] Anonymous. Other meetings of interest. IEEE Spectrum, 17(5):12–15, May 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:OMS

[Ano80-230] Anonymous. Other meetings: Series SE3000 lightweight, instrumentation recorders. IEEE Spectrum, 17(9):10–13, September 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:SIa

[Ano80-231] Anonymous. Scanning the Institute. IEEE Spectrum, 17 (1):26, January 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:SIb

[Ano80-232] Anonymous. Scanning the Institute. IEEE Spectrum, 17 (2):22, February 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:SIc

[Ano80-233] Anonymous. Scanning the Institute. IEEE Spectrum, 17(3):20, March 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:SId

[Ano80-234] Anonymous. Scanning the Institute. IEEE Spectrum, 17(4):14, April 1980. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:SIe

[Ano80-235] Anonymous. Scanning the Institute. IEEE Spectrum, 17(5):23, May 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:SIf

[Ano80-236] Anonymous. Scanning the Institute. IEEE Spectrum, 17(6):18, June 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:SIg

[Ano80-237] Anonymous. Scanning the Institute. IEEE Spectrum, 17(7):14, July 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:SIh

[Ano80-238] Anonymous. Scanning the Institute. IEEE Spectrum, 17(8):14, August 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:SIi

[Ano80-239] Anonymous. Scanning the Institute. IEEE Spectrum, 17 (9):22, September 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:SIj

[Ano80-240] Anonymous. Scanning the Institute. IEEE Spectrum, 17 (10):22, October 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:SIk

[Ano80-241] Anonymous. Scanning the Institute. IEEE Spectrum, 17 (11):18, November 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:SIl

[Ano80-242] Anonymous. Scanning the Institute. IEEE Spectrum, 17 (12):15, December 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:SES

[Ano80-243] Anonymous. Sharper ‘eyes’ for star gazers [a new submillimeterwavelength radio receiver]. IEEE Spectrum, 17(8):26–29, August 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:TSa

[Ano80-244] Anonymous. Technically speaking. IEEE Spectrum, 17(9): 26–28, September 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:TPS

[Ano80-245] Anonymous. Technology and policy: Spectrum asks questions of Carter, Reagan, and Anderson: This year’s contenders for the U.S. Presidency give their views on technology and engineering issues. IEEE Spectrum, 17(10):53–59, October 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:TSb

[Ano80-246] Anonymous. Technology speaking. IEEE Spectrum, 17(11): 23–24, November 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:WEHa

[Ano80-247] Anonymous. What ever happened to … ? IEEE Spectrum, 17(8):16–18, August 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:WEHb

[Ano80-248] Anonymous. What ever happened to … ? IEEE Spectrum, 17(10):24–25, October 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1980:WEHc

[Ano80-249] Anonymous. What ever happened to … ? [ovonic devices]. IEEE Spectrum, 17(12): 16–18, December 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Aa

[Ano81a] Anonymous. [advertisement].

IEEE Spectrum, 18(1):1–3, January 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Ab

[Ano81b] Anonymous. [advertisement].

IEEE Spectrum, 18(1):2–7, January 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Ac

[Ano81c] Anonymous. [advertisement]. IEEE Spectrum, 18(1):1–4, January 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Ad

[Ano81d] Anonymous. [advertisement]. IEEE Spectrum, 18(2):1–2, February 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Ae

[Ano81e] Anonymous. [advertisement]. IEEE Spectrum, 18(2):2–7, February 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Af

[Ano81f] Anonymous. [advertisement]. IEEE Spectrum, 18(2):1–4, February 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Ag

[Ano81g] Anonymous. [advertisement]. IEEE Spectrum, 18(3):1–2, March 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Ah

[Ano81h] Anonymous. [advertisement]. IEEE Spectrum, 18(3):2–5, March 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Ai

[Ano81i] Anonymous. [advertisement].

IEEE Spectrum, 18(3):1–4, March 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Aj

[Ano81j] Anonymous. [advertisement].

IEEE Spectrum, 18(4):1–2, April 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Ak

[Ano81k] Anonymous. [advertisement].

IEEE Spectrum, 18(4):2–3, April 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Al

[Ano81l] Anonymous. [advertisement].

IEEE Spectrum, 18(4):1–5, April 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Am

[Ano81m] Anonymous. [advertisement].

IEEE Spectrum, 18(5):1, May 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:An

[Ano81n] Anonymous. [advertisement].

IEEE Spectrum, 18(5):2–3, May 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Ao

[Ano81o] Anonymous. [advertisement]. IEEE Spectrum, 18(5):1–4, May 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Ap

[Ano81p] Anonymous. [advertisement]. IEEE Spectrum, 18(6):1–2, June 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Aq

[Ano81q] Anonymous. [advertisement]. IEEE Spectrum, 18(6):2–5, June 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Ar

[Ano81r] Anonymous. [advertisement]. IEEE Spectrum, 18(6):1–4, June 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:As

[Ano81s] Anonymous. [advertisement]. IEEE Spectrum, 18(7):1–2, July 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:At

[Ano81t] Anonymous. [advertisement]. IEEE Spectrum, 18(7):2–5, July 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Au

[Ano81u] Anonymous. [advertisement].

IEEE Spectrum, 18(7):1–3, July 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Av

[Ano81v] Anonymous. [advertisement].

IEEE Spectrum, 18(8):1–2, August 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Aw

[Ano81w] Anonymous. [advertisement].

IEEE Spectrum, 18(8):2–5, August 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Ax

[Ano81x] Anonymous. [advertisement].

IEEE Spectrum, 18(8):1–3, August 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Ay

[Ano81y] Anonymous. [advertisement].

IEEE Spectrum, 18(9):1–2, September 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Az

[Ano81z] Anonymous. [advertisement].

IEEE Spectrum, 18(9):2–3, September 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Aaa

[Ano81-27] Anonymous. [advertisement]. IEEE Spectrum, 18(9):1–4, September 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Aab

[Ano81-28] Anonymous. [advertisement]. IEEE Spectrum, 18(10):1–2, October 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Aac

[Ano81-29] Anonymous. [advertisement]. IEEE Spectrum, 18(10):2–3, October 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Aae

[Ano81-30] Anonymous. [advertisement]. IEEE Spectrum, 18(10):1–4, October 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Aaf

[Ano81-31] Anonymous. [advertisement]. IEEE Spectrum, 18(11):1–2, November 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Aag

[Ano81-32] Anonymous. [advertisement]. IEEE Spectrum, 18(11):2–3, November 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Aah

[Ano81-33] Anonymous. [advertisement].

IEEE Spectrum, 18(11):1–4, November 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Aai

[Ano81-34] Anonymous. [advertisement].

IEEE Spectrum, 18(12):1–2, December 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Aaj

[Ano81-35] Anonymous. [advertisement].

IEEE Spectrum, 18(12):2–3, December 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Aak

[Ano81-36] Anonymous. [advertisement].

IEEE Spectrum, 18(12):1–4, December 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Aad

[Ano81-37] Anonymous. The authors. IEEE Spectrum, 18(10):104–107, October 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:AIF

[Ano81-38] Anonymous. Awards: IEEE Field, Service, and Prize Paper Award winners of 1981. IEEE Spectrum, 18(8):54–55, August 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:AIM

[Ano81-39] Anonymous. Awards: IEEE major medalists for 1981. IEEE Spectrum, 18(4):70–71, April 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:BBAc

[Ano81-40] Anonymous. Best bits: Applications of microprocessors. IEEE Spectrum, 18(5):16–17, May 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:BBAf

[Ano81-41] Anonymous. Best bits: Applications of microprocessors. IEEE Spectrum, 18(11):18–21, November 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:BBAe

[Ano81-42] Anonymous. Best bits: Applications of microprocessors: Chip makes drip coffee. IEEE Spectrum, 18(9):19, September 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:BBAb

[Ano81-43] Anonymous. Best bits: Applications of microprocessors: Mechanics on a chip. IEEE Spectrum, 18(3):24–26, March 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:BBAa

[Ano81-44] Anonymous. Best bits: Applications of microprocessors: ‘smart cash cards. IEEE Spectrum, 18 (1):30–31, January 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:BBAd

[Ano81-45] Anonymous. Best bits: Applications of microprocessors: Smart small-craft radar. IEEE Spectrum, 18(7):17, July 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:BRa

[Ano81-46] Anonymous. Book reviews.

IEEE Spectrum, 18(1):101– 105, January 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:BRb

[Ano81-47] Anonymous. Book reviews.

IEEE Spectrum, 18(2):64–71, February 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:BRc

[Ano81-48] Anonymous. Book reviews.

IEEE Spectrum, 18(3):92–103, March 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:BRd

[Ano81-49] Anonymous. Book reviews. IEEE Spectrum, 18(4):88–91, April 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:BRe

[Ano81-50] Anonymous. Book reviews. IEEE Spectrum, 18(5):71–77, May 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:BRf

[Ano81-51] Anonymous. Book reviews. IEEE Spectrum, 18(6):75–85, June 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:BRg

[Ano81-52] Anonymous. Book reviews. IEEE Spectrum, 18(7):70–79, July 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:BRh

[Ano81-53] Anonymous. Book reviews. IEEE Spectrum, 18(8):73–79, August 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:BRi

[Ano81-54] Anonymous. Book reviews. IEEE Spectrum, 18(9):80–81, September 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:BRj

[Ano81-55] Anonymous. Book reviews.

IEEE Spectrum, 18(10):126– 127, October 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:BRk

[Ano81-56] Anonymous. Book reviews.

IEEE Spectrum, 18(11):81–82, November 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:BRl

[Ano81-57] Anonymous. Book reviews.

IEEE Spectrum, 18(12):82–85, December 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:BS

[Ano81-58] Anonymous. Book-shelf. IEEE Spectrum, 18(10):102–103, October 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Cb

[Ano81-59] Anonymous. Calendar. IEEE Spectrum, 18(1):8–13, January 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Cd

[Ano81-60] Anonymous. Calendar. IEEE Spectrum, 18(2):8–10, February 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Cf

[Ano81-61] Anonymous. Calendar. IEEE Spectrum, 18(3):6–15, March 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Ch

[Ano81-62] Anonymous. Calendar. IEEE Spectrum, 18(4):4–9, April 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Cj

[Ano81-63] Anonymous. Calendar. IEEE Spectrum, 18(5):4–7, May 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Cl

[Ano81-64] Anonymous. Calendar. IEEE Spectrum, 18(6):6–13, June 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Cn

[Ano81-65] Anonymous. Calendar. IEEE Spectrum, 18(7):6–13, July 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Cp

[Ano81-66] Anonymous. Calendar. IEEE Spectrum, 18(8):6–7, August 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Cr

[Ano81-67] Anonymous. Calendar. IEEE Spectrum, 18(9):4–11, September 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Cu

[Ano81-68] Anonymous. Calendar. IEEE Spectrum, 18(10):4–9, October 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Cw

[Ano81-69] Anonymous. Calendar. IEEE Spectrum, 18(11):4–9, November 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Cy

[Ano81-70] Anonymous. Calendar. IEEE Spectrum, 18(12):4–9, December 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:CSa

[Ano81-71] Anonymous. Coming in Spectrum. IEEE Spectrum, 18(1): 28–29, January 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:CSb

[Ano81-72] Anonymous. Coming in Spectrum. IEEE Spectrum, 18(2): 24–25, February 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:CSc

[Ano81-73] Anonymous. Coming in Spectrum. IEEE Spectrum, 18 (3):22–23, March 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:CSd

[Ano81-74] Anonymous. Coming in Spectrum. IEEE Spectrum, 18 (4):16–19, April 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:CSe

[Ano81-75] Anonymous. Coming in Spectrum. IEEE Spectrum, 18 (5):20–21, May 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:CSf

[Ano81-76] Anonymous. Coming in Spectrum. IEEE Spectrum, 18 (6):18–19, June 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:CSg

[Ano81-77] Anonymous. Coming in Spectrum. IEEE Spectrum, 18 (7):20–22, July 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:CSh

[Ano81-78] Anonymous. Coming in Spectrum. IEEE Spectrum, 18(8):16, August 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:CSi

[Ano81-79] Anonymous. Coming in Spectrum. IEEE Spectrum, 18(9): 22–24, September 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:CSj

[Ano81-80] Anonymous. Coming in Spectrum. IEEE Spectrum, 18 (10):24, October 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:CSk

[Ano81-81] Anonymous. Coming in Spectrum. IEEE Spectrum, 18(11): 24–26, November 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:CSl

[Ano81-82] Anonymous. Coming in Spectrum. IEEE Spectrum, 18(12): 14–15, December 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Ca

[Ano81-83] Anonymous. Contents. IEEE Spectrum, 18(1):1, January 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Cc

[Ano81-84] Anonymous. Contents. IEEE Spectrum, 18(2):1, February 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Ce

[Ano81-85] Anonymous. Contents. IEEE Spectrum, 18(3):1, March 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Cg

[Ano81-86] Anonymous. Contents. IEEE Spectrum, 18(4):1, April 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Ci

[Ano81-87] Anonymous. Contents. IEEE Spectrum, 18(5):1, May 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Ck

[Ano81-88] Anonymous. Contents. IEEE Spectrum, 18(6):1, June 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Cm

[Ano81-89] Anonymous. Contents. IEEE Spectrum, 18(7):1, July 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Co

[Ano81-90] Anonymous. Contents. IEEE Spectrum, 18(8):1, August 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Cq

[Ano81-91] Anonymous. Contents. IEEE Spectrum, 18(9):1, September 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Ct

[Ano81-92] Anonymous. Contents. IEEE Spectrum, 18(10):1, October 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Cv

[Ano81-93] Anonymous. Contents. IEEE Spectrum, 18(11):1, November 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Cx

[Ano81-94] Anonymous. Contents. IEEE Spectrum, 18(12):1, December 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:ETTa

[Ano81-95] Anonymous. EEs’ tools toys.

IEEE Spectrum, 18(2):20–21, February 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:ETTb

[Ano81-96] Anonymous. EEs’ tools toys.

IEEE Spectrum, 18(4):14–15, April 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:ETTc

[Ano81-97] Anonymous. EEs’ tools toys. IEEE Spectrum, 18(6):20–21, June 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:ETTd

[Ano81-98] Anonymous. EEs’ tools toys. IEEE Spectrum, 18(8):17, August 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:ETTe

[Ano81-99] Anonymous. EEs’ tools toys. IEEE Spectrum, 18(10):25, October 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:ETTf

[Ano81-100] Anonymous. EEs’ tools toys. IEEE Spectrum, 18(12):16, December 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:ERa

[Ano81-101] Anonymous. Energy report. IEEE Spectrum, 18(1): 24–25, January 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:ERb

[Ano81-102] Anonymous. Energy report. IEEE Spectrum, 18(2):22–23, February 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:ERc

[Ano81-103] Anonymous. Energy report.

IEEE Spectrum, 18(3):20–21, March 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:ERd

[Ano81-104] Anonymous. Energy report.

IEEE Spectrum, 18(4):12–13, April 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:ERe

[Ano81-105] Anonymous. Energy report.

IEEE Spectrum, 18(5):18–19, May 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:ERf

[Ano81-106] Anonymous. Energy report.

IEEE Spectrum, 18(6):17, June 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:ERg

[Ano81-107] Anonymous. Energy report.

IEEE Spectrum, 18(7):18–19, July 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:ERh

[Ano81-108] Anonymous. Energy report.

IEEE Spectrum, 18(8):15, August 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:ERi

[Ano81-109] Anonymous. Energy report. IEEE Spectrum, 18(9):20–21, September 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:ERj

[Ano81-110] Anonymous. Energy report. IEEE Spectrum, 18(10):22– 23, October 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:ERk

[Ano81-111] Anonymous. Energy report. IEEE Spectrum, 18(11):22–23, November 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:ERl

[Ano81-112] Anonymous. Energy report. IEEE Spectrum, 18(12):12–13, December 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:ELa

[Ano81-113] Anonymous. The engineer at large. IEEE Spectrum, 18(1): 26–27, January 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:ELb

[Ano81-114] Anonymous. The engineer at large. IEEE Spectrum, 18(3):19, March 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:ELc

[Ano81-115] Anonymous. The engineer at large. IEEE Spectrum, 18 (5):14–15, May 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:ELd

[Ano81-116] Anonymous. The engineer at large. IEEE Spectrum, 18(7):16, July 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:ELe

[Ano81-117] Anonymous. The engineer at large. IEEE Spectrum, 18 (9):18, September 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:ELf

[Ano81-118] Anonymous. The engineer at large. IEEE Spectrum, 18(11): 16–17, November 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:FR

[Ano81-119] Anonymous. For further reading. IEEE Spectrum, 18(1): 83–84, January 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Fa

[Ano81-120] Anonymous. Forum. IEEE Spectrum, 18(1):14–23, January 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Fb

[Ano81-121] Anonymous. Forum. IEEE Spectrum, 18(2):11–19, February 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Fc

[Ano81-122] Anonymous. Forum. IEEE Spectrum, 18(3):16–18, March 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Fd

[Ano81-123] Anonymous. Forum. IEEE Spectrum, 18(4):10–11, April 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Fe

[Ano81-124] Anonymous. Forum. IEEE Spectrum, 18(5):8–13, May 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Ff

[Ano81-125] Anonymous. Forum. IEEE Spectrum, 18(6):14–16, June 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Fg

[Ano81-126] Anonymous. Forum. IEEE Spectrum, 18(7):14–15, July 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Fh

[Ano81-127] Anonymous. Forum. IEEE Spectrum, 18(8):8–14, August 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Fi

[Ano81-128] Anonymous. Forum. IEEE Spectrum, 18(9):12–17, September 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Fj

[Ano81-129] Anonymous. Forum. IEEE Spectrum, 18(10):10–21, October 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Fk

[Ano81-130] Anonymous. Forum. IEEE Spectrum, 18(11):10–15, November 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:Fl

[Ano81-131] Anonymous. Forum. IEEE Spectrum, 18(12):10–11, December 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:GA

[Ano81-132] Anonymous. Guide to authors. IEEE Spectrum, 18(1): 85–87, January 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:ECN

[Ano81-133] Anonymous. I. Engineering changes: New and improved systems and hardware. IEEE Spectrum, 18(4):29–33, April 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:IPPa

[Ano81-134] Anonymous. IEEE publications paper calls. IEEE Spectrum, 18(1):120, January 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:IPPb

[Ano81-135] Anonymous. IEEE publications paper calls. IEEE Spectrum, 18(3):91, March 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:IPPc

[Ano81-136] Anonymous. IEEE publications paper calls. IEEE Spectrum, 18(5):22–24, May 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:IPPd

[Ano81-137] Anonymous. IEEE publications paper calls. IEEE Spectrum, 18 (11):85–87, November 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:ITCa

[Ano81-138] Anonymous. IEEE tables of contents. IEEE Spectrum, 18 (1):90–97, January 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:ITCb

[Ano81-139] Anonymous. IEEE tables of contents. IEEE Spectrum, 18 (2):62–63, February 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:ITCc

[Ano81-140] Anonymous. IEEE tables of contents. IEEE Spectrum, 18 (3):66–75, March 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:ITCd

[Ano81-141] Anonymous. IEEE tables of contents. IEEE Spectrum, 18 (4):83–87, April 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:ITCj

[Ano81-142] Anonymous. IEEE tables of contents. IEEE Spectrum, 18(10): 128–137, October 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:ITCk

[Ano81-143] Anonymous. IEEE tables of contents. IEEE Spectrum, 18(11): 83–85, November 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:ITCl

[Ano81-144] Anonymous. IEEE tables of contents. IEEE Spectrum, 18(12):

86–97, December 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:ITCe

[Ano81-145] Anonymous. IEEE tables of contents for current and future publications. IEEE Spectrum, 18(5):58–63, May 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:ITCf

[Ano81-146] Anonymous. IEEE tables of contents for current and future publications. IEEE Spectrum, 18(6):64–66, June 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:ITCg

[Ano81-147] Anonymous. IEEE tables of contents for current and future publications. IEEE Spectrum, 18(7):64–67, July 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:ITCh

[Ano81-148] Anonymous. IEEE tables of contents for current and future publications. IEEE Spectrum, 18 (8):66–72, August 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:ITCi

[Ano81-149] Anonymous. IEEE tables of contents for current and future publications. IEEE Spectrum, 18 (9):82–91, September 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:IHS

[Ano81-150] Anonymous. II. The human side: Training and tools to help the operator. IEEE Spectrum, 18(4):33–37, April 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:IMF

[Ano81-151] Anonymous. III. Management faces the facts: Damaging accidents are possible. IEEE Spectrum, 18(4):37–40, April 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:INN

[Ano81-152] Anonymous. IV. NRC’s new zeal, a face-lift, and increased regulation. IEEE Spectrum, 18 (4):40–42, April 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:MPb

[Ano81-153] Anonymous. Meeting previews. IEEE Spectrum, 18(2): 77–87, February 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:MPc

[Ano81-154] Anonymous. Meeting previews.

IEEE Spectrum, 18(3):76–81, March 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:MPd

[Ano81-155] Anonymous. Meeting previews.

IEEE Spectrum, 18(4):92–101, April 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:MPe

[Ano81-156] Anonymous. Meeting previews. IEEE Spectrum, 18(5):66–70, May 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:MPf

[Ano81-157] Anonymous. Meeting previews. IEEE Spectrum, 18(6):67–71, June 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:MPg

[Ano81-158] Anonymous. Meeting previews. IEEE Spectrum, 18(7):68–69, July 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:MPh

[Ano81-159] Anonymous. Meeting previews. IEEE Spectrum, 18(8):58–63, August 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:MPi

[Ano81-160] Anonymous. Meeting previews. IEEE Spectrum, 18(9): 70–73, September 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:MPj

[Ano81-161] Anonymous. Meeting previews. IEEE Spectrum, 18(10):

110–121, October 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:MPk

[Ano81-162] Anonymous. Meeting previews. IEEE Spectrum, 18(11): 74–80, November 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:MPl

[Ano81-163] Anonymous. Meeting previews. IEEE Spectrum, 18(12): 72–79, December 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:MPa

[Ano81-164] Anonymous. Meting previews. IEEE Spectrum, 18(1): 106–119, January 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:NRIa

[Ano81-165] Anonymous. New and recent IEEE publications. IEEE Spectrum, 18(1):88–89, January 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:NRIb

[Ano81-166] Anonymous. New and recent IEEE publications. IEEE Spectrum, 18(2):60–61, February 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:NRIc

[Ano81-167] Anonymous. New and recent IEEE publications. IEEE Spectrum, 18(3):64–65, March 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:NRId

[Ano81-168] Anonymous. New and recent IEEE publications. IEEE Spectrum, 18(4):72–79, April 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:NRIe

[Ano81-169] Anonymous. New and recent IEEE publications. IEEE Spectrum, 18(5):56–57, May 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:NRIf

[Ano81-170] Anonymous. New and recent IEEE publications. IEEE Spectrum, 18(6):62–63, June 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:NRIg

[Ano81-171] Anonymous. New and recent IEEE publications. IEEE Spectrum, 18(7):60–63, July 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:NRIh

[Ano81-172] Anonymous. New and recent IEEE publications. IEEE Spectrum, 18(8):56–57, August 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:NRIi

[Ano81-173] Anonymous. New and recent IEEE publications. IEEE Spectrum, 18(9):68–69, September 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:NRIj

[Ano81-174] Anonymous. New and recent IEEE publications. IEEE Spectrum, 18(10):108–109, October 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:NRIk

[Ano81-175] Anonymous. New and recent IEEE publications. IEEE Spectrum, 18(11):72–73, November 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:NRIl

[Ano81-176] Anonymous. New and recent IEEE publications. IEEE Spectrum, 18(12):70–71, December 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:NPAb

[Ano81-177] Anonymous. New product application: Microprocessorcontrolled D-size plotter is sheetfed, automates color and linewidth selection. IEEE Spectrum, 18(2):72–76, February 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:NPAa

[Ano81-178] Anonymous. New product applications. IEEE Spectrum, 18 (1):98–100, January 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:NPAg

[Ano81-179] Anonymous. New product applications. IEEE Spectrum, 18 (8):64–65, August 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:NPAk

[Ano81-180] Anonymous. New product applications: 50-MHz oscilloscopes have 2-mV sensitivity, 10-ns/ division sweep, and dual timebase versions. IEEE Spectrum, 18(12):80–81, December 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:NPAi

[Ano81-181] Anonymous. New product applications: Digital analysis system features up to 104 channels of acquisition, sample rates to 330 MHz, resolution to 1.5 ns. IEEE Spectrum, 18(10):122–125, October 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:NPAj

[Ano81-182] Anonymous. New product applications: Digital waveform storage option permits repetitive sampling to 100 MHz at time resolutions as low as 100 ps. IEEE Spectrum, 18(11): 88–104, November 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:NPAh

[Ano81-183] Anonymous. New product applications: Microprocessor development system has 64 kilobytes of memory; companion emulator has 8-kilobyte static RAM. IEEE Spectrum, 18(9):74–79, September 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:NPAd

[Ano81-184] Anonymous. New product applications: Multibus-compatible, all-CMOS microcomputer can have 32-kiloword memory, consumes less than 0.5 W. IEEE Spectrum, 18(4):80–82, April 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:NPAf

[Ano81-185] Anonymous. New product applications: Peripherals enhance performance of MC68000 VLSI 16-bit microprocessor. IEEE Spectrum, 18(6):72–74, June 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:NPAc

[Ano81-186] Anonymous. New product applications: Single-board computer based on 86/10 CPU contains 8 kilobytes GF RAM, expandable to 16 kilobytes. IEEE Spectrum, 18(3):82–90, March 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:NPAe

[Ano81-187] Anonymous. New product applications: Three-chip, 32-bit mainframe has 1 terabyte virtual memory address space. IEEE Spectrum, 18(5):64–65, May 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:NIa

[Ano81-188] Anonymous. News from industry. IEEE Spectrum, 18(1): 120–122, January 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:NIb

[Ano81-189] Anonymous. News from industry. IEEE Spectrum, 18(2): 88–90, February 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:NIc

[Ano81-190] Anonymous. News from industry. IEEE Spectrum, 18 (3):104–106, March 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:NId

[Ano81-191] Anonymous. News from industry. IEEE Spectrum, 18 (4):102–103, April 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:NIe

[Ano81-192] Anonymous. News from industry. IEEE Spectrum, 18 (5):78–80, May 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:NIf

[Ano81-193] Anonymous. News from industry. IEEE Spectrum, 18 (6):85–86, June 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:NIg

[Ano81-194] Anonymous. News from industry. IEEE Spectrum, 18 (7):79–80, July 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:NIh

[Ano81-195] Anonymous. News from industry. IEEE Spectrum, 18 (8):79–80, August 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:NIi

[Ano81-196] Anonymous. News from industry. IEEE Spectrum, 18(9):

91–92, September 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:NIj

[Ano81-197] Anonymous. News from industry. IEEE Spectrum, 18(10): 137–138, October 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:NIk

[Ano81-198] Anonymous. News from industry. IEEE Spectrum, 18(11): 105–106, November 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:NIl

[Ano81-199] Anonymous. News from industry. IEEE Spectrum, 18(12): 97–98, December 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:NWa

[Ano81-200] Anonymous. News from Washington. IEEE Spectrum, 18 (1):24, January 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:NWb

[Ano81-201] Anonymous. News from Washington. IEEE Spectrum, 18 (2):22, February 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:NWc

[Ano81-202] Anonymous. News from Washington. IEEE Spectrum, 18 (3):20, March 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:NWd

[Ano81-203] Anonymous. News from Washington. IEEE Spectrum, 18 (4):12, April 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:NWe

[Ano81-204] Anonymous. News from Washington. IEEE Spectrum, 18 (5):18, May 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:NWf

[Ano81-205] Anonymous. News from Washington. IEEE Spectrum, 18 (6):17, June 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:NWg

[Ano81-206] Anonymous. News from Washington. IEEE Spectrum, 18 (7):18, July 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:NWh

[Ano81-207] Anonymous. News from Washington. IEEE Spectrum, 18 (8):15, August 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:NWi

[Ano81-208] Anonymous. News from Washington. IEEE Spectrum, 18 (9):20, September 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:NWj

[Ano81-209] Anonymous. News from Washington. IEEE Spectrum, 18 (10):22, October 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:NWk

[Ano81-210] Anonymous. News from Washington. IEEE Spectrum, 18 (11):22, November 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:NWl

[Ano81-211] Anonymous. News from Washington. IEEE Spectrum, 18 (12):12, December 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:SIa

[Ano81-212] Anonymous. Scanning the Institute. IEEE Spectrum, 18 (1):28, January 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:SIb

[Ano81-213] Anonymous. Scanning the Institute. IEEE Spectrum, 18 (2):24, February 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:SIc

[Ano81-214] Anonymous. Scanning the Institute. IEEE Spectrum, 18(3):22, March 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:SId

[Ano81-215] Anonymous. Scanning the Institute. IEEE Spectrum, 18(4):16, April 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:SIe

[Ano81-216] Anonymous. Scanning the Institute. IEEE Spectrum, 18(5):20, May 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:SIf

[Ano81-217] Anonymous. Scanning the Institute. IEEE Spectrum, 18(6):18, June 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:SIg

[Ano81-218] Anonymous. Scanning the Institute. IEEE Spectrum, 18(7):20, July 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:SIh

[Ano81-219] Anonymous. Scanning the Institute. IEEE Spectrum, 18(8):16, August 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:SIi

[Ano81-220] Anonymous. Scanning the Institute. IEEE Spectrum, 18 (9):22, September 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:SIj

[Ano81-221] Anonymous. Scanning the Institute. IEEE Spectrum, 18 (10):24, October 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:SIk

[Ano81-222] Anonymous. Scanning the Institute. IEEE Spectrum, 18 (11):24, November 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:SIl

[Ano81-223] Anonymous. Scanning the Institute. IEEE Spectrum, 18 (12):14, December 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:TSa

[Ano81-224] Anonymous. Technically speaking. IEEE Spectrum, 18(1): 32–34, January 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:TSb

[Ano81-225] Anonymous. Technically speaking. IEEE Spectrum, 18(4): 24–26, April 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:TSc

[Ano81-226] Anonymous. Technically speaking. IEEE Spectrum, 18(6):

23–24, June 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:TSd

[Ano81-227] Anonymous. Technically speaking. IEEE Spectrum, 18(8): 19–20, August 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:TSe

[Ano81-228] Anonymous. Technically speaking. IEEE Spectrum, 18(10): 30–32, October 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:TSf

[Ano81-229] Anonymous. Technically speaking. IEEE Spectrum, 18(12): 20–22, December 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:VLA

[Ano81-230] Anonymous. V. A look ahead:

More plants to come; computer’s role on rise. IEEE Spectrum, 18(4):43–44, April 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:WEHa

[Ano81-231] Anonymous. What ever happened to … ? IEEE Spectrum, 18(2):26–28, February 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:WEHb

[Ano81-232] Anonymous. What ever happened to … ? IEEE Spectrum, 18(4):20–22, April 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:WEHc

[Ano81-233] Anonymous. What ever happened to … ELF? IEEE Spectrum, 18(6):22, June 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:WEHd

[Ano81-234] Anonymous. What ever happened to uncol? IEEE Spectrum, 18(8):18, August 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:WHC

[Ano81-235] Anonymous. Whatever happened to CCDs? IEEE Spectrum, 18(10):26–29, October 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1981:WHM

[Ano81-236] Anonymous. Whatever happened to molecular electronics? IEEE Spectrum, 18(12): 17–19, December 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:IGA

[Ano82a] Anonymous. 1982 IEEE guide for authors. IEEE Spectrum, 19 (1):78–83, January 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:AAd

[Ano82b] Anonymous. About the authors. IEEE Spectrum, 19(10): 114–115, October 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Aa

[Ano82c] Anonymous. [advertisement].

IEEE Spectrum, 19(1):1–3, January 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Ab

[Ano82d] Anonymous. [advertisement].

IEEE Spectrum, 19(1):2, January 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Ac

[Ano82e] Anonymous. [advertisement].

IEEE Spectrum, 19(1):4, January 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Ad

[Ano82f] Anonymous. [advertisement].

IEEE Spectrum, 19(2):1–3, February 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Ae

[Ano82g] Anonymous. [advertisement].

IEEE Spectrum, 19(2):2, February 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Af

[Ano82h] Anonymous. [advertisement]. IEEE Spectrum, 19(2):4, February 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Ag

[Ano82i] Anonymous. [advertisement]. IEEE Spectrum, 19(2):1, February 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Ah

[Ano82j] Anonymous. [advertisement]. IEEE Spectrum, 19(3):1–3, March 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Ai

[Ano82k] Anonymous. [advertisement]. IEEE Spectrum, 19(3):2, March 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Aj

[Ano82l] Anonymous. [advertisement]. IEEE Spectrum, 19(3):4–7, March 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Ak

[Ano82m] Anonymous. [advertisement]. IEEE Spectrum, 19(3):1–2, March 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Al

[Ano82n] Anonymous. [advertisement].

IEEE Spectrum, 19(4):1–3, April 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Am

[Ano82o] Anonymous. [advertisement].

IEEE Spectrum, 19(4):2, April 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:An

[Ano82p] Anonymous. [advertisement].

IEEE Spectrum, 19(4):4–7, April 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Ao

[Ano82q] Anonymous. [advertisement].

IEEE Spectrum, 19(4):1–2, April 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Ap

[Ano82r] Anonymous. [advertisement].

IEEE Spectrum, 19(5):1–4, May 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Aq

[Ano82s] Anonymous. [advertisement].

IEEE Spectrum, 19(5):4–7, May 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Ar

[Ano82t] Anonymous. [advertisement]. IEEE Spectrum, 19(6):1–4, June 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:As

[Ano82u] Anonymous. [advertisement]. IEEE Spectrum, 19(6):4–5, June 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:At

[Ano82v] Anonymous. [advertisement]. IEEE Spectrum, 19(7):1–3, July 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Au

[Ano82w] Anonymous. [advertisement]. IEEE Spectrum, 19(7):2, July 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Av

[Ano82x] Anonymous. [advertisement]. IEEE Spectrum, 19(7):4–5, July 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Aw

[Ano82y] Anonymous. [advertisement]. IEEE Spectrum, 19(8):1–3, August 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Ax

[Ano82z] Anonymous. [advertisement].

IEEE Spectrum, 19(8):2, August 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Ay

[Ano82-27] Anonymous. [advertisement].

IEEE Spectrum, 19(8):4–5, August 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Az

[Ano82-28] Anonymous. [advertisement].

IEEE Spectrum, 19(9):1–3, September 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Aaa

[Ano82-29] Anonymous. [advertisement].

IEEE Spectrum, 19(9):2, September 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Aab

[Ano82-30] Anonymous. [advertisement].

IEEE Spectrum, 19(9):4–7, September 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Aac

[Ano82-31] Anonymous. [advertisement].

IEEE Spectrum, 19(9):1–2, September 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Aae

[Ano82-32] Anonymous. [advertisement]. IEEE Spectrum, 19(11):1–3, November 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Aaf

[Ano82-33] Anonymous. [advertisement]. IEEE Spectrum, 19(11):2, November 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Aag

[Ano82-34] Anonymous. [advertisement]. IEEE Spectrum, 19(11):4–7, November 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Aah

[Ano82-35] Anonymous. [advertisement]. IEEE Spectrum, 19(11):1, November 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Aai

[Ano82-36] Anonymous. [advertisement]. IEEE Spectrum, 19(12):1–3, December 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Aaj

[Ano82-37] Anonymous. [advertisement]. IEEE Spectrum, 19(12):2, December 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Aak

[Ano82-38] Anonymous. [advertisement].

IEEE Spectrum, 19(12):4–5, December 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Aal

[Ano82-39] Anonymous. [advertisement].

IEEE Spectrum, 19(12):1–2, December 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:AIa

[Ano82-40] Anonymous. Author index.

IEEE Spectrum, 19(12):76–84, December 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:AIM

[Ano82-41] Anonymous. Awards: IEEE major medalists for 1982. IEEE Spectrum, 19(5):76–77, May 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:BC

[Ano82-42] Anonymous. [back cover]. IEEE Spectrum, 19(10):c4, October 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:BBAb

[Ano82-43] Anonymous. Best bits: Applications of microprocessors. IEEE Spectrum, 19(3):24–25, March 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:BBAc

[Ano82-44] Anonymous. Best bits: Applications of microprocessors. IEEE Spectrum, 19(5):26–29, May 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:BBAf

[Ano82-45] Anonymous. Best bits: Applications of microprocessors. IEEE Spectrum, 19(11):17, November 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:BBAd

[Ano82-46] Anonymous. Best bits: Applications of microprocessors: ‘flying with microprocessors. IEEE Spectrum, 19(7):17–19, July 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:BBAe

[Ano82-47] Anonymous. Best bits: Applications of microprocessors: Making the Shuttle reliable. IEEE Spectrum, 19(9):22–23, September 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:BBAa

[Ano82-48] Anonymous. Best bits: Applications of microprocessors: Smart heart. IEEE Spectrum, 19(1): 20–23, January 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:BRa

[Ano82-49] Anonymous. Book reviews.

IEEE Spectrum, 19(1):98–103, January 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:BRb

[Ano82-50] Anonymous. Book reviews.

IEEE Spectrum, 19(2):78–81, February 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:BRc

[Ano82-51] Anonymous. Book reviews.

IEEE Spectrum, 19(3):82–83, March 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:BRd

[Ano82-52] Anonymous. Book reviews.

IEEE Spectrum, 19(4):82–83, April 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:BRe

[Ano82-53] Anonymous. Book reviews.

IEEE Spectrum, 19(5):85–88, May 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:BRf

[Ano82-54] Anonymous. Book reviews.

IEEE Spectrum, 19(6):78–79, June 1982. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:BRg

[Ano82-55] Anonymous. Book reviews. IEEE Spectrum, 19(7):65–66, July 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:BRh

[Ano82-56] Anonymous. Book reviews. IEEE Spectrum, 19(8):66–67, August 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:BRi

[Ano82-57] Anonymous. Book reviews. IEEE Spectrum, 19(9):86–88, September 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:BRj

[Ano82-58] Anonymous. Book reviews. IEEE Spectrum, 19(10):128– 137, October 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:BRk

[Ano82-59] Anonymous. Book reviews. IEEE Spectrum, 19(11):60–63, November 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:BRl

[Ano82-60] Anonymous. Book reviews. IEEE Spectrum, 19(12):62–73, December 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:SHR

[Ano82-61] Anonymous. Spectrum’s hardware review. IEEE Spectrum, 19(2):83, February 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Cb

[Ano82-62] Anonymous. Calendar. IEEE Spectrum, 19(1):6–9, January 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Cd

[Ano82-63] Anonymous. Calendar. IEEE Spectrum, 19(2):5–8, February 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Cf

[Ano82-64] Anonymous. Calendar. IEEE Spectrum, 19(3):8–9, March 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Ch

[Ano82-65] Anonymous. Calendar. IEEE Spectrum, 19(4):8–11, April 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Cj

[Ano82-66] Anonymous. Calendar. IEEE Spectrum, 19(5):8–11, May 1982.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Cl

[Ano82-67] Anonymous. Calendar. IEEE Spectrum, 19(6):6–7, June 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Cn

[Ano82-68] Anonymous. Calendar. IEEE Spectrum, 19(7):6–7, July 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Cp

[Ano82-69] Anonymous. Calendar. IEEE Spectrum, 19(8):6–7, August 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Cr

[Ano82-70] Anonymous. Calendar. IEEE Spectrum, 19(9):8–16, September 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Cu

[Ano82-71] Anonymous. Calendar. IEEE Spectrum, 19(10):8–15, October 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Cw

[Ano82-72] Anonymous. Calendar. IEEE Spectrum, 19(11):8–11, November 1982. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Cy

[Ano82-73] Anonymous. Calendar. IEEE Spectrum, 19(12):6–8, December 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:CSa

[Ano82-74] Anonymous. Coming in Spectrum. IEEE Spectrum, 19(1): 26–28, January 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:CSb

[Ano82-75] Anonymous. Coming in Spectrum. IEEE Spectrum, 19(2): 20–21, February 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:CSc

[Ano82-76] Anonymous. Coming in Spectrum. IEEE Spectrum, 19 (3):28–30, March 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:CSd

[Ano82-77] Anonymous. Coming in Spectrum. IEEE Spectrum, 19 (4):17–19, April 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:CSe

[Ano82-78] Anonymous. Coming in Spectrum. IEEE Spectrum, 19 (5):32–36, May 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:CSf

[Ano82-79] Anonymous. Coming in Spectrum. IEEE Spectrum, 19 (6):16–16P, June 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:CSg

[Ano82-80] Anonymous. Coming in Spectrum. IEEE Spectrum, 19 (7):22–24, July 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:CSh

[Ano82-81] Anonymous. Coming in Spectrum. IEEE Spectrum, 19 (8):14–15, August 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:CSi

[Ano82-82] Anonymous. Coming in Spectrum. IEEE Spectrum, 19(9): 26–28, September 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:CSj

[Ano82-83] Anonymous. Coming in Spectrum. IEEE Spectrum, 19(10): 22–23, October 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:CSk

[Ano82-84] Anonymous. Coming in Spectrum. IEEE Spectrum, 19(11):

20–22, November 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:CSl

[Ano82-85] Anonymous. Coming in Spectrum. IEEE Spectrum, 19 (12):14, December 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Ca

[Ano82-86] Anonymous. Contents. IEEE Spectrum, 19(1):3, January 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Cc

[Ano82-87] Anonymous. Contents. IEEE Spectrum, 19(2):3, February 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Ce

[Ano82-88] Anonymous. Contents. IEEE Spectrum, 19(3):3, March 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Cg

[Ano82-89] Anonymous. Contents. IEEE Spectrum, 19(4):3, April 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Ci

[Ano82-90] Anonymous. Contents. IEEE Spectrum, 19(5):3, May 1982.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Ck

[Ano82-91] Anonymous. Contents. IEEE Spectrum, 19(6):3, June 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Cm

[Ano82-92] Anonymous. Contents. IEEE Spectrum, 19(7):3, July 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Co

[Ano82-93] Anonymous. Contents. IEEE Spectrum, 19(8):3, August 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Cq

[Ano82-94] Anonymous. Contents. IEEE Spectrum, 19(9):3, September 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Ct

[Ano82-95] Anonymous. Contents. IEEE Spectrum, 19(10):3, October 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Cv

[Ano82-96] Anonymous. Contents. IEEE Spectrum, 19(11):3, November 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Cx

[Ano82-97] Anonymous. Contents. IEEE Spectrum, 19(12):3, December 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:ETTa

[Ano82-98] Anonymous. EEs’ tools toys.

IEEE Spectrum, 19(2):22–23, February 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:ETTb

[Ano82-99] Anonymous. EEs’ tools toys.

IEEE Spectrum, 19(4):20–23, April 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:ETTc

[Ano82-100] Anonymous. EEs’ tools toys.

IEEE Spectrum, 19(6):17, June 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:ETTd

[Ano82-101] Anonymous. EEs’ tools toys.

IEEE Spectrum, 19(8):16, August 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:ETTe

[Ano82-102] Anonymous. EEs’ tools toys.

IEEE Spectrum, 19(11):16, November 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:ERa

[Ano82-103] Anonymous. Energy report. IEEE Spectrum, 19(1): 24–25, January 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:ERb

[Ano82-104] Anonymous. Energy report. IEEE Spectrum, 19(2):16–19, February 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:ERc

[Ano82-105] Anonymous. Energy report. IEEE Spectrum, 19(3):26–27, March 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:ERd

[Ano82-106] Anonymous. Energy report. IEEE Spectrum, 19(4):16, April 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:ERe

[Ano82-107] Anonymous. Energy report. IEEE Spectrum, 19(5):30–31, May 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:ERf

[Ano82-108] Anonymous. Energy report. IEEE Spectrum, 19(6):14–15, June 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:ERg

[Ano82-109] Anonymous. Energy report.

IEEE Spectrum, 19(7):20–21, July 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:ERh

[Ano82-110] Anonymous. Energy report.

IEEE Spectrum, 19(8):12–13, August 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:ERi

[Ano82-111] Anonymous. Energy report.

IEEE Spectrum, 19(9):24–25, September 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:ERj

[Ano82-112] Anonymous. Energy report.

IEEE Spectrum, 19(10):16– 16P, October 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:ERk

[Ano82-113] Anonymous. Energy report.

IEEE Spectrum, 19(11):18–19, November 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:ERl

[Ano82-114] Anonymous. Energy report.

IEEE Spectrum, 19(12):12–13, December 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:ELa

[Ano82-115] Anonymous. The engineer at large. IEEE Spectrum, 19(1): 18–19, January 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:ELb

[Ano82-116] Anonymous. The engineer at large. IEEE Spectrum, 19 (3):20–23, March 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:ELc

[Ano82-117] Anonymous. The engineer at large. IEEE Spectrum, 19 (5):22–25, May 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:ELd

[Ano82-118] Anonymous. The engineer at large. IEEE Spectrum, 19(7):16, July 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:ELe

[Ano82-119] Anonymous. The engineer at large. IEEE Spectrum, 19(9): 20–21, September 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:ELf

[Ano82-120] Anonymous. The engineer at large. IEEE Spectrum, 19 (12):15, December 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:FR

[Ano82-121] Anonymous. For further reading. IEEE Spectrum, 19(1): 74–75, January 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Fa

[Ano82-122] Anonymous. Forum. IEEE Spectrum, 19(1):10–16, January 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Fb

[Ano82-123] Anonymous. Forum. IEEE Spectrum, 19(2):9–15, February 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Fc

[Ano82-124] Anonymous. Forum. IEEE Spectrum, 19(3):10–17, March 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Fd

[Ano82-125] Anonymous. Forum. IEEE Spectrum, 19(4):12–15, April 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Fe

[Ano82-126] Anonymous. Forum. IEEE Spectrum, 19(5):12–17, May 1982.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Ff

[Ano82-127] Anonymous. Forum. IEEE Spectrum, 19(6):8–13, June 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Fg

[Ano82-128] Anonymous. Forum. IEEE Spectrum, 19(7):8–13, July 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Fh

[Ano82-129] Anonymous. Forum. IEEE Spectrum, 19(8):8–11, August 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Fi

[Ano82-130] Anonymous. Forum. IEEE Spectrum, 19(9):17, September 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Fj

[Ano82-131] Anonymous. Forum. IEEE Spectrum, 19(10):17–21, October 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Fk

[Ano82-132] Anonymous. Forum. IEEE Spectrum, 19(11):12–13, November 1982. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Fl

[Ano82-133] Anonymous. Forum. IEEE Spectrum, 19(12):9–11, December 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:FCa

[Ano82-134] Anonymous. [front cover]. IEEE Spectrum, 19(10):c1, October 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:IFS

[Ano82-135] Anonymous. IEEE Field, Service, and Prize Paper Award winners. IEEE Spectrum, 19 (8):58–59, August 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:IPPa

[Ano82-136] Anonymous. IEEE publications paper calls. IEEE Spectrum, 19(1):97, January 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:IPPb

[Ano82-137] Anonymous. IEEE publications paper calls. IEEE Spectrum, 19(6):71, June 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:IPPc

[Ano82-138] Anonymous. IEEE publications paper calls. IEEE Spectrum, 19(8):68, August 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:IPPd

[Ano82-139] Anonymous. IEEE publications paper calls. IEEE Spectrum, 19 (10):125, October 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:ITCa

[Ano82-140] Anonymous. IEEE tables of contents. IEEE Spectrum, 19 (1):86–92, January 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:ITCb

[Ano82-141] Anonymous. IEEE tables of contents. IEEE Spectrum, 19 (2):84–96, February 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:ITCe

[Ano82-142] Anonymous. IEEE tables of contents. IEEE Spectrum, 19 (5):89–95, May 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:ITCf

[Ano82-143] Anonymous. IEEE tables of contents. IEEE Spectrum, 19 (6):80–89, June 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:ITCg

[Ano82-144] Anonymous. IEEE tables of contents. IEEE Spectrum, 19 (7):67–71, July 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:ITCh

[Ano82-145] Anonymous. IEEE tables of contents. IEEE Spectrum, 19 (8):69–71, August 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:ITCi

[Ano82-146] Anonymous. IEEE tables of contents. IEEE Spectrum, 19(9): 82–85, September 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:ITCk

[Ano82-147] Anonymous. IEEE tables of contents. IEEE Spectrum, 19(11): 74–89, November 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:ITCl

[Ano82-148] Anonymous. IEEE tables of contents. IEEE Spectrum, 19(12): 59–61, December 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:ITCc

[Ano82-149] Anonymous. IEEE tables of contents for current and future publications. IEEE Spectrum, 19 (3):84–95, March 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:ITCd

[Ano82-150] Anonymous. IEEE tables of contents for current and future publications. IEEE Spectrum, 19(4):84–95, April 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:ITCj

[Ano82-151] Anonymous. IEEE tables of contents for current and future publications. IEEE Spectrum, 19 (10):126–127, October 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:IDH

[Ano82-152] Anonymous. II. Design: High technology, low cost. IEEE Spectrum, 19(12):22–30, December 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:IPM

[Ano82-153] Anonymous. III. Psychology: Mind and matter. IEEE Spectrum, 19(12):30–33, December 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:IBC

[Ano82-154] Anonymous. [inside back cover]. IEEE Spectrum, 19(10):c3, October 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:IFC

[Ano82-155] Anonymous. [inside front cover]. IEEE Spectrum, 19(10):c2, October 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:IAT

[Ano82-156] Anonymous. International arms transfers: Imports and exports. IEEE Spectrum, 19(10): 76–79, October 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:LDC

[Ano82-157] Anonymous. “… Diapers, carriage, bottle-warmer, supplemental group life insurance” [advertisement]. IEEE Spectrum, 19 (10):28–29, October 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:LS

[Ano82-158] Anonymous. List of staff.

IEEE Spectrum, 19(10):4, October 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:MPa

[Ano82-159] Anonymous. Meeting previews. IEEE Spectrum, 19(1): 93–97, January 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:MPb

[Ano82-160] Anonymous. Meeting previews. IEEE Spectrum, 19(2): 72–77, February 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:MPc

[Ano82-161] Anonymous. Meeting previews.

IEEE Spectrum, 19(3):76–79, March 1982. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:MPd

[Ano82-162] Anonymous. Meeting previews. IEEE Spectrum, 19(4):74–79, April 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:MPe

[Ano82-163] Anonymous. Meeting previews. IEEE Spectrum, 19(5):81–84, May 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:MPf

[Ano82-164] Anonymous. Meeting previews. IEEE Spectrum, 19(6):68–70, June 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:MPg

[Ano82-165] Anonymous. Meeting previews. IEEE Spectrum, 19(7):63–64, July 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:MPh

[Ano82-166] Anonymous. Meeting previews. IEEE Spectrum, 19(8):62–65, August 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:MPi

[Ano82-167] Anonymous. Meeting previews. IEEE Spectrum, 19(9):

76–79, September 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:MPj

[Ano82-168] Anonymous. Meeting previews. IEEE Spectrum, 19(10): 118–121, October 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:MPk

[Ano82-169] Anonymous. Meeting previews. IEEE Spectrum, 19(11): 70–73, November 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:MPl

[Ano82-170] Anonymous. Meeting previews. IEEE Spectrum, 19(12): 54–55, December 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:MTE

[Ano82-171] Anonymous. More than ever, Atari home computers are speaking your language [advertisement]. IEEE Spectrum, 19 (10):1, October 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:NRIa

[Ano82-172] Anonymous. New and recent IEEE publications. IEEE Spectrum, 19(1):84–85, January 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:NRIb

[Ano82-173] Anonymous. New and recent IEEE publications. IEEE Spectrum, 19(2):70–71, February 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:NRIc

[Ano82-174] Anonymous. New and recent IEEE publications. IEEE Spectrum, 19(3):74–75, March 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:NRId

[Ano82-175] Anonymous. New and recent IEEE publications. IEEE Spectrum, 19(4):72–73, April 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:NRIe

[Ano82-176] Anonymous. New and recent IEEE publications. IEEE Spectrum, 19(5):78–80, May 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:NRIf

[Ano82-177] Anonymous. New and recent IEEE publications. IEEE Spectrum, 19(6):66–67, June 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:NRIg

[Ano82-178] Anonymous. New and recent IEEE publications. IEEE Spectrum, 19(7):61–62, July 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:NRIh

[Ano82-179] Anonymous. New and recent IEEE publications. IEEE Spectrum, 19(8):60–61, August 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:NRIi

[Ano82-180] Anonymous. New and recent IEEE publications. IEEE Spectrum, 19(9):73–75, September 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:NRIj

[Ano82-181] Anonymous. New and recent IEEE publications. IEEE Spectrum, 19(10):116–117, October 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:NRIk

[Ano82-182] Anonymous. New and recent IEEE publications. IEEE Spectrum, 19(11):58–59, November 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:NRIl

[Ano82-183] Anonymous. New and recent IEEE publications. IEEE Spectrum, 19(12):52–53, December 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:NPAj

[Ano82-184] Anonymous. New product applications. IEEE Spectrum, 19 (10):122–124, October 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:NPAa

[Ano82-185] Anonymous. New product applications: Array processor performs 5 million floating-point operations per second, has 200-ns operation time. IEEE Spectrum, 19(1):104–118, January 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:NPAk

[Ano82-186] Anonymous. New product applications: Controller supports up to eight array processors, each transferring at the maximum 3megabyte-per-second rate. IEEE Spectrum, 19(11):64–69, November 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:NPAi

[Ano82-187] Anonymous. New product applications: EJAO-size drafting plotter permits programmed selection of up to eight pens, operates at speeds to 60 cm/s. IEEE Spectrum, 19(9):80–81, September 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:NPAl

[Ano82-188] Anonymous. New product applications: Interactive graphics design system uses station based on 8086 microprocessor and 68000 cluster controllvi. IEEE Spectrum, 19(12):56–58, December 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:NPAg

[Ano82-189] Anonymous. New product applications: Logic analyzer displays data in three colors, offers up to 104 channels of data acquisition. IEEE Spectrum, 19(7):72–79, July 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:NPAf

[Ano82-190] Anonymous. New product applications: Monolithic 16-bit D/A converter available in two accuracy grades for 0^{∘}-to-75^{∘}C range. IEEE Spectrum, 19 (6):72–77, June 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:NPAd

[Ano82-191] Anonymous. New product applications: Plotter draws multicolored graphics at 25-micrometer resolution; ‘microgrip drive uses no stepper motors. IEEE Spectrum, 19(4):80–81, April 1982.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:NPAe

[Ano82-192] Anonymous. New product applications: Pocket-size computer terminal stores 4096 characters in RAM, has 4 kilobytes of ROM. IEEE Spectrum, 19(5):96–103, May 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:NPAb

[Ano82-193] Anonymous. New product applications: Single-board computer with 16-bit wordlength measures 13.2 by 22.8 cm, offers up to 32 kilobytes of EPROM. IEEE Spectrum, 19(2):82, February 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:NPAh

[Ano82-194] Anonymous. New product applications: Single-chip microprocessor comprises 16-bit CPU clock generator, two DMA channels, and programmable interrupt controller. IEEE Spectrum, 19(8):72–79, August 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:NPAc

[Ano82-195] Anonymous. New product applications: Waveform recorder samples input signals at rates to 20 MHz and stores 10bit codes in up to 16-kiloword memory. IEEE Spectrum, 19 (3):80–81, March 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:NIa

[Ano82-196] Anonymous. News from industry. IEEE Spectrum, 19(1): 119–120, January 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:NIb

[Ano82-197] Anonymous. News from industry. IEEE Spectrum, 19(2): 97–98, February 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:NIc

[Ano82-198] Anonymous. News from industry. IEEE Spectrum, 19 (3):95–96, March 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:NId

[Ano82-199] Anonymous. News from industry. IEEE Spectrum, 19 (4):95–96, April 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:NIe

[Ano82-200] Anonymous. News from industry. IEEE Spectrum, 19 (5):103–104, May 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:NIf

[Ano82-201] Anonymous. News from industry. IEEE Spectrum, 19 (6):89–90, June 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:NIg

[Ano82-202] Anonymous. News from industry. IEEE Spectrum, 19 (7):79–80, July 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:NIh

[Ano82-203] Anonymous. News from industry. IEEE Spectrum, 19 (8):79–80, August 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:NIi

[Ano82-204] Anonymous. News from industry. IEEE Spectrum, 19(9): 89–95, September 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:NIj

[Ano82-205] Anonymous. News from industry. IEEE Spectrum, 19 (10):138, October 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:NIk

[Ano82-206] Anonymous. News from industry. IEEE Spectrum, 19 (11):90, November 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:NIl

[Ano82-207] Anonymous. News from industry. IEEE Spectrum, 19(12):

74–75, December 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:NWa

[Ano82-208] Anonymous. News from Washington. IEEE Spectrum, 19 (1):24, January 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:NWb

[Ano82-209] Anonymous. News from Washington. IEEE Spectrum, 19 (2):16, February 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:NWc

[Ano82-210] Anonymous. News from Washington. IEEE Spectrum, 19 (3):26, March 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:NWd

[Ano82-211] Anonymous. News from Washington. IEEE Spectrum, 19 (4):16, April 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:NWe

[Ano82-212] Anonymous. News from Washington. IEEE Spectrum, 19 (5):30, May 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:NWf

[Ano82-213] Anonymous. News from Washington. IEEE Spectrum, 19 (6):14, June 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:NWg

[Ano82-214] Anonymous. News from Washington. IEEE Spectrum, 19 (7):20, July 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:NWh

[Ano82-215] Anonymous. News from Washington. IEEE Spectrum, 19 (8):12, August 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:NWi

[Ano82-216] Anonymous. News from Washington. IEEE Spectrum, 19 (9):24, September 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:NWj

[Ano82-217] Anonymous. News from Washington. IEEE Spectrum, 19 (10):16, October 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:NWk

[Ano82-218] Anonymous. News from Washington. IEEE Spectrum, 19 (11):18, November 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:NWl

[Ano82-219] Anonymous. News from Washington. IEEE Spectrum, 19 (12):12, December 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:QWD

[Ano82-220] Anonymous. Questionnaire:

What do you think? Spectrum solicits your advice. IEEE Spectrum, 19(1):76–77, January 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Ra

[Ano82-221] Anonymous. Reflections. IEEE Spectrum, 19(1):17, January 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Rb

[Ano82-222] Anonymous. Reflections. IEEE Spectrum, 19(3):18–19, March 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Rc

[Ano82-223] Anonymous. Reflections. IEEE Spectrum, 19(5):18–21, May 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Rd

[Ano82-224] Anonymous. Reflections. IEEE Spectrum, 19(7):14–15, July 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Re

[Ano82-225] Anonymous. Reflections. IEEE Spectrum, 19(9):18–19, September 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Rf

[Ano82-226] Anonymous. Reflections. IEEE Spectrum, 19(11):14–15, November 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:SIa

[Ano82-227] Anonymous. Scanning the Institute. IEEE Spectrum, 19 (1):26, January 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:SIb

[Ano82-228] Anonymous. Scanning the Institute. IEEE Spectrum, 19 (2):20, February 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:SIc

[Ano82-229] Anonymous. Scanning the Institute. IEEE Spectrum, 19(3):28, March 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:SId

[Ano82-230] Anonymous. Scanning the Institute. IEEE Spectrum, 19(4):17, April 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:SIe

[Ano82-231] Anonymous. Scanning the Institute. IEEE Spectrum, 19(5):32, May 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:SIf

[Ano82-232] Anonymous. Scanning the Institute. IEEE Spectrum, 19(6):16, June 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:SIg

[Ano82-233] Anonymous. Scanning the Institute. IEEE Spectrum, 19(7):22, July 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:SIh

[Ano82-234] Anonymous. Scanning the Institute. IEEE Spectrum, 19(8):14, August 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:SIi

[Ano82-235] Anonymous. Scanning the Institute. IEEE Spectrum, 19 (9):26, September 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:SIj

[Ano82-236] Anonymous. Scanning the Institute. IEEE Spectrum, 19 (10):22, October 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:SIk

[Ano82-237] Anonymous. Scanning the Institute. IEEE Spectrum, 19 (11):20, November 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:SIl

[Ano82-238] Anonymous. Scanning the Institute. IEEE Spectrum, 19 (12):14, December 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Sb

[Ano82-239] Anonymous. Speakout. IEEE Spectrum, 19(12):17–18, December 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:Sa

[Ano82-240] Anonymous. Staff. IEEE Spectrum, 19(10):2, October 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:SGA

[Ano82-241] Anonymous. Stop guessing at accuracy [advertisement]. IEEE Spectrum, 19(10):5–7, October 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:TSa

[Ano82-242] Anonymous. Technically speaking. IEEE Spectrum, 19(2): 25–26, February 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:TSb

[Ano82-243] Anonymous. Technically speaking. IEEE Spectrum, 19(4): 26–28, April 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:TSc

[Ano82-244] Anonymous. Technically speaking. IEEE Spectrum, 19(6): 20–22, June 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:TSd

[Ano82-245] Anonymous. Technically speaking. IEEE Spectrum, 19(8): 18–20, August 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:TSe

[Ano82-246] Anonymous. Technically speaking. IEEE Spectrum, 19(10): 30–32, October 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:PF

[Ano82-247] Anonymous. To probe further. IEEE Spectrum, 19(10): 112–113, October 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:WDY

[Ano82-248] Anonymous. What do you think?

Spectrum solicits your opinions. IEEE Spectrum, 19(6):65, June 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:WHC

[Ano82-249] Anonymous. Whatever happened to cybernetics? IEEE Spectrum, 19(4):24–25, April 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:WHF

[Ano82-250] Anonymous. Whatever happened to fluid logic? IEEE Spectrum, 19(12):16, December 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:WHM

[Ano82-251] Anonymous. Whatever happened to maglev? IEEE Spectrum, 19(8):17, August 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:WHP

[Ano82-252] Anonymous. Whatever happened to the picturephone? IEEE Spectrum, 19(2):24, February 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1982:WHW

[Ano82-253] Anonymous. Whatever happened to water-scale integration? IEEE Spectrum, 19 (6):18–19, June 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:IGA

[Ano83a] Anonymous. 1983 IEEE guide for authors. IEEE Spectrum, 20 (1):86–89, January 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:AAh

[Ano83b] Anonymous. About the authors. IEEE Spectrum, 20(9): 98–100, September 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:AAp

[Ano83c] Anonymous. About the authors. IEEE Spectrum, 20(11): 120, November 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Aa

[Ano83d] Anonymous. [advertisement].

IEEE Spectrum, 20(1):1–2, January 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Ab

[Ano83e] Anonymous. [advertisement].

IEEE Spectrum, 20(1):2, January 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Ac

[Ano83f] Anonymous. [advertisement].

IEEE Spectrum, 20(1):4–5, January 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Ad

[Ano83g] Anonymous. [advertisement]. IEEE Spectrum, 20(1):1–2, January 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Ae

[Ano83h] Anonymous. [advertisement]. IEEE Spectrum, 20(2):1–3, February 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Af

[Ano83i] Anonymous. [advertisement]. IEEE Spectrum, 20(2):2, February 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Ag

[Ano83j] Anonymous. [advertisement]. IEEE Spectrum, 20(2):4–7, February 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Ah

[Ano83k] Anonymous. [advertisement]. IEEE Spectrum, 20(2):1–2, February 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Ai

[Ano83l] Anonymous. [advertisement]. IEEE Spectrum, 20(3):1–3, March 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Aj

[Ano83m] Anonymous. [advertisement].

IEEE Spectrum, 20(3):2, March 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Ak

[Ano83n] Anonymous. [advertisement].

IEEE Spectrum, 20(3):4–9, March 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Al

[Ano83o] Anonymous. [advertisement].

IEEE Spectrum, 20(3):1–2, March 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Am

[Ano83p] Anonymous. [advertisement].

IEEE Spectrum, 20(4):1–3, April 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:An

[Ano83q] Anonymous. [advertisement].

IEEE Spectrum, 20(4):2, April 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Ao

[Ano83r] Anonymous. [advertisement].

IEEE Spectrum, 20(4):4–7, April 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Ap

[Ano83s] Anonymous. [advertisement]. IEEE Spectrum, 20(4):1–2, April 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Aq

[Ano83t] Anonymous. [advertisement]. IEEE Spectrum, 20(5):1–3, May 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Ar

[Ano83u] Anonymous. [advertisement]. IEEE Spectrum, 20(5):2, May 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:As

[Ano83v] Anonymous. [advertisement]. IEEE Spectrum, 20(5):4–7, May 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:At

[Ano83w] Anonymous. [advertisement]. IEEE Spectrum, 20(5):1–2, May 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Au

[Ano83x] Anonymous. [advertisement]. IEEE Spectrum, 20(6):1–2, June 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Av

[Ano83y] Anonymous. [advertisement].

IEEE Spectrum, 20(6):1–2, June 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Aw

[Ano83z] Anonymous. [advertisement].

IEEE Spectrum, 20(6):4–5, June 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Ax

[Ano83-27] Anonymous. [advertisement].

IEEE Spectrum, 20(6):1–2, June 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Ay

[Ano83-28] Anonymous. [advertisement].

IEEE Spectrum, 20(7):1–3, July 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Az

[Ano83-29] Anonymous. [advertisement].

IEEE Spectrum, 20(7):2, July 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Aaa

[Ano83-30] Anonymous. [advertisement].

IEEE Spectrum, 20(7):4–9, July 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Aab

[Ano83-31] Anonymous. [advertisement]. IEEE Spectrum, 20(7):1–2, July 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Aac

[Ano83-32] Anonymous. [advertisement]. IEEE Spectrum, 20(8):1–3, August 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Aad

[Ano83-33] Anonymous. [advertisement]. IEEE Spectrum, 20(8):2, August 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Aae

[Ano83-34] Anonymous. [advertisement]. IEEE Spectrum, 20(8):4–7, August 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Aag

[Ano83-35] Anonymous. [advertisement]. IEEE Spectrum, 20(8):1–2, August 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Aai

[Ano83-36] Anonymous. [advertisement]. IEEE Spectrum, 20(10):1–3, October 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Aaj

[Ano83-37] Anonymous. [advertisement].

IEEE Spectrum, 20(10):2, October 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Aak

[Ano83-38] Anonymous. [advertisement].

IEEE Spectrum, 20(10):4–7, October 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Aal

[Ano83-39] Anonymous. [advertisement].

IEEE Spectrum, 20(10):1–2, October 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Aam

[Ano83-40] Anonymous. [advertisement].

IEEE Spectrum, 20(11):1–3, November 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Aan

[Ano83-41] Anonymous. [advertisement].

IEEE Spectrum, 20(11):2, November 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Aao

[Ano83-42] Anonymous. [advertisement].

IEEE Spectrum, 20(11):4–7, November 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Aaq

[Ano83-43] Anonymous. [advertisement]. IEEE Spectrum, 20(11):1–2, November 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Aar

[Ano83-44] Anonymous. [advertisement]. IEEE Spectrum, 20(12):1–2, December 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Aas

[Ano83-45] Anonymous. [advertisement]. IEEE Spectrum, 20(12):1–2, December 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Aat

[Ano83-46] Anonymous. [advertisement]. IEEE Spectrum, 20(12):4–9, December 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Aau

[Ano83-47] Anonymous. [advertisement]. IEEE Spectrum, 20(12):1–2, December 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:APP

[Ano83-48] Anonymous. Award/83 Prize Paper and Service Awards [4 awards]. IEEE Spectrum, 20 (10):76–77, October 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Aaf

[Ano83-49] Anonymous. Awards/83. IEEE Spectrum, 20(8):74–75, August 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:BC

[Ano83-50] Anonymous. [back cover]. IEEE Spectrum, 20(9):c4, September 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:BAM

[Ano83-51] Anonymous. Best bist: Applications of microprocessors: Probing transmission lines. IEEE Spectrum, 20(1):14–15, January 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:BBAd

[Ano83-52] Anonymous. Best bits: Applications of microprocessors. IEEE Spectrum, 20(9):20–21, September 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:BBAe

[Ano83-53] Anonymous. Best bits: Applications of microprocessors. IEEE Spectrum, 20(11):26–27, November 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:BBAb

[Ano83-54] Anonymous. Best bits: Applications of microprocessors: a robot armed with micros. IEEE Spectrum, 20(5):18–21, May 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:BBAa

[Ano83-55] Anonymous. Best bits: Applications of microprocessors: Micros replace hydraulics. IEEE Spectrum, 20(3):17, March 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:BBAc

[Ano83-56] Anonymous. Best bits: Applications of microprocessors: Outguessing gases. IEEE Spectrum, 20(7):26–28, July 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:BRa

[Ano83-57] Anonymous. Book reviews. IEEE Spectrum, 20(1):98–111, January 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:BRb

[Ano83-58] Anonymous. Book reviews. IEEE Spectrum, 20(2):80–81, February 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:BRc

[Ano83-59] Anonymous. Book reviews. IEEE Spectrum, 20(3):82–92, March 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:BRd

[Ano83-60] Anonymous. Book reviews.

IEEE Spectrum, 20(4):83–95, April 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:BRe

[Ano83-61] Anonymous. Book reviews.

IEEE Spectrum, 20(5):114–127, May 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:BRf

[Ano83-62] Anonymous. Book reviews.

IEEE Spectrum, 20(6):70–71, June 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:BRg

[Ano83-63] Anonymous. Book reviews.

IEEE Spectrum, 20(7):17–20, July 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:BRh

[Ano83-64] Anonymous. Book reviews.

IEEE Spectrum, 20(8):14–17, August 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:BRi

[Ano83-65] Anonymous. Book reviews.

IEEE Spectrum, 20(9):9–16P, September 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:BRj

[Ano83-66] Anonymous. Book reviews. IEEE Spectrum, 20(10):14– 16, October 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:BRk

[Ano83-67] Anonymous. Book reviews. IEEE Spectrum, 20(11):16–21, November 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:BRl

[Ano83-68] Anonymous. Book reviews. IEEE Spectrum, 20(12):16–18, December 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Cb

[Ano83-69] Anonymous. Calendar. IEEE Spectrum, 20(1):6–9, January 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Cd

[Ano83-70] Anonymous. Calendar. IEEE Spectrum, 20(2):8–11, February 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Cf

[Ano83-71] Anonymous. Calendar. IEEE Spectrum, 20(3):10–13, March 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Ch

[Ano83-72] Anonymous. Calendar. IEEE Spectrum, 20(4):8–13, April 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Cj

[Ano83-73] Anonymous. Calendar. IEEE Spectrum, 20(5):8–16, May 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Cl

[Ano83-74] Anonymous. Calendar. IEEE Spectrum, 20(6):6–13, June 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Cn

[Ano83-75] Anonymous. Calendar. IEEE Spectrum, 20(7):74–75, July 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Cq

[Ano83-76] Anonymous. Calendar. IEEE Spectrum, 20(8):81–83, August 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Ct

[Ano83-77] Anonymous. Calendar. IEEE Spectrum, 20(9):104–107, September 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Cv

[Ano83-78] Anonymous. Calendar. IEEE Spectrum, 20(10):89–94, October 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Cx

[Ano83-79] Anonymous. Calendar. IEEE Spectrum, 20(11):132–135, November 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Caa

[Ano83-80] Anonymous. Calendar. IEEE Spectrum, 20(12):76–80, December 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:CSa

[Ano83-81] Anonymous. Coming in Spectrum. IEEE Spectrum, 20(1): 22–24, January 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:CSb

[Ano83-82] Anonymous. Coming in Spectrum. IEEE Spectrum, 20(2): 17–19, February 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:CSc

[Ano83-83] Anonymous. Coming in Spectrum. IEEE Spectrum, 20 (3):26–28, March 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:CSd

[Ano83-84] Anonymous. Coming in Spectrum. IEEE Spectrum, 20 (4):20–21, April 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:CSe

[Ano83-85] Anonymous. Coming in Spectrum. IEEE Spectrum, 20 (5):31–32, May 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:CSf

[Ano83-86] Anonymous. Coming in Spectrum. IEEE Spectrum, 20 (6):18–19, June 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:CSg

[Ano83-87] Anonymous. Coming in Spectrum. IEEE Spectrum, 20(7):88, July 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:CSh

[Ano83-88] Anonymous. Coming in Spectrum. IEEE Spectrum, 20(8):98, August 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:CSi

[Ano83-89] Anonymous. Coming in Spectrum. IEEE Spectrum, 20(9): 130, September 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:CSj

[Ano83-90] Anonymous. Coming in Spectrum. IEEE Spectrum, 20 (10):114, October 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:CSk

[Ano83-91] Anonymous. Coming in Spectrum. IEEE Spectrum, 20(11): 154, November 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:CSl

[Ano83-92] Anonymous. Coming in Spectrum. IEEE Spectrum, 20(12): 101–103, December 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Ca

[Ano83-93] Anonymous. Contents. IEEE Spectrum, 20(1):3, January 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Cc

[Ano83-94] Anonymous. Contents. IEEE Spectrum, 20(2):3, February 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Ce

[Ano83-95] Anonymous. Contents. IEEE Spectrum, 20(3):3, March 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Cg

[Ano83-96] Anonymous. Contents. IEEE Spectrum, 20(4):3, April 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Ci

[Ano83-97] Anonymous. Contents. IEEE Spectrum, 20(5):3, May 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Ck

[Ano83-98] Anonymous. Contents. IEEE Spectrum, 20(6):3, June 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Cm

[Ano83-99] Anonymous. Contents. IEEE Spectrum, 20(7):3, July 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Co

[Ano83-100] Anonymous. Contents. IEEE Spectrum, 20(8):3, August 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Cr

[Ano83-101] Anonymous. Contents. IEEE Spectrum, 20(9):3, September 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Cu

[Ano83-102] Anonymous. Contents. IEEE Spectrum, 20(10):3, October 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Cw

[Ano83-103] Anonymous. Contents. IEEE Spectrum, 20(11):3, November 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Cy

[Ano83-104] Anonymous. Contents. IEEE Spectrum, 20(12):3, December 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Cp

[Ano83-105] Anonymous. Continuum. IEEE Spectrum, 20(8):20–23, August 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Cz

[Ano83-106] Anonymous. Continuum. IEEE Spectrum, 20(12):22–23, December 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:DDA

[Ano83-107] Anonymous. Data-driven automation 8 to probe further: Computers in manufacturing. IEEE Spectrum, 20(5):95–99, May 1983. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:ETTa

[Ano83-108] Anonymous. EEs’ tools toys.

IEEE Spectrum, 20(1):18–19, January 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:ETTb

[Ano83-109] Anonymous. EEs’ tools toys.

IEEE Spectrum, 20(3):22–23, March 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:ETTc

[Ano83-110] Anonymous. EEs’ tools toys.

IEEE Spectrum, 20(5):26–29, May 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:ETTd

[Ano83-111] Anonymous. EEs’ tools toys.

IEEE Spectrum, 20(7):70–73, July 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:ETTe

[Ano83-112] Anonymous. EEs’ tools toys.

IEEE Spectrum, 20(8):76–80, August 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:ETTf

[Ano83-113] Anonymous. EEs’ tools toys.

IEEE Spectrum, 20(9):101–103, September 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:ETTg

[Ano83-114] Anonymous. EEs’ tools toys. IEEE Spectrum, 20(10):78– 85, October 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:ETTh

[Ano83-115] Anonymous. EEs’ tools toys. IEEE Spectrum, 20(11):124– 131, November 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:ETTi

[Ano83-116] Anonymous. EEs’ tools toys. IEEE Spectrum, 20(12):81–83, December 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:ERa

[Ano83-117] Anonymous. Energy report. IEEE Spectrum, 20(1): 20–21, January 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:ERb

[Ano83-118] Anonymous. Energy report. IEEE Spectrum, 20(2):16, February 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:ERc

[Ano83-119] Anonymous. Energy report. IEEE Spectrum, 20(3):24–25, March 1983. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:ERd

[Ano83-120] Anonymous. Energy report.

IEEE Spectrum, 20(4):17–19, April 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:ERe

[Ano83-121] Anonymous. Energy report.

IEEE Spectrum, 20(5):30, May 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:ERf

[Ano83-122] Anonymous. Energy report.

IEEE Spectrum, 20(6):17, June 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:EAL

[Ano83-123] Anonymous. The engineer al large. IEEE Spectrum, 20(4):24, April 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:ELa

[Ano83-124] Anonymous. The engineer at large. IEEE Spectrum, 20(2): 24–27, February 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:ELb

[Ano83-125] Anonymous. The engineer at large. IEEE Spectrum, 20 (6):22–25, June 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:ELc

[Ano83-126] Anonymous. The engineer at large. IEEE Spectrum, 20 (8):18–19, August 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:ELd

[Ano83-127] Anonymous. The engineer at large. IEEE Spectrum, 20(10): 17–19, October 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:ELe

[Ano83-128] Anonymous. The engineer at large. IEEE Spectrum, 20(12): 19–21, December 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Fa

[Ano83-129] Anonymous. Forum. IEEE Spectrum, 20(1):10–13, January 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Fb

[Ano83-130] Anonymous. Forum. IEEE Spectrum, 20(2):12–15, February 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Fc

[Ano83-131] Anonymous. Forum. IEEE Spectrum, 20(3):14–16, March 1983.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Fd

[Ano83-132] Anonymous. Forum. IEEE Spectrum, 20(4):14–16, April 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Fe

[Ano83-133] Anonymous. Forum. IEEE Spectrum, 20(5):17, May 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Ff

[Ano83-134] Anonymous. Forum. IEEE Spectrum, 20(6):14–16, June 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Fg

[Ano83-135] Anonymous. Forum. IEEE Spectrum, 20(7):10–15, July 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Fh

[Ano83-136] Anonymous. Forum. IEEE Spectrum, 20(8):8–11, August 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Fi

[Ano83-137] Anonymous. Forum. IEEE Spectrum, 20(9):6–8H, September 1983. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Fj

[Ano83-138] Anonymous. Forum. IEEE Spectrum, 20(10):8–13, October 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Fk

[Ano83-139] Anonymous. Forum. IEEE Spectrum, 20(11):8–15, November 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Fl

[Ano83-140] Anonymous. Forum. IEEE Spectrum, 20(12):10–15, December 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:FGF

[Ano83-141] Anonymous. The four generations of fiber-optic systems. IEEE Spectrum, 20(12):39–45, December 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:FCa

[Ano83-142] Anonymous. [front cover]. IEEE Spectrum, 20(9):c1, September 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:HHC

[Ano83-143] Anonymous. The hand-held computer with powerful connections [advertisement]. IEEE Spectrum, 20(9):5, September 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:IA

[Ano83-144] Anonymous. IEEE awards:

1983. IEEE Spectrum, 20 (4):66–67, April 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:IPPa

[Ano83-145] Anonymous. IEEE publications paper calls. IEEE Spectrum, 20(3):79, March 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:IPPb

[Ano83-146] Anonymous. IEEE publications paper calls. IEEE Spectrum, 20(4):76, April 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:ITCa

[Ano83-147] Anonymous. IEEE tables of contents. IEEE Spectrum, 20 (1):92–93, January 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:ITCb

[Ano83-148] Anonymous. IEEE tables of contents. IEEE Spectrum, 20 (2):74–75, February 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:ITCc

[Ano83-149] Anonymous. IEEE tables of contents. IEEE Spectrum, 20 (3):72–75, March 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:ITCd

[Ano83-150] Anonymous. IEEE tables of contents. IEEE Spectrum, 20 (4):76–79, April 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:ITCe

[Ano83-151] Anonymous. IEEE tables of contents. IEEE Spectrum, 20 (5):103–105, May 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:ITCf

[Ano83-152] Anonymous. IEEE tables of contents. IEEE Spectrum, 20 (6):64–67, June 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:ITCg

[Ano83-153] Anonymous. IEEE tables of contents. IEEE Spectrum, 20 (7):79–87, July 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:ITCh

[Ano83-154] Anonymous. IEEE tables of contents. IEEE Spectrum, 20 (8):86–88, August 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:ITCj

[Ano83-155] Anonymous. IEEE tables of contents. IEEE Spectrum, 20 (10):99–113, October 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:ITCk

[Ano83-156] Anonymous. IEEE tables of contents. IEEE Spectrum, 20(11): 141–153, November 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:ITCl

[Ano83-157] Anonymous. IEEE tables of contents. IEEE Spectrum, 20(12): 86–88, December 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:ITCi

[Ano83-158] Anonymous. IEEE tables of contents for current and future publications. IEEE Spectrum, 20(9): 118–129, September 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:IBC

[Ano83-159] Anonymous. [inside back cover]. IEEE Spectrum, 20 (9):c3, September 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:IFC

[Ano83-160] Anonymous. [inside front cover]. IEEE Spectrum, 20 (9):c2, September 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:LS

[Ano83-161] Anonymous. List of staff. IEEE Spectrum, 20(9):4, September 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:MPW

[Ano83-162] Anonymous. Meeting preview: Wescon ’83: VLSI in the spotlight: Rapidly maturing technology of very large-scale integration creates a prime attraction at California meeting. IEEE Spectrum, 20(11):121–123, November 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:MPa

[Ano83-163] Anonymous. Meeting previews. IEEE Spectrum, 20(1): 94–95, January 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:MPb

[Ano83-164] Anonymous. Meeting previews. IEEE Spectrum, 20(2): 76–79, February 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:MPc

[Ano83-165] Anonymous. Meeting previews. IEEE Spectrum, 20(3):76–79, March 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:MPd

[Ano83-166] Anonymous. Meeting previews. IEEE Spectrum, 20(4):70–75, April 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:MPe

[Ano83-167] Anonymous. Meeting previews.

IEEE Spectrum, 20(5):106–111, May 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:MPf

[Ano83-168] Anonymous. Meeting previews.

IEEE Spectrum, 20(6):68–69, June 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:MSE

[Ano83-169] Anonymous. The most successful engine transplant in automotive history [advertisement]. IEEE Spectrum, 20(9):1, September 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:NRIa

[Ano83-170] Anonymous. New and recent IEEE publications. IEEE Spectrum, 20(1):90–91, January 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:NRIb

[Ano83-171] Anonymous. New and recent IEEE publications. IEEE Spectrum, 20(2):72–73, February 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:NRIc

[Ano83-172] Anonymous. New and recent IEEE publications. IEEE Spectrum, 20(3):70–71, March 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:NRId

[Ano83-173] Anonymous. New and recent IEEE publications. IEEE Spectrum, 20(4):68–69, April 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:NRIe

[Ano83-174] Anonymous. New and recent IEEE publications. IEEE Spectrum, 20(5):100–102, May 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:NRIf

[Ano83-175] Anonymous. New and recent IEEE publications. IEEE Spectrum, 20(6):62–63, June 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:NRIg

[Ano83-176] Anonymous. New and recent IEEE publications. IEEE Spectrum, 20(7):76–78, July 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:NRIh

[Ano83-177] Anonymous. New and recent IEEE publications. IEEE Spectrum, 20(8):89–97, August 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:NRIi

[Ano83-178] Anonymous. New and recent IEEE publications. IEEE Spectrum, 20(9):116–117, September 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:NRIj

[Ano83-179] Anonymous. New and recent IEEE publications. IEEE Spectrum, 20(10):87–88, October 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:NRIk

[Ano83-180] Anonymous. New and recent IEEE publications. IEEE Spectrum, 20(11):139–140, November 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:NRIl

[Ano83-181] Anonymous. New and recent IEEE publications. IEEE Spectrum, 20(12):84–85, December 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:NPAb

[Ano83-182] Anonymous. New product applications. IEEE Spectrum, 20 (2):80–97, February 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:NPAe

[Ano83-183] Anonymous. New product applications. IEEE Spectrum, 20 (5):112–113, May 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:NPAd

[Ano83-184] Anonymous. New product applications: Ethernet front-end processor includes 16-bit 8088, 64- or 128-kilobyte RAM, and 8-kilobyte PROM. IEEE Spectrum, 20(4):80–82, April 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:NPAc

[Ano83-185] Anonymous. New product applications: Measurement and control system for microcomputers accepts up to 12 I/O modules, uses Apple-compatible language. IEEE Spectrum, 20 (3):80–81, March 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:NPAf

[Ano83-186] Anonymous. New product applications: Universal logic-design simulator software validates IC tests at gate and functional levels. IEEE Spectrum, 20(6):72–79, June 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:NPAa

[Ano83-187] Anonymous. New product applications: VLSI development system uses two 12-MHz 32-bit processors, offers up to 4 megabytes of physical memory in Multibus card cage. IEEE Spectrum, 20 (1):96–97, January 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:NIa

[Ano83-188] Anonymous. News from industry. IEEE Spectrum, 20(1): 112–114, January 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:NIb

[Ano83-189] Anonymous. News from industry. IEEE Spectrum, 20 (2):98, February 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:NIc

[Ano83-190] Anonymous. News from industry. IEEE Spectrum, 20 (3):92–96, March 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:NId

[Ano83-191] Anonymous. News from industry. IEEE Spectrum, 20(4):96, April 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:NIe

[Ano83-192] Anonymous. News from industry. IEEE Spectrum, 20 (5):128–130, May 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:NIf

[Ano83-193] Anonymous. News from industry. IEEE Spectrum, 20(6):80, June 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:NWa

[Ano83-194] Anonymous. News from Washington. IEEE Spectrum, 20 (1):20, January 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:NWb

[Ano83-195] Anonymous. News from Washington. IEEE Spectrum, 20 (2):16, February 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:NWc

[Ano83-196] Anonymous. News from Washington. IEEE Spectrum, 20 (3):24, March 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:NWd

[Ano83-197] Anonymous. News from Washington. IEEE Spectrum, 20 (4):17, April 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:NWe

[Ano83-198] Anonymous. News from Washington. IEEE Spectrum, 20 (5):30, May 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:NWf

[Ano83-199] Anonymous. News from Washington. IEEE Spectrum, 20 (6):17, June 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:PIa

[Ano83-200] Anonymous. Papers are invited. IEEE Spectrum, 20 (8):84–85, August 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:PIb

[Ano83-201] Anonymous. Papers are invited. IEEE Spectrum, 20(10): 95–98, October 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:PIc

[Ano83-202] Anonymous. Papers are invited. IEEE Spectrum, 20(11): 136–138, November 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:PId

[Ano83-203] Anonymous. Papers are invited. IEEE Spectrum, 20(12): 88–100, December 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:PIV

[Ano83-204] Anonymous. Papers are invited [various calls for papers]. IEEE Spectrum, 20(9): 108–115, September 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:PNRa

[Ano83-205] Anonymous. Program notes [reports of new computer programs]. IEEE Spectrum, 20 (8):24–30, August 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:PNRb

[Ano83-206] Anonymous. Program notes [reports of new computer programs]. IEEE Spectrum, 20 (10):20–24, October 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:PNRc

[Ano83-207] Anonymous. Program notes [reports of new computer programs]. IEEE Spectrum, 20(12): 24–25, December 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Ra

[Ano83-208] Anonymous. Reflections. IEEE Spectrum, 20(1):16, January 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Rb

[Ano83-209] Anonymous. Reflections. IEEE Spectrum, 20(3):18–19, March 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Rc

[Ano83-210] Anonymous. Reflections. IEEE Spectrum, 20(5):22–23, May 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Rd

[Ano83-211] Anonymous. Reflections. IEEE Spectrum, 20(7):16, July 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Re

[Ano83-212] Anonymous. Reflections. IEEE Spectrum, 20(11):22–23, November 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:SIa

[Ano83-213] Anonymous. Scanning the Institute. IEEE Spectrum, 20 (1):22, January 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:SIb

[Ano83-214] Anonymous. Scanning the Institute. IEEE Spectrum, 20 (2):17, February 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:SIc

[Ano83-215] Anonymous. Scanning the Institute. IEEE Spectrum, 20(3):26, March 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:SId

[Ano83-216] Anonymous. Scanning the Institute. IEEE Spectrum, 20(4):20, April 1983. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:SIe

[Ano83-217] Anonymous. Scanning the Institute. IEEE Spectrum, 20(5):31, May 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:SIf

[Ano83-218] Anonymous. Scanning the Institute. IEEE Spectrum, 20(6):18, June 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:SIg

[Ano83-219] Anonymous. Scanning the Institute. IEEE Spectrum, 20(7):88, July 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:SIh

[Ano83-220] Anonymous. Scanning the Institute. IEEE Spectrum, 20(8):98, August 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:SIi

[Ano83-221] Anonymous. Scanning the Institute. IEEE Spectrum, 20 (9):130, September 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:SIj

[Ano83-222] Anonymous. Scanning the Institute. IEEE Spectrum, 20 (10):114, October 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:SIk

[Ano83-223] Anonymous. Scanning the Institute. IEEE Spectrum, 20 (11):154, November 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:SIl

[Ano83-224] Anonymous. Scanning the Institute. IEEE Spectrum, 20 (12):101, December 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Sa

[Ano83-225] Anonymous. Speakout. IEEE Spectrum, 20(2):28–30, February 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Sb

[Ano83-226] Anonymous. Speakout. IEEE Spectrum, 20(4):25, April 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Sc

[Ano83-227] Anonymous. Speakout. IEEE Spectrum, 20(6):26–28, June 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Sd

[Ano83-228] Anonymous. Speakout. IEEE Spectrum, 20(8):12–13, August 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:SOa

[Ano83-229] Anonymous. Spin-offs. IEEE Spectrum, 20(7):25, July 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:SOb

[Ano83-230] Anonymous. Spin-offs. IEEE Spectrum, 20(9):22–24, September 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:SOc

[Ano83-231] Anonymous. Spin-offs. IEEE Spectrum, 20(11):28–29, November 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:Se

[Ano83-232] Anonymous. Staff. IEEE Spectrum, 20(9):2, September 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:TSa

[Ano83-233] Anonymous. Technically speaking. IEEE Spectrum, 20(1): 17, January 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:TSb

[Ano83-234] Anonymous. Technically speaking. IEEE Spectrum, 20(3):

20–21, March 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:TSc

[Ano83-235] Anonymous. Technically speaking. IEEE Spectrum, 20(5): 24–25, May 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:TSd

[Ano83-236] Anonymous. Technically speaking. IEEE Spectrum, 20(7):21, July 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:TSe

[Ano83-237] Anonymous. Technically speaking. IEEE Spectrum, 20(9): 18–19, September 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:TSf

[Ano83-238] Anonymous. Technically speaking. IEEE Spectrum, 20(11): 24–25, November 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:TPF

[Ano83-239] Anonymous. Technology ’83 to probe further: Minis and mainframes. IEEE Spectrum, 20 (1):84–85, January 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:PF

[Ano83-240] Anonymous. To probe further. IEEE Spectrum, 20(9):

96–97, September 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:PFQ

[Ano83-241] Anonymous. To probe further: The quest. IEEE Spectrum, 20(11):118–119, November 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:V

[Ano83-242] Anonymous. Video. IEEE Spectrum, 20(12):26–28, December 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:WHD

[Ano83-243] Anonymous. Whatever happened to distributed computation? IEEE Spectrum, 20 (6):20–21, June 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:WHE

[Ano83-244] Anonymous. Whatever happened to energy parts? IEEE Spectrum, 20(2):20–23, February 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:WHF

[Ano83-245] Anonymous. Whatever happened to four-channel stereo? IEEE Spectrum, 20(11):30–32, November 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:WHSb

[Ano83-246] Anonymous. Whatever happened to solar-power satellites? IEEE Spectrum, 20(7):22–24, July 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:WHSa

[Ano83-247] Anonymous. Whatever happened to the search for life in space? IEEE Spectrum, 20 (4):22–23, April 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1983:WTC

[Ano83-248] Anonymous. What’s up there?

[chart of objects in Earth orbit]. IEEE Spectrum, 20(9): 48–49, September 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:IGA

[Ano84a] Anonymous. 1984 IEEE guide for authors. IEEE Spectrum, 21 (1):99–102, January 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:AAa

[Ano84b] Anonymous. About the authors.

IEEE Spectrum, 21(6):135–137, June 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Aa

[Ano84c] Anonymous. [advertisement].

IEEE Spectrum, 21(1):1–3, January 1984. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Ab

[Ano84d] Anonymous. [advertisement]. IEEE Spectrum, 21(1):2, January 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Ac

[Ano84e] Anonymous. [advertisement]. IEEE Spectrum, 21(1):4–7, January 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Ad

[Ano84f] Anonymous. [advertisement]. IEEE Spectrum, 21(1):1–4, January 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Ae

[Ano84g] Anonymous. [advertisement]. IEEE Spectrum, 21(2):1, February 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Af

[Ano84h] Anonymous. [advertisement]. IEEE Spectrum, 21(2):2, February 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Ag

[Ano84i] Anonymous. [advertisement]. IEEE Spectrum, 21(2):4–7, February 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Ah

[Ano84j] Anonymous. [advertisement].

IEEE Spectrum, 21(2):1–5, February 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Ai

[Ano84k] Anonymous. [advertisement].

IEEE Spectrum, 21(3):1–2, March 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Aj

[Ano84l] Anonymous. [advertisement].

IEEE Spectrum, 21(3):2, March 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Ak

[Ano84m] Anonymous. [advertisement].

IEEE Spectrum, 21(3):4–5, March 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Al

[Ano84n] Anonymous. [advertisement].

IEEE Spectrum, 21(3):1–2, March 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Am

[Ano84o] Anonymous. [advertisement].

IEEE Spectrum, 21(4):1, April 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:An

[Ano84p] Anonymous. [advertisement]. IEEE Spectrum, 21(4):2, April 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Ao

[Ano84q] Anonymous. [advertisement]. IEEE Spectrum, 21(4):4–7, April 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Ap

[Ano84r] Anonymous. [advertisement]. IEEE Spectrum, 21(4):1–3, April 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Aq

[Ano84s] Anonymous. [advertisement]. IEEE Spectrum, 21(5):1, May 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Ar

[Ano84t] Anonymous. [advertisement]. IEEE Spectrum, 21(5):2, May 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:As

[Ano84u] Anonymous. [advertisement]. IEEE Spectrum, 21(5):4, May 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:At

[Ano84v] Anonymous. [advertisement].

IEEE Spectrum, 21(5):1–2, May 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Au

[Ano84w] Anonymous. [advertisement].

IEEE Spectrum, 21(6):1, June 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Av

[Ano84x] Anonymous. [advertisement].

IEEE Spectrum, 21(6):4–7, June 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Aw

[Ano84y] Anonymous. [advertisement].

IEEE Spectrum, 21(6):1–4, June 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Ax

[Ano84z] Anonymous. [advertisement].

IEEE Spectrum, 21(7):1–3, July 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Ay

[Ano84-27] Anonymous. [advertisement].

IEEE Spectrum, 21(7):2, July 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Az

[Ano84-28] Anonymous. [advertisement]. IEEE Spectrum, 21(7):4–7, July 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Aab

[Ano84-29] Anonymous. [advertisement]. IEEE Spectrum, 21(7):1–2, July 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Aac

[Ano84-30] Anonymous. [advertisement]. IEEE Spectrum, 21(8):1, August 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Aad

[Ano84-31] Anonymous. [advertisement]. IEEE Spectrum, 21(8):1–2, August 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Aae

[Ano84-32] Anonymous. [advertisement]. IEEE Spectrum, 21(8):4, August 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Aaf

[Ano84-33] Anonymous. [advertisement]. IEEE Spectrum, 21(8):1–2, August 1984. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Aag

[Ano84-34] Anonymous. [advertisement].

IEEE Spectrum, 21(9):1, September 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Aah

[Ano84-35] Anonymous. [advertisement].

IEEE Spectrum, 21(9):2, September 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Aai

[Ano84-36] Anonymous. [advertisement].

IEEE Spectrum, 21(9):4–7, September 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Aaj

[Ano84-37] Anonymous. [advertisement].

IEEE Spectrum, 21(9):1–2, September 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Aak

[Ano84-38] Anonymous. [advertisement].

IEEE Spectrum, 21(10):1, October 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Aal

[Ano84-39] Anonymous. [advertisement].

IEEE Spectrum, 21(10):2, October 1984. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Aam

[Ano84-40] Anonymous. [advertisement]. IEEE Spectrum, 21(10):4–7, October 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Aan

[Ano84-41] Anonymous. [advertisement]. IEEE Spectrum, 21(10):1–3, October 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Aao

[Ano84-42] Anonymous. [advertisement]. IEEE Spectrum, 21(11):1, November 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Aap

[Ano84-43] Anonymous. [advertisement]. IEEE Spectrum, 21(11):4–7, November 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Aaq

[Ano84-44] Anonymous. [advertisement]. IEEE Spectrum, 21(11):4–7, November 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Aar

[Ano84-45] Anonymous. [advertisement]. IEEE Spectrum, 21(11):1–3, November 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Aas

[Ano84-46] Anonymous. [advertisement].

IEEE Spectrum, 21(12):1, December 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Aat

[Ano84-47] Anonymous. [advertisement].

IEEE Spectrum, 21(12):2, December 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Aau

[Ano84-48] Anonymous. [advertisement].

IEEE Spectrum, 21(12):4–7, December 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Aav

[Ano84-49] Anonymous. [advertisement].

IEEE Spectrum, 21(12):1–4, December 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:AEI

[Ano84-50] Anonymous. Art and electrotechnology: The IEEE centennial art contest winners. IEEE Spectrum, 21(2):56–57, February 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:AIF

[Ano84-51] Anonymous. Awards: IEEE Field Award winners. IEEE Spectrum, 21(8):68–69, August 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:APP

[Ano84-52] Anonymous. Awards/84: Prize Papers and Services Awards. IEEE Spectrum, 21(10):80– 80Q, October 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:BER

[Ano84-53] Anonymous. Before electronics: Room for error: Computer graphics has revolutionized television studios. IEEE Spectrum, 21(10):70–72, October 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:BBAa

[Ano84-54] Anonymous. Best bits: Applications of microprocessors. IEEE Spectrum, 21(1):26–27, January 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:BBAb

[Ano84-55] Anonymous. Best bits: Applications of microprocessors. IEEE Spectrum, 21(3):20–23, March 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:BBAc

[Ano84-56] Anonymous. Best bits: Applications of microprocessors. IEEE Spectrum, 21(5):22–24, May 1984. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:BBAd

[Ano84-57] Anonymous. Best bits: Applications of microprocessors. IEEE Spectrum, 21(7):22–24, July 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:BBAe

[Ano84-58] Anonymous. Best bits: Applications of microprocessors. IEEE Spectrum, 21(9):24–26, September 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:BBAf

[Ano84-59] Anonymous. Best bits: Applications of microprocessors. IEEE Spectrum, 21(11):32, November 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:BRa

[Ano84-60] Anonymous. Book reviews.

IEEE Spectrum, 21(1):11–17, January 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:BRb

[Ano84-61] Anonymous. Book reviews.

IEEE Spectrum, 21(2):14–16, February 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:BRc

[Ano84-62] Anonymous. Book reviews. IEEE Spectrum, 21(3):10–13, March 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:BRd

[Ano84-63] Anonymous. Book reviews. IEEE Spectrum, 21(4):10–15, April 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:BRe

[Ano84-64] Anonymous. Book reviews. IEEE Spectrum, 21(5):14–15, May 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:BRf

[Ano84-65] Anonymous. Book reviews. IEEE Spectrum, 21(6):12, June 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:BRg

[Ano84-66] Anonymous. Book reviews. IEEE Spectrum, 21(7):12–15, July 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:BRh

[Ano84-67] Anonymous. Book reviews. IEEE Spectrum, 21(8):12–15, August 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:BRi

[Ano84-68] Anonymous. Book reviews.

IEEE Spectrum, 21(9):17–21, September 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:BRj

[Ano84-69] Anonymous. Book reviews.

IEEE Spectrum, 21(10):17– 21, October 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:BRk

[Ano84-70] Anonymous. Book reviews.

IEEE Spectrum, 21(11):17–23, November 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:BRl

[Ano84-71] Anonymous. Book reviews.

IEEE Spectrum, 21(12):17–21, December 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:ISA

[Ano84-72] Anonymous. IEEE Spectrum annual index. IEEE Spectrum, 21(12):110–121, December 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Cb

[Ano84-73] Anonymous. Calendar. IEEE Spectrum, 21(1):108–115, January 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Cd

[Ano84-74] Anonymous. Calendar. IEEE Spectrum, 21(2):87–88NN, February 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Cf

[Ano84-75] Anonymous. Calendar. IEEE Spectrum, 21(3):78–81, March 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Ci

[Ano84-76] Anonymous. Calendar. IEEE Spectrum, 21(4):80–83, April 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Ck

[Ano84-77] Anonymous. Calendar. IEEE Spectrum, 21(5):92–97, May 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Cn

[Ano84-78] Anonymous. Calendar. IEEE Spectrum, 21(6):142–151, June 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Cp

[Ano84-79] Anonymous. Calendar. IEEE Spectrum, 21(7):73–79, July 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Ct

[Ano84-80] Anonymous. Calendar. IEEE Spectrum, 21(8):74–75, August 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Cv

[Ano84-81] Anonymous. Calendar. IEEE Spectrum, 21(9):103–106QQ, September 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Cy

[Ano84-82] Anonymous. Calendar. IEEE Spectrum, 21(10):85–93, October 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Caa

[Ano84-83] Anonymous. Calendar. IEEE Spectrum, 21(11):126–129, November 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Cad

[Ano84-84] Anonymous. Calendar. IEEE Spectrum, 21(12):82–84, December 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:CBI

[Ano84-85] Anonymous. The cat battery: Ill-starred and shortlived. IEEE Spectrum, 21(8): 65–67, August 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:CHF

[Ano84-86] Anonymous. Centennial hall of fame: Today’s electrical engineers select the major contributors — technical, industrial, and academic — to the profession and society. IEEE Spectrum, 21(4):64–66, April 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:CSa

[Ano84-87] Anonymous. Coming in Spectrum. IEEE Spectrum, 21(1): 136–138, January 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:CSb

[Ano84-88] Anonymous. Coming in Spectrum. IEEE Spectrum, 21 (2):112, February 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:CSc

[Ano84-89] Anonymous. Coming in Spectrum. IEEE Spectrum, 21(3): 106–108, March 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:CSd

[Ano84-90] Anonymous. Coming in Spectrum. IEEE Spectrum, 21(4): 105–109, April 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:CSe

[Ano84-91] Anonymous. Coming in Spectrum. IEEE Spectrum, 21 (5):114–116, May 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:CSf

[Ano84-92] Anonymous. Coming in Spectrum. IEEE Spectrum, 21(6):168, June 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:CSg

[Ano84-93] Anonymous. Coming in Spectrum. IEEE Spectrum, 21 (7):96–98, July 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:CSh

[Ano84-94] Anonymous. Coming in Spectrum. IEEE Spectrum, 21(8):90, August 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:CSi

[Ano84-95] Anonymous. Coming in Spectrum. IEEE Spectrum, 21(9): 130–132, September 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:CSj

[Ano84-96] Anonymous. Coming in Spectrum. IEEE Spectrum, 21(10): 112–113, October 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:CSk

[Ano84-97] Anonymous. Coming in Spectrum. IEEE Spectrum, 21(11):

154–155, November 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:CSl

[Ano84-98] Anonymous. Coming in Spectrum. IEEE Spectrum, 21(12): 109, December 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Ca

[Ano84-99] Anonymous. Contents. IEEE Spectrum, 21(1):3, January 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Cc

[Ano84-100] Anonymous. Contents. IEEE Spectrum, 21(2):3, February 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Ce

[Ano84-101] Anonymous. Contents. IEEE Spectrum, 21(3):3, March 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Cg

[Ano84-102] Anonymous. Contents. IEEE Spectrum, 21(4):3, April 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Cj

[Ano84-103] Anonymous. Contents. IEEE Spectrum, 21(5):3, May 1984.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Cl

[Ano84-104] Anonymous. Contents. IEEE Spectrum, 21(6):2–3, June 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Co

[Ano84-105] Anonymous. Contents. IEEE Spectrum, 21(7):3, July 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Cq

[Ano84-106] Anonymous. Contents. IEEE Spectrum, 21(8):3, August 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Cu

[Ano84-107] Anonymous. Contents. IEEE Spectrum, 21(9):2–3, September 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Cw

[Ano84-108] Anonymous. Contents. IEEE Spectrum, 21(10):3, October 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Cz

[Ano84-109] Anonymous. Contents. IEEE Spectrum, 21(11):2–3, November 1984. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Cab

[Ano84-110] Anonymous. Contents. IEEE Spectrum, 21(12):3, December 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Ch

[Ano84-111] Anonymous. Continuum. IEEE Spectrum, 21(4):20–21, April 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Cm

[Ano84-112] Anonymous. Continuum. IEEE Spectrum, 21(6):16–16G, June 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Cr

[Ano84-113] Anonymous. Continuum. IEEE Spectrum, 21(8):17, August 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Cx

[Ano84-114] Anonymous. Continuum. IEEE Spectrum, 21(10):22, October 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Cac

[Ano84-115] Anonymous. Continuum. IEEE Spectrum, 21(12):23–24, December 1984. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:ETTa

[Ano84-116] Anonymous. EEs’ tools toys.

IEEE Spectrum, 21(1):103– 107, January 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:ETTb

[Ano84-117] Anonymous. EEs’ tools toys.

IEEE Spectrum, 21(2):91–93, February 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:ETTc

[Ano84-118] Anonymous. EEs’ tools toys.

IEEE Spectrum, 21(3):74–77, March 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:ETTd

[Ano84-119] Anonymous. EEs’ tools toys.

IEEE Spectrum, 21(4):76–79, April 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:ETTe

[Ano84-120] Anonymous. EEs’ tools toys.

IEEE Spectrum, 21(5):86–91, May 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:ETTf

[Ano84-121] Anonymous. EEs’ tools toys.

IEEE Spectrum, 21(6):138–141, June 1984. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:ETTg

[Ano84-122] Anonymous. EEs’ tools toys. IEEE Spectrum, 21(7):69–72, July 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:ETTh

[Ano84-123] Anonymous. EEs’ tools toys. IEEE Spectrum, 21(8):70–73, August 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:ETTi

[Ano84-124] Anonymous. EEs’ tools toys. IEEE Spectrum, 21(9):96–102, September 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:ETTj

[Ano84-125] Anonymous. EEs’ tools toys. IEEE Spectrum, 21(10):82– 84, October 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:ETTk

[Ano84-126] Anonymous. EEs’ tools toys. IEEE Spectrum, 21(11):123– 125, November 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:ETTl

[Ano84-127] Anonymous. EEs’ tools toys. IEEE Spectrum, 21(12):80–81, December 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:ELa

[Ano84-128] Anonymous. The engineer at large. IEEE Spectrum, 21 (2):17, February 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:ELb

[Ano84-129] Anonymous. The engineer at large. IEEE Spectrum, 21(4):16, April 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:ELc

[Ano84-130] Anonymous. The engineer at large. IEEE Spectrum, 21 (6):13–14, June 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:ELd

[Ano84-131] Anonymous. The engineer at large. IEEE Spectrum, 21(8): 16–16O, August 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:ELe

[Ano84-132] Anonymous. The engineer at large. IEEE Spectrum, 21(10): 16–16P, October 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:ELf

[Ano84-133] Anonymous. The engineer at large. IEEE Spectrum, 21 (12):22, December 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Fa

[Ano84-134] Anonymous. Forum. IEEE Spectrum, 21(1):8–10, January 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Fb

[Ano84-135] Anonymous. Forum. IEEE Spectrum, 21(2):8–11, February 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Fc

[Ano84-136] Anonymous. Forum. IEEE Spectrum, 21(3):6–9, March 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Fd

[Ano84-137] Anonymous. Forum. IEEE Spectrum, 21(4):8, April 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Fe

[Ano84-138] Anonymous. Forum. IEEE Spectrum, 21(5):5–11, May 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Ff

[Ano84-139] Anonymous. Forum. IEEE Spectrum, 21(6):8–9, June 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Fg

[Ano84-140] Anonymous. Forum. IEEE Spectrum, 21(7):8–9, July 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Fh

[Ano84-141] Anonymous. Forum. IEEE Spectrum, 21(8):5–8, August 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Fi

[Ano84-142] Anonymous. Forum. IEEE Spectrum, 21(9):8–13, September 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Fj

[Ano84-143] Anonymous. Forum. IEEE Spectrum, 21(10):9–15, October 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Fk

[Ano84-144] Anonymous. Forum. IEEE Spectrum, 21(11):8–15, November 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Fl

[Ano84-145] Anonymous. Forum. IEEE Spectrum, 21(12):9–16, December 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:GDM

[Ano84-146] Anonymous. Gems from the Deutsches Museum: Noteworthy historical artifacts from this Munich showplace are described. IEEE Spectrum, 21(12): 72–73, December 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:GSM

[Ano84-147] Anonymous. Gems from the Science Museum: Five electrotechnological artifacts are selected as representative of the outstanding historical collection of London’s Science Museum. IEEE Spectrum, 21(10):76–77, October 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:GSC

[Ano84-148] Anonymous. Gems from the Smithsonian: From its collection of more than 17 000 historical electrical and electronic artifacts, the curator of the Smithsonian’s electrical section selected these as seminal. IEEE Spectrum, 21 (9):81–83, September 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:IAM

[Ano84-149] Anonymous. IEEE awards: 1984 major medalists. IEEE Spectrum, 21(4):2, April 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:ITCa

[Ano84-150] Anonymous. IEEE tables of contents. IEEE Spectrum, 21(1): 121–135, January 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:ITCb

[Ano84-151] Anonymous. IEEE tables of contents. IEEE Spectrum, 21 (2):97–98, February 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:ITCc

[Ano84-152] Anonymous. IEEE tables of contents. IEEE Spectrum, 21 (3):89–105, March 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:ITCd

[Ano84-153] Anonymous. IEEE tables of contents. IEEE Spectrum, 21 (4):92–104, April 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:ITCe

[Ano84-154] Anonymous. IEEE tables of contents. IEEE Spectrum, 21 (5):103–113, May 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:ITCf

[Ano84-155] Anonymous. IEEE tables of contents. IEEE Spectrum, 21(6):156, June 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:ITCg

[Ano84-156] Anonymous. IEEE tables of contents. IEEE Spectrum, 21 (7):84–95, July 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:ITCh

[Ano84-157] Anonymous. IEEE tables of contents. IEEE Spectrum, 21 (8):76–77, August 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:ITCi

[Ano84-158] Anonymous. IEEE tables of contents. IEEE Spectrum, 21(9): 110–113, September 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:ITCj

[Ano84-159] Anonymous. IEEE tables of contents. IEEE Spectrum, 21 (10):96–99, October 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:ITCk

[Ano84-160] Anonymous. IEEE tables of contents. IEEE Spectrum, 21(11): 134–137, November 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:ITCl

[Ano84-161] Anonymous. IEEE tables of contents. IEEE Spectrum, 21(12): 85–88FF, December 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Ia

[Ano84-162] Anonymous. Innovations. IEEE Spectrum, 21(2):18–21, February 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Ib

[Ano84-163] Anonymous. Innovations. IEEE Spectrum, 21(4):17, April 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Ic

[Ano84-164] Anonymous. Innovations. IEEE Spectrum, 21(6):17–19, June 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Id

[Ano84-165] Anonymous. Innovations. IEEE Spectrum, 21(8):18–19, August 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Ie

[Ano84-166] Anonymous. Innovations. IEEE Spectrum, 21(10):23–25, October 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:If

[Ano84-167] Anonymous. Innovations. IEEE Spectrum, 21(12):25–26, December 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:MPN

[Ano84-168] Anonymous. The machine as partner of the new professional: Today’s instruments are providing more data than the mind can process; expert systems are coming to the rescue but may soon usurp human roles. IEEE Spectrum, 21(6):28–31, June 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:NRIa

[Ano84-169] Anonymous. New and recent IEEE publications. IEEE Spectrum, 21(1):119–120, January 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:NRIb

[Ano84-170] Anonymous. New and recent IEEE publications. IEEE Spectrum, 21(2):99–111, February 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:NRIc

[Ano84-171] Anonymous. New and recent IEEE publications. IEEE Spectrum, 21(3):86–88, March 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:NRId

[Ano84-172] Anonymous. New and recent IEEE publications. IEEE Spectrum, 21(4):84–87, April 1984.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:NRIe

[Ano84-173] Anonymous. New and recent IEEE publications. IEEE Spectrum, 21(5):101–102, May 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:NRIf

[Ano84-174] Anonymous. New and recent IEEE publications. IEEE Spectrum, 21(6):158–167, June 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:NRIg

[Ano84-175] Anonymous. New and recent IEEE publications. IEEE Spectrum, 21(7):83–84, July 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:NRIh

[Ano84-176] Anonymous. New and recent IEEE publications. IEEE Spectrum, 21(8):78–79, August 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:NRIi

[Ano84-177] Anonymous. New and recent IEEE publications. IEEE Spectrum, 21(9):118–129, September 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:NRIj

[Ano84-178] Anonymous. New and recent IEEE publications. IEEE Spectrum, 21(10):94–95, October 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:NRIk

[Ano84-179] Anonymous. New and recent IEEE publications. IEEE Spectrum, 21(11):138–153, November 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:NRIl

[Ano84-180] Anonymous. New and recent IEEE publications. IEEE Spectrum, 21(12):92–108, December 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:PIa

[Ano84-181] Anonymous. Papers are invited. IEEE Spectrum, 21(1): 116–118, January 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:PIb

[Ano84-182] Anonymous. Papers are invited. IEEE Spectrum, 21(2): 94–96, February 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:PIc

[Ano84-183] Anonymous. Papers are invited. IEEE Spectrum, 21 (3):82–85, March 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:PId

[Ano84-184] Anonymous. Papers are invited. IEEE Spectrum, 21 (4):88–91, April 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:PIe

[Ano84-185] Anonymous. Papers are invited.

IEEE Spectrum, 21(5):98–98VV, May 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:PIf

[Ano84-186] Anonymous. Papers are invited.

IEEE Spectrum, 21(6):152, June 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:PIg

[Ano84-187] Anonymous. Papers are invited.

IEEE Spectrum, 21(7):80–80VV, July 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:PIh

[Ano84-188] Anonymous. Papers are invited. IEEE Spectrum, 21 (8):80–89, August 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:PIi

[Ano84-189] Anonymous. Papers are invited. IEEE Spectrum, 21(9):

110–113, September 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:PIj

[Ano84-190] Anonymous. Papers are invited. IEEE Spectrum, 21(10): 100–111, October 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:PIk

[Ano84-191] Anonymous. Papers are invited. IEEE Spectrum, 21(11): 130–133, November 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:PIl

[Ano84-192] Anonymous. Papers are invited. IEEE Spectrum, 21(12):91, December 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:PEW

[Ano84-193] Anonymous. Photo-electric 100: Winners of the IEEE Spectrum centennial photography contest. IEEE Spectrum, 21(8):43–47, August 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:PNRa

[Ano84-194] Anonymous. Program notes [reports of new computer programs]. IEEE Spectrum, 21 (2):24–26, February 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:PNRb

[Ano84-195] Anonymous. Program notes [reports of new computer programs]. IEEE Spectrum, 21 (4):22–24, April 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:PNRc

[Ano84-196] Anonymous. Program notes [reports of new computer programs]. IEEE Spectrum, 21 (6):21–24, June 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:PNRd

[Ano84-197] Anonymous. Program notes [reports of new computer programs]. IEEE Spectrum, 21 (8):22–23, August 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:PNRe

[Ano84-198] Anonymous. Program notes [reports of new computer programs]. IEEE Spectrum, 21 (10):28–30, October 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:PNRf

[Ano84-199] Anonymous. Program notes [reports of new computer programs]. IEEE Spectrum, 21(12): 29–32, December 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Ra

[Ano84-200] Anonymous. Reflections. IEEE Spectrum, 21(1):18–21, January 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Rb

[Ano84-201] Anonymous. Reflections. IEEE Spectrum, 21(3):14–15, March 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Rc

[Ano84-202] Anonymous. Reflections. IEEE Spectrum, 21(5):12–13, May 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Rd

[Ano84-203] Anonymous. Reflections. IEEE Spectrum, 21(7):10–11, July 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Re

[Ano84-204] Anonymous. Reflections. IEEE Spectrum, 21(9):14–15, September 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Rf

[Ano84-205] Anonymous. Reflections. IEEE Spectrum, 21(11):16, November 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:SIa

[Ano84-206] Anonymous. Scanning the Institute. IEEE Spectrum, 21 (1):136, January 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:SIb

[Ano84-207] Anonymous. Scanning the Institute. IEEE Spectrum, 21 (2):112, February 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:SIc

[Ano84-208] Anonymous. Scanning the Institute. IEEE Spectrum, 21(3):106, March 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:SId

[Ano84-209] Anonymous. Scanning the Institute. IEEE Spectrum, 21(4):105, April 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:SIe

[Ano84-210] Anonymous. Scanning the Institute. IEEE Spectrum, 21(5):114, May 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:SIf

[Ano84-211] Anonymous. Scanning the Institute. IEEE Spectrum, 21(6):168, June 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:SIg

[Ano84-212] Anonymous. Scanning the Institute. IEEE Spectrum, 21(7):96, July 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:SIh

[Ano84-213] Anonymous. Scanning the Institute. IEEE Spectrum, 21(8):90, August 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:SIi

[Ano84-214] Anonymous. Scanning the Institute. IEEE Spectrum, 21 (9):130, September 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:SIj

[Ano84-215] Anonymous. Scanning the Institute. IEEE Spectrum, 21 (10):112, October 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:SIk

[Ano84-216] Anonymous. Scanning the Institute. IEEE Spectrum, 21 (11):154, November 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:SIl

[Ano84-217] Anonymous. Scanning the Institute. IEEE Spectrum, 21 (12):109, December 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Sa

[Ano84-218] Anonymous. Speakout. IEEE Spectrum, 21(2):12–13, February 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Sb

[Ano84-219] Anonymous. Speakout. IEEE Spectrum, 21(4):9, April 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Sc

[Ano84-220] Anonymous. Speakout. IEEE Spectrum, 21(6):10–11, June 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Sd

[Ano84-221] Anonymous. Speakout. IEEE Spectrum, 21(8):9–11, August 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Se

[Ano84-222] Anonymous. Speakout. IEEE Spectrum, 21(10):8–8H, October 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Sf

[Ano84-223] Anonymous. Speakout. IEEE Spectrum, 21(12):8, December 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:SOa

[Ano84-224] Anonymous. Spin-offs. IEEE Spectrum, 21(1):28–29, January 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:SOb

[Ano84-225] Anonymous. Spin-offs. IEEE Spectrum, 21(3):24, March 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:SOc

[Ano84-226] Anonymous. Spin-offs. IEEE Spectrum, 21(5):18–21, May 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:SOd

[Ano84-227] Anonymous. Spin-offs. IEEE Spectrum, 21(7):18–21, July 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:SOe

[Ano84-228] Anonymous. Spin-offs. IEEE Spectrum, 21(9):23, September 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:SOf

[Ano84-229] Anonymous. Spin-offs. IEEE Spectrum, 21(11):30–31, November 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:TSa

[Ano84-230] Anonymous. Technically speaking. IEEE Spectrum, 21(1): 22–25, January 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:TSb

[Ano84-231] Anonymous. Technically speaking. IEEE Spectrum, 21(3): 16–19, March 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:TSc

[Ano84-232] Anonymous. Technically speaking. IEEE Spectrum, 21(5): 16–16P, May 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:TSd

[Ano84-233] Anonymous. Technically speaking. IEEE Spectrum, 21(7):16, July 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:TSe

[Ano84-234] Anonymous. Technically speaking. IEEE Spectrum, 21(9): 16, September 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:TSf

[Ano84-235] Anonymous. Technically speaking. IEEE Spectrum, 21(11): 24–27, November 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:TGV

[Ano84-236] Anonymous. Technology’s global village fragments community life: Transportation and communications technologies widen human horizons but splinter human relationships; computer networks could reverse this trend. IEEE Spectrum, 21(6):80–84, June 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:TGH

[Ano84-237] Anonymous. Television graphics hardware: Speed is the key: Up to 2 megabytes must be manipulated within 1/30 of a second. IEEE Spectrum, 21 (10):72–75, October 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:PFa

[Ano84-238] Anonymous. To probe further. IEEE Spectrum, 21(1): 97–98, January 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:PFb

[Ano84-239] Anonymous. To probe further. IEEE Spectrum, 21(11): 121–127, November 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:PFG

[Ano84-240] Anonymous. To probe further: General. IEEE Spectrum, 21 (6):133–134, June 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Va

[Ano84-241] Anonymous. Video. IEEE Spectrum, 21(2):22–23, February 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Vb

[Ano84-242] Anonymous. Video. IEEE Spectrum, 21(4):18–19, April 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Vc

[Ano84-243] Anonymous. Video. IEEE Spectrum, 21(6):20, June 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Vd

[Ano84-244] Anonymous. Video. IEEE Spectrum, 21(8):20–21, August 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Ve

[Ano84-245] Anonymous. Video. IEEE Spectrum, 21(10):26–27, October 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:Vf

[Ano84-246] Anonymous. Video. IEEE Spectrum, 21(12):28, December 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:WFL

[Ano84-247] Anonymous. Wescon ’84: Focus on local area networks: New standards for LANs — and ways of conforming to them — are a prime topic at the Anaheim, Calif., meeting Oct. 30– Nov. 2. IEEE Spectrum, 21 (10):78–79, October 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:WHa

[Ano84-248] Anonymous. Whatever happened to. IEEE Spectrum, 21(5):17, May 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:WHb

[Ano84-249] Anonymous. Whatever happened to. IEEE Spectrum, 21(7):17, July 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:WHc

[Ano84-250] Anonymous. Whatever happened to. IEEE Spectrum, 21 (9):22, September 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:WHd

[Ano84-251] Anonymous. Whatever happened to. IEEE Spectrum, 21 (11):28–29, November 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:WHO

[Ano84-252] Anonymous. Whatever happened to ocean thermal? IEEE Spectrum, 21(1):30–32, January 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1984:WHT

[Ano84-253] Anonymous. Whatever happened to the tunnel diode? IEEE Spectrum, 21(3):25–26, March 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:AII

[Ano85a] Anonymous. 1985 annual index to IEEE Spectrum. IEEE Spectrum, 22(12):100–109, December 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:IGA

[Ano85b] Anonymous. 1985 IEEE guide for authors. IEEE Spectrum, 22 (1):97–100, January 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Aa

[Ano85c] Anonymous. [advertisement].

IEEE Spectrum, 22(1):1–3, January 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Ab

[Ano85d] Anonymous. [advertisement].

IEEE Spectrum, 22(1):2, January 1985. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Ac

[Ano85e] Anonymous. [advertisement]. IEEE Spectrum, 22(1):4–7, January 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Ad

[Ano85f] Anonymous. [advertisement]. IEEE Spectrum, 22(1):1–3, January 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Ae

[Ano85g] Anonymous. [advertisement]. IEEE Spectrum, 22(2):1–3, February 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Af

[Ano85h] Anonymous. [advertisement]. IEEE Spectrum, 22(2):2, February 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Ag

[Ano85i] Anonymous. [advertisement]. IEEE Spectrum, 22(2):4–7, February 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Ah

[Ano85j] Anonymous. [advertisement]. IEEE Spectrum, 22(3):1–3, March 1985. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Ai

[Ano85k] Anonymous. [advertisement].

IEEE Spectrum, 22(3):2, March 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Aj

[Ano85l] Anonymous. [advertisement].

IEEE Spectrum, 22(3):4–5, March 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Ak

[Ano85m] Anonymous. [advertisement].

IEEE Spectrum, 22(4):1–3, April 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Al

[Ano85n] Anonymous. [advertisement].

IEEE Spectrum, 22(4):2, April 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Am

[Ano85o] Anonymous. [advertisement].

IEEE Spectrum, 22(4):4–5, April 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:An

[Ano85p] Anonymous. [advertisement].

IEEE Spectrum, 22(5):1–3, May 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Ao

[Ano85q] Anonymous. [advertisement]. IEEE Spectrum, 22(5):2, May 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Ap

[Ano85r] Anonymous. [advertisement]. IEEE Spectrum, 22(5):4–5, May 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Aq

[Ano85s] Anonymous. [advertisement]. IEEE Spectrum, 22(6):1–3, June 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Ar

[Ano85t] Anonymous. [advertisement]. IEEE Spectrum, 22(6):2, June 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:As

[Ano85u] Anonymous. [advertisement]. IEEE Spectrum, 22(6):4–7, June 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:At

[Ano85v] Anonymous. [advertisement]. IEEE Spectrum, 22(7):1–3, July 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Au

[Ano85w] Anonymous. [advertisement].

IEEE Spectrum, 22(7):2, July 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Av

[Ano85x] Anonymous. [advertisement].

IEEE Spectrum, 22(7):4–5, July 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Aw

[Ano85y] Anonymous. [advertisement].

IEEE Spectrum, 22(8):1–3, August 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Ax

[Ano85z] Anonymous. [advertisement].

IEEE Spectrum, 22(8):2, August 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Ay

[Ano85-27] Anonymous. [advertisement].

IEEE Spectrum, 22(8):4–7, August 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Az

[Ano85-28] Anonymous. [advertisement].

IEEE Spectrum, 22(8):1–2, August 1985. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Aaa

[Ano85-29] Anonymous. [advertisement]. IEEE Spectrum, 22(9):1–3, September 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Aab

[Ano85-30] Anonymous. [advertisement]. IEEE Spectrum, 22(9):2, September 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Aac

[Ano85-31] Anonymous. [advertisement]. IEEE Spectrum, 22(9):4–7, September 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Aad

[Ano85-32] Anonymous. [advertisement]. IEEE Spectrum, 22(9):1, September 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Aae

[Ano85-33] Anonymous. [advertisement]. IEEE Spectrum, 22(10):1–3, October 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Aaf

[Ano85-34] Anonymous. [advertisement]. IEEE Spectrum, 22(10):2, October 1985. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Aag

[Ano85-35] Anonymous. [advertisement].

IEEE Spectrum, 22(10):4–7, October 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Aah

[Ano85-36] Anonymous. [advertisement].

IEEE Spectrum, 22(10):1–4, October 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Aai

[Ano85-37] Anonymous. [advertisement].

IEEE Spectrum, 22(11):1–3, November 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Aaj

[Ano85-38] Anonymous. [advertisement].

IEEE Spectrum, 22(11):2, November 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Aak

[Ano85-39] Anonymous. [advertisement].

IEEE Spectrum, 22(11):4–7, November 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Aal

[Ano85-40] Anonymous. [advertisement].

IEEE Spectrum, 22(11):1–2, November 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Aam

[Ano85-41] Anonymous. [advertisement]. IEEE Spectrum, 22(12):1–3, December 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Aan

[Ano85-42] Anonymous. [advertisement]. IEEE Spectrum, 22(12):2, December 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Aao

[Ano85-43] Anonymous. [advertisement]. IEEE Spectrum, 22(12):4–7, December 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Aap

[Ano85-44] Anonymous. [advertisement]. IEEE Spectrum, 22(12):1–2, December 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:AIF

[Ano85-45] Anonymous. Awards/85: IEEE Field Award winners. IEEE Spectrum, 22(8):68–69, August 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:APP

[Ano85-46] Anonymous. Awards/85 Prize Papers and Service Awards [4 awards]. IEEE Spectrum, 22 (10):84, October 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:BMa

[Ano85-47] Anonymous. [back matter]. IEEE Spectrum, 22(2): 1–3, February 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:BMb

[Ano85-48] Anonymous. [back matter].

IEEE Spectrum, 22(3):1–2, March 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:BMc

[Ano85-49] Anonymous. [back matter].

IEEE Spectrum, 22(4):1–2, April 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:BMd

[Ano85-50] Anonymous. [back matter].

IEEE Spectrum, 22(5):1–2, May 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:BMe

[Ano85-51] Anonymous. [back matter].

IEEE Spectrum, 22(6):1–2, June 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:BMf

[Ano85-52] Anonymous. [back matter].

IEEE Spectrum, 22(7):1–3, July 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:BBAa

[Ano85-53] Anonymous. Best bits: Applications of microprocessors. IEEE Spectrum, 22(1):30–32, January 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:BBAb

[Ano85-54] Anonymous. Best bits: Applications of microprocessors. IEEE Spectrum, 22(3):34–36, March 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:BBAc

[Ano85-55] Anonymous. Best bits: Applications of microprocessors. IEEE Spectrum, 22(5):26–29, May 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:BBAd

[Ano85-56] Anonymous. Best bits: Applications of microprocessors. IEEE Spectrum, 22(7):22–25, July 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:BBAe

[Ano85-57] Anonymous. Best bits: Applications of microprocessors. IEEE Spectrum, 22(9):25–28, September 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:BBAf

[Ano85-58] Anonymous. Best bits: Applications of microprocessors. IEEE Spectrum, 22(11):30–37, November 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:BRa

[Ano85-59] Anonymous. Book reviews.

IEEE Spectrum, 22(1):18–23, January 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:BRb

[Ano85-60] Anonymous. Book reviews.

IEEE Spectrum, 22(2):17–21, February 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:BRc

[Ano85-61] Anonymous. Book reviews.

IEEE Spectrum, 22(3):16–22, March 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:BRd

[Ano85-62] Anonymous. Book reviews.

IEEE Spectrum, 22(4):14–16, April 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:BRe

[Ano85-63] Anonymous. Book reviews.

IEEE Spectrum, 22(5):8–15, May 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:BRf

[Ano85-64] Anonymous. Book reviews. IEEE Spectrum, 22(6):14–16, June 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:BRg

[Ano85-65] Anonymous. Book reviews. IEEE Spectrum, 22(7):9–15, July 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:BRh

[Ano85-66] Anonymous. Book reviews. IEEE Spectrum, 22(8):10–15, August 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:BRi

[Ano85-67] Anonymous. Book reviews. IEEE Spectrum, 22(9):14–16, September 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:BRj

[Ano85-68] Anonymous. Book reviews. IEEE Spectrum, 22(10):10– 15, October 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:BRk

[Ano85-69] Anonymous. Book reviews. IEEE Spectrum, 22(11):14–21, November 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:BRl

[Ano85-70] Anonymous. Book reviews.

IEEE Spectrum, 22(12):14–16, December 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Cb

[Ano85-71] Anonymous. Calendar. IEEE Spectrum, 22(1):103–106, January 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Ce

[Ano85-72] Anonymous. Calendar. IEEE Spectrum, 22(2):76–77, February 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Cg

[Ano85-73] Anonymous. Calendar. IEEE Spectrum, 22(3):92–95, March 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Cj

[Ano85-74] Anonymous. Calendar. IEEE Spectrum, 22(4):86–90, April 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Cl

[Ano85-75] Anonymous. Calendar. IEEE Spectrum, 22(5):117–119, May 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Co

[Ano85-76] Anonymous. Calendar. IEEE Spectrum, 22(6):79–82, June 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Cq

[Ano85-77] Anonymous. Calendar. IEEE Spectrum, 22(7):84–87, July 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Cu

[Ano85-78] Anonymous. Calendar. IEEE Spectrum, 22(8):73–75, August 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Cw

[Ano85-79] Anonymous. Calendar. IEEE Spectrum, 22(9):88–93, September 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Cz

[Ano85-80] Anonymous. Calendar. IEEE Spectrum, 22(10):87–88, October 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Cab

[Ano85-81] Anonymous. Calendar. IEEE Spectrum, 22(11):118–120, November 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Cae

[Ano85-82] Anonymous. Calendar. IEEE Spectrum, 22(12):72–75, December 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:CSa

[Ano85-83] Anonymous. Coming in Spectrum. IEEE Spectrum, 22(1): 128–129, January 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:CSb

[Ano85-84] Anonymous. Coming in Spectrum. IEEE Spectrum, 22(2): 100–101, February 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:CSc

[Ano85-85] Anonymous. Coming in Spectrum. IEEE Spectrum, 22(3): 114–116, March 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:CSd

[Ano85-86] Anonymous. Coming in Spectrum. IEEE Spectrum, 22(4): 103–106, April 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:CSe

[Ano85-87] Anonymous. Coming in Spectrum. IEEE Spectrum, 22 (5):137–138, May 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:CSf

[Ano85-88] Anonymous. Coming in Spectrum. IEEE Spectrum, 22(6):98, June 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:CSg

[Ano85-89] Anonymous. Coming in Spectrum. IEEE Spectrum, 22 (7):104–105, July 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:CSh

[Ano85-90] Anonymous. Coming in Spectrum. IEEE Spectrum, 22 (8):87–90, August 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:CSi

[Ano85-91] Anonymous. Coming in Spectrum. IEEE Spectrum, 22(9): 114–116, September 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:CSj

[Ano85-92] Anonymous. Coming in Spectrum. IEEE Spectrum, 22 (10):104, October 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:CSk

[Ano85-93] Anonymous. Coming in Spectrum. IEEE Spectrum, 22(11): 137–139, November 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:CSl

[Ano85-94] Anonymous. Coming in Spectrum. IEEE Spectrum, 22(12): 97–99, December 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Ca

[Ano85-95] Anonymous. Contents. IEEE Spectrum, 22(1):3, January 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Cc

[Ano85-96] Anonymous. Contents. IEEE Spectrum, 22(2):3, February 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Cf

[Ano85-97] Anonymous. Contents. IEEE Spectrum, 22(3):3, March 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Ch

[Ano85-98] Anonymous. Contents. IEEE Spectrum, 22(4):3, April 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Ck

[Ano85-99] Anonymous. Contents. IEEE Spectrum, 22(5):3, May 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Cm

[Ano85-100] Anonymous. Contents. IEEE Spectrum, 22(6):3, June 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Cp

[Ano85-101] Anonymous. Contents. IEEE Spectrum, 22(7):3, July 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Cr

[Ano85-102] Anonymous. Contents. IEEE Spectrum, 22(8):3, August 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Cv

[Ano85-103] Anonymous. Contents. IEEE Spectrum, 22(9):3, September 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Cx

[Ano85-104] Anonymous. Contents. IEEE Spectrum, 22(10):3, October 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Caa

[Ano85-105] Anonymous. Contents. IEEE Spectrum, 22(11):3, November 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Cac

[Ano85-106] Anonymous. Contents. IEEE Spectrum, 22(12):3, December 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Cd

[Ano85-107] Anonymous. Continuum. IEEE Spectrum, 22(2):27–29, February 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Ci

[Ano85-108] Anonymous. Continuum. IEEE Spectrum, 22(4):22–23, April 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Cn

[Ano85-109] Anonymous. Continuum. IEEE Spectrum, 22(6):18–21, June 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Ct

[Ano85-110] Anonymous. Continuum. IEEE Spectrum, 22(8):17, August 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Cy

[Ano85-111] Anonymous. Continuum. IEEE Spectrum, 22(10):20–21, October 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Cad

[Ano85-112] Anonymous. Continuum. IEEE Spectrum, 22(12):19–21, December 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:ETTa

[Ano85-113] Anonymous. EEs’ tools toys. IEEE Spectrum, 22(1):101– 102, January 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:ETTb

[Ano85-114] Anonymous. EEs’ tools toys. IEEE Spectrum, 22(2):74–75, February 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:ETTc

[Ano85-115] Anonymous. EEs’ tools toys. IEEE Spectrum, 22(3):90–91, March 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:ETTd

[Ano85-116] Anonymous. EEs’ tools toys. IEEE Spectrum, 22(4):82–85, April 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:ETTe

[Ano85-117] Anonymous. EEs’ tools toys. IEEE Spectrum, 22(5):113–116, May 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:ETTf

[Ano85-118] Anonymous. EEs’ tools toys.

IEEE Spectrum, 22(6):77–78, June 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:ETTg

[Ano85-119] Anonymous. EEs’ tools toys.

IEEE Spectrum, 22(7):81–83, July 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:ETTh

[Ano85-120] Anonymous. EEs’ tools toys.

IEEE Spectrum, 22(8):70–72, August 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:ETTi

[Ano85-121] Anonymous. EEs’ tools toys.

IEEE Spectrum, 22(9):84–87, September 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:ETTj

[Ano85-122] Anonymous. EEs’ tools toys.

IEEE Spectrum, 22(10):85– 86, October 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:ETTk

[Ano85-123] Anonymous. EEs’ tools toys.

IEEE Spectrum, 22(11):115– 117, November 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:ETTl

[Ano85-124] Anonymous. EEs’ tools toys. IEEE Spectrum, 22(12):70–71, December 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:ETM

[Ano85-125] Anonymous. The electric talking machine: Electric motors were first used to drive early phonographs in the late 1880s. IEEE Spectrum, 22(2):60–61, February 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:ELa

[Ano85-126] Anonymous. The engineer at large. IEEE Spectrum, 22 (2):26, February 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:ELb

[Ano85-127] Anonymous. The engineer at large. IEEE Spectrum, 22 (4):20–21, April 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:ELc

[Ano85-128] Anonymous. The engineer at large. IEEE Spectrum, 22(6):17, June 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:ELd

[Ano85-129] Anonymous. The engineer at large. IEEE Spectrum, 22(8):16, August 1985. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:ELe

[Ano85-130] Anonymous. The engineer at large. IEEE Spectrum, 22(10): 17–19, October 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:ELf

[Ano85-131] Anonymous. The engineer at large. IEEE Spectrum, 22(12): 17–18, December 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Fa

[Ano85-132] Anonymous. Forum. IEEE Spectrum, 22(1):8–15, January 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Fb

[Ano85-133] Anonymous. Forum. IEEE Spectrum, 22(2):8–16, February 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Fc

[Ano85-134] Anonymous. Forum. IEEE Spectrum, 22(3):6–13, March 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Fd

[Ano85-135] Anonymous. Forum. IEEE Spectrum, 22(4):6–13, April 1985.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Fe

[Ano85-136] Anonymous. Forum. IEEE Spectrum, 22(5):6–7, May 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Ff

[Ano85-137] Anonymous. Forum. IEEE Spectrum, 22(6):9, June 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Fg

[Ano85-138] Anonymous. Forum. IEEE Spectrum, 22(7):6–8, July 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Fh

[Ano85-139] Anonymous. Forum. IEEE Spectrum, 22(8):9, August 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Fi

[Ano85-140] Anonymous. Forum. IEEE Spectrum, 22(9):10–13, September 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Fj

[Ano85-141] Anonymous. Forum. IEEE Spectrum, 22(10):8–9, October 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Fk

[Ano85-142] Anonymous. Forum. IEEE Spectrum, 22(11):10–13, November 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Fl

[Ano85-143] Anonymous. Forum. IEEE Spectrum, 22(12):9–13, December 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:IAM

[Ano85-144] Anonymous. IEEE awards: 1985 major medalists. IEEE Spectrum, 22(4):80–81, April 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:ITCa

[Ano85-145] Anonymous. IEEE tables of contents. IEEE Spectrum, 22(1): 108–109, January 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:ITCb

[Ano85-146] Anonymous. IEEE tables of contents. IEEE Spectrum, 22 (2):83–99, February 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:ITCc

[Ano85-147] Anonymous. IEEE tables of contents. IEEE Spectrum, 22 (3):101–113, March 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:ITCd

[Ano85-148] Anonymous. IEEE tables of contents. IEEE Spectrum, 22 (4):94–102, April 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:ITCe

[Ano85-149] Anonymous. IEEE tables of contents. IEEE Spectrum, 22 (5):124–125, May 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:ITCf

[Ano85-150] Anonymous. IEEE tables of contents. IEEE Spectrum, 22 (6):86–97, June 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:ITCg

[Ano85-151] Anonymous. IEEE tables of contents. IEEE Spectrum, 22 (7):94–103, July 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:ITCh

[Ano85-152] Anonymous. IEEE tables of contents. IEEE Spectrum, 22 (8):81–86, August 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:ITCi

[Ano85-153] Anonymous. IEEE tables of contents. IEEE Spectrum, 22(9): 102–113, September 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:ITCj

[Ano85-154] Anonymous. IEEE tables of contents. IEEE Spectrum, 22 (10):93, October 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:ITCk

[Ano85-155] Anonymous. IEEE tables of contents. IEEE Spectrum, 22(11): 126–136, November 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:ITCl

[Ano85-156] Anonymous. IEEE tables of contents. IEEE Spectrum, 22 (12):79, December 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Ia

[Ano85-157] Anonymous. Innovations. IEEE Spectrum, 22(2):30–31, February 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Ib

[Ano85-158] Anonymous. Innovations. IEEE Spectrum, 22(4):24–25, April 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Ic

[Ano85-159] Anonymous. Innovations. IEEE Spectrum, 22(6):22–23, June 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Id

[Ano85-160] Anonymous. Innovations. IEEE Spectrum, 22(8):18–21, August 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Ie

[Ano85-161] Anonymous. Innovations. IEEE Spectrum, 22(10):22–25, October 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:If

[Ano85-162] Anonymous. Innovations. IEEE Spectrum, 22(12):22–23, December 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:MTa

[Ano85-163] Anonymous. Managing technology. IEEE Spectrum, 22 (5):30–32, May 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:MTb

[Ano85-164] Anonymous. Managing technology. IEEE Spectrum, 22 (7):26–28, July 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:MTc

[Ano85-165] Anonymous. Managing technology. IEEE Spectrum, 22(9): 29–32, September 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:MTd

[Ano85-166] Anonymous. Managing technology. IEEE Spectrum, 22(11): 38–42, November 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:NRIa

[Ano85-167] Anonymous. New and recent IEEE publications. IEEE Spectrum, 22(1):110–127, January 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:NRIb

[Ano85-168] Anonymous. New and recent IEEE publications. IEEE Spectrum, 22(2):78–79, February 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:NRIc

[Ano85-169] Anonymous. New and recent IEEE publications. IEEE Spectrum, 22(3):96–97, March 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:NRId

[Ano85-170] Anonymous. New and recent IEEE publications. IEEE Spectrum, 22(4):92–93, April 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:NRIe

[Ano85-171] Anonymous. New and recent IEEE publications. IEEE Spectrum, 22(5):126–136, May 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:NRIf

[Ano85-172] Anonymous. New and recent IEEE publications. IEEE Spectrum, 22(6):84–85, June 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:NRIg

[Ano85-173] Anonymous. New and recent IEEE publications. IEEE Spectrum, 22(7):92–93, July 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:NRIh

[Ano85-174] Anonymous. New and recent IEEE publications. IEEE Spectrum, 22(8):79–80, August 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:NRIi

[Ano85-175] Anonymous. New and recent IEEE publications. IEEE Spectrum, 22(9):94–97, September 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:NRIj

[Ano85-176] Anonymous. New and recent IEEE publications. IEEE Spectrum, 22(10):94–103, October 1985. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:NRIk

[Ano85-177] Anonymous. New and recent IEEE publications. IEEE Spectrum, 22(11):121–124, November 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:NRIl

[Ano85-178] Anonymous. New and recent IEEE publications. IEEE Spectrum, 22(12):80–96, December 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:PIa

[Ano85-179] Anonymous. Papers are invited.

IEEE Spectrum, 22(1):107, January 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:PIb

[Ano85-180] Anonymous. Papers are invited. IEEE Spectrum, 22(2): 80–82, February 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:PIc

[Ano85-181] Anonymous. Papers are invited.

IEEE Spectrum, 22(3):98–100, March 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:PId

[Ano85-182] Anonymous. Papers are invited. IEEE Spectrum, 22(4):91, April 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:PIe

[Ano85-183] Anonymous. Papers are invited. IEEE Spectrum, 22(5):120–123, May 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:PIf

[Ano85-184] Anonymous. Papers are invited. IEEE Spectrum, 22(6):83, June 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:PIg

[Ano85-185] Anonymous. Papers are invited. IEEE Spectrum, 22 (7):88–91, July 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:PIh

[Ano85-186] Anonymous. Papers are invited. IEEE Spectrum, 22 (8):76–78, August 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:PIi

[Ano85-187] Anonymous. Papers are invited. IEEE Spectrum, 22(9): 98–101, September 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:PIj

[Ano85-188] Anonymous. Papers are invited. IEEE Spectrum, 22(10):

91–92, October 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:PIk

[Ano85-189] Anonymous. Papers are invited. IEEE Spectrum, 22(11): 125, November 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:PIl

[Ano85-190] Anonymous. Papers are invited. IEEE Spectrum, 22(12): 76–78, December 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:PNRa

[Ano85-191] Anonymous. Program notes [reports of new computer programs]. IEEE Spectrum, 22 (2):33–34, February 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:PNRb

[Ano85-192] Anonymous. Program notes [reports of new computer programs]. IEEE Spectrum, 22 (4):27–30, April 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:PNRc

[Ano85-193] Anonymous. Program notes [reports of new computer programs]. IEEE Spectrum, 22 (6):26–28, June 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:PNRd

[Ano85-194] Anonymous. Program notes [reports of new computer programs]. IEEE Spectrum, 22 (8):27–28, August 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:PNRe

[Ano85-195] Anonymous. Program notes [reports of new computer programs]. IEEE Spectrum, 22 (10):26–29, October 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:PNRf

[Ano85-196] Anonymous. Program notes [reports of new computer programs]. IEEE Spectrum, 22(12): 24–25, December 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Ra

[Ano85-197] Anonymous. Reflections. IEEE Spectrum, 22(1):16–17, January 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Rb

[Ano85-198] Anonymous. Reflections. IEEE Spectrum, 22(3):14–15, March 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Rc

[Ano85-199] Anonymous. Reflections. IEEE Spectrum, 22(5):16, May 1985.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Rd

[Ano85-200] Anonymous. Reflections. IEEE Spectrum, 22(7):16, July 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Re

[Ano85-201] Anonymous. Reflections. IEEE Spectrum, 22(9):8, September 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Rf

[Ano85-202] Anonymous. Reflections. IEEE Spectrum, 22(11):6–9, November 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:SIa

[Ano85-203] Anonymous. Scanning the Institute. IEEE Spectrum, 22 (1):128, January 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:SIb

[Ano85-204] Anonymous. Scanning the Institute. IEEE Spectrum, 22 (2):100, February 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:SIc

[Ano85-205] Anonymous. Scanning the Institute. IEEE Spectrum, 22(3):114, March 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:SId

[Ano85-206] Anonymous. Scanning the Institute. IEEE Spectrum, 22(4):103, April 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:SIe

[Ano85-207] Anonymous. Scanning the Institute. IEEE Spectrum, 22(5):137, May 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:SIf

[Ano85-208] Anonymous. Scanning the Institute. IEEE Spectrum, 22(6):98, June 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:SIg

[Ano85-209] Anonymous. Scanning the Institute. IEEE Spectrum, 22(7):104, July 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:SIh

[Ano85-210] Anonymous. Scanning the Institute. IEEE Spectrum, 22(8):87, August 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:SIi

[Ano85-211] Anonymous. Scanning the Institute. IEEE Spectrum, 22 (9):114, September 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:SIj

[Ano85-212] Anonymous. Scanning the Institute. IEEE Spectrum, 22 (10):104, October 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:SIk

[Ano85-213] Anonymous. Scanning the Institute. IEEE Spectrum, 22 (11):137, November 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:SIl

[Ano85-214] Anonymous. Scanning the Institute. IEEE Spectrum, 22 (12):97, December 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Sa

[Ano85-215] Anonymous. Speakout. IEEE Spectrum, 22(2):22–25, February 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Sb

[Ano85-216] Anonymous. Speakout. IEEE Spectrum, 22(4):17–19, April 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Sc

[Ano85-217] Anonymous. Speakout. IEEE Spectrum, 22(6):8, June 1985.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Sd

[Ano85-218] Anonymous. Speakout. IEEE Spectrum, 22(8):8, August 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Se

[Ano85-219] Anonymous. Speakout. IEEE Spectrum, 22(10):16, October 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Sf

[Ano85-220] Anonymous. Speakout. IEEE Spectrum, 22(12):8, December 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:SOa

[Ano85-221] Anonymous. Spin-offs. IEEE Spectrum, 22(1):28–29, January 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:SOb

[Ano85-222] Anonymous. Spin-offs. IEEE Spectrum, 22(3):30–33, March 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:SOc

[Ano85-223] Anonymous. Spin-offs. IEEE Spectrum, 22(5):24–25, May 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:SOd

[Ano85-224] Anonymous. Spin-offs. IEEE Spectrum, 22(7):20–21, July 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:SOe

[Ano85-225] Anonymous. Spin-offs. IEEE Spectrum, 22(9):24, September 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:SOf

[Ano85-226] Anonymous. Spin-offs. IEEE Spectrum, 22(11):26–29, November 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:TSa

[Ano85-227] Anonymous. Technically speaking. IEEE Spectrum, 22(1): 24–25, January 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:TSb

[Ano85-228] Anonymous. Technically speaking. IEEE Spectrum, 22(3):26, March 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:TSc

[Ano85-229] Anonymous. Technically speaking. IEEE Spectrum, 22(5):

17–19, May 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:TSd

[Ano85-230] Anonymous. Technically speaking. IEEE Spectrum, 22(7):17, July 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:TSe

[Ano85-231] Anonymous. Technically speaking. IEEE Spectrum, 22(9): 17–19, September 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:TSf

[Ano85-232] Anonymous. Technically speaking. IEEE Spectrum, 22(11): 22–23, November 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:PFa

[Ano85-233] Anonymous. To probe further. IEEE Spectrum, 22(1): 95–96, January 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:PFb

[Ano85-234] Anonymous. To probe further. IEEE Spectrum, 22(5):110–112, May 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:PFc

[Ano85-235] Anonymous. To probe further. IEEE Spectrum, 22(11):

113–114, November 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Va

[Ano85-236] Anonymous. Video. IEEE Spectrum, 22(2):32, February 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Vb

[Ano85-237] Anonymous. Video. IEEE Spectrum, 22(4):26, April 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Vc

[Ano85-238] Anonymous. Video. IEEE Spectrum, 22(6):24–25, June 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Vd

[Ano85-239] Anonymous. Video. IEEE Spectrum, 22(8):22–26, August 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Ve

[Ano85-240] Anonymous. Video. IEEE Spectrum, 22(10):30–32, October 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:Vf

[Ano85-241] Anonymous. Video. IEEE Spectrum, 22(12):26–28, December 1985. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:WHa

[Ano85-242] Anonymous. Whatever happened to. IEEE Spectrum, 22 (1):26–27, January 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:WHb

[Ano85-243] Anonymous. Whatever happened to. IEEE Spectrum, 22 (3):27–29, March 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:WHc

[Ano85-244] Anonymous. Whatever happened to. IEEE Spectrum, 22 (5):20–23, May 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:WHd

[Ano85-245] Anonymous. Whatever happened to. IEEE Spectrum, 22 (7):18–19, July 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:WHe

[Ano85-246] Anonymous. Whatever happened to. IEEE Spectrum, 22 (9):20–23, September 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:WHf

[Ano85-247] Anonymous. Whatever happened to. IEEE Spectrum, 22 (11):24–25, November 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:IGA

[Ano86a] Anonymous. 1986 IEEE guide for authors. IEEE Spectrum, 23 (1):90–93, January 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Aa

[Ano86b] Anonymous. [advertisement].

IEEE Spectrum, 23(1):1–3, January 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Ab

[Ano86c] Anonymous. [advertisement].

IEEE Spectrum, 23(1):2, January 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Ac

[Ano86d] Anonymous. [advertisement].

IEEE Spectrum, 23(1):4–5, January 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Ad

[Ano86e] Anonymous. [advertisement].

IEEE Spectrum, 23(1):1–2, January 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Ae

[Ano86f] Anonymous. [advertisement].

IEEE Spectrum, 23(2):1–3, February 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Af

[Ano86g] Anonymous. [advertisement]. IEEE Spectrum, 23(2):4, February 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Ag

[Ano86h] Anonymous. [advertisement]. IEEE Spectrum, 23(2):4–5, February 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Ah

[Ano86i] Anonymous. [advertisement]. IEEE Spectrum, 23(2):1, February 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Ai

[Ano86j] Anonymous. [advertisement]. IEEE Spectrum, 23(3):1–3, March 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Aj

[Ano86k] Anonymous. [advertisement]. IEEE Spectrum, 23(3):2, March 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Ak

[Ano86l] Anonymous. [advertisement]. IEEE Spectrum, 23(3):4–7, March 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Al

[Ano86m] Anonymous. [advertisement].

IEEE Spectrum, 23(4):1–2, April 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Am

[Ano86n] Anonymous. [advertisement].

IEEE Spectrum, 23(4):1–2, April 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:An

[Ano86o] Anonymous. [advertisement].

IEEE Spectrum, 23(4):4–7, April 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Ao

[Ano86p] Anonymous. [advertisement].

IEEE Spectrum, 23(4):1–2, April 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Ap

[Ano86q] Anonymous. [advertisement].

IEEE Spectrum, 23(5):1–3, May 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Aq

[Ano86r] Anonymous. [advertisement].

IEEE Spectrum, 23(5):2, May 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Ar

[Ano86s] Anonymous. [advertisement]. IEEE Spectrum, 23(5):4–5, May 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:As

[Ano86t] Anonymous. [advertisement]. IEEE Spectrum, 23(5):1–2, May 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:At

[Ano86u] Anonymous. [advertisement]. IEEE Spectrum, 23(6):1–2, June 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Au

[Ano86v] Anonymous. [advertisement]. IEEE Spectrum, 23(6):1–2, June 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Av

[Ano86w] Anonymous. [advertisement]. IEEE Spectrum, 23(6):4–5, June 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Aw

[Ano86x] Anonymous. [advertisement]. IEEE Spectrum, 23(6):1–2, June 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Ax

[Ano86y] Anonymous. [advertisement].

IEEE Spectrum, 23(7):1–3, July 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Ay

[Ano86z] Anonymous. [advertisement].

IEEE Spectrum, 23(7):2, July 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Az

[Ano86-27] Anonymous. [advertisement].

IEEE Spectrum, 23(7):4–7, July 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Aaa

[Ano86-28] Anonymous. [advertisement].

IEEE Spectrum, 23(7):1–2, July 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Aab

[Ano86-29] Anonymous. [advertisement].

IEEE Spectrum, 23(8):1–2, August 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Aac

[Ano86-30] Anonymous. [advertisement].

IEEE Spectrum, 23(8):1–2, August 1986. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Aad

[Ano86-31] Anonymous. [advertisement]. IEEE Spectrum, 23(8):4–5, August 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Aae

[Ano86-32] Anonymous. [advertisement]. IEEE Spectrum, 23(8):1–2, August 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Aaf

[Ano86-33] Anonymous. [advertisement]. IEEE Spectrum, 23(9):1–3, September 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Aag

[Ano86-34] Anonymous. [advertisement]. IEEE Spectrum, 23(9):2, September 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Aah

[Ano86-35] Anonymous. [advertisement]. IEEE Spectrum, 23(9):4–7, September 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Aai

[Ano86-36] Anonymous. [advertisement]. IEEE Spectrum, 23(9):1, September 1986. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Aaj

[Ano86-37] Anonymous. [advertisement].

IEEE Spectrum, 23(10):1–2, October 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Aak

[Ano86-38] Anonymous. [advertisement].

IEEE Spectrum, 23(10):1–2, October 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Aal

[Ano86-39] Anonymous. [advertisement].

IEEE Spectrum, 23(10):4–5, October 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Aam

[Ano86-40] Anonymous. [advertisement].

IEEE Spectrum, 23(11):1–3, November 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Aan

[Ano86-41] Anonymous. [advertisement].

IEEE Spectrum, 23(11):1–2, November 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Aao

[Ano86-42] Anonymous. [advertisement].

IEEE Spectrum, 23(11):4–7, November 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Aap

[Ano86-43] Anonymous. [advertisement]. IEEE Spectrum, 23(11):1, November 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Aaq

[Ano86-44] Anonymous. [advertisement]. IEEE Spectrum, 23(12):1–2, December 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Aar

[Ano86-45] Anonymous. [advertisement]. IEEE Spectrum, 23(12):1–2, December 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Aas

[Ano86-46] Anonymous. [advertisement]. IEEE Spectrum, 23(12):4–7, December 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Aat

[Ano86-47] Anonymous. [advertisement]. IEEE Spectrum, 23(12):1–2, December 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:AIF

[Ano86-48] Anonymous. Awards/86: IEEE Field Award winners. IEEE Spectrum, 23(8):70–71, August 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:APP

[Ano86-49] Anonymous. Awards/86: Prize Papers and Service Awards. IEEE Spectrum, 23(9):66, September 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:BBAa

[Ano86-50] Anonymous. Best bits: Applications of microprocessors. IEEE Spectrum, 23(1):26–27, January 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:BBAb

[Ano86-51] Anonymous. Best bits: Applications of microprocessors. IEEE Spectrum, 23(3):18–21, March 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:BBAc

[Ano86-52] Anonymous. Best bits: Applications of microprocessors. IEEE Spectrum, 23(5):28–29, May 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:BRa

[Ano86-53] Anonymous. Book reviews.

IEEE Spectrum, 23(1):18–23, January 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:BRb

[Ano86-54] Anonymous. Book reviews. IEEE Spectrum, 23(2):12–15, February 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:BRc

[Ano86-55] Anonymous. Book reviews. IEEE Spectrum, 23(3):12–15, March 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:BRd

[Ano86-56] Anonymous. Book reviews. IEEE Spectrum, 23(4):17–23, April 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:BRe

[Ano86-57] Anonymous. Book reviews. IEEE Spectrum, 23(5):14–19, May 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:BRf

[Ano86-58] Anonymous. Book reviews. IEEE Spectrum, 23(6):13–25, June 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:BRg

[Ano86-59] Anonymous. Book reviews. IEEE Spectrum, 23(7):18–23, July 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:BRh

[Ano86-60] Anonymous. Book reviews.

IEEE Spectrum, 23(8):10–16, August 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:BRi

[Ano86-61] Anonymous. Book reviews.

IEEE Spectrum, 23(9):16–19, September 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:BRj

[Ano86-62] Anonymous. Book reviews.

IEEE Spectrum, 23(10):10– 16, October 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:BRk

[Ano86-63] Anonymous. Book reviews.

IEEE Spectrum, 23(11):12–15, November 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:BRl

[Ano86-64] Anonymous. Book reviews.

IEEE Spectrum, 23(12):10–15, December 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:ISA

[Ano86-65] Anonymous. IEEE Spectrum annual index. IEEE Spectrum, 23 (12):89–98, December 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:BCM

[Ano86-66] Anonymous. Burning coal more cleanly and efficiently: Faced with tougher antipollution laws and inflationary costs, electric utilities turn to advanced coal combustion technologies. IEEE Spectrum, 23(8):64–69, August 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Cb

[Ano86-67] Anonymous. Calendar. IEEE Spectrum, 23(1):99–101, January 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Ce

[Ano86-68] Anonymous. Calendar. IEEE Spectrum, 23(2):75–84, February 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Cg

[Ano86-69] Anonymous. Calendar. IEEE Spectrum, 23(3):87–92, March 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Ci

[Ano86-70] Anonymous. Calendar. IEEE Spectrum, 23(4):81–82, April 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Ck

[Ano86-71] Anonymous. Calendar. IEEE Spectrum, 23(5):92–95, May 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Cm

[Ano86-72] Anonymous. Calendar. IEEE Spectrum, 23(6):74–79, June 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Co

[Ano86-73] Anonymous. Calendar. IEEE Spectrum, 23(7):83–84, July 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Cq

[Ano86-74] Anonymous. Calendar. IEEE Spectrum, 23(8):75–76, August 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Ct

[Ano86-75] Anonymous. Calendar. IEEE Spectrum, 23(9):69–70, September 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Cv

[Ano86-76] Anonymous. Calendar. IEEE Spectrum, 23(10):78–82, October 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Cx

[Ano86-77] Anonymous. Calendar. IEEE Spectrum, 23(11):88–90NN, November 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Cz

[Ano86-78] Anonymous. Calendar. IEEE Spectrum, 23(12):60–65, December 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:CSa

[Ano86-79] Anonymous. Coming in Spectrum. IEEE Spectrum, 23(1): 119–122, January 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:CSb

[Ano86-80] Anonymous. Coming in Spectrum. IEEE Spectrum, 23(2): 99–102, February 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:CSc

[Ano86-81] Anonymous. Coming in Spectrum. IEEE Spectrum, 23(3):103, March 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:CSd

[Ano86-82] Anonymous. Coming in Spectrum. IEEE Spectrum, 23(4):98, April 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:CSe

[Ano86-83] Anonymous. Coming in Spectrum. IEEE Spectrum, 23(5):112, May 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:CSf

[Ano86-84] Anonymous. Coming in Spectrum. IEEE Spectrum, 23(6):90, June 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:CSg

[Ano86-85] Anonymous. Coming in Spectrum. IEEE Spectrum, 23 (7):96–98, July 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:CSh

[Ano86-86] Anonymous. Coming in Spectrum. IEEE Spectrum, 23(8):90, August 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:CSi

[Ano86-87] Anonymous. Coming in Spectrum. IEEE Spectrum, 23 (9):84, September 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:CSj

[Ano86-88] Anonymous. Coming in Spectrum. IEEE Spectrum, 23 (10):98, October 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:CSk

[Ano86-89] Anonymous. Coming in Spectrum. IEEE Spectrum, 23(11): 108, November 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:CSl

[Ano86-90] Anonymous. Coming in Spectrum. IEEE Spectrum, 23(12): 72–88, December 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Ca

[Ano86-91] Anonymous. Contents. IEEE Spectrum, 23(1):3, January 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Cc

[Ano86-92] Anonymous. Contents. IEEE Spectrum, 23(2):3, February 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Cf

[Ano86-93] Anonymous. Contents. IEEE Spectrum, 23(3):3, March 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Ch

[Ano86-94] Anonymous. Contents. IEEE Spectrum, 23(4):3, April 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Cj

[Ano86-95] Anonymous. Contents. IEEE Spectrum, 23(5):3, May 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Cl

[Ano86-96] Anonymous. Contents. IEEE Spectrum, 23(6):3, June 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Cn

[Ano86-97] Anonymous. Contents. IEEE Spectrum, 23(7):3, July 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Cp

[Ano86-98] Anonymous. Contents. IEEE Spectrum, 23(8):3, August 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Cr

[Ano86-99] Anonymous. Contents. IEEE Spectrum, 23(9):3, September 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Cu

[Ano86-100] Anonymous. Contents. IEEE Spectrum, 23(10):3, October 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Cw

[Ano86-101] Anonymous. Contents. IEEE Spectrum, 23(11):3, November 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Cy

[Ano86-102] Anonymous. Contents. IEEE Spectrum, 23(12):3, December 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Cd

[Ano86-103] Anonymous. Continuum. IEEE Spectrum, 23(2):17, February 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:ETTa

[Ano86-104] Anonymous. EEs’ tools toys. IEEE Spectrum, 23(1):94–98, January 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:ETTb

[Ano86-105] Anonymous. EEs’ tools toys. IEEE Spectrum, 23(2):72–74, February 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:ETTc

[Ano86-106] Anonymous. EEs’ tools toys. IEEE Spectrum, 23(3):83–86, March 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:ETTd

[Ano86-107] Anonymous. EEs’ tools toys.

IEEE Spectrum, 23(4):79–80, April 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:ETTe

[Ano86-108] Anonymous. EEs’ tools toys.

IEEE Spectrum, 23(5):88–91, May 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:ETTf

[Ano86-109] Anonymous. EEs’ tools toys.

IEEE Spectrum, 23(6):72–73, June 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:ETTg

[Ano86-110] Anonymous. EEs’ tools toys.

IEEE Spectrum, 23(7):81–82, July 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:ETTh

[Ano86-111] Anonymous. EEs’ tools toys.

IEEE Spectrum, 23(8):72–74, August 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:ETTi

[Ano86-112] Anonymous. EEs’ tools toys.

IEEE Spectrum, 23(10):74– 77, October 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:ETTj

[Ano86-113] Anonymous. EEs’ tools toys. IEEE Spectrum, 23(11):84–85, November 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:ETTk

[Ano86-114] Anonymous. EEs’ tools toys. IEEE Spectrum, 23(12):57–59, December 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:ELa

[Ano86-115] Anonymous. The engineer at large. IEEE Spectrum, 23 (2):16, February 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:ELb

[Ano86-116] Anonymous. The engineer at large. IEEE Spectrum, 23(4):24, April 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:ELc

[Ano86-117] Anonymous. The engineer at large. IEEE Spectrum, 23 (6):20–23, June 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:ELd

[Ano86-118] Anonymous. The engineer at large. IEEE Spectrum, 23 (8):20–21, August 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:ELe

[Ano86-119] Anonymous. The engineer at large. IEEE Spectrum, 23(10): 20–23, October 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:ELf

[Ano86-120] Anonymous. The engineer at large. IEEE Spectrum, 23 (12):17, December 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Fa

[Ano86-121] Anonymous. Forum. IEEE Spectrum, 23(1):8–15, January 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Fb

[Ano86-122] Anonymous. Forum. IEEE Spectrum, 23(2):6–8, February 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Fc

[Ano86-123] Anonymous. Forum. IEEE Spectrum, 23(3):9–11, March 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Fd

[Ano86-124] Anonymous. Forum. IEEE Spectrum, 23(4):8–15, April 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Fe

[Ano86-125] Anonymous. Forum. IEEE Spectrum, 23(5):8–13, May 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Ff

[Ano86-126] Anonymous. Forum. IEEE Spectrum, 23(6):6–11, June 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Fg

[Ano86-127] Anonymous. Forum. IEEE Spectrum, 23(7):12–17, July 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Fh

[Ano86-128] Anonymous. Forum. IEEE Spectrum, 23(8):6–9, August 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Fi

[Ano86-129] Anonymous. Forum. IEEE Spectrum, 23(9):12–15, September 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Fj

[Ano86-130] Anonymous. Forum. IEEE Spectrum, 23(10):6–9, October 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Fk

[Ano86-131] Anonymous. Forum. IEEE Spectrum, 23(11):9–11, November 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Fl

[Ano86-132] Anonymous. Forum. IEEE Spectrum, 23(12):8–9, December 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:IAM

[Ano86-133] Anonymous. IEEE awards: 1986 major medalists. IEEE Spectrum, 23(5):86–87, May 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:ITCa

[Ano86-134] Anonymous. IEEE tables of contents. IEEE Spectrum, 23 (2):88–98, February 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:ITCb

[Ano86-135] Anonymous. IEEE tables of contents. IEEE Spectrum, 23 (3):94–102, March 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:ITCc

[Ano86-136] Anonymous. IEEE tables of contents. IEEE Spectrum, 23 (4):86–97, April 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:ITCd

[Ano86-137] Anonymous. IEEE tables of contents. IEEE Spectrum, 23 (5):102–111, May 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:ITCe

[Ano86-138] Anonymous. IEEE tables of contents. IEEE Spectrum, 23 (6):83–85, June 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:ITCf

[Ano86-139] Anonymous. IEEE tables of contents. IEEE Spectrum, 23 (7):88–95, July 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:ITCg

[Ano86-140] Anonymous. IEEE tables of contents. IEEE Spectrum, 23 (8):77–78, August 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:ITCh

[Ano86-141] Anonymous. IEEE tables of contents. IEEE Spectrum, 23(9): 73–76, September 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:ITCi

[Ano86-142] Anonymous. IEEE tables of contents. IEEE Spectrum, 23 (10):83–85, October 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:ITCj

[Ano86-143] Anonymous. IEEE tables of contents. IEEE Spectrum, 23(11): 95–107, November 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:ITCk

[Ano86-144] Anonymous. IEEE tables of contents. IEEE Spectrum, 23(12): 66–68, December 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:SOB

[Ano86-145] Anonymous. In search of an optical ‘brain’. IEEE Spectrum, 23(8):49–51, August 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Ia

[Ano86-146] Anonymous. Innovations. IEEE Spectrum, 23(2):18–21, February 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Ib

[Ano86-147] Anonymous. Innovations. IEEE Spectrum, 23(4):25, April 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Ic

[Ano86-148] Anonymous. Innovations. IEEE Spectrum, 23(6):24–25, June 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Id

[Ano86-149] Anonymous. Innovations. IEEE Spectrum, 23(8):22–25, August 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Ie

[Ano86-150] Anonymous. Innovations. IEEE Spectrum, 23(10):24, October 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:If

[Ano86-151] Anonymous. Innovations. IEEE Spectrum, 23(12):18–19, December 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:LLC

[Ano86-152] Anonymous. Lessons learned: Classic debates aired first by mainframe developers, then by minicomputer makers are discovered anew by the PC industry. IEEE Spectrum, 23 (5):46–49, May 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:LAC

[Ano86-153] Anonymous. The lure of analog computing. IEEE Spectrum, 23(8):40–44, August 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:MTa

[Ano86-154] Anonymous. Managing technology. IEEE Spectrum, 23 (1):28–29, January 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:MTb

[Ano86-155] Anonymous. Managing technology. IEEE Spectrum, 23 (3):22–23, March 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:MTc

[Ano86-156] Anonymous. Managing technology. IEEE Spectrum, 23 (5):30–32, May 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:MTd

[Ano86-157] Anonymous. Managing technology. IEEE Spectrum, 23 (7):30–32, July 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:MTe

[Ano86-158] Anonymous. Managing technology. IEEE Spectrum, 23(9): 28–30, September 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:MTf

[Ano86-159] Anonymous. Managing technology. IEEE Spectrum, 23(11): 24–34, November 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:NRIa

[Ano86-160] Anonymous. New and recent IEEE publications. IEEE Spectrum, 23(1):102–103, January 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:NRIb

[Ano86-161] Anonymous. New and recent IEEE publications. IEEE Spectrum, 23(2):86–87, February 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:NRIc

[Ano86-162] Anonymous. New and recent IEEE publications. IEEE Spectrum, 23(3):104–10, March 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:NRId

[Ano86-163] Anonymous. New and recent IEEE publications. IEEE Spectrum, 23(4):84–85, April 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:NRIe

[Ano86-164] Anonymous. New and recent IEEE publications. IEEE Spectrum, 23(5):96–97, May 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:NRIf

[Ano86-165] Anonymous. New and recent IEEE publications. IEEE Spectrum, 23(6):86–89, June 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:NRIg

[Ano86-166] Anonymous. New and recent IEEE publications. IEEE Spectrum, 23(7):86–87, July 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:NRIh

[Ano86-167] Anonymous. New and recent IEEE publications. IEEE Spectrum, 23(8):79–82, August 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:NRIi

[Ano86-168] Anonymous. New and recent IEEE publications. IEEE Spectrum, 23(9):71–72, September 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:NRIj

[Ano86-169] Anonymous. New and recent IEEE publications. IEEE Spectrum, 23(10):86–87, October 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:NRIk

[Ano86-170] Anonymous. New and recent IEEE publications. IEEE Spectrum, 23(11):93–94, November 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:NRIl

[Ano86-171] Anonymous. New and recent IEEE publications. IEEE Spectrum, 23(12):69–70, December 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:PIa

[Ano86-172] Anonymous. Papers are invited. IEEE Spectrum, 23(1): 104–118, January 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:PIb

[Ano86-173] Anonymous. Papers are invited. IEEE Spectrum, 23(2):85, February 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:PIc

[Ano86-174] Anonymous. Papers are invited. IEEE Spectrum, 23(3):93, March 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:PId

[Ano86-175] Anonymous. Papers are invited. IEEE Spectrum, 23(4):83, April 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:PIe

[Ano86-176] Anonymous. Papers are invited. IEEE Spectrum, 23(5):98–101, May 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:PIf

[Ano86-177] Anonymous. Papers are invited. IEEE Spectrum, 23 (6):80–82, June 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:PIg

[Ano86-178] Anonymous. Papers are invited. IEEE Spectrum, 23(7):85, July 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:PIh

[Ano86-179] Anonymous. Papers are invited. IEEE Spectrum, 23 (8):83–89, August 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:PIi

[Ano86-180] Anonymous. Papers are invited. IEEE Spectrum, 23(9): 77–83, September 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:PIj

[Ano86-181] Anonymous. Papers are invited. IEEE Spectrum, 23(10): 88–97, October 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:PIk

[Ano86-182] Anonymous. Papers are invited. IEEE Spectrum, 23(11): 86–87, November 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:PIl

[Ano86-183] Anonymous. Papers are invited. IEEE Spectrum, 23(12):71, December 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:PSE

[Ano86-184] Anonymous. Photons in the service of electrons. IEEE Spectrum, 23(8):38–40, August 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:PNRa

[Ano86-185] Anonymous. Program notes [reports of new computer programs]. IEEE Spectrum, 23 (2):22–23, February 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:PNRb

[Ano86-186] Anonymous. Program notes [reports of new computer programs]. IEEE Spectrum, 23 (4):26–27, April 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:PNRc

[Ano86-187] Anonymous. Program notes [reports of new computer programs]. IEEE Spectrum, 23 (6):26–27, June 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:PNRd

[Ano86-188] Anonymous. Program notes [reports of new computer programs]. IEEE Spectrum, 23 (8):26–29, August 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:PNRe

[Ano86-189] Anonymous. Program notes [reports of new computer programs]. IEEE Spectrum, 23 (10):25, October 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:PNRf

[Ano86-190] Anonymous. Program notes [reports of new computer programs]. IEEE Spectrum, 23(12): 20–21, December 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:PP

[Ano86-191] Anonymous. Prospects and pitfalls. IEEE Spectrum, 23 (8):55–57, August 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:POS

[Ano86-192] Anonymous. Prospects for an optical supercomputer. IEEE Spectrum, 23(8):44–49, August 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Ra

[Ano86-193] Anonymous. Reflections. IEEE Spectrum, 23(1):6–7, January 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Rb

[Ano86-194] Anonymous. Reflections. IEEE Spectrum, 23(3):8, March 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Rc

[Ano86-195] Anonymous. Reflections. IEEE Spectrum, 23(5):6–7, May 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Rd

[Ano86-196] Anonymous. Reflections. IEEE Spectrum, 23(7):8–11, July 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Re

[Ano86-197] Anonymous. Reflections. IEEE Spectrum, 23(9):8–11, September 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Rf

[Ano86-198] Anonymous. Reflections. IEEE Spectrum, 23(11):8, November 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:SIa

[Ano86-199] Anonymous. Scanning the Institute. IEEE Spectrum, 23 (1):119, January 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:SIb

[Ano86-200] Anonymous. Scanning the Institute. IEEE Spectrum, 23 (2):99, February 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:SIc

[Ano86-201] Anonymous. Scanning the Institute. IEEE Spectrum, 23(3):103, March 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:SId

[Ano86-202] Anonymous. Scanning the Institute. IEEE Spectrum, 23(4):98, April 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:SIe

[Ano86-203] Anonymous. Scanning the Institute. IEEE Spectrum, 23(5):112, May 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:SIf

[Ano86-204] Anonymous. Scanning the Institute. IEEE Spectrum, 23(6):90, June 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:SIg

[Ano86-205] Anonymous. Scanning the Institute. IEEE Spectrum, 23(7):96, July 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:SIh

[Ano86-206] Anonymous. Scanning the Institute. IEEE Spectrum, 23(8):90, August 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:SIi

[Ano86-207] Anonymous. Scanning the Institute. IEEE Spectrum, 23 (9):84, September 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:SIj

[Ano86-208] Anonymous. Scanning the Institute. IEEE Spectrum, 23 (10):98, October 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:SIk

[Ano86-209] Anonymous. Scanning the Institute. IEEE Spectrum, 23 (11):108, November 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:SIl

[Ano86-210] Anonymous. Scanning the Institute. IEEE Spectrum, 23 (12):72, December 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:SLR

[Ano86-211] Anonymous. A shopping list of requirements. IEEE Spectrum, 23(8):51–55, August 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Sb

[Ano86-212] Anonymous. Speakout. IEEE Spectrum, 23(2):9–11, February 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Sc

[Ano86-213] Anonymous. Speakout. IEEE Spectrum, 23(4):16, April 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Se

[Ano86-214] Anonymous. Speakout. IEEE Spectrum, 23(6):18–19, June 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Sg

[Ano86-215] Anonymous. Speakout. IEEE Spectrum, 23(8):17–19, August 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Sj

[Ano86-216] Anonymous. Speakout. IEEE Spectrum, 23(10):17–19, October 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Sm

[Ano86-217] Anonymous. Speakout. IEEE Spectrum, 23(12):16–16H, December 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:SL

[Ano86-218] Anonymous. Spectral lines. IEEE Spectrum, 23(5):33, May 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:SO

[Ano86-219] Anonymous. Spin-offs. IEEE Spectrum, 23(3):17, March 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Sa

[Ano86-220] Anonymous. Spinoffs. IEEE Spectrum, 23(1):24–25, January 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Sd

[Ano86-221] Anonymous. Spinoffs. IEEE Spectrum, 23(5):22–25, May 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Sf

[Ano86-222] Anonymous. Spinoffs. IEEE Spectrum, 23(7):26–27, July 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Sh

[Ano86-223] Anonymous. Spinoffs. IEEE Spectrum, 23(9):26–27, September 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Sk

[Ano86-224] Anonymous. Spinoffs. IEEE Spectrum, 23(11):20–23, November 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:TSa

[Ano86-225] Anonymous. Technically speaking. IEEE Spectrum, 23(1): 16–17, January 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:TSb

[Ano86-226] Anonymous. Technically speaking. IEEE Spectrum, 23(3):16, March 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:TSc

[Ano86-227] Anonymous. Technically speaking. IEEE Spectrum, 23(5): 20–21, May 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:TSd

[Ano86-228] Anonymous. Technically speaking. IEEE Spectrum, 23(7): 24–25, July 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:TSe

[Ano86-229] Anonymous. Technically speaking. IEEE Spectrum, 23(9): 20–23, September 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:TSf

[Ano86-230] Anonymous. Technically speaking. IEEE Spectrum, 23(11): 16, November 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:PFa

[Ano86-231] Anonymous. To probe further. IEEE Spectrum, 23(1): 88–89, January 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:PFb

[Ano86-232] Anonymous. To probe further. IEEE Spectrum, 23(8):57, August 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:PFc

[Ano86-233] Anonymous. To probe further. IEEE Spectrum, 23(11): 80–83, November 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Va

[Ano86-234] Anonymous. Video. IEEE Spectrum, 23(2):24–26, February 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Vb

[Ano86-235] Anonymous. Video. IEEE Spectrum, 23(4):28–30, April 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Vc

[Ano86-236] Anonymous. Video. IEEE Spectrum, 23(6):28–30, June 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Vd

[Ano86-237] Anonymous. Video. IEEE Spectrum, 23(8):30–32, August 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Ve

[Ano86-238] Anonymous. Video. IEEE Spectrum, 23(10):26–28, October 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Vf

[Ano86-239] Anonymous. Video. IEEE Spectrum, 23(12):22–24, December 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:WHa

[Ano86-240] Anonymous. Whatever happened to. IEEE Spectrum, 23 (1):30–31, January 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:WHb

[Ano86-241] Anonymous. Whatever happened to. IEEE Spectrum, 23 (3):24–30, March 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:WHc

[Ano86-242] Anonymous. Whatever happened to. IEEE Spectrum, 23 (5):26–26I, May 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:WHd

[Ano86-243] Anonymous. Whatever happened to. IEEE Spectrum, 23 (7):28–29, July 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:WHe

[Ano86-244] Anonymous. Whatever happened to. IEEE Spectrum, 23 (9):24–25, September 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:WHf

[Ano86-245] Anonymous. Whatever happened to. IEEE Spectrum, 23 (11):17–19, November 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:IGA

[Ano87a] Anonymous. 1987 IEEE guide for authors. IEEE Spectrum, 24 (1):73–76, January 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:Aa

[Ano87b] Anonymous. [advertisements]. IEEE Spectrum, 24(1):1–2, January 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:Ab

[Ano87c] Anonymous. [advertisements].

IEEE Spectrum, 24(1):4–7, January 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:Ac

[Ano87d] Anonymous. [advertisements].

IEEE Spectrum, 24(2):1–2, February 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:Ad

[Ano87e] Anonymous. [advertisements].

IEEE Spectrum, 24(2):4–5, February 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:Ae

[Ano87f] Anonymous. [advertisements].

IEEE Spectrum, 24(2):94, February 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:Af

[Ano87g] Anonymous. [advertisements].

IEEE Spectrum, 24(3):1–2, March 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:Ag

[Ano87h] Anonymous. [advertisements].

IEEE Spectrum, 24(3):4–5, March 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:Ah

[Ano87i] Anonymous. [advertisements]. IEEE Spectrum, 24(4):1–2, April 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:Ai

[Ano87j] Anonymous. [advertisements]. IEEE Spectrum, 24(4):4–7, April 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:Aj

[Ano87k] Anonymous. [advertisements]. IEEE Spectrum, 24(5):1–2, May 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:Ak

[Ano87l] Anonymous. [advertisements]. IEEE Spectrum, 24(5):6–9, May 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:Al

[Ano87m] Anonymous. [advertisements]. IEEE Spectrum, 24(6):1–2, June 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:Am

[Ano87n] Anonymous. [advertisements]. IEEE Spectrum, 24(6):6–9, June 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:An

[Ano87o] Anonymous. [advertisements].

IEEE Spectrum, 24(7):1–2, July 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:Ao

[Ano87p] Anonymous. [advertisements].

IEEE Spectrum, 24(7):6–7, July 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:Ap

[Ano87q] Anonymous. [advertisements].

IEEE Spectrum, 24(8):1–2, August 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:Aq

[Ano87r] Anonymous. [advertisements].

IEEE Spectrum, 24(8):6, August 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:Ar

[Ano87s] Anonymous. [advertisements].

IEEE Spectrum, 24(9):1–2, September 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:As

[Ano87t] Anonymous. [advertisements].

IEEE Spectrum, 24(9):6–7, September 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:Au

[Ano87u] Anonymous. [advertisements]. IEEE Spectrum, 24(10):1–2, October 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:Av

[Ano87v] Anonymous. [advertisements]. IEEE Spectrum, 24(10):6–7, October 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:Aw

[Ano87w] Anonymous. [advertisements]. IEEE Spectrum, 24(11):1–2, November 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:Ax

[Ano87x] Anonymous. [advertisements]. IEEE Spectrum, 24(11):6–7, November 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:Ay

[Ano87y] Anonymous. [advertisements]. IEEE Spectrum, 24(12):1–2, December 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:Az

[Ano87z] Anonymous. [advertisements]. IEEE Spectrum, 24(12):6–9, December 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:Aaa

[Ano87-27] Anonymous. [advertisements].

IEEE Spectrum, 24(12):98, December 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:AT

[Ano87-28] Anonymous. Anatomy of a tragedy. IEEE Spectrum, 24 (2):44–51, February 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:BCa

[Ano87-29] Anonymous. [back cover]. IEEE Spectrum, 24(1):c4, January 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:BCb

[Ano87-30] Anonymous. [back cover]. IEEE Spectrum, 24(2):c4, February 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:BC

[Ano87-31] Anonymous. [back cover]. IEEE Spectrum, 24(3):c4, March 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:BCc

[Ano87-32] Anonymous. [back cover]. IEEE Spectrum, 24(4):c4, April 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:BCd

[Ano87-33] Anonymous. [back cover]. IEEE Spectrum, 24(5):c4, May 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:BCe

[Ano87-34] Anonymous. [back cover]. IEEE Spectrum, 24(6):c4, June 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:BCf

[Ano87-35] Anonymous. [back cover]. IEEE Spectrum, 24(7):c4, July 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:BCg

[Ano87-36] Anonymous. [back cover]. IEEE Spectrum, 24(8):c4, August 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:BCh

[Ano87-37] Anonymous. [back cover]. IEEE Spectrum, 24(9):c4, September 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:BCi

[Ano87-38] Anonymous. [back cover]. IEEE Spectrum, 24(10):c4, October 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:BCj

[Ano87-39] Anonymous. [back cover]. IEEE Spectrum, 24(11):c4, November 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:BCk

[Ano87-40] Anonymous. [back cover]. IEEE Spectrum, 24(12):c4, December 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:BICa

[Ano87-41] Anonymous. [back inside cover].

IEEE Spectrum, 24(1):c3, January 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:BICb

[Ano87-42] Anonymous. [back inside cover].

IEEE Spectrum, 24(2):c3, February 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:BIC

[Ano87-43] Anonymous. [back inside cover].

IEEE Spectrum, 24(3):c3, March 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:BICc

[Ano87-44] Anonymous. [back inside cover].

IEEE Spectrum, 24(4):c3, April 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:BICd

[Ano87-45] Anonymous. [back inside cover]. IEEE Spectrum, 24(5):c3, May 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:BICe

[Ano87-46] Anonymous. [back inside cover]. IEEE Spectrum, 24(6):c3, June 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:BICf

[Ano87-47] Anonymous. [back inside cover]. IEEE Spectrum, 24(7):c3, July 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:BICg

[Ano87-48] Anonymous. [back inside cover]. IEEE Spectrum, 24(8):c3, August 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:BICh

[Ano87-49] Anonymous. [back inside cover]. IEEE Spectrum, 24 (9):c3, September 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:BICi

[Ano87-50] Anonymous. [back inside cover]. IEEE Spectrum, 24(10):c3, October 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:BICj

[Ano87-51] Anonymous. [back inside cover]. IEEE Spectrum, 24 (11):c3, November 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:BICk

[Ano87-52] Anonymous. [back inside cover].

IEEE Spectrum, 24(12):c3, December 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:BRa

[Ano87-53] Anonymous. Book reviews.

IEEE Spectrum, 24(2):14–16, February 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:BRb

[Ano87-54] Anonymous. Book reviews.

IEEE Spectrum, 24(3):12–15, March 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:BRc

[Ano87-55] Anonymous. Book reviews.

IEEE Spectrum, 24(4):12–15, April 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:BRd

[Ano87-56] Anonymous. Book reviews.

IEEE Spectrum, 24(5):13–17, May 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:BRe

[Ano87-57] Anonymous. Book reviews. IEEE Spectrum, 24(6):10–15, June 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:BRf

[Ano87-58] Anonymous. Book reviews. IEEE Spectrum, 24(9):14–17, September 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:BRg

[Ano87-59] Anonymous. Book reviews. IEEE Spectrum, 24(10):12– 15, October 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:BRh

[Ano87-60] Anonymous. Book reviews. IEEE Spectrum, 24(12):13–15, December 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:ISA

[Ano87-61] Anonymous. IEEE Spectrum annual index. IEEE Spectrum, 24 (12):86–96, December 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:Cb

[Ano87-62] Anonymous. Calendar. IEEE Spectrum, 24(1):82–84, January 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:Cd

[Ano87-63] Anonymous. Calendar. IEEE Spectrum, 24(2):70–75, February 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:Cf

[Ano87-64] Anonymous. Calendar. IEEE Spectrum, 24(3):56, March 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:Ch

[Ano87-65] Anonymous. Calendar. IEEE Spectrum, 24(4):62–63, April 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:Cj

[Ano87-66] Anonymous. Calendar. IEEE Spectrum, 24(5):76–77, May 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:Cl

[Ano87-67] Anonymous. Calendar. IEEE Spectrum, 24(6):63–66, June 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:Cn

[Ano87-68] Anonymous. Calendar. IEEE Spectrum, 24(7):56–58, July 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:Cp

[Ano87-69] Anonymous. Calendar. IEEE Spectrum, 24(8):61–62, August 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:CMCa

[Ano87-70] Anonymous. Calendar meetings conferences and conventions. IEEE Spectrum, 24(9): 58–60, September 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:CMCb

[Ano87-71] Anonymous. Calendar: Meetings conferences and conventions. IEEE Spectrum, 24(10): 82–83, October 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:CMCc

[Ano87-72] Anonymous. Calendar: Meetings conferences and conventions. IEEE Spectrum, 24 (11):70, November 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:CMCd

[Ano87-73] Anonymous. Calendar: Meetings conferences and conventions. IEEE Spectrum, 24(12): 65–82, December 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:CSJ

[Ano87-74] Anonymous. Case study: JVC videomovie: Compact and lightweight. IEEE Spectrum, 24(5):35–39, May 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:CFH

[Ano87-75] Anonymous. Challenger’s final hours. IEEE Spectrum, 24(2): 37–39, February 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:CSa

[Ano87-76] Anonymous. Coming in Spectrum. IEEE Spectrum, 24 (1):106, January 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:CSb

[Ano87-77] Anonymous. Coming in Spectrum. IEEE Spectrum, 24 (2):93, February 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:CSc

[Ano87-78] Anonymous. Coming in Spectrum. IEEE Spectrum, 24(3):76, March 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:CSd

[Ano87-79] Anonymous. Coming in Spectrum. IEEE Spectrum, 24(4):78, April 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:CSe

[Ano87-80] Anonymous. Coming in Spectrum. IEEE Spectrum, 24(5):90, May 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:CSf

[Ano87-81] Anonymous. Coming in Spectrum. IEEE Spectrum, 24(6):78, June 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:CSg

[Ano87-82] Anonymous. Coming in Spectrum. IEEE Spectrum, 24(7):70, July 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:CSh

[Ano87-83] Anonymous. Coming in Spectrum. IEEE Spectrum, 24(8):74, August 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:CSi

[Ano87-84] Anonymous. Coming in Spectrum. IEEE Spectrum, 24 (9):76, September 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:CSk

[Ano87-85] Anonymous. Coming in Spectrum. IEEE Spectrum, 24 (10):100, October 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:CSl

[Ano87-86] Anonymous. Coming in Spectrum. IEEE Spectrum, 24 (11):92, November 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:CSm

[Ano87-87] Anonymous. Coming in Spectrum. IEEE Spectrum, 24 (12):97, December 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:Ca

[Ano87-88] Anonymous. Contents. IEEE Spectrum, 24(1):3, January 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:Cc

[Ano87-89] Anonymous. Contents. IEEE Spectrum, 24(2):3, February 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:Ce

[Ano87-90] Anonymous. Contents. IEEE Spectrum, 24(3):3, March 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:Cg

[Ano87-91] Anonymous. Contents. IEEE Spectrum, 24(4):3, April 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:Ci

[Ano87-92] Anonymous. Contents. IEEE Spectrum, 24(5):5, May 1987.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:Ck

[Ano87-93] Anonymous. Contents. IEEE Spectrum, 24(6):5, June 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:Cm

[Ano87-94] Anonymous. Contents. IEEE Spectrum, 24(7):5, July 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:Co

[Ano87-95] Anonymous. Contents. IEEE Spectrum, 24(8):4–5, August 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:Cq

[Ano87-96] Anonymous. Contents. IEEE Spectrum, 24(9):4–5, September 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:Cr

[Ano87-97] Anonymous. Contents. IEEE Spectrum, 24(10):4–5, October 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:Ct

[Ano87-98] Anonymous. Contents. IEEE Spectrum, 24(11):4–5, November 1987. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:Cu

[Ano87-99] Anonymous. Contents. IEEE Spectrum, 24(12):4–5, December 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:CMR

[Ano87-100] Anonymous. Corrections: [misspelling of Richard Feynman’s name]. IEEE Spectrum, 24(4):10, April 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:ETTa

[Ano87-101] Anonymous. EEs’ tools toys.

IEEE Spectrum, 24(1):77–79, January 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:ETTb

[Ano87-102] Anonymous. EEs’ tools toys.

IEEE Spectrum, 24(2):68–69, February 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:ETTc

[Ano87-103] Anonymous. EEs’ tools toys.

IEEE Spectrum, 24(3):53–54, March 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:ETTd

[Ano87-104] Anonymous. EEs’ tools toys.

IEEE Spectrum, 24(4):58–60, April 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:ETTe

[Ano87-105] Anonymous. EEs’ tools toys. IEEE Spectrum, 24(5):73–74, May 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:ETTf

[Ano87-106] Anonymous. EEs’ tools toys. IEEE Spectrum, 24(6):60–62, June 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:ETTg

[Ano87-107] Anonymous. EEs’ tools toys. IEEE Spectrum, 24(7):54–55, July 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:ETTh

[Ano87-108] Anonymous. EEs’ tools toys. IEEE Spectrum, 24(8):58–59, August 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:ETTi

[Ano87-109] Anonymous. EEs’ tools toys. IEEE Spectrum, 24(9):56–57, September 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:ETTj

[Ano87-110] Anonymous. EEs’ tools toys. IEEE Spectrum, 24(10):77– 78, October 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:ETTk

[Ano87-111] Anonymous. EE’s tools toys.

IEEE Spectrum, 24(11):68–69, November 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:ETTl

[Ano87-112] Anonymous. EEs’ tools toys.

IEEE Spectrum, 24(12):61–62, December 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:ELa

[Ano87-113] Anonymous. The engineer at large. IEEE Spectrum, 24(2): 18–19, February 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:ELb

[Ano87-114] Anonymous. The engineer at large. IEEE Spectrum, 24(4):16, April 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:ELc

[Ano87-115] Anonymous. The engineer at large. IEEE Spectrum, 24 (10):20, October 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:ELd

[Ano87-116] Anonymous. The engineer at large. IEEE Spectrum, 24(12):

17–19, December 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:FFa

[Ano87-117] Anonymous. Faults failures. IEEE Spectrum, 24(10):16– 19, October 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:FFb

[Ano87-118] Anonymous. Faults failures. IEEE Spectrum, 24(12):16, December 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:Fa

[Ano87-119] Anonymous. Forum. IEEE Spectrum, 24(1):10, January 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:Fb

[Ano87-120] Anonymous. Forum. IEEE Spectrum, 24(2):6–13, February 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:Fd

[Ano87-121] Anonymous. Forum. IEEE Spectrum, 24(4):10–11, April 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:Fe

[Ano87-122] Anonymous. Forum. IEEE Spectrum, 24(6):16–16H, June 1987.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:Ff

[Ano87-123] Anonymous. Forum. IEEE Spectrum, 24(7):10–11, July 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:Fg

[Ano87-124] Anonymous. Forum. IEEE Spectrum, 24(8):7, August 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:Fh

[Ano87-125] Anonymous. Forum. IEEE Spectrum, 24(10):8–9, October 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:Fi

[Ano87-126] Anonymous. Forum. IEEE Spectrum, 24(12):10–11, December 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:FCa

[Ano87-127] Anonymous. [front cover].

IEEE Spectrum, 24(1):c1, January 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:FCb

[Ano87-128] Anonymous. [front cover]. IEEE Spectrum, 24(2):c1, February 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:FCc

[Ano87-129] Anonymous. [front cover]. IEEE Spectrum, 24(3):c1, March 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:FCd

[Ano87-130] Anonymous. [front cover]. IEEE Spectrum, 24(4):c1, April 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:FCe

[Ano87-131] Anonymous. [front cover]. IEEE Spectrum, 24(5):c1, May 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:FCf

[Ano87-132] Anonymous. [front cover]. IEEE Spectrum, 24(6):c1, June 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:FCg

[Ano87-133] Anonymous. [front cover]. IEEE Spectrum, 24(7):c1, July 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:FCh

[Ano87-134] Anonymous. [front cover]. IEEE Spectrum, 24(8):c1, August 1987.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:FCi

[Ano87-135] Anonymous. [front cover]. IEEE Spectrum, 24(9):c1, September 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:FCj

[Ano87-136] Anonymous. [front cover]. IEEE Spectrum, 24(10):c1, October 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:FCk

[Ano87-137] Anonymous. [front cover]. IEEE Spectrum, 24(11):c1, November 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:FCl

[Ano87-138] Anonymous. [front cover]. IEEE Spectrum, 24(12):c1, December 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:FICa

[Ano87-139] Anonymous. [front inside cover].

IEEE Spectrum, 24(1):c2, January 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:FICb

[Ano87-140] Anonymous. [front inside cover].

IEEE Spectrum, 24(2):c2, February 1987. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:FICc

[Ano87-141] Anonymous. [front inside cover]. IEEE Spectrum, 24(3):c2, March 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:FICd

[Ano87-142] Anonymous. [front inside cover]. IEEE Spectrum, 24(4):c2, April 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:FICe

[Ano87-143] Anonymous. [front inside cover]. IEEE Spectrum, 24(5):c2, May 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:FICf

[Ano87-144] Anonymous. [front inside cover]. IEEE Spectrum, 24(6):c2, June 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:FICg

[Ano87-145] Anonymous. [front inside cover]. IEEE Spectrum, 24(7):c2, July 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:FICh

[Ano87-146] Anonymous. [front inside cover]. IEEE Spectrum, 24(8):c2, August 1987. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:FICi

[Ano87-147] Anonymous. [front inside cover]. IEEE Spectrum, 24 (9):c2, September 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:FICj

[Ano87-148] Anonymous. [front inside cover].

IEEE Spectrum, 24(10):c2, October 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:FICk

[Ano87-149] Anonymous. [front inside cover]. IEEE Spectrum, 24 (11):c2, November 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:FICl

[Ano87-150] Anonymous. [front inside cover].

IEEE Spectrum, 24(12):c2, December 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:HFJ

[Ano87-151] Anonymous. The history of the flawed joint. IEEE Spectrum, 24 (2):39–44, February 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:IAM

[Ano87-152] Anonymous. IEEE awards: 1987 major medalists. IEEE Spectrum, 24(7):52–53, July 1987.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:ITCa

[Ano87-153] Anonymous. IEEE tables of contents. IEEE Spectrum, 24 (1):87–105, January 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:ITCe

[Ano87-154] Anonymous. IEEE tables of contents. IEEE Spectrum, 24 (6):70–77, June 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:ITCb

[Ano87-155] Anonymous. IEEE tables of contents for current and future publications. IEEE Spectrum, 24 (2):76–77, February 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:ITCc

[Ano87-156] Anonymous. IEEE tables of contents for current and future publications. IEEE Spectrum, 24 (3):73–75, March 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:ITCd

[Ano87-157] Anonymous. IEEE tables of contents for current and future publications. IEEE Spectrum, 24(4):64–66, April 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:ITCf

[Ano87-158] Anonymous. IEEE tables of contents for current and future publications. IEEE Spectrum, 24(7):62–69, July 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:ITCg

[Ano87-159] Anonymous. IEEE tables of contents: For current and future publications. IEEE Spectrum, 24 (8):67–73, August 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:ITCh

[Ano87-160] Anonymous. IEEE tables of contents: For current and future publications. IEEE Spectrum, 24 (9):64–66, September 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:ITCi

[Ano87-161] Anonymous. IEEE tables of contents for current and future publications. IEEE Spectrum, 24 (10):86–99, October 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:ITCj

[Ano87-162] Anonymous. IEEE tables of contents for current and future publications. IEEE Spectrum, 24 (11):72–74, November 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:ITCk

[Ano87-163] Anonymous. IEEE tables of contents for current and future publications. IEEE Spectrum, 24 (12):85, December 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:IPPa

[Ano87-164] Anonymous. Innovations patents, processes, and products. IEEE Spectrum, 24(2):20–21, February 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:IPPb

[Ano87-165] Anonymous. Innovations: Patents, processes, and products. IEEE Spectrum, 24(4): 17–19, April 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:IPPc

[Ano87-166] Anonymous. Innovations patents, processes, and products. IEEE Spectrum, 24(10):21, October 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:IPPd

[Ano87-167] Anonymous. Innovations: Patents, processes, and products. IEEE Spectrum, 24(12): 20–21, December 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:MTa

[Ano87-168] Anonymous. Managing technology. IEEE Spectrum, 24 (3):20–22, March 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:MTb

[Ano87-169] Anonymous. Managing technology. IEEE Spectrum, 24(9): 22–24, September 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:NRIa

[Ano87-170] Anonymous. New and recent IEEE publications. IEEE Spectrum, 24(1):85–86, January 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:NRIb

[Ano87-171] Anonymous. New and recent IEEE publications. IEEE Spectrum, 24(2):79–92, February 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:NRIc

[Ano87-172] Anonymous. New and recent IEEE publications. IEEE Spectrum, 24(3):57–72, March 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:NRId

[Ano87-173] Anonymous. New and recent IEEE publications. IEEE Spectrum, 24(4):67–77, April 1987.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:NRIe

[Ano87-174] Anonymous. New and recent IEEE publications. IEEE Spectrum, 24(5):78–89, May 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:NRIf

[Ano87-175] Anonymous. New and recent IEEE publications. IEEE Spectrum, 24(6):68–69, June 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:NRIg

[Ano87-176] Anonymous. New and recent IEEE publications. IEEE Spectrum, 24(7):60–61, July 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:NRIh

[Ano87-177] Anonymous. New and recent IEEE publications. IEEE Spectrum, 24(8):63–66, August 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:NRIi

[Ano87-178] Anonymous. New and recent IEEE publications. IEEE Spectrum, 24(9):61–63, September 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:NRIj

[Ano87-179] Anonymous. New and recent IEEE publications. IEEE Spectrum, 24(10):84–85, October 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:NRIk

[Ano87-180] Anonymous. New and recent IEEE publications. IEEE Spectrum, 24(11):75–91, November 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:NRIl

[Ano87-181] Anonymous. New and recent IEEE publications. IEEE Spectrum, 24(12):83–84, December 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:NLa

[Ano87-182] Anonymous. News log. IEEE Spectrum, 24(5):3–4, May 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:NLb

[Ano87-183] Anonymous. News log. IEEE Spectrum, 24(6):3–4, June 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:NLc

[Ano87-184] Anonymous. News log. IEEE Spectrum, 24(7):3–4, July 1987.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:NLd

[Ano87-185] Anonymous. News log. IEEE Spectrum, 24(8):3, August 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:NLe

[Ano87-186] Anonymous. News log. IEEE Spectrum, 24(9):3, September 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:NLf

[Ano87-187] Anonymous. News log. IEEE Spectrum, 24(10):3, October 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:NLg

[Ano87-188] Anonymous. News log. IEEE Spectrum, 24(11):3, November 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:NLh

[Ano87-189] Anonymous. News log. IEEE Spectrum, 24(12):3, December 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:PIa

[Ano87-190] Anonymous. Papers are invited. IEEE Spectrum, 24(1):

80–81, January 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:PIb

[Ano87-191] Anonymous. Papers are invited.

IEEE Spectrum, 24(2):78, February 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:PIc

[Ano87-192] Anonymous. Papers are invited. IEEE Spectrum, 24(3):55, March 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:PId

[Ano87-193] Anonymous. Papers are invited. IEEE Spectrum, 24(4):61, April 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:PIe

[Ano87-194] Anonymous. Papers are invited. IEEE Spectrum, 24(5):75, May 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:PIf

[Ano87-195] Anonymous. Papers are invited. IEEE Spectrum, 24(6):67, June 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:PIg

[Ano87-196] Anonymous. Papers are invited. IEEE Spectrum, 24(7):59, July 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:PIh

[Ano87-197] Anonymous. Papers are invited. IEEE Spectrum, 24(8):60, August 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:PIi

[Ano87-198] Anonymous. Papers are invited. IEEE Spectrum, 24(9): 67–75, September 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:PIj

[Ano87-199] Anonymous. Papers are invited. IEEE Spectrum, 24(10):81, October 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:PIk

[Ano87-200] Anonymous. Papers are invited. IEEE Spectrum, 24 (11):71, November 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:PIl

[Ano87-201] Anonymous. Papers are invited. IEEE Spectrum, 24(12): 63–64, December 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:PWC

[Ano87-202] Anonymous. A portfolio of wartime components. IEEE Spectrum, 24(11):52–53, November 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:PNSa

[Ano87-203] Anonymous. Program notes software news. IEEE Spectrum, 24 (2):22–23, February 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:PNSb

[Ano87-204] Anonymous. Program notes:

Software topics. IEEE Spectrum, 24(4):20–23, April 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:PNSc

[Ano87-205] Anonymous. Program notes:

Software topics. IEEE Spectrum, 24(10):22–27, October 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:PNSd

[Ano87-206] Anonymous. Program notes:

Software topics. IEEE Spectrum, 24(12):22–25, December 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:R

[Ano87-207] Anonymous. Reflections. IEEE Spectrum, 24(7):8–9, July 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:RCE

[Ano87-208] Anonymous. Reflections: Celebrities as engineers. IEEE Spectrum, 24(9):8–11, September 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:SIa

[Ano87-209] Anonymous. Scanning the Institute. IEEE Spectrum, 24 (1):106, January 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:SIb

[Ano87-210] Anonymous. Scanning the Institute. IEEE Spectrum, 24 (2):93, February 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:SIc

[Ano87-211] Anonymous. Scanning the Institute. IEEE Spectrum, 24(3):76, March 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:SId

[Ano87-212] Anonymous. Scanning the Institute. IEEE Spectrum, 24(4):78, April 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:SIe

[Ano87-213] Anonymous. Scanning the Institute. IEEE Spectrum, 24(5):90, May 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:SIf

[Ano87-214] Anonymous. Scanning the Institute. IEEE Spectrum, 24(6):78, June 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:SIg

[Ano87-215] Anonymous. Scanning the Institute. IEEE Spectrum, 24(7):70, July 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:SIh

[Ano87-216] Anonymous. Scanning the Institute. IEEE Spectrum, 24(8):74, August 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:SIi

[Ano87-217] Anonymous. Scanning the Institute. IEEE Spectrum, 24 (9):76, September 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:SIj

[Ano87-218] Anonymous. Scanning the Institute. IEEE Spectrum, 24 (10):100, October 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:SIk

[Ano87-219] Anonymous. Scanning the Institute. IEEE Spectrum, 24 (11):92, November 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:SIl

[Ano87-220] Anonymous. Scanning the Institute. IEEE Spectrum, 24 (12):97, December 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:SRa

[Ano87-221] Anonymous. Software reviews. IEEE Spectrum, 24(8):16, August 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:SRb

[Ano87-222] Anonymous. Software reviews. IEEE Spectrum, 24(11): 14–15, November 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:Sa

[Ano87-223] Anonymous. Speakout. IEEE Spectrum, 24(2):17, February 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:Sb

[Ano87-224] Anonymous. Speakout. IEEE Spectrum, 24(4):8–9, April 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:STC

[Ano87-225] Anonymous. Speakout: a time for cooperation. IEEE Spectrum, 24(10):10–11, October 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:SMB

[Ano87-226] Anonymous. Speakout: Missing the big view. IEEE Spectrum, 24(12):12, December 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:SO

[Ano87-227] Anonymous. Spin-offs. IEEE Spectrum, 24(3):19, March 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:SON

[Ano87-228] Anonymous. Spin-offs: Navy sensor on the brain. IEEE Spectrum, 24(9):20–21, September 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:TSa

[Ano87-229] Anonymous. Technically speaking. IEEE Spectrum, 24(3): 16–17, March 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:TSb

[Ano87-230] Anonymous. Technically speaking. IEEE Spectrum, 24(9): 18, September 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:T

[Ano87-231] Anonymous. Technology ’87.

IEEE Spectrum, 24(1):26–27, January 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:PFa

[Ano87-232] Anonymous. To probe further. IEEE Spectrum, 24(1): 71–72, January 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:PFb

[Ano87-233] Anonymous. To probe further. IEEE Spectrum, 24 (5):71–72, May 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:TPA

[Ano87-234] Anonymous. Twelve prestigious awards are announced: Winners of the IEEE’s specialist awards cover the fields of precision measurements, power transmission and distribution, electromagnetics, and supercomputer compilers. IEEE Spectrum, 24 (8):56–57, August 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:VPTa

[Ano87-235] Anonymous. Video programs and technology. IEEE Spectrum, 24(2):23–26, February 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:VPTb

[Ano87-236] Anonymous. Video: Programs and technology. IEEE Spectrum, 24(4):24–26, April 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:VPTc

[Ano87-237] Anonymous. Video programs and technology. IEEE Spectrum, 24(10):28–30, October 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:VPTd

[Ano87-238] Anonymous. Video: Programs and technology. IEEE Spectrum, 24(12):26–28, December 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:WHA

[Ano87-239] Anonymous. Whatever happened to analog computers? IEEE Spectrum, 24(3):18, March 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:WHM

[Ano87-240] Anonymous. Whatever happened to: Microwave powered aircraft? IEEE Spectrum, 24 (9):19, September 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1988:AR

[Ano88a] Anonymous. Ads to remember. IEEE Spectrum, 25(11): 124–125, November 1988. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1988:Aa

[Ano88b] Anonymous. [advertisement].

IEEE Spectrum, 25(11):1, November 1988. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1988:Ab

[Ano88c] Anonymous. [advertisement]. IEEE Spectrum, 25(11):5, November 1988. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1988:ACS

[Ano88d] Anonymous. Among the classics: Salient events, inventions, discoveries, successes, and triumphs of man over matter in this exciting quarter century. IEEE Spectrum, 25(11):76, November 1988. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1988:BC

[Ano88e] Anonymous. [back cover]. IEEE Spectrum, 25(11):c4, November 1988. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1988:BIC

[Ano88f]  Anonymous.  [back inside

cover]. IEEE Spectrum, 25 (11):c3, November 1988. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1988:BBSa

[Ano88g] Anonymous. Bits bites of Spectrum. IEEE Spectrum, 25(11): 6–54, November 1988. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1988:BBSb

[Ano88h] Anonymous. Bits bites of Spectrum. IEEE Spectrum, 25(11): 142–202, November 1988. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1988:C

[Ano88i] Anonymous. Contents. IEEE Spectrum, 25(11):2, November 1988. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1988:FUE

[Ano88j] Anonymous. A few uncommon engineers: They include giants, heroes, millionaires, and mavericks some in the public eye and others who are not. IEEE Spectrum, 25(11):56–69, November 1988. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1988:FWS

[Ano88k] Anonymous. Firsts and worsts:

Some news was good, some bad, some mixed: Remember these? IEEE Spectrum, 25(11): 126–131, November 1988. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1988:FIC

[Ano88l]  Anonymous.  [front inside

cover]. IEEE Spectrum, 25 (11):c2, November 1988. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1988:MSE

[Ano88m] Anonymous. Milestones: Salient events, inventions, discoveries, successes, and triumphs of man over matter in this exciting quarter century. IEEE Spectrum, 25 (11):70–75, November 1988. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1988:S

[Ano88n] Anonymous.  Staff.  IEEE

Spectrum, 25(11):3–4, November 1988. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1988:APW

[Ano88o] Anonymous. Up in arms: a profile of world military trade. IEEE Spectrum, 25(12):36–37, November 1988. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1988:WIW

[Ano88p] Anonymous. What’s UP and down: Insidious as well as blatant changes have occurred in 25 years; our inventory reveals some surprises. IEEE Spectrum, 25(11):120–123, November 1988. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Antinone:1987:HPC

[Ant87] R. J. Antinone. How to prevent circuit zapping: Modern chips are extrasensitive to electrical overstress, but the underlying physics is well understood, giving designers a growing arsenal of protective techniques. IEEE Spectrum, 24 (4):34–38, April 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Apt:1985:PPG

[Apt85] C. M. Apt. Perfecting the picture: a guide is offered for handicapping the race for first place in flat-panel displays; none can yet match CRTs for price and performance. IEEE Spectrum, 22(7):60–66, July 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Asal:1987:CSM

[Asa87] T. Asal. Case study: Miniature compact disk players: Help from the factory. IEEE Spectrum, 24(5):44–46, May 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Augustine:1982:BMH

[Aug82] N. R. Augustine. Brilliant missiles on the horizon. IEEE Spectrum, 19(10):96–99, October 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1985:MBC

[AW85] J. A. Adam and P. Wallich.

Mind-boggling complexity: The challenge is to design an intelligent system that finds and destroys targets in a nuclear battlefield. IEEE Spectrum, 22(9): 36–46, September 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bradt:1983:MHS

[BA83] L. J. Bradt and J. K. Allred. Materials-handling systems: Automating of materials handling can integrate otherwise isolated islands of automated machine tools. IEEE Spectrum, 20 (5):74–77, May 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1986:CCO

[BA86] T. E. Bell and F. T. Andrews. Communications: Coherent optical communication shows promise, the FCC continues on its path of deregulation, and satellite communications go high-frequency. IEEE Spectrum, 23(1):49–52, January 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1989:TAM

[BA89] T. E. Bell and J. A. Adam. Technology ’89: aerospace and military. IEEE Spectrum, 26 (1):64–66, January 1989. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Backus:1980:PIC

[Bac80] C. E. Backus. Photovoltaics III: Concentrators: Focusing the Sun’s rays, using cheap, longlived optics, will be exploited in direct sunlight regions. IEEE Spectrum, 17(2):34–36, February 1980. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Backus:1982:SFL

[Bac82] J. Backus. Software I: Functionlevel computing: a new programming method, linked to radically different architectures, may greatly simplify software development. IEEE Spectrum, 19 (8):22–27, August 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bagdikian:1984:EHT

[Bag84] B. H. Bagdikian. Education:

How technology influences journalism: Minicameras and microwave dishes have facilitated the gathering and transmission of news, but have the city editors and TV news producers raised the quality of the product they deliver? IEEE Spectrum, 21 (6):109–111, June 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Baliga:1981:PSS

[Bal81] B. J. Baliga. Power semiconductors: Switching lots of watts at high speeds: State-of-the-art development yields solid-state devices capable of handling over 10 million watts. IEEE Spectrum, 18(12):42–48, December 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Baldwin:1984:EMI

[Bal84a] L. V. Baldwin. The education marketplace: Instructional television: Last year 70 000 engineers enrolled in 2000 courses delivered by video technology; over 3000 master of science degrees were earned via TV in the last two decades. IEEE Spectrum, 21(11):110–114, November 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Baldwin:1984:EUE

[Bal84b] L. V. Baldwin. An electronic university: Engineers across the United States can earn master’s degrees by enrolling in a new television-based institution. IEEE Spectrum, 21(11): 108–110, November 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Baldauf:1985:WCN

[Bal85] D. R. Baldauf. The workhorse CRT: New life: This old standby is still first in monochrome text displays, and color CRTs are widely chosen for graphics. IEEE Spectrum, 22(7):67–73, July 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Barth:1981:SSS

[Bar81] P. W. Barth. Solid state: Silicon sensors meet integrated circuits: The combination of sensing and circuit technology in monolithic form is making transducers cheaper and more versatile. IEEE Spectrum, 18(9): 33–39, September 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1987:IPPb

[BBJ87] T. E. Bell, R. H. Baer, and H. Jensen. Innovations: patents, processes, and products. IEEE Spectrum, 24(8):20–21, August 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bar-Chalm:1982:SSI

[BCUY82] N. Bar-Chalm, I. Ury, and A. Yariv. Solid state: Integrated optoelectronics: The marriage of lasers, detectors, and transistors in a single monolithic package promises fact, reliable data transmission. IEEE Spectrum, 19(5):38–45, May 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bollinger:1983:SAL

[BD83] J. G. Bollinger and N. A. Duffie.

Sensors and actuators: Lowcost sensors linearize, condition, and preprocess inputs from measured parameters. New actuators provide precise position control. IEEE Spectrum, 20 (5):70–73, May 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1987:FFF

[BE87] T. E. Bell and K. Esch. The fatal flaw in flight 511: Events leading up to the ill-fated Challenger launch proved more of a surprise than the disaster itself for what they revealed of NASA’s inability to correct obvious design errors. IEEE Spectrum, 24 (2):36–37, February 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1989:SSC

[BE89] T. E. Bell and K. Esch. The Space Shuttle: a case of subjective engineering. IEEE Spectrum, 26(6):42–46, June 1989. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Beeby:1983:HIS

[Bee83] W. P. Beeby. The heart of integration: a sound data base: Establishing a computer-accessible set of facts about design and manufacturing is a major undertaking with potentially great rewards. IEEE Spectrum, 20 (5):44–48, May 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Beggs:1983:STA

[Beg83] J. Beggs. Space tomorrow: The Antarctica model. IEEE Spectrum, 20(9):89–90, September 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Belohoubek:1981:CAM

[Bel81] E. F. Belohoubek. Communications: Advanced microwave circuits: Computer-aided design and innovative physical configurations help increase performance and reduce costs. IEEE Spectrum, 18(2):44–47, February 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1984:RBP

[Bel84a] R. A. Bell. Robert A. Bell on power and energy: Uncertainties and perceptions of risk bold back energy technology exploitation. IEEE Spectrum, 21 (1):39–40, January 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1984:CSG

[Bel84b] T. E. Bell. Communicating:

Students give oral presentations, create visual aids, and learn engineering by writing. IEEE Spectrum, 21(11):59–60, November 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1984:CLD

[Bel84c] T. E. Bell. Communications:

Long-distance fiber-optic networks, direct-broadcast satellites, and low-cost PBXs bring increased communications capacity to customers. IEEE Spectrum, 21(1):53–57, January 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1984:NTLb

[Bel84d] T. E. Bell. The new television:

Looking behind the tube: In this second part of a two-part series, the outlook for emerging directbroadcast satellite technology and entrenched cable television is examined. IEEE Spectrum, 21 (9):52–62, September 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1984:NTLa

[Bel84e] T. E. Bell. The new television: Looking behind the tube: The viewer gets a host of new options with the advent of new frequency allocations. The likely impact on the traditional broadcaster is discussed in this first part of a twopart series. IEEE Spectrum, 21 (8):48–56, August 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1984:SAP

[Bel84f] T. E. Bell. Societal awareness: Programs stress the interaction of engineering, environmental issues, and policy. IEEE Spectrum, 21(11):60–61, November 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1985:CHI

[Bel85a] T. E. Bell. Communications: Highlights include the AT t divestiture, a new long-distance record in fiber-optic communications, local-area networks, and cellular radio. IEEE Spectrum, 22(1):56–59, January 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1985:DDI

[Bel85b] T. E. Bell. The decision to divest: Incredible or inevitable? The contest that ultimately fractured the Bell System was as much a clash between the participants’ deeply held philosophies as it was about antitrust laws. IEEE Spectrum, 22(11):46–55, November 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1985:JRB

[Bel85c] T. E. Bell. Japan reaches beyond silicon: Intent on dominating high-speed computing and optical communications, Japanese laboratories are aggressively studying materials in the gallium arsenide family. IEEE Spectrum, 22(10):46–53, October 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1985:RHP

[Bel85d] T. E. Bell. Robots in the home:

Promises, promises: While great expectations are held for certain robot types, the robots for fun and educational purposes are limited in their adaptability to useful tasks. IEEE Spectrum, 22(5):51–55, May 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1985:WWW

[Bel85e] T. E. Bell. The way we wash.

IEEE Spectrum, 22(5):106–108, May 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1985:WAC

[Bel85f] T. E. Bell. What’s ahead: Choices for tomorrow: Sobering institutional and technical issues will determine the shape of future telecommunications systems. IEEE Spectrum, 22(11): 111–112, November 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1986:SSS

[Bel86a]  C. G. Bell.  A surge for

solid state: Revitalization of the U.S. semiconductor industry may depend on supplying unconnected transistors and letting the buyer specify on-chip interconnections. IEEE Spectrum, 23(4):71–74, April 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1986:OCF

[Bel86b] T. E. Bell. Optical computing: a field in flux. IEEE Spectrum, 23(8):34–38, August 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1986:QBA

[Bel86c] T. E. Bell. The quest for ballistic action: Avoiding collisions during electron transport to increase switching speeds is the goal of the ultimate transistor. IEEE Spectrum, 23(2):36–38, February 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1987:CSV

[Bel87a] C. G. Bell. Case study: VAX:

Offshoot of a sturdy family tree. IEEE Spectrum, 24(5):30, May 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1987:CIS

[Bel87b]  T. E. Bell.  Communica-

tions: ISDN services are introduced worldwide. Coherent optical communications components are developed in Japan. Fiber-optic submarine cable systems to span the Pacific. Additional spectrum space is allocated for mobile communications. IEEE Spectrum, 24 (1):42–44, January 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1987:DDM

[Bel87c]  T. E. Bell.  Defensive de-

signs minimize hazards: To prevent customer injuries and avoid lawsuits, engineers must anticipate bizarre or unexpected ways in which products might be misused. IEEE Spectrum, 24(5):39–40, May 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1987:IPPa

[Bel87d] T. E. Bell. Innovations patents, processes, and products. IEEE Spectrum, 24(6):20–21, June 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1987:IWD

[Bel87e] T. E. Bell. It was the driver and the driven: Electrotechnology helped win the war; in turn, the pressures of battle sharply accelerated electrotechnical developments. IEEE Spectrum, 24 (8):32–36, August 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1987:QSY

[Bel87f] T. E. Bell. Quitting to start your own company: Tensions run high when creative employees leave to set up a business based on knowledge gained and skills honed at their previous employer. IEEE Spectrum, 24(9):52–55, September 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1988:GBC

[Bel88a] C. G. Bell. Gordon Bell calls for a US research network. IEEE Spectrum, 25(2):54–57, February 1988. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1988:FOL

[Bel88b] T. E. Bell. 1970 fiber optics: Low-loss fibers, reliable lasers, and millimeter-wave research joined to create optical communication. IEEE Spectrum, 25(11):97–102, November 1988. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1988:BBP

[Bel88c] T. E. Bell. Bell break-up plus five: mixed reviews. IEEE Spectrum, 25(13):26–31, December 1988. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1988:CBB

[Bel88d]  T. E. Bell.  Choosing the

best battery for portable equipment. IEEE Spectrum, 25 (3):30–35, March 1988. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1988:EPK

[Bel88e] T. E. Bell. Earthquake prediction: keeping an ear to the ground. IEEE Spectrum, 25 (5):47–52, May 1988. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1988:TC

[Bel88f] T. E. Bell. Technology ’88: communications. IEEE Spectrum, 25 (1):41–43, January 1988. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1989:BFP

[Bel89a] T. E. Bell. Back to the future:

a perspective (defense spending). IEEE Spectrum, 26(11): 42–44, November 1989. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1989:DEE

[Bel89b] T. E. Bell. ’Doc’ Edgerton:

an EE for all seasons. IEEE Spectrum, 26(9):52–57, September 1989. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1989:MML

[Bel89c] T. E. Bell. Managing Murphy’s Law: engineering a minimumrisk system. IEEE Spectrum, 26(6):24–27, June 1989. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1989:NGS

[Bel89d] T. E. Bell. Next-generation spacecraft control. IEEE Spectrum, 26(12):34–38, December 1989. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1989:TT

[Bel89e] T. E. Bell. Technology ’89: telecommunications. IEEE Spectrum, 26(1):41–43, January 1989. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bernhard:1980:AAD

[Ber80a] R. Bernhard. Avionics: All digital jets on the horizon: The Boeing 767 is the most talked about new airliner in years because of its ambitious use of digital avionics. IEEE Spectrum, 17 (10):38–41, October 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bernhard:1980:CBT

[Ber80b] R. Bernhard. Computers: Bubbles take on disks: Application advances for bubble memories depend upon increased density and speed, standardization, and support circuitry. IEEE Spectrum, 17(5):30–33, May 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bernhard:1980:CES

[Ber80c] R. Bernhard. Computers: emphasis on software: The microprocessor becomes more versatile while peripheral circuits get smart. IEEE Spectrum, 17 (1):32–37, January 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bernhard:1980:EAC

[Ber80d] R. Bernhard. An electronic advisor/companion: Computers able to function as personal consultants require breakthroughs in the field of artificial intelligence. IEEE Spectrum, 17(9): 39–44, September 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bernhard:1980:NDC

[Ber80e] R. Bernhard. The ‘no-downtime’ computer. IEEE Spectrum, 17 (9):33–38, September 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bernhard:1980:SRA

[Ber80f] R. Bernhard. Software: The Russian algorithm: a second look: What was labeled a breakthrough in linear programming seems now to be largely of theoretical importance. IEEE Spectrum, 17(5):56–57, May 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bernhard:1980:SSL

[Ber80g] R. Bernhard. Solid state looks to VLSI: Meanwhile, uncommitted logic arrays enjoy a revival and new bubble memories are introduced. IEEE Spectrum, 17 (1):44–49, January 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bergen:1981:PES

[Ber81a] L. Bergen. Power/energy: Satellite control of electric power distribution: a single satellite can manage loads, automate distribution, and read all power meters from coast to coast. IEEE Spectrum, 18 (6):43–47, June 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bernhard:1981:CMS

[Ber81b] R. Bernhard. Computers. I. Micros and software. IEEE Spectrum, 18(1):38–41, January 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bernhard:1981:CIM

[Ber81c] R. Bernhard. Computers. II. Minis and mainframes. IEEE Spectrum, 18(1):42–44, January 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bernhard:1981:CIS

[Ber81d] R. Bernhard. Computers III.

Small business machines: Wireless terminals, new computers, and an electronic publishing setup activate small business systems. IEEE Spectrum, 18 (1):45–46, January 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bernhard:1981:LMA

[Ber81e] R. Bernhard. Lessons from the military: Applications, environments, and servicemen pose reliability challenges not found in commerce or industry. IEEE Spectrum, 18(10):73–78, October 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bernhard:1981:MHM

[Ber81f] R. Bernhard. More hardware means less software. IEEE Spectrum, 18(12):30–37, December 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bernhard:1981:OOL

[Ber81g] R. Bernhard. Overlooking the obvious: In large, complex systems, even the most careful designer can lose sight of an obvious error. IEEE Spectrum, 18 (10):85–89, October 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bernhard:1981:RTP

[Ber81h] R. Bernhard. Reliability test procedures: LSI circuits: Fault models make high-volume testing practical. IEEE Spectrum, 18 (10):67–72, October 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bernhard:1981:SSV

[Ber81i] R. Bernhard. Solid state. I. VLSI/LSI. IEEE Spectrum, 18 (1):57–61, January 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bernhard:1981:SSI

[Ber81j] R. Bernhard. Solid state: II. Linear ICs: Bipolar technology still leads CMOS and NMCS in linear integrated circuits, but CMOS is catching up. IEEE Spectrum, 18 (1):62–63, January 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bernhard:1981:SSR

[Ber81k] R. Bernhard. Solid state: The 64-Rkb RAM teaches a VLSI lessen: Solutions to the problems in developing 64-kb dynamic RAMs will hasten the arrival of new VLSI memory technology. IEEE Spectrum, 18 (6):38–42, June 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bernhard:1982:BSS

[Ber82a] R. Bernhard. Breaching system security. IEEE Spectrum, 19(6):24–31, June 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bernhard:1982:CQO

[Ber82b] R. Bernhard. Communications:

The quandary of office automation: Systems designers are heatedly debating the most efficient communication hardware for the electronic office of the future. IEEE Spectrum, 19(9): 34–39, September 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bernhard:1982:CCS

[Ber82c] R. Bernhard. Computers: Computing at the speed limit: Computers 1000 times faster than today’s supercomputers would benefit vital scientific applications. IEEE Spectrum, 19 (7):26–31, July 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bernhard:1982:EC

[Ber82d] R. Bernhard. Electronic countermeasures. IEEE Spectrum, 19 (10):59–62, October 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bernhard:1982:GSP

[Ber82e]  R. Bernhard.  Giants in

small packages [array processors]. IEEE Spectrum, 19(2): 39–44, February 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bernhard:1982:LIC

[Ber82f]  R. Bernhard.  Linear ICs:

Converters, with system-level components, are more sophisticated; op amps feature low noise. IEEE Spectrum, 19(1): 54–55, January 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bernhard:1982:SBM

[Ber82g] R. Bernhard. Small-business machines: a management-oriented software package and standalone information graphics systems are among major new offerings. IEEE Spectrum, 19 (1):37–38, January 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bernhard:1982:SSR

[Ber82h] R. Bernhard. Solid state: Rethinking the 256-kb RAM: Creativity in the factory may be more important than innovation in product design. IEEE Spectrum, 19(5):46–51, May 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bernhard:1982:TMM

[Ber82i] R. Bernhard. Technology ’82. Minis and mainframes. IEEE Spectrum, 19(1):34–33, January 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bernhard:1982:VLC

[Ber82j] R. Bernhard. VLSI/LSI components: ICs are denser and faster than was thought achievable through state-of-the-art processing and computer-aided design. IEEE Spectrum, 19 (1):49–53, January 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bereiter:1983:EED

[Ber83a] S. Bereiter. Education: Engineers with a difference: University students learn to deal with issues and problems that arise from the interaction of technology and society. IEEE Spectrum, 20(2):63–66, February 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bereiter:1983:PCI

[Ber83b]  S. Bereiter.  The personal

computer invades higher education. IEEE Spectrum, 20 (6):59–61, June 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bristol:1987:OHF

[BFB87] R. Bristol, G. Fladstol, and C. Brannon. Oscilloscopes: How to find the one that fits: New options, from scopes on PC cards to glitch detection, may at first bewilder users but will ultimately help in matching instruments to applications as never before. IEEE Spectrum, 24 (6):34–39, June 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bondurant:1989:FNR

[BG89] D. W. Bondurant and F. P. Gnadinger. Ferroelectrics for nonvolatile RAMs. IEEE Spectrum, 26(7):30–33, July 1989. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bhasavanich:1985:ATJ

[Bha85] D. Bhasavanich. An American in Tokyo: Jumping to the Japanese beat: During a year’s internship, an engineer from Westinghouse learns to move in sync with the hard-driving members of the Mitsubishi team. IEEE Spectrum, 22(9):72–83, September 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1988:IBH

[BJ88] T. E. Bell and P. Janowski. The image benders: How the arts and media depict the engineer: Is the public deceived? IEEE Spectrum, 25(11):132–136, November 1988. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Borriello:1985:VSD

[BKBC85] G. Borriello, R. H. Katz, A. G. Bell, and L. Conway. VLSI system design by the numbers: Design frames and system kits promise to simplify the development of VLSI-based systems. IEEE Spectrum, 22(2): 44–50, February 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Booth:1983:CCA

[BKW83] K. S. Booth, D. H. Kochanek, and M. Wein. Computers: Computers animate films and video: Computers permit spectacular visual effects and take the drudgery out of film cartooning. IEEE Spectrum, 20 (2):44–51, February 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Blumfield:1980:PBF

[Blu80]  C. W. Blumfield.  Power:

Britain’s fast breeder: The Dounreay Test Reactor experience has instilled confidence in its designers to press for commercialization. IEEE Spectrum, 17(4):35–39, April 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bernhard:1981:MP

[BM81] R. Bernhard and N. Mokhoff.

The mission profile. IEEE Spectrum, 18(10):47–49, October 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1983:TPJ

[BMoF^{+}83] T. E. Bell, T. Moto-oka, M. A.

Fischetti, R. S. Cooper, R. E. Kahn, E. Bloch, J. D. Meindl, R. B. Fair, B. W. Oakley, and H. Nasko. The teams and the players: Japan, the United States, and Western Europe provide striking contrasts in national interests and techniques: Strategies. IEEE Spectrum, 20 (11):45–46, November 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1988:SO

[BMR88] C. G. Bell, G. S. Miranker, and J. J. Rubinstein. Supercomputing for one. IEEE Spectrum, 25(4):46–50, April 1988. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Borgatta:1983:COC

[Bor83] L. S. Borgatta. Chips oust clips. IEEE Spectrum, 20 (4):42–47, April 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Borrill:1989:HSB

[Bor89] P. L. Borrill. High-speed 32-bit buses for forward-looking computers. IEEE Spectrum, 26 (7):34–37, July 1989. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bottomley:1983:NMR

[Bot83] P. A. Bottomley. Nuclear magnetic resonance: Beyond physical imaging: a powerful new diagnostic tool that uses magnetic fields and radio waves creates pictures of the body’s internal chemistry. IEEE Spectrum, 20(2):32–38, February 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Botez:1985:LDP

[Bot85] D. Botez. Laser diodes are power-packed: Single-mode laser diodes have a rosy future, both as individual sources of power up to 100 milliwatts and in arrays promising up to half a watt. IEEE Spectrum, 22 (6):43–53, June 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bowlby:1985:DMT

[Bow85] R. Bowlby. The DIP may take its final bows: The dualin-line package, the reigning IC package for several generations, is losing position to newcomers for packaging advanced chips. IEEE Spectrum, 22 (6):37–42, June 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Boxleitner:1989:HDE

[Box89] W. Boxleitner. How to defeat electrostatic discharge. IEEE Spectrum, 26(8):36–40, August 1989. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Booth:1984:TPT

[BPEB84] T. L. Booth, T. L. Powers, J. M.

Early, and R. A. Bell. Technology ’84: the penetrating technologies: Am analysis by the experts: Taylor L. Booth on computers: Software is the limitation. IEEE Spectrum, 21 (1):36–37, January 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bronicki:1981:PES

[Bro81a] Y. L. Bronicki. Power/energy:

a solar-pond power plant: Israel reviews an experiment which got under way more then a year ago: Will it prove a model for more like it? IEEE Spectrum, 18 (2):56–59, February 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Brown:1981:SSS

[Bro81b] W. L. Brown. Solid state: Superfast annealing: Application of laser and electron-beam annealing portends three-dimensional ICs and economical photoelectric cells. IEEE Spectrum, 18 (4):50–56, April 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Brookner:1988:SRR

[Bro88] E. Brookner. Spaceborne radar remote sensing: applications and techniques [book review]. IEEE Spectrum, 25(12):17, November 1988. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bruno:1987:MCD

[Bru87] R. Bruno. Making compact disks interactive. IEEE Spectrum, 24 (11):40–46, November 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bryson:1984:FTE

[Bry84] E. C. Bryson. Frederick E. Terman: Educator and mentor: This popular professor’s influence extended far beyond the Stanford University campus; it reached into industry, too. IEEE Spectrum, 21 (3):71–73, March 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Burns:1981:HSD

[BS81] D. E. Burns and T. C. Spang.

How systems differ: Telephone system: Electronic switching, redundancy, and automatic recovery reduce outages. IEEE Spectrum, 18(10):55–67, October 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Brusil:1989:TUT

[BS89] P. J. Brusil and D. P. Stokesberry. Toward a unified theory of managing large networks. IEEE Spectrum, 26 (4):39–42, April 1989. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Buchheim:1982:TMV

[Buc82] R. W. Buchheim. Treaty making and verification. IEEE Spectrum, 19(10):105–108, October 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Buchheim:1983:MMS

[Buc83] R. W. Buchheim. Minimizing the misuse of space. IEEE Spectrum, 20(9):38–41, September 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bugliarello:1984:LHC

[Bug84] G. Bugliarello. Lifestyle: Healthcare costs: Technology to the rescue? The medical establishment encourages only those new tools that it controls while impeding the implementation of systems that could minimize doctor and hospital costs. IEEE Spectrum, 21(6):97–100, June 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cadogan:1982:PEW

[CA82] J. B. Cadogan and D. F. Ancona. Power/energy: Wind turbines at work: High-capacity wind machines are being integrated with the electric power grid in oil-dependent, wind-rich areas. IEEE Spectrum, 19 (8):46–51, August 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Callaham:1982:MLW

[Cal82] M. B. Callaham. Military: Laser weapons: a variety of targets might be neutralized by laser beams, but only if weapons employing them can be utilized with great finesse. IEEE Spectrum, 19(3):51–55, March 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Carlson:1980:PVA

[Car80] D. E. Carlson. Photovoltaics V: amorphous silicon cells: a highly homogeneous amorphous semiconductor material promises lowcost solar cells. IEEE Spectrum, 17(2):39–41, February 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Carlson:1983:AGE

[Car83] W. B. Carlson. Among the giants: Elihu Thomson: Man of many facets: This foremost inventor and industrialist also directed a corporate engineering and research department and was president of MIT. IEEE Spectrum, 20(10):72–75, October 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Carmichael:1984:DEC

[Car84] D. Carmichael. Do EE careers suffer from poor management? IEEE Spectrum, 21 (7):50–54, July 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Casasent:1981:DPP

[Cas81] D. Casasent. Data processing:

Pattern recognition: a review: Optical techniques developed to discern military targets are being studied for industrial and commercial applications. IEEE Spectrum, 18(3):28–33, March 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Caulfield:1982:THR

[Cau82] H. J. Caulfield. Technology:

Holography: a reassessment: Holography enjoys a modest but secure and steadily growing position in areas, among others, such as nondestructive testing. IEEE Spectrum, 19 (8):39–45, August 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cortes-Comerer:1985:CRA

[CC85] N. Cortes-Comerer. The consumer reels in the aftermath: Changing rules, rising rates, questionable quality, and overwhelming options challenge the residential subscriber. IEEE Spectrum, 22(11):70–77, November 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cortes-Comerer:1986:JMO

[CC86a] N. Cortes-Comerer. JIT is made to order: Beginning with the obvious goal of cutting costs by reducing inventories, just-in-time production now embraces an assortment of sound industrial practices. IEEE Spectrum, 23 (9):57–62, September 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cortes-Comerer:1986:VGS

[CC86b] N. Cortes-Comerer. A venerable giant sharpens its claws: General Electric’s drive to increase profits and market share means drastic restructuring in product lines, employment, and manpower utilization. IEEE Spectrum, 23 (2):54–65, February 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cortes-Comerer:1987:MSG

[CC87] N. Cortes-Comerer. Motto for specialists: Give some, get some: Parallel development of products and their manufacturing processes, within ever shorter lead times, is a formidable challenge for both designers and management. IEEE Spectrum, 24(5):41, May 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cendes:1989:UMM

[Cen89] Z. J. Cendes. Unlocking the magic of Maxwell’s equations. IEEE Spectrum, 26(4):29–33, April 1989. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Chain:1982:UAF

[Cha82] J. T. Chain. The U.S. air force: Is it vulnerable? IEEE Spectrum, 19(10):85–87, October 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Chen:1982:PEA

[Che82]  A. C. M. Chen.  Power/

energy: Automated power distribution: Increasingly diverse and complex power operation and distribution systems will mean a larger role for microprocessor and communications technologies. IEEE Spectrum, 19 (4):55–60, April 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Chen:1986:CDR

[Che86] P. P. Chen. The compact disk ROM: How it works: an offshoot of the compact digital audio disk, this multimegabyte storage technique is revolutionizing database technology. IEEE Spectrum, 23 (4):43–49, April 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Chen:1987:PSF

[Che87] D. Y. Chen. Power semiconductors: Fast, tough, and compact: MOS field-effect transistors and gate turnoff thyristors are bringing new levels of performance in switching frequencies and current and voltage ratings. IEEE Spectrum, 24(9):30–35, September 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Chen:1989:SAP

[Che89a] K. Chen. Sonar array processing borrows from geophysics. IEEE Spectrum, 26(9):58–59, September 1989. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Chen:1989:ETW

[Che89b] K. T. Chen. Education: trouble waiting in the wings. IEEE Spectrum, 26(11):60–64, November 1989. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Chen:1989:IAS

[Che89c] K. T. Chen. International arms sales sag. IEEE Spectrum, 26 (11):56–60, November 1989. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Chen:1989:SJM

[Che89d] K. T. Chen. The state of Japan’s military art. IEEE Spectrum, 26 (9):28–33, September 1989. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1983:T

[CHG^{+}83] D. Christiansen, E. Herbert, F. Guterl, G. Kaplan, P. Wallich, M. A. Fischetti, C. Truxal, and Tekla S. Perry. Technology ’83. IEEE Spectrum, 20 (1):26–27, January 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Chi:1981:CHD

[Chi81] C. S. Chi. Computers: Higher densities for disk memories: Despite rumors of their impending demise, disk memories are alive and well, and with good reason. IEEE Spectrum, 18 (3):39–43, March 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1980:HFA

[Chr80a] D. Christiansen. Hiring, firing, and age discrimination. IEEE Spectrum, 17(1):29, January 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1980:SLSb

[Chr80b] D. Christiansen. Spectral lines:

a secret that’s no surprise. IEEE Spectrum, 17(7):21, July 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1980:SLA

[Chr80c] D. Christiansen. Spectral lines: Advice on how to run a railroad. IEEE Spectrum, 17(6):21, June 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1980:SLEa

[Chr80d] D. Christiansen. Spectral lines: Engineering education under fire. IEEE Spectrum, 17(8):19, August 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1980:SLEb

[Chr80e] D. Christiansen. Spectral lines: Engineering education under fire: II. IEEE Spectrum, 17 (10):27, October 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1980:SLGa

[Chr80f] D. Christiansen. Spectral lines: Good news amid the bad. IEEE Spectrum, 17(3):23, March 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1980:SLGb

[Chr80g] D. Christiansen. Spectral lines: Good news amid the bad, part II. IEEE Spectrum, 17(5):29, May 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1980:SLH

[Chr80h] D. Christiansen. Spectral lines: Helping the younger world. IEEE Spectrum, 17 (11):25, November 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1980:SLF

[Chr80i] D. Christiansen. Spectral lines:

On falsifying data. IEEE Spectrum, 17(12):19, December 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1980:SLSa

[Chr80j] D. Christiansen. Spectral lines:

Saving energy by design. IEEE Spectrum, 17(4):17, April 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1980:SLTa

[Chr80k] D. Christiansen. Spectral lines:

The tape (information) gap. IEEE Spectrum, 17(2):25, February 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1980:SLTb

[Chr80l] D. Christiansen. Spectral lines:

They said it couldn’t be done … or can it? IEEE Spectrum, 17(9):29, September 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1980:T

[Chr80m] D. Christiansen. Technology ’80. IEEE Spectrum, 17(1): 30–31, January 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1981:I

[Chr81a]  D. Christiansen.  Introduc-

tion. IEEE Spectrum, 18(10): 34–35, October 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1981:SLM

[Chr81b] D. Christiansen. Spectral lines: a measure of class. IEEE Spectrum, 18(6):25, June 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1981:SLAb

[Chr81c] D. Christiansen. Spectral lines: Advice to authors. IEEE Spectrum, 18(11):27, November 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1981:SLE

[Chr81d] D. Christiansen. Spectral lines: an exercise in ethics. IEEE Spectrum, 18(1):35, January 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1981:SLG

[Chr81e] D. Christiansen. Spectral lines: Getting the story straight. IEEE Spectrum, 18(5):25, May 1981.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1981:SLI

[Chr81f] D. Christiansen. Spectral lines:

IEEE examines its heritage. IEEE Spectrum, 18(12):23, December 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1981:SLPa

[Chr81g]  D. Christiansen.  Spectral

lines: Peer review reviewed. IEEE Spectrum, 18(8):21, August 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1981:SLPb

[Chr81h] D. Christiansen. Spectral lines:

Publication policies. IEEE Spectrum, 18(9):25, September 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1981:SLRb

[Chr81i] D. Christiansen. Spectral lines:

Reliability vignettes. IEEE Spectrum, 18(10):33, October 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1981:SLS

[Chr81j] D. Christiansen. Spectral lines:

Science, technology, and the government. IEEE Spectrum, 18(7):23, July 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1981:SLAa

[Chr81k] D. Christiansen. Spectral lines: The anonymity of mice and men. IEEE Spectrum, 18(2):29, February 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1981:SLD

[Chr81l] D. Christiansen. Spectral lines: The density (and frailty) of our history. IEEE Spectrum, 18(3):27, March 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1981:SLRa

[Chr81m] D. Christiansen. Spectral lines: The rebirth of nuclear power? IEEE Spectrum, 18(4):27, April 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1981:T

[Chr81n] D. Christiansen. Technology’81. IEEE Spectrum, 18 (1):36–37, January 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1982:BPS

[Chr82a] D. Christiansen. Beyond peacekeeping: Systematic approaches. IEEE Spectrum, 19(10):110– 111, October 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1982:SLAa

[Chr82b] D. Christiansen. Spectral lines: Arms control. IEEE Spectrum, 19(6):23, June 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1982:SLAb

[Chr82c] D. Christiansen. Spectral lines:

Arms control II. IEEE Spectrum, 19(7):25, July 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1982:SLAc

[Chr82d] D. Christiansen. Spectral lines:

Arms control III. IEEE Spectrum, 19(8):21, August 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1982:SLB

[Chr82e] D. Christiansen. Spectral lines:

Behind the issue. IEEE Spectrum, 19(10):33, October 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1982:SLC

[Chr82f]  D. Christiansen.  Spectral

lines: Computer and math anxiety. IEEE Spectrum, 19 (11):23, November 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1982:SLD

[Chr82g] D. Christiansen. Spectral lines:

Dead end — whose fault? IEEE Spectrum, 19(2):27, February 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1982:SLH

[Chr82h] D. Christiansen. Spectral lines: Hear no data, see no data, speak no data. IEEE Spectrum, 19(5):37, May 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1982:SLSa

[Chr82i] D. Christiansen. Spectral lines: Shortage or underutilization? IEEE Spectrum, 19(3):31, March 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1982:SLSb

[Chr82j] D. Christiansen. Spectral lines: The space shuffle. IEEE Spectrum, 19(9):29, September 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1982:SLWb

[Chr82k] D. Christiansen. Spectral lines: Walking the technology-transfer tightrope. IEEE Spectrum, 19 (12):19, December 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1982:SLWa

[Chr82l] D. Christiansen. Spectral lines: What readers liked. IEEE Spectrum, 19(1):29, January 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1982:SLY

[Chr82m] D. Christiansen. Spectral lines:

You said it, we didn’t. IEEE Spectrum, 19(4):29, April 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1982:T

[Chr82n] D. Christiansen. Technology ’82. IEEE Spectrum, 19(1): 30–31, January 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1983:MEM

[Chr83a] D. Christiansen. Multiple exposure: Multiple exposure. IEEE Spectrum, 20(2):31, February 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1983:SLCa

[Chr83b] D. Christiansen. Spectral lines:

a centennial hall of fame. IEEE Spectrum, 20(4):27, April 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1983:SLL

[Chr83c] D. Christiansen. Spectral lines:

a little more knowledge. IEEE Spectrum, 20(6):29, June 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1983:SLS

[Chr83d] D. Christiansen. Spectral lines:

a space tribute in covers. IEEE Spectrum, 20(9):25, September 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1983:SLCc

[Chr83e] D. Christiansen. Spectral lines: Computers and society. IEEE Spectrum, 20(11):33, November 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1983:SLCb

[Chr83f] D. Christiansen. Spectral lines: Concerned engineers. IEEE Spectrum, 20(7):29, July 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1983:SLE

[Chr83g]  D. Christiansen.  Spectral

lines: Evaluating the competition. IEEE Spectrum, 20(3):29, March 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1983:SLG

[Chr83h] D. Christiansen. Spectral lines: Going on 20. IEEE Spectrum, 20(1):25, January 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1983:SLTa

[Chr83i] D. Christiansen. Spectral lines: Technical illiteracy. IEEE Spectrum, 20(10):25, October 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1983:SLA

[Chr83j] D. Christiansen. Spectral lines:

The automatic factory. IEEE Spectrum, 20(5):33, May 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1983:SLTb

[Chr83k] D. Christiansen. Spectral lines:

Toward a technically literate public. IEEE Spectrum, 20 (12):29, December 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1983:SLW

[Chr83l] D. Christiansen. Spectral lines:

What’s wrong with this picture? IEEE Spectrum, 20(8):31, August 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1984:SLCb

[Chr84a] D. Christiansen. Spectral lines:

Campus vignettes. IEEE Spectrum, 21(11):35, November 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1984:SLCa

[Chr84b] D. Christiansen. Spectral lines:

Caveat lector. IEEE Spectrum, 21(8):25, August 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1984:SLH

[Chr84c] D. Christiansen. Spectral lines:

High-tech publishing. IEEE Spectrum, 21(12):33, December 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1984:SLL

[Chr84d] D. Christiansen. Spectral lines: Lost skill. IEEE Spectrum, 21(5):25, May 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1984:SLN

[Chr84e] D. Christiansen. Spectral lines: Nuclear trials and tribulations. IEEE Spectrum, 21(4):25, April 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1984:SLO

[Chr84f] D. Christiansen. Spectral lines: Opening a time capsule. IEEE Spectrum, 21(2):27, February 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1984:SLSa

[Chr84g] D. Christiansen. Spectral lines: Saving the best. IEEE Spectrum, 21(7):25, July 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1984:SLSb

[Chr84h] D. Christiansen. Spectral lines: Science for life or death. IEEE Spectrum, 21(10):31, October 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1984:SLTb

[Chr84i]  D. Christiansen.  Spectral

lines: Technology transfer at issue. IEEE Spectrum, 21 (9):27, September 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1984:SLI

[Chr84j] D. Christiansen. Spectral lines:

The issues we avoid. IEEE Spectrum, 21(6):25, June 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1984:SLJ

[Chr84k] D. Christiansen. Spectral lines:

The joys and hazards of looking ahead. IEEE Spectrum, 21(1):33, January 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1984:SLTa

[Chr84l] D. Christiansen. Spectral lines:

Through the looking glass. IEEE Spectrum, 21(3):27, March 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1984:T

[Chr84m] D. Christiansen. Technology ’84. IEEE Spectrum, 21(1): 34–35, January 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1985:SLBb

[Chr85a] D. Christiansen. Spectral lines:

Behind the MX. IEEE Spectrum, 22(5):33–34, May 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1985:SLBa

[Chr85b] D. Christiansen. Spectral lines: Beyond divestiture. IEEE Spectrum, 22(1):33, January 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1985:SLD

[Chr85c] D. Christiansen. Spectral lines: Don’t hang up yet. IEEE Spectrum, 22(11):43–44, November 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1985:SLE

[Chr85d] D. Christiansen. Spectral lines: Education: Good news and bad. IEEE Spectrum, 22(6):29, June 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1985:SLYa

[Chr85e] D. Christiansen. Spectral lines: Is the 4-year-baccalaureate obsolete? IEEE Spectrum, 22(8):29, August 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1985:SLJ

[Chr85f] D. Christiansen. Spectral lines: Just for the fun of it. IEEE Spectrum, 22(4):31, April 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1985:SLM

[Chr85g] D. Christiansen. Spectral lines:

Mixed signals. IEEE Spectrum, 22(2):35, February 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1985:SLC

[Chr85h] D. Christiansen. Spectral lines:

‘the challenge of all time’. IEEE Spectrum, 22(9):33, September 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1985:SLWc

[Chr85i] D. Christiansen. Spectral lines:

Wasted talents. IEEE Spectrum, 22(10):33, October 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1985:SLWa

[Chr85j] D. Christiansen. Spectral lines:

What we print. IEEE Spectrum, 22(3):37, March 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1985:SLWb

[Chr85k] D. Christiansen. Spectral lines:

Where are our heroes? IEEE Spectrum, 22(7):29, July 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1985:SLYb

[Chr85l] D. Christiansen. Spectral lines:

You said it. IEEE Spectrum, 22 (12):29, December 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1985:T

[Chr85m] D. Christiansen. Technology ’85. IEEE Spectrum, 22(1): 34–35, January 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1986:SLH

[Chr86a] D. Christiansen. Spectral lines: a helping hand. IEEE Spectrum, 23(6):31, June 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1986:SLN

[Chr86b] D. Christiansen. Spectral lines: a NASA design defect. IEEE Spectrum, 23(4):31, April 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1986:SLS

[Chr86c] D. Christiansen. Spectral lines: a stirring giant. IEEE Spectrum, 23(2):27, February 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1986:SLEb

[Chr86d] D. Christiansen. Spectral lines: Engineering anonymity. IEEE Spectrum, 23(8):33, August 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1986:SLEa

[Chr86e] D. Christiansen. Spectral lines: Enlarging the circle. IEEE Spectrum, 23(1):33, January 1986.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1986:SLL

[Chr86f] D. Christiansen. Spectral lines:

In the line of duty. IEEE Spectrum, 23(3):31, March 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1986:SLG

[Chr86g] D. Christiansen. Spectral lines:

On good bosses. IEEE Spectrum, 23(9):31, September 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1986:SLOb

[Chr86h] D. Christiansen. Spectral lines:

One world. IEEE Spectrum, 23 (12):25, December 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1986:SLOa

[Chr86i] D. Christiansen. Spectral lines:

Our expansive past and explosive future. IEEE Spectrum, 23(10):29, October 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1986:SLV

[Chr86j] D. Christiansen. Spectral lines:

Verification and arms control. IEEE Spectrum, 23(7):33, July 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1986:SLJ

[Chr86k] Donald Christiansen. Spectral lines: From Jennys to jets. IEEE Spectrum, 23(11):35, November 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1987:APE

[Chr87a] D. Christiansen. Age as a pejorative: Ethics: More talk, less action: The technological war. IEEE Spectrum, 24(8):27, August 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1987:SLa

[Chr87b] D. Christiansen. Spectral lines. IEEE Spectrum, 24(1):25, January 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1987:SLb

[Chr87c] D. Christiansen. Spectral lines. IEEE Spectrum, 24(5):25–26, May 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1987:SLc

[Chr87d] D. Christiansen. Spectral lines. IEEE Spectrum, 24(6):27, June 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1987:SLd

[Chr87e] D. Christiansen. Spectral lines. IEEE Spectrum, 24(7):21, July 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1987:SLf

[Chr87f]  D. Christiansen.  Spectral

lines. IEEE Spectrum, 24 (11):23, November 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1987:SLP

[Chr87g] D. Christiansen. Spectral lines: a pervasive machine. IEEE Spectrum, 24(10):31, October 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1987:SLS

[Chr87h] D. Christiansen. Spectral lines:

a system gone awry. IEEE Spectrum, 24(3):23, March 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1987:SLEb

[Chr87i] D. Christiansen. Spectral lines:

an electronic nation. IEEE Spectrum, 24(12):29, December 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1987:SLEa

[Chr87j] D. Christiansen. Spectral lines:

Electromagnetic hazards. IEEE Spectrum, 24(9):25, September 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1987:SLO

[Chr87k] D. Christiansen. Spectral lines: One flight too many. IEEE Spectrum, 24(2):27, February 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1987:SLR

[Chr87l] D. Christiansen. Spectral lines: Radio days. IEEE Spectrum, 24(4):27, April 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1988:BBR

[Chr88a] D. Christiansen. Bigger bangs for the right price. IEEE Spectrum, 25(12):23, November 1988. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1988:PT

[Chr88b] D. Christiansen. A powerful time. IEEE Spectrum, 25 (11):55, November 1988. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Chu:1988:PDS

[Chu88]  D. Chu.  Phase digitizing

sharpens timing measurements. IEEE Spectrum, 25(7):28–32, July 1988. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Clark:1980:CHD

[Cla80a] T. Clark. Communications: How Diana touched the Moon: an impromptu band of Army radio experimenters established Man’s first contact with outer space in 1946. IEEE Spectrum, 17 (5):44–48, May 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Clarke:1980:HCI

[Cla80b] K. E. Clarke. Home communications III: Videotext standards: International experts are hard at work defining standards to interface new services with divergent systems. IEEE Spectrum, 17(3):33–35, March 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Clemens:1982:CEV

[Cle82] J. K. Clemens. Consumer electronics: Video disks: Three choices: Three incompatible approaches to video recording for playback are compared. IEEE Spectrum, 19(3):38–42, March 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cooper:1983:CCM

[CN83] G. R. Cooper and R. W. Nettleton. Communications: Cellular mobile technology: The great multiplier: It can greatly expand mobile communications, but some technical and regulatory issues remain unsolved. IEEE Spectrum, 20(6):30–37, June 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Coggeshall:1984:ECV

[Cog84] I. S. Coggeshall. An electric centura: Variations on a theme by Oppenheimer: Looking both ways from 1984, this celebration of the IEEE’s Centennial addresses the give-and-take between electrotechnology and civilization. IEEE Spectrum, 21 (2):70–86, February 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Conrad:1986:LMC

[Con86] M. Conrad. The lure of molecular computing: While marketable products seem decades away, researchers are crystallizing theories and devices that will give biological organisms the power to compute. IEEE Spectrum, 23(10):55–60, October 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Corwin:1984:CNS

[Cor84] W. L. Corwin. A communications network for the summer olympics: Behind-the-scenes communicating at the games in los angeles seemed a complicated and thankless task, yet AT t took it on. IEEE Spectrum, 21(7):38–44, July 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Corcoran:1987:CCA

[Cor87a] E. Corcoran. Cooperating to compete: Allied companies share expertise, distribute risk, acquire technology, and provide products for niche markets. IEEE Spectrum, 24(10):53–56, October 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Corcoran:1987:MEP

[Cor87b] E. Corcoran. Medical electronics:

Pacemaker responds to changes in body motion worldwide use of artificial hearts increases threedimensional diagnostic imaging matures. IEEE Spectrum, 24 (1):66–68, January 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Corcoran:1987:SIR

[Cor87c] E. Corcoran. She incites revolutions with chips and networks: From chip designs by college students to intelligent computers for the military, Lynn Conway has pioneered some hot technology programs. And she’s at it again! IEEE Spectrum, 24(12): 46–52, December 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Corcoran:1987:SCS

[Cor87d] E. Corcoran. Strategic computing: a status report: In the fouryear-old program that the DoD launched to create intelligent machines, the emphasis is now on technology that will give U.S. industry an edge. IEEE Spectrum, 24(4):50–54, April 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Corcoran:1987:TSa

[Cor87e] E. Corcoran. Technically speaking. IEEE Spectrum, 24(1): 16–17, January 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Corcoran:1987:TSb

[Cor87f] E. Corcoran. Technically speaking. IEEE Spectrum, 24(5): 18–19, May 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Corcoran:1987:TSc

[Cor87g] E. Corcoran. Technically speaking. IEEE Spectrum, 24(7): 14–15, July 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Corcoran:1987:TSd

[Cor87h] E. Corcoran. Technically speaking. IEEE Spectrum, 24(11): 16–24, November 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Corcoran:1988:TME

[Cor88a] E. Corcoran. Technology ’88: medical electronics. IEEE Spectrum, 25(1):65–67, January 1988. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Correspondent:1988:CDA

[Cor88b] F. G. Correspondent. 1983 compact disc: Arch rivals Philips and Sony combined forces to produce an audio breakthrough.

IEEE Spectrum, 25(11):102– 108, November 1988. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Costello:1988:LA

[Cos88] R. B. Costello. A look ahead.

IEEE Spectrum, 25(12):68–69, November 1988. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Corcoran:1986:MEI

[CP86] E. Corcoran and A. R. Potvin.

Medical electronics: Implantable defibrillators, cochlear devices, and lithotripters lead the list of recent developments. IEEE Spectrum, 23(1):82–85, January 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Carley:1988:HAA

[CR88] L. R. Carley and R. A. Rutenbar. How to automate analog IC designs. IEEE Spectrum, 25 (8):26–30, August 1988. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Chou:1987:CVW

[CS87] S. Chou and C. Simonsen. Chip voltage: Why less is better: a reduced power-supply standard is in the offing as manufacturers and designers grapple with the tradeoffs of submicrometer devices. IEEE Spectrum, 24 (4):39–43, April 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cunningham:1980:MUL

[Cun80]  J. A. Cunningham.  Man-

agement: Using the learning curve as a management tool: The learning curve can help in preparing cost reduction programs, pricing forecasts, and product development goals. IEEE Spectrum, 17(6):45–48, June 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Curien:1983:PFN

[Cur83] H. Curien. The pride of France: a national commitment. IEEE Spectrum, 20(9):74–76, September 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cyert:1984:ENT

[Cye84] R. M. Cyert. Education: New teacher’s pet: The computer: Microcomputers are being imaginatively introduced into liberalarts curricula, thereby offering hitherto unavailable data and exercises. IEEE Spectrum, 21 (6):120–122, June 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Davidson:1987:S

[DA87] R. Davidson and A. E. Alcorn. Spinoffs. IEEE Spectrum, 24 (11):18–19, November 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Daniel:1982:WUR

[Dan82] C. D. Daniel. War 5: U.S. readiness: Preparedness of the U.S. military is discussed in essays by an incumbent and three former military leaders; a fifth essay analyzes industry’s role: The U.S Army: Is it ready? IEEE Spectrum, 19(10):80–83, October 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Davies:1983:SSC

[Dav83a] R. D. Davies. Solid state:

The case for CMOS: Its low power consumption and many design choices mean that by the decade’s end, half of all ICs may be made by this technique. IEEE Spectrum, 20(10):26–32, October 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Davis:1983:CA

[Dav83b] A. L. Davis. Computer architecture. IEEE Spectrum, 20(11): 94–99, November 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Davidson:1987:PDP

[Dav87a] R. Davidson. Part I. Defining the product: Product planners assess users’ needs: Designers of commercial military, and consumer systems rely on the customer more than ever to help set design parameters. IEEE Spectrum, 24(5):29–30, May 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Davidson:1987:SOa

[Dav87b] R. Davidson. Spin-offs. IEEE Spectrum, 24(1):18–19, January 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Davidson:1987:SOb

[Dav87c] R. Davidson. Spin-offs. IEEE Spectrum, 24(5):22–23, May 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Davidson:1987:SOc

[Dav87d] R. Davidson. Spin-offs. IEEE Spectrum, 24(7):17, July 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dario:1985:TSG

[DD85] P. Dario and D. De Rossi. Tactile sensors and the gripping challenge: Increasing the performance of sensors over a wide range of force is a first step toward robotry that can hold and manipulate objects as humans do. IEEE Spectrum, 22(8):46–53, August 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dawson:1986:FCD

[DD86] W. K. Dawson and R. W. Dobinson. A framework for computer design: Instead of starting off with a microprocessor, equipment designers must now select from among the many available bus standards before choosing hardware and software. IEEE Spectrum, 23(10):49–54, October 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

deFranco:1982:PES

[de 82] N. de Franco. Power/energy:

The superelectrification of Brazil: a huge hydroelectric complex under construction at Itaipu plus planned long-distance transmission lines signal the country’s ambitions. IEEE Spectrum, 19(5):57–63, May 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

deNeuf:1983:EHH

[de 83] D. K. de Neuf. EE heroes and happenings: a philatelic review: The impending issuance of four new stamps prompts a retrospective look at electrotechnology from the philatelist’s viewpoint. IEEE Spectrum, 20 (8):60–69, August 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

deGeus:1989:LSS

[de 89] A. J. de Geus. Logic synthesis speeds ASIC design. IEEE Spectrum, 26(8):27–31, August 1989. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Delogne:1980:CRG

[Del80] P. Delogne. Communications:

Radio goes underground: Communicating in mine tunnels has been taken seriously by radio and microwave engineers only in the last decade. IEEE Spectrum, 17(7):26–29, July 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

DeLauer:1982:WFM

[DeL82] R. D. DeLauer. War 1: The force multiplier: Advanced technology, and particularly electronics, provides leverage in modern military defense systems. IEEE Spectrum, 19(10):35–37, October 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Denaro:1981:ANA

[Den81] R. P. Denaro. Aerospace: Navstar: The all-purpose satellite: The military sees Navstar as a way of keeping wide-ranging forces on target, and potential civilian users are showing interest too. IEEE Spectrum, 18(5):35–40, May 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

DeVine:1981:NPP

[DeV81] J. C. DeVine. Nuclear power: a progress report: Cleaning up TMI: On the second anniversary of the accident at Three Mile Island’s nuclear reactor, decontamination is progressing slowly. IEEE Spectrum, 18 (3):44–49, March 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Drummond:1988:QTS

[DGZ88] T. J. Drummond, P. L. Gourley, and T. E. Zipperian. Quantum-tailored solid-state devices. IEEE Spectrum, 25 (6):33–37, June 1988. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dinneen:1982:CGV

[Din82] G. P. Dinneen. C3 gets a vote of confidence. IEEE Spectrum, 19 (10):58–59, October 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Draper:1985:PPV

[DJ85] R. F. Draper and T. V. Johnson. Planning a probing voyage to Jupiter: The adroit Galileo spacecraft, crammed with advanced instrumentation, should answer some age-old questions about the mysterious planet. IEEE Spectrum, 22(6):70–76, June 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Daniels:1985:MTM

[DM85] H. A. Daniels and N. Mayur.

More than mainframes: Power utilities are finding that the mixand-match trend that is evident in computer systems is for them a matter of survival. IEEE Spectrum, 22(8):54–61, August 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dongarra:1987:CBP

[DMW87] J. Dongarra, J. L. Martin, and J. Worlton. Computer benchmarking: Paths and pitfalls: The most popular way of rating computer performance can confuse as well as inform; avoid misunderstanding by asking just what the benchmark is measuring. IEEE Spectrum, 24 (7):38–43, July 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dorros:1983:CTN

[Dor83] I. Dorros. Communications: Telephone nets go digital: Digital communication links promise flexible, economical telephone services for both home and office. IEEE Spectrum, 20 (4):48–53, April 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Douglas:1981:CCD

[Dou81] A. Douglas. Communications: The crystal detector: By 1920, G. W. Pickard had tested 31,250 possible combinations of materials in search of a practical detector. IEEE Spectrum, 18 (4):64–69, April 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Douglass:1983:DLN

[Dou83a] R. J. Douglass. A day in the life of a new-generation computer user. IEEE Spectrum, 20 (11):74–75, November 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Douglass:1983:NUT

[Dou83b] R. J. Douglass. Needs and uses: Technology applied in research settings suggests that virtually everyone will use the new-generation computing systems. IEEE Spectrum, 20(11): 41–45, November 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dougherty:1985:JJD

[Dou85]  J. J. Dougherty.  John J.

Dougherty on power and energy: How much power from what sources? IEEE Spectrum, 22(1):41, January 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Doyle:1981:HPF

[Doy81] E. A. Doyle. How parts fail:

a fundamental explanation for any part failure exists, and it’s up to the failure analyst to find the cause with modern sleuthing. IEEE Spectrum, 18 (10):36–43, October 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Doddington:1981:CSR

[DS81] G. R. Doddington and T. B.

Schalk. Computers: Speech recognition: Turning theory to practice: New ICs have brought the requisite computer power to speech technology; an evaluation of equipment shows where it stands today. IEEE Spectrum, 18 (9):26–32, September 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

deMoraes:1983:CMI

[dS83] J. de Moraes and V. F. Salatko. Coming: 12600 megawatts at Itaipu Island. IEEE Spectrum, 20(8):46–52, August 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dalton:1986:HEH

[DTS86] G. W. Dalton, P. H. Thompson, and W. N. Smallwood. Helping engineers help themselves: Although largely an individual responsibility, professional development affects a company’s ability to retain top-notch professionals. IEEE Spectrum, 23(12): 43–47, December 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dupuy:1982:TFS

[Dup82]  J. Dupuy.  Transportation:

France’s superfast train: The French National Railroad’s fleet of 260-km/h electrics competes favorably with the airlines between Paris and Lyons. IEEE Spectrum, 19(7):38–43, July 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Duttweiler:1980:CBE

[Dut80] D. L. Duttweiler. Communications: Bell’s echo-killer chip: Annoying echoes in two-way conversations via satellite are eliminated by using an inexpensive silicon circuit. IEEE Spectrum, 17(10):34–37, October 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Early:1984:JME

[Ear84] J. M. Early. James M. Early on solid state: The theoretical ideals come ever closer. IEEE Spectrum, 21(1):38–39, January 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Eastman:1986:BEC

[Eas86] L. F. Eastman. Ballistic electrons in compound semiconductors. IEEE Spectrum, 23(2): 42–45, February 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Elachi:1982:SIR

[EG82] C. Elachi and J. Granger. Spaceborne imaging radars probe ‘in depth’. IEEE Spectrum, 19(11): 24–29, November 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Eidson:1981:SSF

[Eid81]  J. C. Eidson.  Solid state:

Fast electron-beam lithography: High blanking speeds may make this new system a serious challenger in producing submicrometer ICs. IEEE Spectrum, 18 (7):24–28, July 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Eisenhut:1984:NRV

[Eis84] D. G. Eisenhut. NRC as referee: Voluminous studies of TMI have helped the Nuclear Regulatory Commission generate an action plan to alter existing nuclear reactors. IEEE Spectrum, 21(4):33–39, April 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Eden:1983:SSI

[ELW83] R. C. Eden, A. R. Livingston, and B. M. Welch. Solid state: Integrated circuits: The case for gallium arsenide: This superfast semiconductor holds great promise for high-speed computers, and manufacturing difficulties are now beginning to be overcome. IEEE Spectrum, 20 (12):30–37, December 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Edelson:1982:CGM

[EMM82] B. I. Edelson, R. B. Marsten, and W. L. Morgan. Communications: Greater message capacity for satellites: Techniques such as the use of new frequencies, digital encoding, and bandwidth compression will increase mission efficiency. IEEE Spectrum, 19(3):56–64, March 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Emmett:1985:KAT

[Emm85] A. Emmett. Karmarkar’s algorithm: a threat to simplex? An elegant new method of solving linear equations has yet to be thoroughly tested. IEEE Spectrum, 22(12):54–55, December 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Esch:1986:HNP

[Esc86] K. Esch. How NASA prepared to cope with disaster: Extensive plans were in place to mitigate the effects of a Shuttle failure, whether on the launch pad or aloft. IEEE Spectrum, 23 (3):32–36, March 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fischetti:1984:TPN

[FAEH84] M. A. Fischetti, J. A. Adam, D. G. Eisenhut, and J. Horgan. TMI plus 5: Nuclear power on the ropes: The 1979 accident at the Three Mile Island nuclear plant triggered a domino reaction that threatens the future of the U.S. nuclear power industry. IEEE Spectrum, 21(4):26–27, April 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fagenbaum:1980:CEA

[Fag80a] J. Fagenbaum. Computers: Electronics aids the athlete: Biomechanical systems are useful in human factors engineering, and in the design of sports equipment and prosthetic devices. IEEE Spectrum, 17(7):36–39, July 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fagenbaum:1980:IFE

[Fag80b] J. Fagenbaum. Innovation in the factory: Energy-efficient processes are sought, and the computer is used to merge design and fabrication functions. IEEE Spectrum, 17(1):54–57, January 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fagenbaum:1980:ICS

[Fag80c] J. Fagenbaum. Instrumentation: Controlling the synfuel process: Sophisticated electronic instrumentation and safety controls for coal gasification and liquefaction systems are under development. IEEE Spectrum, 17(11): 31–36, November 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fagenbaum:1980:MPR

[Fag80d] J. Fagenbaum. Magnetohydrodynamic power: The road to an operational magnetohydrodynamic power plant is paved with formidable engineering obstacles. IEEE Spectrum, 17(9): 67–72, September 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fagenbaum:1980:PEC

[Fag80e] J. Fagenbaum. Power/energy: Cogeneration: an energy saver: This old and proven practice is even more attractive as the costs of primary fuels soar. IEEE Spectrum, 17(8):30–34, August 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fagenbaum:1980:TPH

[Fag80f]  J. Fagenbaum.  A tribute:

Patrick E. Haggerty: Engineer and visionary: His colleagues remember him with affection and respect, a persistent, persuasive manager who was an inspiration to his associates. IEEE Spectrum, 17(12):20–23, December 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fagenbaum:1981:CNB

[Fag81a] J. Fagenbaum. Computers: a new breed: The software engineer: As more and better software designers are needed, universities and industry are under pressure to establish programs to train them. IEEE Spectrum, 18 (9):62–67, September 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fagenbaum:1981:MIC

[Fag81b] J. Fagenbaum. Manufacturing. I. Industrial control. IEEE Spectrum, 18(1):70–73, January 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fagenbaum:1981:SSI

[Fag81c] J. Fagenbaum. Solid state.

III. Discrete devices. IEEE Spectrum, 18(1):64–65, January 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fagenbaum:1982:CWL

[Fag82a] J. Fagenbaum. Careers: We look at ourselves: The manufacturing engineer: The best ones are good planners and organizers, with a knack for interfacing with multidisciplined individuals. IEEE Spectrum, 19 (2):64–69, February 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fagenbaum:1982:DDF

[Fag82b] J. Fagenbaum. Discrete devices: Fabrication and processing techniques, borrowed from LSI technology, play major roles in new component designs. IEEE Spectrum, 19(1):56–57, January 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fagenbaum:1982:ICC

[Fag82c] J. Fagenbaum. Industrial controls: Computers and microprocessors lead to improvements in batch-process control, motor inverter systems, and industrial laser equipment. IEEE Spectrum, 19(1):61–64, January 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fan:1989:NAS

[Fan89] J. C. C. Fan. New anatomies for semiconductor wafers. IEEE Spectrum, 26(4):34–38, April 1989. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Farowich:1986:CTO

[Far86] S. A. Farowich. Communicating in the technical office: The Technical and Office Protocols specification, released in November, will be tempered in the fires of realistic applications. IEEE Spectrum, 23(4):63–67, April 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Feinstein:1982:WFW

[Fei82] J. Feinstein. War 6: Future weaponry: Trends in military R&D are reviewed by present and former U.S. defense officials; an expert comments on communications in space: Research thrusts of the U.S. DoD. IEEE Spectrum, 19(10):91–93, October 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Frey:1980:SST

[FEK80] H. H. Frey, P. Etter, and P. Knapp. Solid state: Thyristors on the move: New applications are seen in power converters and inverters for trains, industrial motors, and HVDC. IEEE Spectrum, 17(3):57–62, March 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fischetti:1986:SSE

[FG86] M. A. Fischetti and B. L. Gregory. Solid state: Economics and politics dominate the industry; nonetheless, application-specific and power-logic circuits spark new trends. IEEE Spectrum, 23 (1):53–56, January 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Feigenbaum:1983:BBA

[FHRW^{+}83] E. A. Feigenbaum, F. Hayes-Roth, D. L. Waltz, R. Reddy, V. Zue, and T. Kanade. The building blocks: Artificial intelligence, software engineering, computer architecture, and very large-scale integration will play major roles: Artificial intelligence. IEEE Spectrum, 20(11): 77–91, November 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fischetti:1985:SDH

[FHW85] M. A. Fischetti, J. Horgan, and P. Wallich. The superstructure: Designing for high-tech: The high-tech house is hardly passive; its heating, security, communications, and lighting systems function independently, yet talk with one another at the command of computers. IEEE Spectrum, 22(5):36–40, May 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fisher:1980:LES

[Fis80] P. D. Fisher. Law enforcement: Shortcomings of radar speed measurement: It’s based on sound principles, but present systems have practical limitations and may be misused. IEEE Spectrum, 17(12):28–31, December 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fischetti:1983:ASE

[Fis83a] M. A. Fischetti. The art and science of ergonomics. IEEE Spectrum, 20(9):66–67, September 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fischetti:1983:BEQ

[Fis83b] M. A. Fischetti. Biomedical engineering: The quest for the ultimate artificial heart: The salt lake city achievement is only the beginning; in the offing are smaller, lighter systems, and eventually a fully implantable electrohydraulic heart. IEEE Spectrum, 20(3):39–44, March 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fischetti:1983:PEE

[Fis83c] M. A. Fischetti. Power/energy:

EPRI’s first decade: The power utility industry’s R&D association today concentrates less on long-term research and more on implementable, money-saving technologies. IEEE Spectrum, 20(4):54–57, April 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fischetti:1983:PED

[Fis83d] M. A. Fischetti. Power/energy:

The DOE’s chief on U.S. energy policy: Energy Secretary Donald Hodel explains how laissez-faire and a balanced energy supply are necessary to react quickly in an unstable world. IEEE Spectrum, 20(10):55–59, October 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fischetti:1983:PHB

[Fis83e] M. A. Fischetti. Probing the human body: Imaging allows a safer look inside patients, while the microprocessor aids physicians in rendering diagnoses. IEEE Spectrum, 20(1): 75–78, January 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fischetti:1983:SDE

[Fis83f] M. A. Fischetti. Supplying demand efficiently: Utilities step up direct-current transmission; more alternative energy systems come on line. IEEE Spectrum, 20(1):59–63, January 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fischetti:1983:TCL

[Fis83g] M. A. Fischetti. Telecommunications: Clarke’s legacy. IEEE Spectrum, 20(9):60–62, September 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fischetti:1983:US

[Fis83h] M. A. Fischetti. The United States. IEEE Spectrum, 20(11): 51–69, November 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fischetti:1983:VLC

[Fis83i] M. A. Fischetti. VLSI/LSI components: CMOS gains at the expense of other technologies, as NMOS prepares the way for the 256-kb RAM. IEEE Spectrum, 20(1):43–47, January 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fischetti:1984:BAB

[Fis84a] M. A. Fischetti. Band-aids and better: Hardware, instrumentation, and operations are improving on a piecemeal basis and more systematic studies are under way. IEEE Spectrum, 21 (4):39–43, April 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fischetti:1984:BOD

[Fis84b] M. A. Fischetti. Business outlook: Design work stations, silicon foundries, and a DOD communications network presage a fundamental alteration of the semiconductor industry. IEEE Spectrum, 21(9):47–51, September 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fischetti:1984:CRI

[Fis84c] M. A. Fischetti. But challenges remain: The issues may be clearer and the problems better defined, but solutions may not be simple and will often involve negotiation and compromise. IEEE Spectrum, 21 (4):49–56, April 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fischetti:1984:FCF

[Fis84d] M. A. Fischetti. Fabrication:

chip fabricators run the ‘anchor leg in producing a complex chip in the allotted time: they begin on day 15 and finish on day 31. IEEE Spectrum, 21(9): 35–39, September 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fischetti:1984:PCT

[Fis84e] M. A. Fischetti. Photovoltaiccell technologies joust for position: The fabrication of solar cells has become a bona fide industry with three versions vying to dominate the marketplace. IEEE Spectrum, 21 (3):40–47, March 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fischetti:1984:PES

[Fis84f] M. A. Fischetti. Power and energy: Slowed electricity demand and high interest rates have forced utilities to generate, transmit, and distribute power more efficiently. IEEE Spectrum, 21(1):72–77, January 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fischetti:1984:SSM

[Fis84g] M. A. Fischetti. Solid state: Memories amass more data per chip, while several companies ready the first nonoptical lithography machines. IEEE Spectrum, 21(1):58–63, January 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fischetti:1984:TNE

[Fis84h] M. A. Fischetti. Turning neutrons into electricity: Spurred on by recent scientific triumphs at Princeton and Culham, England, engineers take up the challenge of designing practical fusion reactors that can compete with fossil-fuel power plants. IEEE Spectrum, 21 (8):33–42, August 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fischetti:1985:BBG

[Fis85a] M. A. Fischetti. Body building goes high-tech. IEEE Spectrum, 22(5):102–105, May 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fischetti:1985:BBB

[Fis85b] M. A. Fischetti. Bypassing is big business for business: Both traditional and unconventional suppliers vie for the corporate dollar. IEEE Spectrum, 22(11): 78–83, November 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fischetti:1985:EWL

[Fis85c]  M. A. Fischetti.  Exotic

weaponry: Lasers, particle beams, and high-speed projectiles are expected to be the backbone of the SDI arsenal. IEEE Spectrum, 22(9):47–54, September 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fischetti:1985:HC

[Fis85d] M. A. Fischetti. Heating and cooling. IEEE Spectrum, 22 (5):40–43, May 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fischetti:1985:RDD

[Fis85e] M. A. Fischetti. Robots do the dirty work: Some walk, some roll as they go about their tedious, hazardous chores in nuclearpower plants. IEEE Spectrum, 22(4):65–73, April 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fischetti:1985:SSS

[Fis85f] M. A. Fischetti. Solid state: In smarter design and fabrication of ICs, power and logic are mixed on the same chip and circuits are integrated over an entire wafer. IEEE Spectrum, 22 (1):60–64, January 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fischetti:1986:EUP

[Fis86a] M. A. Fischetti. Electric utilities: Poised for deregulation?: Competitive forces are an outgrowth of technical advances, regulatory constraints, and a slackening in the growth of demand. IEEE Spectrum, 23 (5):34–40, May 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fischetti:1986:HSM

[Fis86b] M. A. Fischetti. Hero or scapegoat? most standard practices of commercial airline pilots are safe, yet pilots are often blamed for accidents. IEEE Spectrum, 23 (11):75–77, November 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fischetti:1986:PCT

[Fis86c] M. A. Fischetti. The puzzle of Chernobyl: Taken by surprise, technologists pose scenarios, map the radiation plume, and call for better emergency notification. IEEE Spectrum, 23(7):34–41, July 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fischetti:1986:QHG

[Fis86d] M. A. Fischetti. Quebec hydro: La grande tour: Three powerhouses, 200 dikes, and a vast flooded area near James Bay, Que., produce more electricity than any other hydroelectric complex in North America. IEEE Spectrum, 23(10): 30–36, October 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fischetti:1986:RPM

[Fis86e] M. A. Fischetti. A review of progress at MCC: Although its current objectives do not include ‘beating Japan,’ this U.S. organization is positioned to help industry do so. IEEE Spectrum, 23(3):76–82, March 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fischetti:1986:WRR

[Fis86f] M. A. Fischetti. When reactors reach old age: Can an obsolete nuclear power plant site he safely and economically restored to its original environmental condition? IEEE Spectrum, 23(2):28–35, February 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fischetti:1987:BNA

[Fis87a] M. A. Fischetti. Banishing the necktie: As automakers discover as much useful knowledge on the factory floor as in the executive suite, respect and productivity follow. IEEE Spectrum, 24 (10):50–52, October 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fischetti:1987:ISR

[Fis87b] M. A. Fischetti. Inherently safe reactors: They’d work if we’d let them: Although they shut down automatically and cannot melt or explode, advanced designs are limited to demonstrations because of economics and institutional inertia. IEEE Spectrum, 24(4):28–33, April 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fischetti:1987:SSB

[Fis87c] M. A. Fischetti. The silver screen blossoms into color: Like it or not, old favorites like the Maltese Falcon are being colorized; here’s how it’s done. IEEE Spectrum, 24(8):50–55, August 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fischetti:1987:SSJ

[Fis87d] M. A. Fischetti. Solid state:

Japan agrees not to undersell U.S. semiconductor makers. Engineers scramble to make application-specific ICs. Digital signal processing enters marketplace. Bipolar and MOS circuits integrated on same chip. IEEE Spectrum, 24(1):45–47, January 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fitzgerald:1987:CDC

[Fit87a]  K. Fitzgerald.  Compressing

the design cycle: CAD and CAE are crucial competitive tools, but honing them to maximum effectiveness is slow work Karen Fitzgerald, Associate Editor. IEEE Spectrum, 24(10): 39–42, October 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fitzgerald:1987:RR

[Fit87b] K. Fitzgerald. Regulation and retrenchment. IEEE Spectrum, 24 (11):30–34, November 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fitzgerald:1987:TER

[Fit87c] K. Fitzgerald. Targeting the enemy: Radar, radio control and inertial-guidance systems helped improve the accuracy of both offensive and defensive weaponry. IEEE Spectrum, 24 (8):42–49, August 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fitzgerald:1988:DSA

[Fit88a] K. Fitzgerald. 1972 digital scopes: Analog-to-digital converters and memories turned a venerable instrument into a remarkable new tool. IEEE Spectrum, 25(11):82–86, November 1988. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fitzgerald:1988:FRW

[Fit88b] K. Fitzgerald. Fielding reliable weapons more quickly. IEEE Spectrum, 25(12):60–64, November 1988. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fitzgerald:1988:GMW

[Fit88c] K. Fitzgerald. The genesis of major weapon systems. IEEE Spectrum, 25(12):41–44, November 1988. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fitzgerald:1988:LCL

[Fit88d]  K. Fitzgerald.  Lilco con-

templates life without Shoreham. IEEE Spectrum, 25(10): 20–23, October 1988. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fitzgerald:1988:SFV

[Fit88e] K. Fitzgerald. Superconductivity: fact vs. fancy. IEEE Spectrum, 25(5):30–41, May 1988. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fitzgerald:1988:TI

[Fit88f]  K. Fitzgerald.  Technology

’88: instrumentation. IEEE Spectrum, 25(1):47–49, January 1988. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fitzgerald:1988:TME

[Fit88g] K. Fitzgerald. Technology ’88:

the main event [superconductivity]. IEEE Spectrum, 25 (1):27–28, January 1988. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fitzgerald:1989:NWR

[Fit89a] K. Fitzgerald. Nuclear weapons reactors: too hot to handle? IEEE Spectrum, 26(6):37–41, June 1989. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fitzgerald:1989:PBC

[Fit89b] K. Fitzgerald. Probing Boeing’s crossed connections (aircraft instrumentation). IEEE Spectrum, 26(5):30–35, May 1989. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fitzgerald:1989:QIP

[Fit89c] K. Fitzgerald. The quest for intruder-proof computer systems. IEEE Spectrum, 26(8): 22–26, August 1989. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fitzgerald:1989:TI

[Fit89d] K. Fitzgerald. Technology ’89 instrumentation. IEEE Spectrum, 26(1):50–52, January 1989. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fitzgerald:1989:TME

[Fit89e]  K. Fitzgerald.  Technology

’89: medical electronics. IEEE Spectrum, 26(1):67–69, January 1989. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fitzgerald:1989:TMT

[Fit89f]  K. Fitzgerald.  Technology

in medicine: too much too soon? IEEE Spectrum, 26(12): 24–29, December 1989. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fletcher:1982:MEI

[Fle82] S. G. Fletcher. Medical electronics II: Seeing speech in real time: The deaf can now view tongue, jaw, and other vocaltract movements on a CRT display. IEEE Spectrum, 19 (4):42–45, April 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fletcher:1985:EHT

[Fle85a] C. Fletcher. Epicurean hightech. IEEE Spectrum, 22 (5):81–85, May 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fletcher:1985:VRE

[Fle85b] C. Fletcher. Videotex: Return engagement: Undaunted by unsuccessful ventures, U.S. corporate giants are planning new investments to give the technology a second chance. IEEE Spectrum, 22(10):34–38, October 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fletcher:1985:WST

[Fle85c] C. Fletcher. Weathering the storm: Terminations and transfers: Uncertainty became the dominant theme on Jan. 1, 1984 for employees of both AT&T and the Bell operating companies. IEEE Spectrum, 22(11): 97–103, November 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Florman:1982:CSV

[Flo82] S. C. Florman. Careers: a skeptic views ethics in engineering: Competing with recession and unemployment for people’s concern, ethics still manages to arouse lively debates. IEEE Spectrum, 19 (8):56–57, August 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Florman:1987:S

[Flo87]  S. C. Florman.  Speakout.

IEEE Spectrum, 24(6):18–19, June 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Florig:1988:FBB

[Flo88]  H. K. Florig.  The future

battlefield: a blast of gigawatts? [microwave-based weapons]. IEEE Spectrum, 25 (3):50–54, March 1988. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Flschettl:1983:SAM

[Fls83] M. A. Flschettl. The space-age materials challenge. IEEE Spectrum, 20(9):65, September 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Flynn:1987:SRP

[Fly87] M. S. Flynn. Out-sourcing rediscovered: Purchasing parts and even complete vehicles from independent suppliers can keep costs down and quality up. IEEE Spectrum, 24(10):47–49, October 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fischetti:1984:FFI

[FM84a] M. A. Fischetti and J. Morgan. A framework for the future: Industry joins government in establishing campus beachheads, as technology begins to transform university life. IEEE Spectrum, 21(11):102–108, November 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fischetti:1984:MPT

[FM84b] M. A. Fischetti and J. Morgan. Model programs: There is a smorgasbord of offerings from universities, industry, professional societies, and commercial firms. IEEE Spectrum, 21(11):98–101, November 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Forrest:1986:ODT

[For86] S. R. Forrest. Optical detectors: Three contenders: Depending on the application, the photoeonductor, p-i-n diode, or avalanche photodiode may prove the best choice. IEEE Spectrum, 23(5):76–85, May 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fischetti:1984:OFH

[FP84] M. A. Fischetti and Tekla S.

Perry. Overcoming formidable hurdles: Engineers who are disabled, female, nondegreed, foreign-born, or retrained and a disadvantaged student reveal how to succeed despite difficulties. IEEE Spectrum, 21(11): 86–91, November 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fischetti:1986:WPS

[FP86] M. A. Fischetti and Tekla S.

Perry. What price safety? Air travelers may be unaware that unless the U.S. Congress acts, increasing the margins of safety may be unaffordable as airlines compete for market share. IEEE Spectrum, 23(11):80, November 1986. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Frey:1984:WDE

[Fre84] C. Frey. What is due the engineer? Why do U.S. engineers have little influence, while Japanese engineers head many companies and Soviet engineers end up in the Politburo? IEEE Spectrum, 21(6):64–68, June 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Freifeld:1986:AAP

[Fre86] K. Freifeld. Art analysis probing beneath the image: Electrotechnology backs up museum conservators in analyzing, conserving, and restoring works of art and in unmasking forgeries. IEEE Spectrum, 23(6):66–71, June 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Freeman:1988:UPG

[Fre88a] R. Freeman. User-programmable gate arrays. IEEE Spectrum, 25 (13):32–35, December 1988. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Freese:1988:ODB

[Fre88b] R. P. Freese. Optical disks become erasable. IEEE Spectrum, 25(2):41–45, February 1988. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedman:1984:WUG

[Fri84] E. A. Friedman. The wired university: In a growing number of institutions, educators are integrating powerful personal computers and networks into a new learning environment. IEEE Spectrum, 21(11): 115–120, November 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Frosch:1983:HBB

[Fro83] R. A. Frosch. Hybrids: The best of both worlds. IEEE Spectrum, 20(9):30–35, September 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Frutkin:1983:PSU

[Fru83] A. Frutkin. The politics of space: U.S. policy: a drama in n acts. IEEE Spectrum, 20(9): 70–74, September 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fischetti:1985:SRS

[FT85] M. A. Fischetti and C. Truxal.

Simulating ‘the right stuff’ [military simulators]. IEEE Spectrum, 22(3):38–47, March 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fullagar:1980:CCA

[Ful80] D. Fullagar. Cmos comes of age. IEEE Spectrum, 17(12): 24–27, December 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Frantz:1982:DCH

[FW82] G. A. Frantz and R. H. Wiggins. Design case history: speak spell learns to talk. IEEE Spectrum, 19(2):45–49, February 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fitzgerald:1987:NGR

[FW87] K. Fitzgerald and P. Wallich. Next-generation race bogs down: Japan, the United States, and Europe are still far from their declared goals for intelligent machines, but Japan’s hidden agenda may already have succeeded. IEEE Spectrum, 24 (6):28–33, June 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fischetti:1986:NTF

[FWVH86] M. A. Fischetti, P. Wallich, J. Voelcker, and J. Horgan. New times at the FAA: Since deregulation, the number of pilots, planes, and airlines has overwhelmed the Federal Aviation Administration. IEEE Spectrum, 23(11):67–69, November 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fischetti:1985:PEU

[FZ85] M. A. Fischetti and Glenn Zorpette. Power and energy: Utilities boost efficiency and save money through energymanagement systems, pumped storage, and fuel cells, while industrial producers pursue cogeneration. IEEE Spectrum, 22 (1):72–76, January 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fischetti:1987:PEC

[FZ87a] M. A. Fischetti and Glenn Zorpette. Power and energy: Cogenerators battle with utilities before regulators nuclear industry undergoes sweeping changes after Chernobyl explosion operating life is extended at old generating units U.S. manufacturers decline, European makers gain. IEEE Spectrum, 24 (1):53–55, January 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fitzgerald:1987:FDG

[FZ87b] K. Fitzgerald and Glenn Zorpette. The first decade: Great expectations. IEEE Spectrum, 24 (11):25–30, November 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fischetti:1986:PEC

[FZL86] M. A. Fischetti, Glenn Zorpette, and W. D. Lang. Power and energy: Computers bring dramatic improvements in power management, permitting an increase in the wheeling of power and spurring the development of HVDC systems. IEEE Spectrum, 23(1):65–68, January 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goutzoulis:1988:DEM

[GA88] A. P. Goutzoulis and I. J.

Abramovitz. Digital electronics meets its match (acoustooptic devices). IEEE Spectrum, 25 (8):21–25, August 1988. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gansler:1982:UDI

[Gan82] J. S. Gansler. U.S. defense industry: Its bottlenecks. IEEE Spectrum, 19(10):89–90, October 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Garrett:1981:ANC

[Gar81] H. B. Garrett. Aerospace: Neutralizing charged-up spacecraft: Electrical charge gathered from space plasma can cause malfunctions or failure, but techniques for minimizing its effects are available. IEEE Spectrum, 18(7):44–47, July 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Garwin:1982:S

[Gar82] R. L. Garwin. Speakout. IEEE Spectrum, 19(10):24–25, October 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gardner:1984:LTE

[Gar84] R. M. Gardner. Lifestyle: Tomorrow’s electronic hospital is here today: Computer use is revolutionizing not only hospital laboratories and record-keeping, but also emergency treatment, diagnostics, and intensive care. IEEE Spectrum, 21(6):101–103, June 1984. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Gaushell:1985:APG

[Gau85] D. J. Gaushell. Automating the power grid: Real-time communication between customers, the electricity grid, and power plants can cut costs for both supplier and user; the technology options are many. IEEE Spectrum, 22 (10):39–45, October 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gwen:1985:DCG

[GB85] Gwen and C. G. Bell. Digging for computer ‘gold’ [Boston Computer Museum]. IEEE Spectrum, 22(12):56–62, December 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Giallorenzi:1986:OFS

[GBDC86] T. G. Giallorenzi, J. A. Bucaro, A. Dandridge, and J. H. Cole. Optical-fiber sensors challenge the competition: Resistance to corrosion and immunity to interference head the list of benefits in detecting stimuli ranging from pressure to magnetism. IEEE Spectrum, 23(9): 44–50, September 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gellings:1981:PED

[Gel81] C. W. Gellings. Power/energy:

Demand-side load management: The rising cost of peak-demand power means that utilities must encourage customers to manage power usage. IEEE Spectrum, 18 (12):49–52, December 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Getting:1982:GSC

[Get82] I. A. Getting. Global survival: Challenge to engineers. IEEE Spectrum, 19(10):109, October 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gevarter:1983:CES

[Gev83]  W. B. Gevarter.  Comput-

ers: Expert systems: Limited but powerful: Using a collection of facts, rules of thumb, and methods of applying those rules and making inferences, a new type of computer system is emerging. IEEE Spectrum, 20 (8):39–45, August 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gelsinger:1989:MC

[GGPY89] P. P. Gelsinger, P. A. Gargini, G. H. Parker, and A. Y. C. Yu. Microprocessors circa 2000. IEEE Spectrum, 26(10): 43–47, October 1989. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Garzia:1986:DES

[GGZ86] R. F. Garzia, M. R. Garzia, and B. P. Zeigler. Discrete-event simulation: When prototyping of large, complex systems is impossible, this technique, based on known behaviors of interacting elements, is useful. IEEE Spectrum, 23(12):32–36, December 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gibson:1980:SCE

[Gib80] R. Gibson. Satellite communications I: ESA at the crossroads: Projects for the 1980s at the space agency include satellite launches and a space platform. IEEE Spectrum, 17 (3):48–53, March 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gibson:1985:TPA

[Gib85] J. M. Gibson. Tools for probing ‘atomic’ action [electron microscopes]. IEEE Spectrum, 22 (12):38–44, December 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Glenn:1983:NAU

[Gle83] J. H. Glenn. The next 25:

Agenda for the U.S. IEEE Spectrum, 20(9):91–95, September 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldman:1981:SMS

[Gol81] D. T. Goldman. Standards:

The metric system: Its status and future: The U.S. creeps reluctantly toward metrication; the electrical engineering profession encounters few problems. IEEE Spectrum, 18 (4):60–63, April 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldman:1987:BR

[Gol87]  J. E. Goldman.  Book re-

views. IEEE Spectrum, 24 (7):12–13, July 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gallini:1981:SNW

[GP81] L. D. Gallini and F. Prokoski. Sociotechnology: Nuclear weapons monitoring: The international community is desperately trying to devise schemes and systems to discourage proliferation. It could be an unending struggle. IEEE Spectrum, 18(7):48–51, July 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Graham:1982:CSS

[Gra82a] A. K. Graham. Computer software: Software design: Breaking the bottleneck: a visionary approach is recommended in place of the piecemeal policy now in effect. IEEE Spectrum, 19 (3):43–50, March 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Graube:1982:CLA

[Gra82b] M. Graube. Communications: Local area nets: a pair of standards: Confronted by marketplace needs, the IEEE 802 committee is proposing standards for two kinds of networks: Contention access and token passing. IEEE Spectrum, 19(6):60–64, June 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gray:1982:ICT

[Gra82c] P. E. Gray. Information control I: Technology transfer at issue: The academic viewpoint: Educators believe efforts to limit transfer of knowledge at the university level are likely to weaken the U.S. lead in innovation. IEEE Spectrum, 19(5):64–68, May 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Grayson:1984:EEA

[Gra84] L. P. Grayson. EE education around the world: The educational systems of the United States, the Soviet Union, Japan, West Germany, France, the United Kingdom, and India are compared. IEEE Spectrum, 21 (11):65–71, November 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Greenberg:1980:PTO

[Gre80] D. S. Greenberg. Politics of technology: OTA: Congress’ own think tank: After seven years of controversy, the Office of Technology Assessment is shunning the spotlight to survive in a political jungle. IEEE Spectrum, 17(7):44–47, July 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Green:1985:PGJ

[Gre85] P. E. Green. Paul E. Green, Jr. on communications: Progress in fiber optics, LANs, and PBXs. IEEE Spectrum, 22 (1):38–39, January 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Groover:1983:FOM

[Gro83] M. P. Groover. Fundamental operations: Machining, forming, joining, assembly, and inspection are basic functions to be integrated and controlled in the automatic factory. IEEE Spectrum, 20(5):65–69, May 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Glatzel:1980:TPM

[GS80] K. Glatzel and H. Schulz. Transportation: The promise of maglev: Tests will continue with a high-speed vehicle and track incorporating a linear motor. IEEE Spectrum, 17(3):63–66, March 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guterl:1982:ECT

[GT82] F. Guterl and C. Truxal. Entertainment/control technology: The wonderful world of Epcot: Walt Disney Productions unveils a futuristic amusement park that may be even more intriguing to engineers than to the public at large. IEEE Spectrum, 19(9): 46–55, September 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guterl:1983:MPW

[GT83] F. Guterl and C. Truxal. Militarization: Peace or war? IEEE Spectrum, 20(9):35–38, September 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guterl:1982:CT

[Gut82a] F. Guterl. Cameras that ‘think’.

IEEE Spectrum, 19(6):32–37, June 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guterl:1982:CRB

[Gut82b] F. Guterl. The clinch river breeder. IEEE Spectrum, 19 (3):65–68, March 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guterl:1982:DCH

[Gut82c] F. Guterl. Design case history: the ‘infinite’ matrix printer. IEEE Spectrum, 19(9): 40–45, September 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guterl:1982:ITB

[Gut82d] F. Guterl. Intelligence: Its technological base. IEEE Spectrum, 19(10):62–64, October 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guterl:1982:MWS

[Gut82e] F. Guterl. Major weapons systems. IEEE Spectrum, 19(10): 40–43, October 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guterl:1982:MPE

[Gut82f] F. Guterl. Meeting preview: Electro ’82: Micros, fiber optics, and robotics: The frontiers of technology and its application to systems will be explored in Boston, May 25 to 27. IEEE Spectrum, 19(5):74–75, May 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guterl:1982:MPW

[Gut82g] F. Guterl. Meeting preview: Wescon ’82: a micro bonanza: Microprocessors and their applications will be among the main attractions at Anaheim in September. IEEE Spectrum, 19 (9):71–72, September 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guterl:1982:MDC

[Gut82h] F. Guterl. Microprocessors: Design case history: Biomation’s logic analyzer: an old lesson is relearned: Utility for users, rather than product sophistication, is key to success. IEEE Spectrum, 19(5):52–56, May 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guterl:1982:TNP

[Gut82i] F. Guterl. Today’s navies: Plying a sea of debate. IEEE Spectrum, 19(10):43–48, October 1982. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Guterl:1983:IHT

[Gut83a] F. Guterl. Internationalization:

Harmony or tension? IEEE Spectrum, 20(9):41–47, September 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guterl:1983:MPE

[Gut83b] F. Guterl. Meeting preview:

Electro ’83 micros, CAD, and telecommunications: a professional program, tutorial seminars, and exhibits entice visitors to the April 19–21 convention in the Big Apple. IEEE Spectrum, 20(4):63, April 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guterl:1983:MCA

[Gut83c]  F. Guterl.  Microprocessors:

Chip architecture: a revolution brewing: Seeking to tailor new micros to challenging applications, designers are scrutinizing entrenched mainframe concepts as never before. IEEE Spectrum, 20(7):30–37, July 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guterl:1983:MEG

[Gut83d] F. Guterl. Microprocessors:

Enhanced general-purpose and number-crunching processors are supported by specialized microcontroller ‘slaves’. IEEE Spectrum, 20(1):34–35, January 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guterl:1983:NWN

[Gut83e] F. Guterl. Networks: Wiring a nation. IEEE Spectrum, 20 (11):74–75, November 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guterl:1983:NGI

[Gut83f] F. Guterl. Next-generation impacts: Technology watchers discuss industrial and office automation, knowledge-based systems, computer privacy, and education. IEEE Spectrum, 20(11): 111–117, November 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guterl:1983:PCN

[Gut83g] F. Guterl. Personal computers: As new systems move up to 16-bit microcomputers, the goal of software compatibility remains as elusive as ever. IEEE Spectrum, 20(1): 36–37, January 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guterl:1983:SES

[Gut83h] F. Guterl. Supersmart equipment: Software for computerbased instruments brings users a plethora of benefits. IEEE Spectrum, 20(1):48–54, January 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guterl:1983:UQA

[Gut83i]  F. Guterl.  An unanswered

question: Automation’s effect on society: Little research has been done and few plans have been laid to ease the transition of developed industrial nations to a highly automated society. IEEE Spectrum, 20 (5):89–92, May 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guterl:1984:SHPd

[Gut84a] F. Guterl. Spectrum/Harris poll:

Education. IEEE Spectrum, 21 (6):128–132, June 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guterl:1984:SHPb

[Gut84b] F. Guterl. Spectrum/Harris poll:

EE specialties. IEEE Spectrum, 21(6):49–52, June 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guterl:1984:SHPc

[Gut84c] F. Guterl. Spectrum/Harris poll:

The career. IEEE Spectrum, 21(6):59–63, June 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guterl:1984:SHPa

[Gut84d] F. Guterl. Spectrum/Harris poll:

The job. IEEE Spectrum, 21 (6):38–43, June 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guterl:1984:DCH

[Gut84e] F. Guterl. Design case history: Apple’s Macintosh: a small team of little-known designers, challenged to produce a low-cost, exceptionally easy-to-use personal computer, turns out a technical milestone. IEEE Spectrum, 21 (12):34–43, December 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guterl:1984:IIO

[Gut84f] F. Guterl. Instrumentation: Industry offers a pragmatic alternative to the dearth of academic programs. IEEE Spectrum, 21 (11):56–57, November 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guterl:1984:IVC

[Gut84g] F. Guterl. Instrumentation: VLSI chips present new challenges for testing, while personal computers and local networks lead to new approaches. IEEE Spectrum, 21(1):64–67, January 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guterl:1984:MMI

[Gut84h] F. Guterl. Microprocessors: The microprocessor industry shifts from hardware as the driving force in chip design to software, with the widespread acceptance of the UNIX operating system. IEEE Spectrum, 21 (1):50–52, January 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guterl:1984:TTT

[Gut84i] F. Guterl. Technologists and technicians: Two- and four-year programs offer hands-on education in electronics and computers to complement the engineer’s skills in activities ranging from research to manufacturing. IEEE Spectrum, 21(11):61–64, November 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guterl:1984:TPC

[Gut84j] F. Guterl. Technology ’84 personal computers: Common sense replaces computerese, steps are taken toward standardization, and interactive graphics gains ground. IEEE Spectrum, 21 (1):41–46, January 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guterl:1984:TDC

[Gut84k] F. Guterl. Telecommunications:

Digital communications and computer science are combined into popular programs. IEEE Spectrum, 21(11):57–55, November 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guterl:1984:TPM

[Gut84l] F. Guterl. Testing: Previously the most neglected area, chip testing must begin during the chip-design phase and continue through fabrication. IEEE Spectrum, 21(9):40–46, September 1984. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Guterl:1985:ALR

[Gut85a] F. Guterl. Audio in the living room. IEEE Spectrum, 22(5):78–80, May 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guterl:1985:IBI

[Gut85b] F. Guterl. Instrumentation: Bench instruments are linked by personal computers and the steps in devising chip tests and designing chips are linked by software. IEEE Spectrum, 22 (1):65–67, January 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guterl:1985:JS

[Gut85c] F. Guterl. Not just for sleeping. IEEE Spectrum, 22(5):90–92, May 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guterl:1985:TPC

[Gut85d] F. Guterl. Technology *85: personal computers. IEEE Spectrum, 22(1):45–49, January 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guterl:1988:FAB

[Gut88] F. Guterl. French arms buying puts technocrats at the helm. IEEE Spectrum, 25(12): 39–40, November 1988. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guterl:1989:DOB

[Gut89]  F. Guterl.  The dual ori-

gins of a bipolar breakthrough (i/sup 2/l). IEEE Spectrum, 26(7):25–29, July 1989. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guterl:1984:POM

[GWF84] F. Guterl, P. Wallich, and M. A.

Fischetti. In pursuit of the onemonth chip: The objective is to describe the functions required of an as yet nonexistent integrated circuit, then have it delivered in 31 days or less. IEEE Spectrum, 21(9):28–29, September 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guterl:1985:FOP

[GZ85] F. Guterl and Glenn Zorpette.

Fiber optics: Poised to displace satellites: Expansion of fiberoptic cable networks will lead to the creation of new communication services and may limit the role of satellites. IEEE Spectrum, 22(8):30–37, August 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hamilton:1986:ZDS

[Ham86] M. H. Hamilton. Zero-defect software: The elusive goal: It is theoretically possible but difficult to achieve; logic and interface errors are most common, but errors in user intent may also occur. IEEE Spectrum, 23 (3):47–53, March 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hanover:1989:NIH

[Han89] G. Hanover. Networking the intelligent home. IEEE Spectrum, 26(10):48–49, October 1989. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Harrison:1983:CWV

[Har83] A. P. Harrison. Communications: The world versus RCA: Circumventing the superhet: Because a clearly superior tuning circuit was denied them by a patent award, manufacturers developed both clumsy and elegant alternatives. IEEE Spectrum, 20 (2):67–71, February 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hecht:1985:VEO

[Hec85] J. Hecht. Victorian experiments and optical communications: Precursors of fiber-optic communications were invented a century ago, but no one at that time attempted to synthesize the fragmented knowledge. IEEE Spectrum, 22(2):69–73, February 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Heffron:1984:TCA

[Hef84] C. Heffron. Teleconferencing comes of age: Electronic meetings enter a new era, thanks to personal computers, video printers, slide projectors, and bright graphics, among other tools. IEEE Spectrum, 21(10): 61–67, October 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Herbert:1982:FFN

[Her82] E. Herbert. Finding the flaws in nuclear power plants. IEEE Spectrum, 19(9):56–61, September 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Herbert:1983:CLD

[Her83a] E. Herbert. Computers: Learning from dinosaurs. IEEE Spectrum, 20(9):58–60, September 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Herbert:1983:MMM

[Her83b] E. Herbert. Minis and mainframes: More powerful versions of existing systems emerge, as reduced hardware costs encourage parallel-architecture development. IEEE Spectrum, 20 (1):28–33, January 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Herbert:1983:STA

[Her83c] E. Herbert. Space technology:

Apollo: The driver and the driven. IEEE Spectrum, 20(9): 56–58, September 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Herbert:1983:SM

[Her83d] E. Herbert. Superproject management. IEEE Spectrum, 20 (9):68–69, September 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Herbert:1984:CEV

[Her84] E. Herbert. Continuing education: a view from the top: How to apply a formal education and continue the learning process is one of the issues nine successful engineers discuss. IEEE Spectrum, 21(11):80–86, November 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Horgan:1985:MWB

[HGF85] J. Horgan, F. Guterl, and C. Fletcher. Making waves in the bath. IEEE Spectrum, 22(5):94–95, May 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hutcheson:1987:OIR

[HHH87] L. D. Hutcheson, P. Hangen, and A. Husain. Optical interconnects replace hardwire: Light promises to relieve bottlenecks in electric interconnections from cabinet to cabinet, board to board, IC to IC and even within chips. IEEE Spectrum, 24 (3):30–35, March 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hicks:1983:PEH

[Hic83] K. L. Hicks. Power/energy: Hybrid load shedding is frequency based: Time at a given frequency is also important, bringing stabilization in a matter of seconds even when as much as two thirds of the load must be shed. IEEE Spectrum, 20(2):52–56, February 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hills:1983:CAG

[Hil83]  A. Hills.  Communications:

Alaska’s giant satellite network: Telephone services and radio and television broadcasts for virtually every Alaskan community are ensured by the state’s satellite. IEEE Spectrum, 20 (7):50–55, July 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hills:1985:SMP

[Hil85] A. Hills. Satellites and mobile phones: Planning a marriage: Dramatic improvements in rural communications may be in the offing as direct service by satellite to mobile phones moves closer to reality. IEEE Spectrum, 22(8):62–67, August 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hills:1986:SEW

[Hil86] A. Hills. Subzero engineering:

At -40^{∘}F, with the wind howling at 60 miles per hour over permanently frozen ground, unconventional engineering techniques were required. IEEE Spectrum, 23(12):52–56, December 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hopcroft:1987:TBC

[HK87] J. E. Hopcroft and D. B. Krafft. Toward better computer science: Dependent on software, the United States needs a national policy on computer-science research: The scientists need more time and money for research, teaching, and leadership. IEEE Spectrum, 24(12):58–60, December 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Howe:1987:HPP

[HM87] C. D. Howe and B. Moxon. How to program parallel processors: Software developers must divide problems into manageable chunks and ensure efficient communications between computing elements. IEEE Spectrum, 24 (9):36–41, September 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Harrop:1980:SCI

[HMDF80] P. Harrop, J. Magarshack, R. Dessert, and J. R. Forrest. Satellite communications II: Television for everyone: Lowcost Earth stations that can select from a broad menu of satellite broadcasts are sought. IEEE Spectrum, 17(3):54–56, March 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hecht-Nielsen:1988:NPH

[HN88] R. Hecht-Nielsen. Neurocomputing: picking the human brain. IEEE Spectrum, 25 (3):36–41, March 1988. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hanes:1982:NPC

[HOD82] L. F. Hanes, J. F. O’Brien, and R. DiSalvo. Nuclear power: Control-room design: Lessons from TMI: Improved instrumentation plus computer-based decision aids under development should help minimize the potential consequences of accidents. IEEE Spectrum, 19 (6):46–53, June 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hondius:1980:CDP

[Hon80]  F. W. Hondius.  Comput-

ers: Data privacy: The European Community grapples with protecting individual rights in the midst of rampant computer progress. IEEE Spectrum, 17 (3):67–71, March 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Horgan:1984:EEI

[Hor84a] J. Horgan. Education: Education initiatives by states, industry, and the Federal government may not help schools that need them most. IEEE Spectrum, 21 (1):94–96, January 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Horgan:1984:GSG

[Hor84b] J. Horgan. A guide to school guides: Selecting the right college or graduate program is aided by both traditional references and emerging computer programs. IEEE Spectrum, 21 (11):76–79, November 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Horgan:1984:LEI

[Hor84c] J. Horgan. Learning from elsewhere: Incidents at sites other than TMI have expanded the nuclear knowledge base and may help contribute to the solutions of new and old problems. IEEE Spectrum, 21 (4):43–49, April 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Horgan:1984:MEB

[Hor84d] J. Horgan. Medical electronics: Biomedical engineers are exploiting electric signals to and from the body to aid rehabilitation of the handicapped. IEEE Spectrum, 21(1):90–93, January 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Horgan:1985:EW

[Hor85a] J. Horgan. Electronic watchdogs. IEEE Spectrum, 22 (5):43–46, May 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Horgan:1985:MEN

[Hor85b] J. Horgan. Medical electronics: NMR reaches the marketplace, biosensors show promise, and progress is made in pacemakers and implantable drug pumps. IEEE Spectrum, 22 (1):89–94, January 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Horgan:1985:MN

[Hor85c] J. Horgan. Monitoring the nursery. IEEE Spectrum, 22(5):93, May 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Horgan:1985:SNS

[Hor85d] J. Horgan. Safeguarding the national security: Can the military and other government functions be confident that the newly fragmented telecommunications system will respond to crises? IEEE Spectrum, 22(11):84–89, November 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Horgan:1985:TIT

[Hor85e] J. Horgan. Thwarting the information thieves: Fear of spying through simple or sophisticated electronics has spawned an industry whose challenge is to block the illegal interception of intelligence. IEEE Spectrum, 22(7):30–41, July 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Horgan:1986:PPP

[Hor86a] J. Horgan. The politics of peace pacts: Technical, monitoring capabilities are but part of the solution to arms control: Attitudes of the superpowers are becoming more crucial. IEEE Spectrum, 23(7):70–80, July 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Horgan:1986:RAA

[Hor86b] J. Horgan. Roboticists aim to ape nature: Engineers are thinking in new ways about sensing, chips, and robots as they study dolphins, slugs, and homo sapiens. IEEE Spectrum, 23 (2):66–71, February 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Horgan:1986:TVT

[Hor86c] J. Horgan. Technology vs. terrorism. IEEE Spectrum, 23(11): 78–79, November 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Horgan:1986:UNW

[Hor86d] J. Horgan. Underground nuclear weapons testing: despite improvements in simulation techniques, weapons designers defend their need for ‘live’ testing to help maintain the U.S. nuclear deterrent. IEEE Spectrum, 23(4):32–43, April 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hosenball:1983:WOW

[Hos83] S. N. Hosenball. Who owns what? Earth law for space. IEEE Spectrum, 20(9):76–79, September 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

House:1985:CHH

[Hou85] C. H. House. Charles H. House on instrumentation: a marriage of computers and test equipment. IEEE Spectrum, 22(1): 40–41, January 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hagelbarger:1983:HFE

[HT83] D. W. Hagelbarger and R. A.

Thompson. Human-factors engineering: Experiments in teleterminal design: The successful marriage of a computer terminal and a telephone depend on the unit’s size and shape, as well as on how it communicates with the user. IEEE Spectrum, 20 (10):40–45, October 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Huson:1984:CGP

[HTW84] F. R. Huson, L. Teng, and H. Wiedemann. The case for giant particle accelerators: Particle collisions can yield a better understanding of the fundamental nature of matter and energy. IEEE Spectrum, 21 (3):53–62, March 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hurwitz:1987:F

[Hur87] H. Hurwitz. Forum. IEEE Spectrum, 24(11):9, November 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hibbs:1983:ASM

[HW83] A. R. Hibbs and W. S. Wilson. Aerospace: Satellites map the oceans: High-flying sensors can measure currents, map the ocean floor, plot winds at sea, and identify prospective fishing grounds. IEEE Spectrum, 20 (10):46–54, October 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Iacono:1984:ORA

[IK84] S. Iacono and R. Kling. Office routine: The automated pink collar: Secretaries and clerks are often enthusiastic about new technologies, until they discover that shortsighted managers use the new tools to isolate them further. IEEE Spectrum, 21 (6):73–76, June 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Inagaki:1989:HLL

[Ina89] T. Inagaki. Hologram lenses lead to compact scanners. IEEE Spectrum, 26(3):39–43, March 1989. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jackson:1980:HCT

[Jac80] R. N. Jackson. Home communications I: Teletext and viewdata: Text and graphics are broadcast or transmitted over telephone wires to the home television receiver. IEEE Spectrum, 17(3):26–28, March 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jacobs:1982:CLP

[Jac82] G. Jacobs. Communications:

Low-power television: Shortrange, low-cost TV stations are in the offing as the FCC prepares to establish broadcast requirements. IEEE Spectrum, 19(6):54–59, June 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jako:1985:LCS

[Jak85] C. J. Jako. Lasers cut a swath in surgical and medical applications: Their early promise is being fulfilled, particularly for the surgeon, while research hints at usefulness in nonsurgical cancer therapy. IEEE Spectrum, 22(3):82–87, March 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jayant:1986:CSL

[Jay86] N. S. Jayant. Coding speech at low bit rates: Advanced algorithms and hardware for voice telecommunications are paring hit rates by at least a factor of four, without losing intelligibility. IEEE Spectrum, 23(8):58–63, August 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1980:ADP

[JCH80] R. K. Jurgen, D. E. Cole, and L. T. Harbeck. Automobiles: Detroit plans a comeback: Billions of dollars are at stake in U.S. auto makers’ catch-up battle with small, fuel-efficient imports. IEEE Spectrum, 17(11):

40–43, November 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jelalian:1981:LRI

[Jel81] A. V. Jelalian. Laser radar improvements. IEEE Spectrum, 18 (11):46–51, November 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jenkins:1984:TEF

[JI84] R. V. Jenkins and P. B. Israel. Thomas A. Edison: Flamboyant inventor: Shrewd and nonconformist in many respects, Edison helped create modern electrical engineering, yet he never became a professional engineer himself. IEEE Spectrum, 21(12): 74–79, December 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jacob:1987:DAS

[JJ87] L. Jacob and D. Johnson. Designing analog systems: Concept to components. IEEE Spectrum, 24(5):62–63, May 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Johnston:1984:WNV

[Joh84] D. Johnston. Work: a newsman’s view of technology: a prizewinning investigative reporter reflects on his past, when he ‘banged out stories’ on a manual typewriter, and on how far his profession has come. IEEE Spectrum, 21(6):112–113, June 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1985:SIH

[JP85] R. A. Jurgen and Tekla S.

Perry. Special issue: The hightech home. IEEE Spectrum, 22(5):34–35, May 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:1981:AEE

[JT81] T. O. Jones and M. U. Trenne.

Automotive electronics; electronic fuel injection: Diesels are next: Auto producers turn to electronics to meet emission standards for diesel engines. It’s tougher going compared with gasoline engines. IEEE Spectrum, 18(5):30–34, May 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1980:EME

[Jur80] R. K. Jurgen. Electronics in medicine: Exploiting microprocessors; new hearing and seeing aids are developed; diagnostic imaging is refined. IEEE Spectrum, 17(1):81–86, January 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1981:AEA

[Jur81a]  R. K. Jurgen.  Automotive

electronics: All-electronic dashboards coming: CRT, liquid crystal, and electroluminescent displays show promise for largearea, multifunction driver information systems in future cars. IEEE Spectrum, 18(6):34–37, June 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1981:AED

[Jur81b] R. K. Jurgen. Automotive electronics: Detroit bets on electronics to stymie Japan: Programmable controllers, robots, lasers, and microwave detectors maintain quality on U.S. auto makers’ production and assembly fines. IEEE Spectrum, 18 (7):29–32, July 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1982:AED

[Jur82a] R. K. Jurgen. Automotive electronics: Drivers get more options in 1983: Car makers embrace electronics in an all-out drive to improve performance, handling, fuel economy, riding quality, driver aids, and entertainment. IEEE Spectrum, 19 (11):30–36, November 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1982:CES

[Jur82b] R. K. Jurgen. Consumer electronics: Stereophonic sound for television: Three transmission systems, vying for selection as the U.S. standard, are compared. IEEE Spectrum, 19(9): 30–33, September 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1982:IEI

[Jur82c] R. K. Jurgen. Industrial electronics: Industry’s workhorse gets smarter: The programmable controller adds intelligent input/ output modules, block programming in ladder diagram logic, self-diagnosis, and data highways. IEEE Spectrum, 19(2): 34–38, February 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1982:IEO

[Jur82d]  R. K. Jurgen.  Industrial

electronics: Our computers aren’t speaking: Innovation in computer-based manufacturing has led to a pressing need for interface standards. IEEE Spectrum, 19(9):62–65, September 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1983:ADU

[Jur83a] R. K. Jurgen. Automotive: Detroit unveils sophisticated electronics: Engines are more completely controlled, an electronic suspension debuts, synthetic voices tell more, and on-board diagnostics proliferate. IEEE Spectrum, 20(10):33–39, October 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1983:CEP

[Jur83b] R. K. Jurgen. Consumer electronics: The problems and promises of high-definition television: Its future looks brighter, yet challenges remain in bandwidth, standards, and compatibility. IEEE Spectrum, 20(12): 46–51, December 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1983:DDA

[Jur83c] R. K. Jurgen. Data-driven automation. IEEE Spectrum, 20 (5):33–35, May 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1984:MED

[Jur84] R. K. Jurgen. More electronics in Detroit’s 1985 models: Bells and whistles take a hack seat to functional, efficient design and to the integration of instrumentation. IEEE Spectrum, 21(10):54–60, October 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1985:IPS

[Jur85a] R. K. Jurgen. Impact 2000: a place in the sun: Sponsored by Boston Edison as an energy-conservation experiment, this home in the suburbs was seen during its construction by public-television viewers across the United States. IEEE Spectrum, 22(5):64–68, May 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1985:LSM

[Jur85b] R. K. Jurgen. Less sizzle, more beef for Detroit in ’86: Increasing electronic content in 1986 models represents serious functions like transmission, engine, and antitheft controls, antilock braking, and diagnostics. IEEE Spectrum, 22(10):77–83, October 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1986:CDN

[Jur86a] R. K. Jurgen. Coming from detroit: Networks on wheels: Multiplexed wiring in new models promises cost-effective control and accurate diagnostics, as well as fewer wires and less redundancy. IEEE Spectrum, 23(6):53–59, June 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1986:DMN

[Jur86b] R. K. Jurgen. Detroit’s 1987 models: New electronic inroads: Suspensions are the latest conquest in the trend toward total electronic control of all major automotive subsystems. IEEE Spectrum, 23(10):68–73, October 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1986:PKC

[Jur86c] R. K. Jurgen. The promise of the Kalina cycle: Using an ammonia-water mixture, the Kalina steam cycle may permit thermal-mechanical-electrical energy conversion efficiencies of 45 percent. IEEE Spectrum, 23(4):68–70, April 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1986:SEI

[Jur86d] R. K. Jurgen. The specialties: Experts identify the most outstanding developments or the most difficult problems in their fields. IEEE Spectrum, 23(1): 86–87, January 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1986:T

[Jur86e]  R. K. Jurgen.  Technology:

’86. IEEE Spectrum, 23(1): 34–35, January 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1987:DPW

[Jur87a] R. K. Jurgen. Desktop publishing: What it can and cannot do: an evolutionary outgrowth of word processing, the low-cost alternative to typesetting services has decided limitations in many applications. IEEE Spectrum, 24(3):50–52, March 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1987:DDF

[Jur87b] R. K. Jurgen. Detroit ’88: Driver-friendly innovations: Car handling, riding comfort, and safety improve with systems that act automatically rather than wait for driver input. IEEE Spectrum, 24(12):53–57, December 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1987:SEH

[Jur87c] R. K. Jurgen. The specialties:

Ergonomics and human factors engineering are redefining the programmer’s art electrical engineering education needs a total curriculum review trellis-coded modulation schemes improve signal-to-noise ratios. IEEE Spectrum, 24(1):69–70, January 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1987:SSG

[Jur87d] R. K. Jurgen. Survival strategy: Go global: International alliances, quality, and efficiency mark the winners in the high-stakes automobilemanufacturing game. IEEE Spectrum, 24(10):34–38, October 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1987:WHA

[Jur87e] R. K. Jurgen. Whatever happened to: Alternative engines, automatic traffic control and continuously variable transmissions? IEEE Spectrum, 24 (10):64–66, October 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1987:WHP

[Jur87f] R. K. Jurgen. Whatever happened to Project Tinkertoy? IEEE Spectrum, 24(5):20–21, May 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1987:WHS

[Jur87g] R. K. Jurgen. Whatever happened to superconducting generators? IEEE Spectrum, 24 (1):17, January 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1988:HDT

[Jur88a] R. K. Jurgen. High-definition television update. IEEE Spectrum, 25(4):56–62, April 1988. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1988:NFD

[Jur88b] R. K. Jurgen. New frontiers for Detroit’s big three. IEEE Spectrum, 25(10):32–34, October 1988. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1988:SLS

[Jur88c] R. K. Jurgen. Sarnoff Labs: ‘still crazy’ but coping. IEEE Spectrum, 25(13):36–39, December 1988. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1988:TCE

[Jur88d] R. K. Jurgen. Technology ’88: consumer electronics. IEEE Spectrum, 25(1):56–58, January 1988. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1988:TSE

[Jur88e] R. K. Jurgen. Technology ’88:

the specialities [electrical engineering]. IEEE Spectrum, 25 (1):70, January 1988. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1989:CJH

[Jur89a] R. K. Jurgen. Chasing Japan in the HDTV race. IEEE Spectrum, 26(10):26–30, October 1989. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1989:GCE

[Jur89b]  R. K. Jurgen.  Global ’90

cars: electronics-aided. IEEE Spectrum, 26(12):45–49, December 1989. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1989:TCE

[Jur89c] R. K. Jurgen. Technology ’89:

Consumer electronics. IEEE Spectrum, 26(1):59–61, January 1989. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1989:TS

[Jur89d] R. K. Jurgen. Technology ’89:

the specialities. IEEE Spectrum, 26(1):70, January 1989. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1987:GA

[JV87] R. K. Jurgen and J. Voelcker.

The global automobile. IEEE Spectrum, 24(10):32–33, October 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1986:TAP

[KA86] Gadi Kaplan and T. A. Adam. Transportation: AC propulsion and the use of microprocessors for locomotive control and diagnostics may help an ailing railroad industry retrack itself. IEEE Spectrum, 23(1): 73–75, January 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kahne:1981:ECE

[Kah81] S. Kahne. Education: a crisis in electrical engineering manpower: a review of the problems facing U.S. industry as a result, and a strategy for corrective action. IEEE Spectrum, 18(6):50–52, June 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kahne:1982:ECI

[Kah82] S. Kahne. Education: cracks in the Ivory Tower: academics are beginning to flee a deteriorating campus environment for jobs in industry — a trend threatening the U.S. technological base. IEEE Spectrum, 19 (3):69–73, March 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kahn:1983:NGC

[Kah83] R. E. Kahn. A new generation in computing: Microelectronics and artificial intelligence may produce advanced computers that are both fast and smart. IEEE Spectrum, 20(11): 36–41, November 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kahn:1984:COS

[Kah84] David Kahn. Cryptology and the origins of spread spectrum: Engineers during World War II developed an unbreakable scrambler to guarantee secure communications between Allied leaders; actress Hedy Lamarr played a role in the technology. IEEE Spectrum, 21(9):70–80, September 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaminski:1986:PCF

[Kam86] M. A. Kaminski. Protocols for communicating in the factory. IEEE Spectrum, 23(4):56–62, April 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1980:CSW

[Kap80a] Gadi Kaplan. Computers/ software: Words into action: I: Spoken advice can be recognized and converted to standard digital commands. IEEE Spectrum, 17(6):22–26, June 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1980:ECE

[Kap80b] Gadi Kaplan. An electric car in every garage: This optimistic goal is far off; progress toward it hinges on high-performance batteries. IEEE Spectrum, 17(9): 79–84, September 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1980:EPM

[Kap80c] Gadi Kaplan. Electrotechnology is a prime mover: Electronic engine controls save fuel; battery cars are tested; automated group rapid transit is funded. IEEE Spectrum, 17 (1):69–74, January 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1980:E

[Kap80d] Gadi Kaplan. Europe ’80. IEEE Spectrum, 17(3):24–25, March 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1980:PED

[Kap80e] Gadi Kaplan. Power/energy developments: High efficiencies and less downtime are sought by utilities as the oil squeeze worsens. IEEE Spectrum, 17 (1):58–63, January 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1980:SPC

[Kap80f] Gadi Kaplan. Superconducting power cables: Both technological push and demand pull are needed to spur progress in superconducting, supercooled cables. IEEE Spectrum, 17(9):

61–65, September 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1981:CWL

[Kap81a] Gadi Kaplan. Careers: We look at ourselves: The researcher: Driven by the thrill of discovery, researchers seek to contribute to society and to enhance their images among peers. IEEE Spectrum, 18(8):46–50, August 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1981:CII

[Kap81b] Gadi Kaplan. Communications. III. Information transfer. IEEE Spectrum, 18(1): 51–52, January 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1981:PEE

[Kap81c] Gadi Kaplan. Power/energy: I.

Efficient devices/systems: Utilities explore new means to cut energy losses during transmission and distribution. IEEE Spectrum, 18(1):74–77, January 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1982:PTP

[Kap82a] Gadi Kaplan. Peace 7: Threats to peace: Scenarios of war that worry experts include accidental nuclear exchanges and escalating brushfire wars. IEEE Spectrum, 19(10):100–101, October 1982. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1982:PPE

[Kap82b] Gadi Kaplan. Peace 8: Peace efforts: Broad-based technical and political involvement seems a prerequisite for regional and global security: International approaches to peacekeeping. IEEE Spectrum, 19(10):102–105, October 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1982:PSH

[Kap82c] Gadi Kaplan. Power systems: Hope is rekindled for nuclear energy; new components can reduce power-system operating costs. IEEE Spectrum, 19(1): 65–68, January 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1982:PET

[Kap82d] Gadi Kaplan. Power/energy: Thyristors: Future workhorses in power transmission: Research is under way to cut energy losses and costs and to improve reliability. IEEE Spectrum, 19(12): 40–45, December 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1983:ARC

[Kap83a] Gadi Kaplan. Automated railways: Computers help run Japan’s second-generation Shinkansen lines and the French Lille urban metro system. IEEE Spectrum, 20(1):68–70, January 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1983:FOI

[Kap83b] Gadi Kaplan. Fiber optics, ICs, and satellites: Advanced optical fibers, integrated circuits for telephony, and solid-state amplifiers for satellites presage a new era. IEEE Spectrum, 20 (1):38–40, January 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1983:PNP

[Kap83c] Gadi Kaplan. Power: Nuclearpower-plant malfunction analysis: High hopes are held for improving nuclear-plant safety through the study of transients using advanced hardware and software tools. IEEE Spectrum, 20(6):53–58, June 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1983:PEN

[Kap83d] Gadi Kaplan. Power/energy:

New York’s fuel-cell power plant: On the verge of success: If all goes well, Manhattan’s 4.8-MW plant should be delivering reliable power this month. IEEE Spectrum, 20(12):60–65, December 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1983:PEP

[Kap83e] Gadi Kaplan. Power/energy:

Prospects for nuclear power: The nuclear power industry is improving plant operation and safety procedures and reducing maintenance costs as it hopes for a brighter political climate. IEEE Spectrum, 20 (3):58–63, March 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1983:RPA

[Kap83f] Gadi Kaplan. Reliable power: Approaching perfection. IEEE Spectrum, 20(9):64–65, September 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1983:RSB

[Kap83g] Gadi Kaplan. Remote sensing: The best view in town. IEEE Spectrum, 20(9):62–64, September 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1984:EID

[Kap84a] Gadi Kaplan. Electromagnetics: Industrial demand spawns a ‘renaissance ’ in an established discipline. IEEE Spectrum, 21 (11):55–56, November 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1984:FEW

[Kap84b] Gadi Kaplan. Formal education: World survey of schools: Number and type of degrees awarded, curricular composition for first EE degree, and a breakdown of curricular activities are tabulated. IEEE Spectrum, 21(11):

71–75, November 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1984:IEP

[Kap84c] Gadi Kaplan. Industrial electronics: Precision robots, vision systems, and expert systems intensify the computer’s role on the production floor. IEEE Spectrum, 21(1):68–71, January 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1984:JIN

[Kap84d] Gadi Kaplan. Japan’s information network system. IEEE Spectrum, 21(5):50–52, May 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1984:RTL

[Kap84e] Gadi Kaplan. Rail transportation: Linear-induction motors, inverters, and microprocessors play a big role in new systems that roll on steel, rubber, or air. IEEE Spectrum, 21 (1):82–85, January 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1984:TWG

[Kap84f]  Gadi Kaplan.  Taming the

world’s greatest tokamak: Diverse electrotechnical disciplines are used to probe, test, measure, and analyze a premier fusion reactor. IEEE Spectrum, 21 (9):64–69, September 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1985:ACH

[Kap85a] Gadi Kaplan. All the comforts of home. IEEE Spectrum, 22(5):74–78, May 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1985:IER

[Kap85b] Gadi Kaplan. Industrial electronics: Robots, vision systems, smart sensors, adaptive controls, and an automation protocol accelerate integrated manufacturing. IEEE Spectrum, 22(1): 68–71, January 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1985:ICT

[Kap85c] Gadi Kaplan. Invasions and counterinvasions: Tariffs, protectionism, and politics affect the struggle for pieces of the huge international telecommunications pie. IEEE Spectrum, 22 (11):64–69, November 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1985:TNC

[Kap85d] Gadi Kaplan. Transportation: New computer controls make travel more comfortable, thyristors cut costs of motor operation, and new steering systems for cars are tested to improve fuel economy. IEEE Spectrum, 22(1):81–84, January 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1985:XIC

[Kap85e] Gadi Kaplan. The X-29: is it coming or going? IEEE Spectrum, 22(6):54–63, June 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kapron:1985:FOS

[Kap85f] F. P. Kapron. Fiber-optic system tradeoffs: With the choice of short-wavelength and longwavelength technologies available, the designer can tailor a fiber-optic system to specific applications. IEEE Spectrum, 22 (3):68–75, March 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1987:IEA

[Kap87a] Gadi Kaplan. Industrial electronics: Automation for increased competitiveness highprecision, low-cost vision system for OEM monitoring controllers’ faults cutting a part at a computer terminal. IEEE Spectrum, 24(1):50–52, January 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1987:SRG

[Kap87b] Gadi Kaplan. Special report: On good design. IEEE Spectrum, 24(5):27–28, May 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1987:TMM

[Kap87c] Gadi Kaplan. Transportation:

Microprocessors monitor rail car systems software optimizes the headway in Miami’s people mover microprocessors replace relays in mechanical switches and signaling AC propulsion systems improve adhesion, simplify maintenance. IEEE Spectrum, 24 (1):59–61, January 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1987:VPTa

[Kap87d] Gadi Kaplan. Video programs and technology. IEEE Spectrum, 24(6):24, June 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1988:TIE

[Kap88] Gadi Kaplan. Technology ’88: industrial electronics. IEEE Spectrum, 25(1):50–52, January 1988. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kato:1980:WAI

[Kat80]  Y. Kato.  Words into ac-

tion III: a commercial system: Dynamic programming handles varying word lengths in this pioneer continuous-speech recognizer. IEEE Spectrum, 17(6):29, June 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Katz:1987:PVT

[Kat87] I. Katz. Part V. Tools for good design: Automating electronic design: From architectural specification through design, verification, and layout, computer-aided engineering is changing the way EEs do their jobs. IEEE Spectrum, 24(5):55–56, May 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kayton:1987:S

[Kay87] M. Kayton. Speakout. IEEE Spectrum, 24(8):8–11, August 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1987:VPTb

[KC87] Gadi Kaplan and W. S. Ciciora. Video programs and technology. IEEE Spectrum, 24 (8):24–26, August 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Keck:1983:CSM

[Kec83] D. B. Keck. Communications:

Single-mode fibers outperform multimode cables: Successful field trials and imminent commercial use stem from high capacity and the need for fewer repeaters. IEEE Spectrum, 20 (3):30–38, March 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Keller:1987:IOF

[Kel87] M. N. Keller. The industry in 1997: Only flexible makers of high-quality efficiently-produced cars will prosper in the fragmented marketplace of the late 20th century. IEEE Spectrum, 24 (10):72–78, October 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Keyworth:1983:FSL

[Key83] G. A. Keyworth. The future of space: Looking ahead in space. IEEE Spectrum, 20(9): 80–84, September 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kutcher:1983:DDA

[KG83] M. Kutcher and E. Gorin. Datadriven automation. Moving data, not paper, enhances productivity. IEEE Spectrum, 20(5):84–88, May 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kidder:1981:CDM

[Kid81] T. Kidder. Computer design: the microkids and the Hardy Boys: an inside look at how a maverick team from Data General ‘rescued the company by designing a competitive 32-bit superminicomputer in record time’. IEEE Spectrum, 18(9):48–55, September 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kirby:1980:CWP

[Kir80] R. C. Kirby. Communications: WARC plots spectrum use: The world administrative radio conference sought accommodating solutions between developed and third world states. IEEE Spectrum, 17(2):61–63, February 1980. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Kirtley:1983:PES

[Kir83] J. L. Kirtley. Power/energy: Supercool generation: More compact and efficient machines are in the offing as superconducting generators are about to reach the field. IEEE Spectrum, 20 (4):28–35, April 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kobayashi:1984:SOA

[KK84] S. Kobayashi and T. Kimura.

Semiconductor optical amplifiers: Beyond their role as transmitters in fiber-optic systems, laser diodes can be made to amplify light, perhaps promising simple all-optical repeaters. IEEE Spectrum, 21(5):26–33, May 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1985:RSS

[KL85] Gadi Kaplan and E. J. Lerner.

Realism in synthetic speech: Synthesized speech may be intelligible, but it often sounds artificial; researchers are solving that problem. IEEE Spectrum, 22(4):32–37, April 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kessler:1980:ACA

[KLCPT80] D. J. Kessler, P. M. Landry, B. G. Cour-Palais, and R. E. Taylor. Aerospace: Collision avoidance in space: Proliferating payloads and space debris prompt action to prevent accidents. IEEE Spectrum, 17 (6):37–41, June 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kline:1984:OIT

[Kli84] R. R. Kline. Origins of the issues: Theory versus practice, industrial participation, the role of the humanities, and four versus more years are the issues examined. IEEE Spectrum, 21(11): 38–39, November 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kline:1987:ROW

[Kli87] R. Kline. R&D: Organizing for war. IEEE Spectrum, 24(11): 54–61, November 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kolm:1982:SAA

[KM82] H. Kolm and P. Mongeau. Space and aeronautics: an alternative launching medium: Accelerating payloads by electromagnetic force may one day prove an alternative to conventional rockets. IEEE Spectrum, 19 (4):30–36, April 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kummer:1987:CSP

[KM87] D. Kummer and T. Moorman. Case study: PCjr: Misreading the market. IEEE Spectrum, 24(5):30–31, May 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1988:TT

[KM88a] Gadi Kaplan and E. E. Murphy. Technology ’88: transportation. IEEE Spectrum, 25 (1):59–60, January 1988. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Knapp:1988:BAS

[KM88b] B. Knapp and K. Mei. The beauty of ‘almost standard’ VLSI. IEEE Spectrum, 25(2): 35–40, February 1988. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1986:IEC

[KP86] Gadi Kaplan and C. L. Phillips.

Industrial electronics: Computerbased manufacturing systems, expert systems, and process controllers continue to draw attention, as does robotry for the assembly of surface-mounted components. IEEE Spectrum, 23 (1):61–64, January 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kent:1986:EAF

[KS86a] E. W. Kent and M. O. Shneier.

Eyes for automatons: Faster, more accurate, and more consistent than humans can hope to be, machine vision systems are on their way to broad application. IEEE Spectrum, 23 (3):37–47, March 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kim:1986:FOG

[KS86b] B. Y. Kim and H. J. Shaw. Fiberoptic gyroscopes: In harsh, confining environments this advanced gyroscope, a close cousin to the ring laser gyro, offers great advantages. IEEE Spectrum, 23 (3):54–60, March 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kallis:1987:PHS

[KSB87] J. M. Kallis, L. A. Strattan, and T. T. Bui. Programs help spot hot spots: Computerized thermal analysis enhances the reliability of electronic products by identifying potential thermal problems at a point where designers can easily fix them. IEEE Spectrum, 24(3):36–43, March 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kutcher:1983:AIA

[Kut83] M. Kutcher. Automating it all: By automating the transfers and transforms of data, as well as of parts and materials, an enterprise can satisfy conflicting objectives. IEEE Spectrum, 20(5):40–43, May 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kuwabara:1984:JTM

[Kuw84] M. Kuwabara. Japan tests its model communications system:

Digital transmission, communications processing, and an innovative bit-based tariff will be studied in the nation’s milestone Mitaka plant. IEEE Spectrum, 21(5):53–56, May 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kashin:1986:HPP

[KW86] S. Kashin and R. G. Welch.

How the promising PCC car was prematurely derailed: The U.S. electric streetcar industry failed to reap the rewards of its 1930s research, aimed at producing a revolutionary trolley car. IEEE Spectrum, 23(12):48–51, December 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1984:SPP

[KZG^{+}84] Gadi Kaplan, Glenn Zorpette, F. Guterl, E. M. Malstrom, and T. E. Bell. Scarce programs: From power engineering and electromagnetics to telecommunications and instrumentation, specialized programs are in short supply. IEEE Spectrum, 21 (11):54–55, November 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lake:1982:UNC

[Lak82] J. S. Lake. The U.S. navy: Can ships survive? IEEE Spectrum, 19(10):83–85, October 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Larson:1985:WHT

[Lar85] C. Larson. When high-tech goes awry. IEEE Spectrum, 22(5):109, May 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Leahy:1985:SIG

[Lea85] M. F. Leahy. Superfine IC geometries: Gas-plasma etching may supplant chemical etching in manufacturing high-performance integrated circuits. IEEE Spectrum, 22(2):36–43, February 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lemer:1981:CFG

[Lem81a] E. J. Lemer. Computers: Fast graphics use parallel techniques: Designers of computer graphics systems exploit parallel processing to provide the speed needed for interactive performance. IEEE Spectrum, 18 (3):34–38, March 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lemer:1981:MEE

[Lem81b] E. J. Lemer. Military electronics: Electromagnetic pulses: Potential crippler: Three bombs, exploded in space over the U.S., could black out the nation, wipe out communications, and make computers useless. IEEE Spectrum, 18(5):41–46, May 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lemer:1981:MKP

[Lem81c] E. J. Lemer. Military: Killer particle beams: Despite continuing research, the feasibility of building particle-beam weapons and the wisdom of using them remain highly uncertain. IEEE Spectrum, 18(9):44–47, September 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lemer:1983:ETE

[Lem83a] E. J. Lemer. Entertainment technology: Electronically synthesized music: Whether deemed an imitation or an art form of its own, electronic music is competing with conventional instruments. IEEE Spectrum, 20 (6):46–52, June 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lemer:1983:TMD

[Lem83b] E. J. Lemer. Technology and the military: DOD’s Darpa at 25. IEEE Spectrum, 20(8):70–73, August 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lepselter:1981:SSX

[Lep81] M. P. Lepselter. Solid state: Xray lithography breaks the submicrometer barrier; a new Xray lithography system allowed Bell Labs engineers to make the smallest, fastest MOSFETs yet reported. IEEE Spectrum, 18(5):26–29, May 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lerner:1980:CCS

[Ler80a] E. J. Lerner. Computers: Computers that see: Solutions to this difficult challenge may reside in hardware rather than software, or a combination of both. IEEE Spectrum, 17(10):28–33, October 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lerner:1980:EHR

[Ler80b] E. J. Lerner. Electromagnetic hazards: RF radiation: Biological effects: Sharp controversy over the effects of electromagnetic radiation may subside in the wake of new understanding and modified standards. IEEE Spectrum, 17(12):51–59, December 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lerner:1980:PEL

[Ler80c] E. J. Lerner. Power/energy: Low-head hydro power: Small may be beautiful but, so far the economics for small-scale hydroelectric sites cloud a bright potential. IEEE Spectrum, 17(11): 37–39, November 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lerner:1980:PEM

[Ler80d] E. J. Lerner. Power/energy: Magnetic fusion power: Controlled thermonuclear fusion power using magnetic confinement nows seems practical, but the best method remains uncertain. IEEE Spectrum, 17(12):

44–50, December 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lerner:1981:CCA

[Ler81a]  E. J. Lerner.  Computers:

Computer-aided manufacturing. IEEE Spectrum, 18(11): 34–39, November 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lerner:1981:CMS

[Ler81b] E. J. Lerner. Computers: Microcomputer standards: Weighing the pros and cons: Standardization of microcomputer hardware and software has been slowed by manufacturers’ vested interests. IEEE Spectrum, 18 (5):47–50, May 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lerner:1981:CCG

[Ler81c] E. J. Lerner. Computers: The computer graphics revolution: Slowly but surely, new software and hardware are making a reality of graphics specialists’ dreams. IEEE Spectrum, 18 (2):35–39, February 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lerner:1981:IFS

[Ler81d] E. J. Lerner. Instrumentation. I.

Field service. IEEE Spectrum, 18(1):66, January 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lerner:1981:IIL

[Ler81e] E. J. Lerner. Instrumentation. II. Laboratory/bench units. IEEE Spectrum, 18(1):67–68, January 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lerner:1981:MEE

[Ler81f] E. J. Lerner. Military electronics: EMPs and nuclear power: The U.S. Nuclear Regulatory Commission is studying whether electromagnetic pulses from H-bombs could trigger nuclear plant accidents. IEEE Spectrum, 18(6):48–49, June 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lerner:1981:MTR

[Ler81g] E. J. Lerner. Modeling techniques: a repetitive cycle of modeling, testing, and redesign through final production specifications is crucial to eliminating a system’s weak links. IEEE Spectrum, 18(10):50–54, October 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lerner:1981:QCU

[Ler81h] E. J. Lerner. Quality control: US and Japan. IEEE Spectrum, 18 (10):97–101, October 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lerner:1981:RSD

[Ler81i] E. J. Lerner. Reliable systems: Design and tests. IEEE Spectrum, 18(10):50, October 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lerner:1982:ATC

[Ler82a] E. J. Lerner. Air traffic control:

Automating U.S. air lanes: a review: The Federal Aviation Administration’s new 20-year plan to modernize air traffic control faces rough weather. IEEE Spectrum, 19(11):46–51, November 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lerner:1982:CPT

[Ler82b]  E. J. Lerner.  Computers:

Products that talk: Speechsynthesis devices are being incorporated into dozens of products as difficult technical problems are solved. IEEE Spectrum, 19(7):32–37, July 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lerner:1982:CGW

[Ler82c]  E. J. Lerner.  Computers:

The great weather network: Weather forecasting can be improved through better models and speedier reporting. IEEE Spectrum, 19(2):50–57, February 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lerner:1982:ENB

[Ler82d] E. J. Lerner. Electronics and the nuclear battlefield. IEEE Spectrum, 19(10):64–65, October 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lerner:1982:ICA

[Ler82e] E. J. Lerner. Instrumentation: Computer-aided troubleshooting becomes easier and faster; automated testing of VLSI circuits poses challenges. IEEE Spectrum, 19(1):58–60, January 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lerner:1982:MWG

[Ler82f] E. J. Lerner. Micros in ‘white goods’ [consumer electronics]. IEEE Spectrum, 19(4):50–54, April 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lerner:1982:SIA

[Ler82g] E. J. Lerner. Software II: Automating programming: Researchers are beginning to shift the growing burden of programming to the computer itself. IEEE Spectrum, 19(8): 28–33, August 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lerner:1982:SIP

[Ler82h] E. J. Lerner. Software III: Programming for nonprogrammers: New languages are helping those who are not professionally trained to communicate with computers. IEEE Spectrum, 19(8):34–38, August 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lerner:1982:TCS

[Ler82i]  E. J. Lerner.  Tactical c3:

Survival under duress. IEEE Spectrum, 19(10):56–57, October 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lerner:1982:WTM

[Ler82j] E. J. Lerner. War 3: Technology in the military: Command, control, and communications along with countermeasures and intelligence offer some solutions: Strategic c3: a goal unreached. IEEE Spectrum, 19 (10):53–55, October 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lerner:1982:WMP

[Ler82k] E. J. Lerner. War 4: Major power strategies: The U.S. debates disarmament strategies, and NATO’s readiness, while the USSR promises no first use: The superpowers: Nuclear at issue. IEEE Spectrum, 19(10): 66–71, October 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lerner:1983:AAC

[Ler83a] E. J. Lerner. Aerospace: The automated cockpit: The digital revolution catches up with commercial airline cockpits and stirs old questions over how many crew members are needed. IEEE Spectrum, 20(2):57–62, February 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lerner:1983:CWL

[Ler83b] E. J. Lerner. Careers: We look at ourselves: The chief engineer: The duties of this all-important engineering executive may run the gamut from central policy decisions to systems design. IEEE Spectrum, 20(10):65–71, October 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lerner:1983:CSH

[Ler83c] E. J. Lerner. Crossroads in space: Humanization: Man or machine? IEEE Spectrum, 20 (9):28–30, September 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lerner:1983:DTM

[Ler83d] E. J. Lerner. Digital TV: makers bet on VLSI. IEEE Spectrum, 20 (2):39–43, February 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lerner:1983:LES

[Ler83e] E. J. Lerner. Law enforcement: Sleuthing by computer: Forensic computer techniques help law-enforcement agencies identify suspects and organize evidence. IEEE Spectrum, 20 (7):44–49, July 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lerner:1984:BEE

[Ler84a] E. J. Lerner. Biological effect of electromagnetic fields: New findings linking changes in organisms to irradiation by weak fields encourage researchers to posit theories, although research is incomplete. IEEE Spectrum, 21(5):57–69, May 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lerner:1984:DFA

[Ler84b] E. J. Lerner. Data-flow architecture: a decentralized structure based on the flow of data will permit future computers to operate at even higher speeds. IEEE Spectrum, 21(4):57–63, April 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lerner:1984:DRE

[Ler84c] E. J. Lerner. The drive to regulate electromagnetic fields: In the United States, states and localities are imposing regulations, and industry is seeking action by the Federal government. IEEE Spectrum, 21 (3):63–70, March 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lerner:1984:EMM

[Ler84d] E. J. Lerner. Electro ’84: Memories and more: an outstanding professional program, tutorial seminars, and exhibits will greet attendees of the May electronics convention in Boston.

IEEE Spectrum, 21(4):71–75, April 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lerner:1984:GCM

[Ler84e] E. J. Lerner. The great culture machine: brainwasher or educator? IEEE Spectrum, 21 (6):104–108, June 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lerner:1984:IDW

[Ler84f] E. J. Lerner. International data wars are brewing: U.S. telecommunications deregulation, Third World protectionism, and European privacy regulations may complicate or curtail data flow across borders. IEEE Spectrum, 21(7):45–49, July 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lerner:1984:LTR

[Ler84g] E. J. Lerner. Low-temperature reactors for space heating: Simply constructed nuclear plants are being used in the USSR to provide heat at a cost substantially below that of oil-fired plants. IEEE Spectrum, 21(2): 50–55, February 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lerner:1984:NEI

[Ler84h] E. J. Lerner. Nurturing energy infants and orphans: One full-time director is responsible for the DOE’s Division of Advanced Energy Projects, overseeing its $10 million budget and caring for projects that no one else wants. IEEE Spectrum, 21(7):55–59, July 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Linke:1987:COD

[LH87] R. A. Linke and P. S. Henry. Coherent optical detection: a thousand calls on one circuit: Dazzling applications for both longdistance and local networks are in the offing as experimental systems adapt long-successful radio techniques. IEEE Spectrum, 24 (2):52–57, February 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Liu:1984:MPM

[Liu84] T. Liu. Maintenance proctors for mainframe computer: Specialpurpose computers supervise the proper operation of their host computer to improve reliability and reduce down time. IEEE Spectrum, 21(2):36–42, February 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Llacer:1981:MEN

[Lla81] J. Llacer. Medical electronics: Nuclear medical imaging: Enhanced system designs provide sharper diagnostic images and better biochemical-activity analysis. IEEE Spectrum, 18 (7):33–37, July 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lombardo:1981:ASF

[LM81] T. G. Lombardo and N. Mokhoff. Aerospace: Shuttle firsts put to the test: After its near-perfect maiden flight, Columbia needs fine tuning for a trip that will highlight its remote arm and its first payload. IEEE Spectrum, 18(8):34–39, August 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Loferski:1980:PSC

[Lof80a] J. J. Loferski. Photovoltaics I: Solar-cell arrays: Clean and silent, solar-cell arrays are targeted to be competitive by 1986. IEEE Spectrum, 17(2): 26–31, February 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Loferski:1980:PIA

[Lof80b] J. J. Loferski. Photovoltaics IV: Advanced materials: Both polycrystalline and tandem cells promise higher efficiencies than silicon. IEEE Spectrum, 17(2): 37–38, February 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Logsdon:1983:NDC

[Log83] J. M. Logsdon. NASA’s dual challenge: Serving yet striving. IEEE Spectrum, 20(9): 86–89, September 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lombardo:1980:AMS

[Lom80a] T. G. Lombardo. Aerospace and military: Space probes score successes; DOD presses highspeed ICs; MLS begins airport tests. IEEE Spectrum, 17(1): 75–80, January 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lombardo:1980:ASF

[Lom80b] T. G. Lombardo. Air safety:

The Federal Aviation Administration under scrutiny: The U.S. Congress and others have ideas about improving air safety that include restructuring the FAA and altering its charter. IEEE Spectrum, 17(11):53–56, November 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lombardo:1980:CWL

[Lom80c] T. G. Lombardo. Careers: We look at ourselves: a leisure-time portrait: Here’s a profile of the IEEE member, constructed from Spectrum interviews and a comprehensive survey. IEEE Spectrum, 17(4):40–43, April 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lombardo:1980:CPA

[Lom80d] T. G. Lombardo. ‘Collisionproof’ airspace: Reducing the risk of midair collisions is the goal of an air-traffic control system proposed by the FAA. IEEE Spectrum, 17(9):85–88, September 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lombardo:1980:NPT

[Lom80e] T. G. Lombardo. Nuclear power: TMI: an insider’s viewpoint: Texas Instrument’s Patrick E. Haggerty, a member of the President’s Commission on TMI, reviews the nuclear outlook. IEEE Spectrum, 17(5):52–55, May 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lombardo:1980:PEN

[Lom80f] T. G. Lombardo. Power/energy: Nuclear waste disposal: Where and how?: How to safely handle one of the most hazardous substances is an urgent problem shared by technologists and the public. IEEE Spectrum, 17 (12):38–43, December 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lombardo:1981:CIM

[Lom81a] T. G. Lombardo. Communications. IV. Military/aerospace. IEEE Spectrum, 18(1):53–56, January 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lombardo:1981:PET

[Lom81b] T. G. Lombardo. Power/energy: TMI plus 2: Two years after the ordeal at Three Mile Island, a restructured, safer nuclear industry is envisaged. IEEE Spectrum, 18(4):28–29, April 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lombardo:1981:TTT

[Lom81c] T. G. Lombardo. Technology transfer: Technology: Dichotomous tool: It can foster trade, but threaten security. Current U.S. policy has tightened controls on exports to the Soviet Union and Eastern Europe. IEEE Spectrum, 18(5):51–55, May 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lombardo:1982:MAS

[Lom82] T. G. Lombardo. Military/ aerospace systems: NASA Shuttle successes mark a year of transition; meanwhile, military programs are revitalized. IEEE Spectrum, 19(1):45–48, January 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lombardo:1980:NPM

[LP80] T. G. Lombardo and Tekla S.

Perry. Nuclear power: Mitigating the effects of a nuclear accident: Regulators, states, and utilities argue about distance, time, communication, and who’s in charge. IEEE Spectrum, 17(7):30–35, July 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lin:1981:CCE

[LS81] Y. S. Lin and I. L. Sanders.

Computer: Contiguous-element memories increase storage tenfold: an overlapping-triangle configuration for bubble memories may displace permalloychevron types. IEEE Spectrum, 18(2):30–34, February 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucas:1980:CTS

[Luc80] P. M. Lucas. Consumer telephone I: a switch to electronics: Telephone systems will go completely electronic as reed relays are displaced by digital electronic switching. IEEE Spectrum, 17(3):36–37, March 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:1983:PR

[Luc83] R. W. Lucky. The papermill [reflections]. IEEE Spectrum, 20 (9):17, September 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:1984:EMT

[Luc84]  R. W. Lucky.  ‘Entertain

me’: Telvision’s omnipresence [lifestyle]. IEEE Spectrum, 21 (6):85–89, June 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:1987:DDR

[Luc87a] R. W. Lucky. Diminishing dinosaurs [reflections]. IEEE Spectrum, 24(11):8, November 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:1987:FGR

[Luc87b] R. W. Lucky. The footsteps of giants [reflections]. IEEE Spectrum, 24(3):6–7, March 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:1987:MPR

[Luc87c] R. W. Lucky. Management power [reflections]. IEEE Spectrum, 24(5):10, May 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:1987:PR

[Luc87d] R. W. Lucky. Plugging in [reflections]. IEEE Spectrum, 24(1):8, January 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lurito:1984:MWE

[Lur84] P. W. Lurito. The message was electric: Images of electricity were first exploited to sell Victorian products; such advertising helped give electricity itself a high profile and promote technological progress. IEEE Spectrum, 21(9):84–95, September 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Luther:1988:YTL

[Lut88] A. C. Luther. You are there … and in control (digital interactive video). IEEE Spectrum, 25 (9):45–50, September 1988. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Maccoby:1984:NWM

[Mac84] M. Maccoby. A new way of managing: Studies comparing U.S. companies with others suggest that U.S. management has failed to grasp that traditional incentives will not motivate an educated work force. IEEE Spectrum, 21(6):69–72, June 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Maccoby:1988:SDW

[Mac88] M. Maccoby. Self-developers: why the new engineers work. IEEE Spectrum, 25(2):50–53, February 1988. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Madan:1986:ASP

[Mad86] A. Madan. Amorphous silicon: from promise to practice. IEEE Spectrum, 23(9):38–43, September 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mains:1983:TSM

[Mai83] S. Mains. Technology and society: The Minnesota powerline wars: Elaborate efforts by the state to mediate a powerindustry public-interest dispute resulted largely in lessons on what not to do. IEEE Spectrum, 20(7):56–62, July 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Malstrom:1984:MCA

[Mal84] E. M. Malstrom. Manufacturing: Computer applications in the factory spur corresponding efforts in the university. IEEE Spectrum, 21(11):58–59, November 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Middelhoek:1980:MGI

[MAN80] S. Middelhoek, J. B. Angell, and D. J. W. Noorlag. Microprocessors get integrated sensors: Sensing devices and signal processing built into one silicon chip portend a new class of smart sensors. IEEE Spectrum, 17 (2):42–46, February 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Martino:1985:LAC

[Mar85] J. P. Martino. Looking ahead with confidence: Cases of technological forecasts that failed are well known, yet careful use of forecasting techniques can help avoid misdirected R&D. IEEE Spectrum, 22(3):76–81, March 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Martin:1986:GMC

[Mar86] G. J. Martin. Gyroscopes may cease spinning: Optical gyroscopes without moving parts are more reliable and more rugged than conventional mechanical systems. IEEE Spectrum, 23 (2):48–53, February 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mason:1980:SWP

[Mas80a] J. Mason. Standards: Why parts fit: The role of IEC: The international electrotechnical commission is busier than ever on the eve of its 75th anniversary. IEEE Spectrum, 17(6):42–44, June 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mason:1980:CDF

[Mas80b] J. F. Mason. Crack-detectives foil aircraft failure: Detecting flaws in aircraft is based on nondestructive testing with new interpretive aids on the way. IEEE Spectrum, 17(2):47–53, February 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mason:1980:SSV

[Mas80c] J. F. Mason. Solid state: VLSI goes to school: In the international struggle for technological leadership, U.S. universities are developing research centers for very-large-scale integrated circuits. IEEE Spectrum, 17(11): 48–52, November 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Masiello:1985:CPW

[Mas85] R. D. Masiello. Computers in power: a welcome invader: Supervisory computer systems cut utility costs by scheduling generating units, planning against outages, and coordinating the sale of power. IEEE Spectrum, 22(2):51–59, February 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

McCoy:1985:CPB

[McC85] M. D. McCoy. Chip pirates:

Beware the law: a new Federal statute is written to protect semiconductor chip designs; the law, however, contains ambiguities that have yet to be tested. IEEE Spectrum, 22 (7):74–80, July 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

MccGwire:1988:SSP

[Mcc88] Michael MccGwire. Soviet state procurement eschews state of the art. IEEE Spectrum, 25(12): 34–38, November 1988. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Molnar:1984:EEH

[MD84] A. R. Molnar and D. K. Deringer. Education: ‘edutainment’: How to laugh and learn: a grass-roots movement that started in the home is putting school systems under pressure to move computers into the classroom. IEEE Spectrum, 21 (6):114–118, June 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Melvin:1980:CLH

[Mel80] D. K. Melvin. Communications: LSI helps telephones go digital: Low-cost, high-density chips accelerate the digital design of switching systems, terminals, and networks. IEEE Spectrum, 17(6):30–33, June 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Merchant:1983:DDA

[Mer83] M. E. Merchant. Data-driven automation. Production: a dynamic challenge. IEEE Spectrum, 20(5):36–39, May 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Meyer:1986:SSG

[Mey86]  S. M. Meyer.  The Soviet

‘spy gaps’. IEEE Spectrum, 23(7):67–69, July 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Morgan:1985:PLF

[MFNL85] M. G. Morgan, H. K. Florig, I. Nair, and D. Lincoln. Powerline fields and human health: Can exposure to 60-Hz fields cause biological problems in humans? The answer may be found through a process of risk assessment. IEEE Spectrum, 22 (2):62–68, February 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Michalasky:1981:QC

[Mic81] J. Michalasky. Quality control. IEEE Spectrum, 18(10):90, October 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Midwinter:1980:OFS

[Mid80] J. E. Midwinter. Optical fiber systems: Low-loss links: Longwavelength intercity telecommunications increase repeater spacing and promise low cost. IEEE Spectrum, 17(3):45–47, March 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mihalasky:1981:TCQ

[Mih81] J. Mihalasky. Tools for controlling quality: What is done prior to production, as well as thereafter, is as important as the production process itself in ensuring the quality of products. IEEE Spectrum, 18(10):90–93, October 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Miller:1981:RNF

[Mil81] M. M. Miller. Rebirth for nuclear fuel. IEEE Spectrum, 18 (11):52–57, November 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Miller:1987:AMD

[Mil87] R. Miller. Automating mechanical design: Software for desktop computers replaces trial and error in modeling and analysis, and sometimes serves as the front-end database for computerbased manufacturing. IEEE Spectrum, 24(5):59, May 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Miyaoka:1984:DAC

[Miy84] S. Miyaoka. Digital audio is compact and rugged: Highdensity pitted disks in conjunction with sharply focused, precision-tracking laser pickup provide quality sound reproduction. IEEE Spectrum, 21 (3):35–39, March 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Miyazawa:1989:HTJ

[Miy89]  M. Miyazawa.  Homegrown

technology for Japan’s heavy-fit launcher. IEEE Spectrum, 26 (3):51–55, March 1989. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

McLean:1983:CAS

[MMB83] C. McLean, M. Mitchell, and E. Barkmeyer. A computer architecture for small-batch manufacturing: Industry will profit from a system that defines the functions of its componentmanufacturing modules and standardizes their interfaces. IEEE Spectrum, 20(5):59–64, May 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

McMahon:1981:CFO

[MNS81] D. H. McMahon, A. R. Nelson, and W. B. Spillman. Communications: Fiber-optic transducers: Sensing and control systems based on fiber-optic conversion of input variables into modulated light signals are in development. IEEE Spectrum, 18 (12):24–29, December 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moto-oka:1983:JFG

[Mo83] T. Moto-oka. Japan: The fifth generation: a quantum jump in friendliness. IEEE Spectrum, 20 (11):46–51, November 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Munshi:1989:FPE

[MO89] M. Z. A. Munshi and B. B.

Owens. Flat polymer electrolytes promise thin-film power. IEEE Spectrum, 26(8):32–35, August 1989. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mokhoff:1980:CMO

[Mok80a] N. Mokhoff. Communications and microwave: Orbit and spectrum congestion; tests for electronic mail; phone by light. IEEE Spectrum, 17(1): 38–43, January 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mokhoff:1980:CCH

[Mok80b] N. Mokhoff. A computer center for the homeowner. IEEE Spectrum, 17(9):73–78, September 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mokhoff:1980:CEI

[Mok80c] N. Mokhoff. Consumer electronics: Intelligent television receivers; personal computers proliferate; cable TV expands. IEEE Spectrum, 17(1):64–68, January 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mokhoff:1980:GVC

[Mok80d] N. Mokhoff. The global video conference: Both user acceptance and economics block the widespread use of video conferencing. IEEE Spectrum, 17(9): 45–47, September 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mokhoff:1980:MBE

[Mok80e] N. Mokhoff. Microprocessors: Beating an electronic opponent is a challenge: But the road from the concept of an intelligent game to the final product is both challenging and tortuous. IEEE Spectrum, 17(11):26–30, November 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mokhoff:1980:SSF

[Mok80f] N. Mokhoff. Solid state: Four targets of VLSI: Medical electronics, telecommunications, terminals, and scientific computers are candidates for its early application. IEEE Spectrum, 17(6):34–36, June 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mokhoff:1981:CFO

[Mok81a] N. Mokhoff. Communications. I. Fiber optics. IEEE Spectrum, 18 (1):47–49, January 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mokhoff:1981:CII

[Mok81b] N. Mokhoff. Communications. II.

Integrated circuits. IEEE Spectrum, 18(1):50, January 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mokhoff:1981:CEV

[Mok81c] N. Mokhoff. Consumer electronics. I. Video products. IEEE Spectrum, 18(1):78–79, January 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mokhoff:1981:CEIa

[Mok81d] N. Mokhoff. Consumer electronics. II. Personal computers. IEEE Spectrum, 18(1):80, January 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mokhoff:1981:CEIb

[Mok81e] N. Mokhoff. Consumer electronics. III. Toys and games. IEEE Spectrum, 18(1):81, January 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mokhoff:1981:CEIc

[Mok81f] N. Mokhoff. Consumer electronics. IV. Audio products. IEEE Spectrum, 18(1):82, January 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mokhoff:1981:CEC

[Mok81g] N. Mokhoff. Consumer electronics: The continuing quest for 3-D television: The solution lies in deceiving the eye; viewers must envision a realistic image where a planar one exists. IEEE Spectrum, 18(2):48–51, February 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mokhoff:1981:DCL

[Mok81h] N. Mokhoff. Data communications: Local data nets: Untying the office knot: Establishing standards for accessing local data networks may facilitate intra-office communications. IEEE Spectrum, 18 (4):57–59, April 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mokhoff:1981:STP

[Mok81i] N. Mokhoff. A step toward ‘perfect’ resolution [high-definition TV]. IEEE Spectrum, 18(7): 56–60, July 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mokhoff:1982:AHF

[Mok82a] N. Mokhoff. Audio: Highfidelity sound emanates from smaller, modular systems; innovations are seen in tape recorders and turntables. IEEE Spectrum, 19(1):73, January 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mokhoff:1982:BCL

[Mok82b] N. Mokhoff. Business communications: Local area networks provide an architecture for the expanded use of automated office equipment. IEEE Spectrum, 19(1):43–44, January 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mokhoff:1982:CID

[Mok82c] N. Mokhoff. Communication ICs: The digital evolution in telephony continues: a signal processor furthers distributed switching and monolithic SLICs decrease component count. IEEE Spectrum, 19(1):41–42, January 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mokhoff:1982:FOS

[Mok82d] N. Mokhoff. Fiber optics:

Single-mode cables and longwavelength transmission are seen as requisite for next-generation systems. IEEE Spectrum, 19 (1):39–40, January 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mokhoff:1982:PCC

[Mok82e] N. Mokhoff. Personal computers/calculators: New units do more and cost less; peripheral interfaces further extend their capabilities. IEEE Spectrum, 19 (1):69–70, January 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mokhoff:1982:VSL

[Mok82f] N. Mokhoff. Video: On the shopping list of equipment are a new video-disk system and an electronic video still camera. IEEE Spectrum, 19(1): 71–72, January 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Morgan:1981:RAC

[Mor81a] M. G. Morgan. Risk assessment: Choosing and managing technology-induced risk: How much risk should we choose to live with? How should we assess and manage the risks we face? IEEE Spectrum, 18(12): 53–60, December 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Morgan:1981:RAP

[Mor81b] M. G. Morgan. Risk assessment: Probing the question of technology-induced risk: What risk is and how we think about it are analyzed by establishing a framework of real and perceived effects. IEEE Spectrum, 18(11): 58–64, November 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Morgan:1985:CHO

[Mor85] J. Morgan. C3I for the home owner. IEEE Spectrum, 22 (5):46–48, May 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mosteller:1981:ESA

[Mos81] W. Mosteller. The elusiveness of software: In the absence of hard models, rules of thumb have tended to govern a field in which it’s more important for a program to be useful than to be totally correct. IEEE Spectrum, 18(10):46, October 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moye:1980:AEG

[Moy80] J. E. Moye. Aerospace: Europe gets aboard the Shuttle: The European Space Agency’s Spacelab is making the dream of research in space come true. IEEE Spectrum, 17(5):39–43, May 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mueller:1982:ISA

[MR82] R. K. Mueller and R. L. Rylander. Instrumentation:’seeing’ acoustically: New nondestructive techniques permit subsurface examination of materials for density and elasticity. IEEE Spectrum, 19(2):28–33, February 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

McDonald:1984:TWS

[MRRS84] J. F. McDonald, E. H. Rogers, K. Rose, and A. J. Steckl. The trials of wafer-scale integration: Although major technical problems have been overcome since WSI was first tried in the 1960s, commercial companies can’t yet make it fly. IEEE Spectrum, 21 (10):32–39, October 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Maycock:1981:SSE

[MS81] P. D. Maycock and E. N. Stirewalt. Solid state/energy: Solarcell systems that work: In applications from irrigation pumps to village power systems, photovoltaic installations are now producing electricity economically. IEEE Spectrum, 18(9): 40–43, September 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Morkoc:1984:HST

[MS84] H. Morkoc and P. M. Solomon. The HEMT: a superfast transistor: an experimental GaAs– AlGoAs device switches in picoseconds and generates little heat. This is just what supercomputers need. IEEE Spectrum, 21 (2):28–35, February 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Murphy:1987:WHG

[MS87] E. E. Murphy and J. J. Suran. Whatever happened to geniponics? IEEE Spectrum, 24 (11):17, November 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Munson:1988:DPS

[Mun88] R. Munson. Deregulation and the power struggle (electricity supply industry). IEEE Spectrum, 25(5):61–63, May 1988. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Murphy:1985:BMS

[Mur85] B. T. Murphy. Bernard T.

Murphy on solid state: Biopolar and MOS on the same chip? IEEE Spectrum, 22 (1):39–40, January 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Murphy:1988:AAS

[Mur88a] E. E. Murphy. All aboard for solid state (train control system). IEEE Spectrum, 25(13): 42–45, December 1988. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Murphy:1988:ESM

[Mur88b] E. E. Murphy. Endangered species: Many traditional engineering tools and techniques have been pushed into obsolescence by technology itself. IEEE Spectrum, 25(11):118– 119, November 1988. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Murphy:1989:AAT

[Mur89a] E. E. Murphy. Aging aircraft:

too old to fly? IEEE Spectrum, 26(6):28–31, June 1989. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Murphy:1989:MBO

[Mur89b] E. E. Murphy. Making and baking those old-time components. IEEE Spectrum, 26 (3):56–58, March 1989. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Murphy:1989:RWT

[Mur89c] E. E. Murphy. A rising war on terrorists. IEEE Spectrum, 26 (11):33–36, November 1989. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Murphy:1989:TDA

[Mur89d] E. E. Murphy. Technology ’89: design automation. IEEE Spectrum, 26(1):34–37, January 1989. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Murphy:1989:TT

[Mur89e] E. E. Murphy. Technology ’89 transportation. IEEE Spectrum, 26(1):62–63, January 1989. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Musa:1985:JDM

[Mus85] J. D. Musa. John D. Musa on software: Productivity, quality, and human factors. IEEE Spectrum, 22(1):37, January 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Musa:1989:TMS

[Mus89] J. D. Musa. Tools for measuring software reliability. IEEE Spectrum, 26(2):39–42, February 1989. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Murphy:1988:TDA

[MW88] E. E. Murphy and K. I. Werner.

Technology ’88: design automation. IEEE Spectrum, 25(1): 35–37, January 1988. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Nagel:1983:RYS

[Nag83] R. N. Nagel. Robots: Not yet smart enough: Improvements are needed in mechanical design, hands and manipulators, sensors, programming methods, kinematic models of behavior, and intelligence. IEEE Spectrum, 20(5):78–83, May 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Nahin:1983:AGO

[Nah83] Paul J. Nahin. Among the giants: Oliver Heaviside: genius and curmudgeon: he called establishment mathematicians ‘woodenheaded’ when his powerful new methods perplexed them. His fellow EEs didn’t know what to make of him either. IEEE Spectrum, 20(7):63–69, July 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Naisbitt:1983:CHT

[Nai83] J. Naisbitt. Computertrends:

The high-tech, knowledge-based society will spawn a counterbalancing concern about personal and humanistic values. IEEE Spectrum, 20(11):110– 111, November 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Nasko:1983:ECM

[Nas83] H. Nasko. European Common Market. IEEE Spectrum, 20(11): 71–72, November 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Nathan:1986:GAB

[NH86] M. I. Nathan and M. Heiblum. A gallium arsenide ballistic transistor? IEEE Spectrum, 23 (2):45–47, February 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Newman:1987:CSM

[NK87] J. Newman and M. Krakinowski. Case study: Matrix printer: No pulleys, belts, or screw. IEEE Spectrum, 24(5):50–51, May 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Nilsson:1980:CTI

[NS80] B. Nilsson and C. G. Svala. Consumer telephone III: Mixing new with old: an interfacing scheme for analog and digital systems features modular architecture. IEEE Spectrum, 17 (3):41–44, March 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Oakley:1983:GB

[Oak83] B. W. Oakley. Great Britain.

IEEE Spectrum, 20(11):69–71, November 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

OConnor:1986:SD

[O’C86] P. R. O’Connor. 10 steps to deregulation. IEEE Spectrum, 23(5):40–43, May 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Oren:1986:CDR

[OK86] T. Oren and G. A. Kildall. The compact disk ROM: Applications software: It optimizes access time, is compatible with various operating systems, and has the potential for multimedia use. IEEE Spectrum, 23 (4):49–55, April 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Oliver:1987:BR

[Oli87]  B. M. Oliver.  Book re-

views. IEEE Spectrum, 24(11): 10–13, November 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Pake:1985:RDX

[Pak85a] G. E. Pake. Research and development: Xerox Palo Alto Research Center (PARC). IEEE Spectrum, 22(10):54, October 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Pake:1985:RXP

[Pak85b] G. E. Pake. Research at Xerox PARC: a founder’s assessment: The first director of the Xerox Palo Alto Research Center tells how his management philosophy worked in launching the center and making it a success. IEEE Spectrum, 22(10):54–61, October 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Pandhi:1987:UDC

[Pan87a] S. N. Pandhi. The universal data connection: The Integrated-Services Digital Network will have a common communications outlet for plugging in telephones, computers, and such almost like household appliances. IEEE Spectrum, 24(7):31–37, July 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Panofsky:1987:BR

[Pan87b] W. K. H. Panofsky. Book reviews. IEEE Spectrum, 24 (1):11–15, January 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Parker:1984:MFD

[Par84] D. B. Parker. The many faces of data vulnerability: Information can be tapped when resident in memory. It can also be intercepted en route, on paper, and elsewhere. IEEE Spectrum, 21(5):46–49, May 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Panish:1980:SSM

[PC80] M. G. Panish and A. Y. Cho.

Solid state: Molecular beam epitaxy: Streams of molecules in a vacuum chamber build devices from the atom up. IEEE Spectrum, 17(4):18–23, April 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1986:CED

[PC86] Tekla S. Perry and W. S. Ciciora.

Consumer electronics [developments in 1985]. IEEE Spectrum, 23(1):69–72, January 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Poate:1986:PTM

[PD86] J. M. Poate and R. C. Dynes.

Progress toward a metal-base transistor. IEEE Spectrum, 23 (2):38–42, February 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Peattie:1981:QCI

[Pea81] G. Peattie. Quality control for ICs: Manufacturing methods and screening procedures that are good enough today will have to be improved dramatically for the next generation of ICs. IEEE Spectrum, 18(10):93–97, October 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1980:ESS

[Per80a] T. Perry. Energy: Solar shift at SERI: The still young Solar Energy Research Institute is training its sights on shortterm objectives. IEEE Spectrum, 17(4):44–45, April 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1980:CWL

[Per80b] Tekla S. Perry. Careers: We look at ourselves: The EE student: a picture of today’s electrical engineering students emerges from interviews with 16 of them, aged 20 to 24, at 10 U.S. colleges. IEEE Spectrum, 17(10): 46–50, October 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1980:TNS

[Per80c] Tekla S. Perry. Telecommunications: NTIA: a small agency tackles large issues: The cartercreated national telecommunications and information administration influences technical and social communications issues. IEEE Spectrum, 17(12): 60–62, December 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1981:CES

[Per81a] Tekla S. Perry. Consumer electronics: Shrinking products pose problems: Small, well-designed audio and video units capture the buyer’s fancy but challenge human-factors engineers. IEEE Spectrum, 18(12):38–41, December 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1981:EEE

[Per81b] Tekla S. Perry. Education: Engineering education: Coping with the crisis: Government, industry, and academic leaders meet to target solutions to the ominous decline in the academic environment. IEEE Spectrum, 18 (11):65–71, November 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1981:EFE

[Per81c] Tekla S. Perry. Ethics: Five ethical dilemmas: Compare your own responses with those of your colleagues who responded to Spectrum’s survey. IEEE Spectrum, 18(6):53–61, June 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1981:EKH

[Per81d] Tekla S. Perry. Ethics: Knowing how to blow the whistle: Speaking out about an employer’s unethical practices may bring public esteem to the whistle blower, but the many possible pitfalls must be considered. IEEE Spectrum, 18(9):56–61, September 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1981:NPI

[Per81e] Tekla S. Perry. Nuclear power:

INPO: Utilities create their own policeman: The burden of helping demonstrate the nuclear industry’s commitment to safety is on the Institute of Nuclear Power Operations. IEEE Spectrum, 18 (3):58–61, March 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1982:MSS

[Per82a] Tekla S. Perry. Management: Six steps to becoming a successful entrepreneur: Starting one’s own business requires both capital and a ‘feeling’ knowing how others did it may help. IEEE Spectrum, 19(12):46–51, December 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1982:NCE

[Per82b] Tekla S. Perry. NATO credibility: an expert assessment. IEEE Spectrum, 19(10):71–75, October 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1982:SAR

[Per82c] Tekla S. Perry. Standards: Antitrust ruling chills standards setting: a U.S. supreme court ruling has clouded standards setting with the threat of antitrust liability. IEEE Spectrum, 19 (8):52–55, August 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1982:WLO

[Per82d] Tekla S. Perry. We look at ourselves: How to make it big: Engineers as entrepreneurs: Abandoning a secure job and regular paycheck may seem like jumping from a plane without a parachute. What makes engineers leap? IEEE Spectrum, 19(7):55–60, July 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1982:SWE

[Per82e] W. J. Perry. Strategic weapons exploit electronics. IEEE Spectrum, 19(10):94–96, October 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1983:CHD

[Per83a] Tekla S. Perry. Careers: How does your office affect your work? Spectrum visited four U.S. companies to help answer the question posed by this article. Do any of them sound familiar? IEEE Spectrum, 20(10):60–64, October 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1983:CHT

[Per83b] Tekla S. Perry. Commercialization: High time for profits? IEEE Spectrum, 20(9): 49–55, September 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1983:CFP

[Per83c] Tekla S. Perry. Company A: a friendly place … but. IEEE Spectrum, 20(12):66–70, December 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1983:CEC

[Per83d] Tekla S. Perry. Consumer electronics: Component television: This systematic approach is designed to help untangle the maze of wires and cables in many U.S. living rooms. IEEE Spectrum, 20(6):38–45, June 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1983:CEV

[Per83e] Tekla S. Perry. Consumer electronics: Video games: The next wave: Interactive videodisk technology lends reality to arcade games, but purists believe the ultimate lies in realtime computer-generated graphics. IEEE Spectrum, 20(12): 52–59, December 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1983:GKM

[Per83f] Tekla S. Perry. Govermnent: Keyworth at mid-term: an exclusive report from Presidential Science Advisor George A. Keyworth II on technology policy in the Reagan administration. IEEE Spectrum, 20(8): 53–59, August 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1983:TGA

[Per83g] Tekla S. Perry. Technology and government: All branches of government are becoming involved with issues once the sole concern of engineers. IEEE Spectrum, 20 (1):81–83, January 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1984:CBC

[Per84a] Tekla S. Perry. Company B:

The competitive edge: Ambition and aggression are the keys to getting ahead. Are technical talents being wasted in the clash? IEEE Spectrum, 21 (2):67–69, February 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1984:CCQ

[Per84b] Tekla S. Perry. Company C:

Quiescent and secure: Engineers here love security and this company provides that. But does security make for a lean and hungry organization? IEEE Spectrum, 21(4):67–71, April 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1984:CDF

[Per84c] Tekla S. Perry. Company D: a fortress undisturbed: Here the research engineers are well insulated from distracting pressures and free to pursue their interests. But is the real world too remote? IEEE Spectrum, 21(5):77–80, May 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1984:CEN

[Per84d] Tekla S. Perry. Consumer electronics: New audio and video developments await their true test in the marketplace; software support may make the difference. IEEE Spectrum, 21 (1):78–81, January 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1984:TNV

[Per84e] Tekla S. Perry. TV networks vie for viewer’s votes: Sophisticated computer graphics will be the weapons as television technologists compete to speedily convert 1984 election results into exciting home images. IEEE Spectrum, 21(10):68–69, October 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1985:CES

[Per85a]  Tekla S. Perry.  Consumer

electronics: Stereo sound for television and new video technologies grabbed and held the consumer-electronics spotlight in 1984. IEEE Spectrum, 22(1): 77–80, January 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1985:ECD

[Per85b] Tekla S. Perry. Eaglecrest: a commuter’s dream: Communication is everything in this hightech aerie in the foothills of the Sierra Nevada; it is designed to permit the professional to work at home, far from the madding crowd. IEEE Spectrum, 22(5):69–73, May 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1985:LLL

[Per85c] Tekla S. Perry. From lab to lap: Liquid-crystal, electroluminescent, and plasma displays are among the major contenders for use in truly portable computers. IEEE Spectrum, 22 (7):53–59, July 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1985:PPP

[Per85d] Tekla S. Perry. Portia’s perfect pad: Superhigh-tech: It may be too high-tech for some tastes, but an IEEE-member husbandand-wife team waits eagerly to move into this state-of-the-art Dallas showplace. IEEE Spectrum, 22(5):56–63, May 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1986:RWC

[Per86] Tekla S. Perry. The real world:

Controllers and technicians warn of a system stretched to its limits, citing mounting delays, overcrowded radio bands, and a rising number of near-misses. IEEE Spectrum, 23(11):44–49, November 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1987:WPH

[Per87a] Tekla S. Perry. At work or play, he’s the captain: Engineermanager Howard Sachs races sailboats and builds machines. From microprocessors to supercomputers under often-daunting conditions. IEEE Spectrum, 24 (6):56–59, June 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1987:CED

[Per87b] Tekla S. Perry. Consumer electronics: Digital audio tape products delayed for fear of hurting CD sales. Debut of a 4-mm camcorder by Samsung. A new 10kHz bandwidth standard for AM radio. Soon standards for home automation marketplace to be released in 1987. IEEE Spectrum, 24(1):56–58, January 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1987:DD

[Per87c] Tekla S. Perry. The designer’s designer. IEEE Spectrum, 24 (5):42–44, May 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1987:FFJ

[Per87d] Tekla S. Perry. Famous first jobs: Seven leading engineers tell about their first professional jobs, their reasons for choosing them, and how those choices affected their careers. IEEE Spectrum, 24(7):44–51, July 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1987:HFH

[Per87e] Tekla S. Perry. Hypermedia: Finally here: Even the primitive products now entering the market under the general name of hypertext are dazzling, and visionaries see lots more to come. IEEE Spectrum, 24(11):38–39, November 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1987:SGK

[Per87f] Tekla S. Perry. Semiconductors go to Korea: Capitalizing on licensing agreements with U.S. manufacturers, Korean electronics companies are working their way into the global marketplace. IEEE Spectrum, 24 (12):34–39, December 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1988:VJS

[Per88a] Tekla S. Perry. 1975 the VCR:

JVC and Sony transformed an ingenious concept pioneered by Ampex into a major industry. IEEE Spectrum, 25(11): 112–117, November 1988. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1988:EBG

[Per88b] Tekla S. Perry. Electronic banking goes to market. IEEE Spectrum, 25(2):46–49, February 1988. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1988:RKI

[Per88c] Tekla S. Perry. For the record:

Kilby and the IC. IEEE Spectrum, 25(13):40–41, December 1988. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1988:LSL

[Per88d] Tekla S. Perry. The longest survivor loses its grip [tv receiver manufacture]. IEEE Spectrum, 25(8):16–20, August 1988. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1988:PPA

[Per88e] Tekla S. Perry. ‘PostScript’ prints anything: a case history. IEEE Spectrum, 25(5):42–46, May 1988. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1989:BSC

[Per89a] Tekla S. Perry. The battle for the SX-70 [camera integrated circuitry]. IEEE Spectrum, 26 (5):45–49, May 1989. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1989:SPP

[Per89b]  Tekla S. Perry.  In search

of peaceful pastures (defense electrical engineers). IEEE Spectrum, 26(11):49–52, November 1989. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1989:IS

[Per89c] Tekla S. Perry. Intel’s secret is out. IEEE Spectrum, 26 (4):22–28, April 1989. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Pitts:1988:AJL

[PFMP88] F. L. Pitts, B. D. Fisher, V. Mazur, and R. A. Perala. Aircraft jolts from lightning bolts (electronic systems protection). IEEE Spectrum, 25(7):34–38, July 1988. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Posner:1986:CES

[PHG86] E. C. Posner, R. L. Horttor, and T. L. Grant. Communicating from the edge of the solar system: Centimeter-band transmissions, three-dimensional data compression, and laser-based systems may launch unmanned missions to deep space. IEEE Spectrum, 23(6):60–65, June 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Pickett:1982:MEH

[Pic82] J. M. Pickett. Medical electronics. I: Hearing through sight and feeling: Microelectronic aids to speech and comprehension in the form of visual and tactile devices may soon be available. IEEE Spectrum, 19 (4):37–41, April 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Pinet:1980:CTI

[Pin80] A. Pinet. Consumer telephone II:

France goes digital: Low noise, switching versatility, and a reduction in A/D and D/A conversions result. IEEE Spectrum, 17(3):38–40, March 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Pinchot:1985:IEP

[Pin85] G. Pinchot. Introducing the ‘entrepreneur’ [product management]. IEEE Spectrum, 22 (4):74–79, April 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Pritchard:1981:CGS

[PK81] W. L. Pritchard and C. A. Kase. Communications: Getting set for direct-broadcast satellites: Three different technical approaches vie for FCC approval; here a generic system is examined. IEEE Spectrum, 18 (8):22–28, August 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Pashley:1989:FMB

[PL89] R. D. Pashley and S. K. Lai. Flash memories: the best of two worlds. IEEE Spectrum, 26(12): 30–33, December 1989. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Powers:1984:TPC

[Pow84] T. L. Powers. Tom L. Powers on communications: VLSI will reduce bandwidth costs and make complex services possible. IEEE Spectrum, 21(1): 37–38, January 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Pardee:1988:STS

[PP88] S. Pardee and T. P. Pennino.

Software tools speed circuit board design. IEEE Spectrum, 25 (9):40–44, September 1988. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Prokoski:1981:WND

[Pro81] F. Prokoski. Watchdog for nuclear development: Technology can help monitor civilian uses of nuclear energy while discouraging military applications. IEEE Spectrum, 18 (7):51–55, July 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1982:CEV

[PTW82] T. Perry, C. Truxal, and P. Wallich. Consumer electronics: Video games: The electronic big bang. IEEE Spectrum, 19(12): 20–21, December 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1989:MMD

[PV89] Tekla S. Perry and J. Voelcker.

Of mice and menus: designing the user-friendly interface. IEEE Spectrum, 26(9):46–51, September 1989. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1983:MDC

[PW83] Tekla S. Perry and P. Wallich. Microprocessors: Design case history: The Atari Video Computer System: By omitting lots of hardware, designers added flexibility and gave videogame programmers room to be creative. IEEE Spectrum, 20 (3):45–51, March 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1984:CCC

[PW84] Tekla S. Perry and P. Wallich. Can computer crime be stopped? The proliferation of microcomputers in today’s information society has brought with it new problems in protecting both computer systems and their resident intelligence. IEEE Spectrum, 21(5):34–45, May 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1985:CDN

[PW85a] Tekla S. Perry and P. Wallich. Computer displays: New choices, new tradeoffs: Advances on several fronts let users and designers juggle lightness, brightness, and price as well as much-sought-after flatness. IEEE Spectrum, 22(7):52–53, July 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1985:DCH

[PW85b] Tekla S. Perry and P. Wallich. Design case history: the Commodore 64. IEEE Spectrum, 22 (3):48–58, March 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1985:IPI

[PW85c] Tekla S. Perry and P. Wallich.

Inside the PARC: the ‘information architects’. IEEE Spectrum, 22(10):62–76, October 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1986:MMA

[PW86] Tekla S. Perry and P. Wallich.

A matter of margins: Aging radios, radars, and computers in a 23 000-element system help controllers route aircraft during their 45 million hours per year in U.S. airspace. IEEE Spectrum, 23(11):38–44, November 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1986:OBS

[PWZ^{+}86] Tekla S. Perry, P. Wallich, Glenn Zorpette, J. Voelcker, and J. Horgan. Our burdened skies. IEEE Spectrum, 23(11): 36–37, November 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Pyko:1987:WMU

[Pyk87] K. Pyko. Why most U.S. students can’t get into engineering school: Math and science courses are poorly taught and sometimes unavailable, but innovative programs may win more college admissions. IEEE Spectrum, 24 (8):28–31, August 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Pyko:1986:CPE

[PZ86] K. Pyko and Glenn Zorpette. Can power engineering education be re-energized? IEEE Spectrum, 23(12):26–31, December 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1989:SEP

[PZ89] Tekla S. Perry and Glenn Zorpette. Supercomputer experts predict expansive growth. IEEE Spectrum, 26(2):26–33, February 1989. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rabinow:1981:CPA

[Rab81] J. Rabinow. Communications: Postal automation: It can be done: The toughest task is to integrate the technology into a system that is archaic, unwieldy, and a victim of its own inertia. IEEE Spectrum, 18 (3):50–57, March 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rackman:1980:PFP

[Rac80]  M. I. Rackman.  Patents:

Firmware patents can be firm: Firmware-based inventions may be patentable, but there are traps for the unwary. IEEE Spectrum, 17(8):35–40, August 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ramo:1982:ARA

[Ram82]  S. Ramo.  Arms reduction:

an ambitious proposal. IEEE Spectrum, 19(10):108–109, October 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rauch:1987:MCH

[Rau87] K. Rauch. Math chips: How they work: Augmenting microprocessors, they speed up math operations while giving systems designers a variety of performance, cost, and integration options. IEEE Spectrum, 24 (7):25–30, July 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Raskin:1981:MMG

[RD81] D. Raskin and L. A. Davis. Materials: Metallic glasses: Researchers are grappling with novel production techniques that would revolutionize transformer manufacturing. IEEE Spectrum, 18(11):28–33, November 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Reddy:1980:WAI

[Red80] R. Reddy. Words into action II: a task-oriented system: Harpy is an experimental, continuous-speech recognition system that exploits a lowcost minicomputer. IEEE Spectrum, 17(6):26–28, June 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Reid:1980:ICM

[Rei80] A. A. Reid. II: Creating a market for Prestel. IEEE Spectrum, 17(3):29–32, March 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Richter:1988:PQL

[Ric88]  H. L. Richter.  Post-quake

lessons for power utilities. IEEE Spectrum, 25(13):46–48, December 1988. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riley:1981:SCE

[Ril81] W. B. Riley. Southern California Edison bets on energy alternatives. IEEE Spectrum, 18 (8):40–45, August 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riley:1987:CSC

[Ril87] R. Riley. Case study: Cordless drill: Rejuvenating a product. IEEE Spectrum, 24(5):32, May 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rivard:1980:AEM

[Riv80] J. G. Rivard. Automotive electronics: Microcomputers hit the road: U.S. automakers are turning to digital electronic controls to boost fuel economy and reduce pollution. IEEE Spectrum, 17 (11):44–47, November 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rivard:1987:ASE

[Riv87] J. G. Rivard. The automobile in 1997: Systems engineering promises integrated performance features, more luxuries, and advanced diagnostics. IEEE Spectrum, 24(10):67–71, October 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ragaller:1980:PEM

[RK80] K. Ragaller and U. Kogelschatz.

Power/energy: Managing switching arcs: Plasma physics research tells us how to dynamically control arcs in high-power circuit breakers. IEEE Spectrum, 17(8):20–25, August 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rogers:1981:PER

[Rog81] T. F. Rogers. Power/energy: Reflector satellites for solar power: Solar-cell arrays in arid zones plus microwave transmission to space and back can send power where needed. IEEE Spectrum, 18(7):38–43, July 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Roizen:1986:DIP

[Roi86]  J. Roizen.  Dubrovnik im-

passe puts high-definition TV on hold: a report on the turmoil in setting program standards concludes with a scenario for future developments. IEEE Spectrum, 23(9):32–37, September 1986. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Rosenstein:1980:PTN

[Ros80] A. B. Rosenstein. Politics of technology: National professions foundation: an idea whose time has come?: With U.S. leadership in technology and the professions declining and the very quality of life in jeopardy, an NPF is recommended to remedy the situation. IEEE Spectrum, 17(8):41–43, August 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rosenbaum:1983:PMC

[Ros83] J. D. Rosenbaum. A propitious marriage: CAD and manufacturing: One typical CAD system integrated with manufacturing allows part design in man-hours instead of man-days and generates numerical-control tapes in a day. IEEE Spectrum, 20(5):49–52, May 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rosenberg:1984:OEE

[Ros84] R. Rosenberg. The origins of EE education: a matter of degree. IEEE Spectrum, 21 (7):60–68, July 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rosenbaum:1988:HIE

[Ros88] A. D. Rosenbaum. How Israeli engineers improve U.S. weaponry. IEEE Spectrum, 25 (12):38–39, November 1988. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rowe:1982:PER

[Row82] W. D. Rowe. Power/energy: Renewable energy: Target for 2050: Massive capital investment, international cooperation, and technical breakthroughs will be key in attaining this goal. IEEE Spectrum, 19(2):58–63, February 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rubinstein:1980:TSN

[Rub80] E. Rubinstein. Technology and society: Nuclear power pervades the public interest while alternate solutions are investigated. IEEE Spectrum, 17(1): 87–90, January 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rubinstein:1982:TWP

[Rub82] E. Rubinstein. Technology in war and peace. IEEE Spectrum, 19(10):34–35, October 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rubinstein:1983:S

[Rub83] E. Rubinstein. Space 25. IEEE Spectrum, 20(9):26–27, September 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rubinstein:1984:BTI

[Rub84] E. Rubinstein. Beyond 1984:

Technology and the individual.

IEEE Spectrum, 21(6):26–27, June 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rumennik:1985:PDC

[Rum85] V. Rumennik. Power devices are in the chips: New power integrated circuits, called PICs, put both power-handling semiconductors and logic on the same IC chip; soon they may be part of every household appliance. IEEE Spectrum, 22 (7):42–51, July 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ryder:1984:WIW

[Ryd84] J. D. Ryder. The way it was: a renowned educator gives a firsthand account of the decades that shaped EE education. IEEE Spectrum, 21(11):39–43, November 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Seth:1985:GCT

[SA85] S. C. Seth and V. D. Agrawal. Gutting chip-testing costs: Designing VLSI circuits for testability is the most efficient way to reduce the relative costs of assuring high chip reliability. IEEE Spectrum, 22(4):38–45, April 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sanchez:1987:F

[San87] H. Sanchez. Forum. IEEE Spectrum, 24(5):11–12, May 1987.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Santo:1988:BDL

[San88a] B. Santo. 1970 1k-bit DRAM:

The LSI era was born in a technical competition between semiconductor and core memories. IEEE Spectrum, 25(11): 108–112, November 1988. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Santo:1988:BIT

[San88b] B. Santo. Bolstering infrastructure through boom and bust. IEEE Spectrum, 25(12): 51–55, November 1988. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Santo:1988:MJC

[San88c]  B. Santo.  From the mir-

ror: a just-concluded poll reveals some unexpected selfperceptions. IEEE Spectrum, 25(11):136–141, November 1988. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Santo:1989:BCB

[San89a] B. Santo. BiCMOS circuitry: the best of both worlds. IEEE Spectrum, 26(5):50–53, May 1989. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Santo:1989:ISS

[San89b] B. Santo. Industry: seeking strength in mergers. IEEE Spectrum, 26(11):64–66, November 1989. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Santo:1989:RSV

[San89c] B. Santo. R&D: salvation in verification. IEEE Spectrum, 26 (11):66–69, November 1989. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Santo:1989:TIE

[San89d] B. Santo. Technology ’89: industrial electronics. IEEE Spectrum, 26(1):53–55, January 1989. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Santo:1989:TSS

[San89e] B. Santo. Technology ’89: solid state. IEEE Spectrum, 26(1): 47–49, January 1989. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Santo:1989:XRL

[San89f] B. Santo. X-ray lithography: the best is yet to come. IEEE Spectrum, 26(2):48–49, February 1989. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Songco:1980:EMH

[SAPM80] D. C. Songco, S. I. Allen, P. S. Plexico, and R. A. Morford. Electronics in medicine: How computers talk to the blind: Sightless computer programmers provide input by keyboard and are responded to by a synthetic speech unit. IEEE Spectrum, 17(5):34–38, May 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sargent:1982:MEI

[Sar82] D. C. Sargent. Medical electronics III: Rhythmic cues aid lip readers: Through knowledge of the frequency and intensity characteristics of syllables, the deaf can learn to relate speech with text. IEEE Spectrum, 19 (4):46–49, April 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Schiecke:1981:CEP

[Sch81] K. Schiecke. Consumer electronics: Projection television: Correcting distortions. IEEE Spectrum, 18(11):40–45, November 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Smith:1981:CEW

[SD81] E. K. Smith and E. W. Davis.

Consumer electronics: Windinduced ions thwart TV reception: High-altitude ionized layers can bounce a signal up to 1500 kilometers outside its normal area of coverage. IEEE Spectrum, 18(2):52–55, February 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Strojwas:1988:VLD

[SD88] A. J. Strojwas and S. W. Director. VLSI: linking design and manufacturing. IEEE Spectrum, 25(10):24–28, October 1988. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Searle:1989:BWN

[Sea89] Lloyd Searle. The bombsight war: Norden vs. Sperry. IEEE Spectrum, 26(9):60–64, September 1989. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Schaefer:1982:CBS

[SF82] D. H. Schaefer and J. R. Fischer. Computers: Beyond the supercomputer: NASA is developing ultrafast machines to process satellite data that would choke a mere supercomputer. IEEE Spectrum, 19 (3):32–37, March 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Stix:1988:FC

[SFV^{+}88] Gary Stix, K. Fitzgerald, J. Voelcker, Glenn Zorpette, T. E. Bell, F. Guterl, B. Santo, and Tekla S. Perry. Front cover. IEEE Spectrum, 25(11):c1, November 1988. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Shape:1985:FFR

[SH85] R. E. Shape and N. B. Hornback. The friendly fields of RF: Encouraged by successes in fusing fractured bones, researchers look to electromagnetic fields for healing wounds and shrinking tumors. IEEE Spectrum, 22(6):64–69, June 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Shackil:1980:CWL

[Sha80a] A. F. Shackil. Careers: We look at ourselves: The design engineer: an introspective look at the personal and professional attributes of today’s designers, gleaned from interviews with several of them. IEEE Spectrum, 17(12):63–67, December 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Shackil:1980:EHE

[Sha80b] A. F. Shackil. An electronic ‘human eye’. IEEE Spectrum, 17 (9):89–91, September 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Shackil:1980:IDS

[Sha80c]  A. F. Shackil.  Instrumen-

tation: Digital storage oscilloscopes: Digital scopes may evolve like the digital multimeter, offering data-capture, transfer, and processing in addition to present capabilities. IEEE Spectrum, 17(7):22–25, July 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Shackil:1980:IGS

[Sha80d] A. F. Shackil. Instruments get smarter: Innovative digital design techniques lead to more intelligent, versatile, and communicative instruments. IEEE Spectrum, 17(1):50–53, January 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Shackil:1980:MPW

[Sha80e] A. F. Shackil. Meeting preview: Wescon ’80: Electronics at Disneyland: Anaheim will be the scene of high technology disclosures. IEEE Spectrum, 17 (8):44–45, August 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Shackil:1980:PPE

[Sha80f] A. F. Shackil. Program preview: Electro ’80: Microprocessors star: Opening May 13 in Boston, the 3-day meeting features three blocks of technical sessions daily. IEEE Spectrum, 17(4):46–47, April 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Shackil:1981:CSS

[Sha81a] A. F. Shackil. Communications shares spotlight with microprocessors. IEEE Spectrum, 18 (3):62–63, March 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Shackil:1981:IIM

[Sha81b] A. F. Shackil. Instrumentation. III. Medical equipment. IEEE Spectrum, 18(1):69, January 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Shackil:1981:MPW

[Sha81c] A. F. Shackil. Meeting preview: Wescon celebrates its 30th in San Francisco: Technical sessions will concentrate on ICs and the microprocessor and business sessions on how to cope in today’s economy. IEEE Spectrum, 18(8):51–53, August 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Shackil:1981:MMM

[Sha81d] A. F. Shackil. Microcomputers: Microprocessors and the M.D.: a new breed of smart medical equipment can diagnose, monitor, analyze, and rehabilitate. IEEE Spectrum, 18 (4):45–49, April 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Shackil:1981:MDCa

[Sha81e]  A. F. Shackil.  Micropro-

cessors: Design case history: Singer’s electronic sewing machine: Putting a microprocessor in a sewing machine surprised the skeptics and put its developers for ahead of the competition. IEEE Spectrum, 18 (2):40–43, February 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Shackil:1981:MDCb

[Sha81f]  A. F. Shackil.  Micropro-

cessors: Design case history: Wang’s word processor: a novel distributed-processing design enabled its creators to revolutionize text-handling systems. IEEE Spectrum, 18(8):29–33, August 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Shackil:1982:MPR

[Sha82] A. P. Shackil. Microprocessors: Progress is reported in programming, speed, complexity, and versatility; new 32-bit devices are announced. IEEE Spectrum, 19(1):32–33, January 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Shaiken:1984:AIB

[Sha84] H. Shaiken. Automation in industry: Bleaching the blue collar: Computerization can increase workers’ output, but it may also distance product designers from assembly and process requirements. IEEE Spectrum, 21(6):77–79, June 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Shapiro:1987:SDL

[Sha87] R. M. Shapiro. Software for designing large systems: How to automate in creating, verifying, and analyzing systems when subelements, components, and interrelationships are too numerous to track. IEEE Spectrum, 24(5):62, May 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Shogase:1988:VSC

[Sho88] H. Shogase. The very smart card: a plastic pocket bank. IEEE Spectrum, 25(10):35–39, October 1988. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Shumate:1989:OFR

[Shu89] P. W. Shumate. Optical fibers reach into homes. IEEE Spectrum, 26(2):43–47, February 1989. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Silverstein:1989:HSF

[Sil89] M. F. Silverstein. How to select a flat-panel display. IEEE Spectrum, 26(9):41–45, September 1989. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Simpson:1983:SSS

[Sim83] T. R. Simpson. Solar system sorties: Ending the hiatus. IEEE Spectrum, 20(9):84–86, September 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Singh:1987:MPP

[Sin87] M. Singh. Modeling power plants. IEEE Spectrum, 24 (5):56–58, May 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Shibata:1989:OEL

[SK89] J. Shibata and T. Kajiwara. Optics and electronics are living together. IEEE Spectrum, 26 (2):34–38, February 1989. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Stiffler:1981:HCF

[SM81] J. J. Stiffler and W. Mosteller. How computers fail. IEEE Spectrum, 18(10):44, October 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Smith:1987:CMD

[Smi87] A. J. Smith. Cache memory design: an evolving art: Designers are looking to line size, degree of associativity, and virtual addresses as important parameters in speeding up the operation. IEEE Spectrum, 24(12): 40–45, December 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Solomon:1988:SCF

[Sol88] C. W. Solomon. Switchedcapacitor filters: precise, compact, inexpensive. IEEE Spectrum, 25(6):28–32, June 1988. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Spitzer:1986:ADJ

[Spi86] C. R. Spitzer. All-digital jets are taking off: Aboard developmental commercial and military aircraft, digital electronics score high in compactness, control flexibility, and reliability. IEEE Spectrum, 23(9): 51–56, September 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sterzer:1980:EMR

[SPM^{+}80] F. Sterzer, R. W. Paglione, J. Mendecki, E. Friedenthal, and C. Botstein. Electronics in medicine: RF therapy for malignancy: Heating of malignant tissues (hyperthermia) by RF radiation presents a new tool in the arsenal of weapons against cancer. IEEE Spectrum, 17(12): 32–37, December 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Scales:1984:ASR

[SS84] W. C. Scales and R. Swanson.

Air and sea rescue via satellite systems: Even experimental systems have helped survivors of air and sea accidents. Two different approaches are discussed. IEEE Spectrum, 21 (3):48–52, March 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Smith:1981:SSB

[ST81] P. W. Smith and W. J. Tomlinson. Solid state: Bistable optical devices promise subpicosecond switching: Extensive research in materials and phenomena could lead to their ultimate use in optical communications, despite high power dissipation. IEEE Spectrum, 18(6):26–33, June 1981. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Staff:1986:AJR

[Sta86] S. Staff. Assessing Japan’s role in telecommunications: Building on successes in devices and small systems, Japan’s R&D turns to making the large systems in which the United States has excelled. IEEE Spectrum, 23(6):47–52, June 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Staff:1987:HCH

[Sta87a]  S. Staff.  How computers

helped stampede the stock market: Many analysts blame automated program trading for the worldwide plunge. But the machines only did what they were told. IEEE Spectrum, 24(12): 30–33, December 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Staff:1987:TMT

[Sta87b] S. Staff. Too much, too soon: Information overload: As designers strive to build more powerful hardware and software systems, the risk is that the complexity will overwhelm humans who use them. IEEE Spectrum, 24(6):51–55, June 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Stern:1981:CEU

[Ste81] N. Stern. Computers: From Eniac to Univac: As inventors, Eckert and Mauchly were clear successes, but as entrepreneurs they had some difficult times. IEEE Spectrum, 18 (12):61–69, December 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Stiffler:1981:VCM

[Sti81] J. J. Stiffler. The vulnerability of computers: Malfunctions may be due to “illegal” operations, to hardware failures, or to combinations of hardware and software failures that simply elude pinpointing. IEEE Spectrum, 18 (10):44–45, October 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Stix:1988:MLO

[Sti88a] Gary Stix. 1969 Moon Lander: One small step for a man was made possible by one giant step for engineering. IEEE Spectrum, 25(11):76–82, November 1988. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Stix:1988:CAE

[Sti88b] Gary Stix. Competition alters equation for industry. IEEE Spectrum, 25(12):47–51, November 1988. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Stix:1988:DSA

[Sti88c] Gary Stix. ‘Digitized sunshine’ aids architects. IEEE Spectrum, 25(3):46–49, March 1988. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Stix:1988:TWC

[Sti88d] Gary Stix. Telephone wiring:

a conduit for networking standards. IEEE Spectrum, 25 (6):38–41, June 1988. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Stix:1988:WHL

[Sti88e] Gary Stix. Working hot: life at 765 kV. IEEE Spectrum, 25 (9):54–56, September 1988. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Stix:1989:BTW

[Sti89a] Gary Stix. Bhopal: a tragedy in waiting. IEEE Spectrum, 26(6):47–50, June 1989. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Stix:1989:SPD

[Sti89b] Gary Stix. From swords to plowshares (defense spending). IEEE Spectrum, 26(11):45–49, November 1989. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Stix:1989:MHC

[Sti89c] Gary Stix. Manufacturing hurdles challenge large-LCD developers. IEEE Spectrum, 26(9): 36–40, September 1989. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Stix:1989:TDC

[Sti89d] Gary Stix. Technology ’89: data communications. IEEE Spectrum, 26(1):43–46, January 1989. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Stix:1989:TP

[Sti89e] Gary Stix. Technology ’89: peripherals. IEEE Spectrum, 26 (1):40, January 1989. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Schweppe:1982:PEH

[STK82] F. C. Schweppe, R. D. Tabors, and J. L. Kirtley. Power/energy: Homeostatic control for electric power usage: a new scheme for putting the customer in the control loop would exploit microprocessors to deliver energy more efficiently. IEEE Spectrum, 19(7):44–48, July 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sugarman:1980:BCR

[Sug80a] R. Sugarman. The blue-collar robot: The machine shop and the assembly line will be early beneficiaries of advanced robotic systems. IEEE Spectrum, 17(9): 53–60, September 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sugarman:1980:CSC

[Sug80b] R. Sugarman. Computers: Superpower computers: The omnipresent microprocessor can hardly supplant the large highspeed computer in modeling complex systems and phenomena. IEEE Spectrum, 17(4): 28–34, April 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sugarman:1980:TOS

[Sug80c] R. Sugarman. The technologist’s own ‘supercomputer’. IEEE Spectrum, 17(9):49–52, September 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sumney:1980:SSV

[Sum80] L. W. Sumney. Solid state: VLSI with a vengeance: The Pentagon’s very-high-speed integrated circuits project requires revolutionary advances in systems and IC technology. IEEE Spectrum, 17(4):24–27, April 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sumney:1982:SSV

[Sum82a] L. W. Sumney. Solid state: VH-SIC: a status report: The first two years of the program’s lifespan is viewed by some as gestation and others as incubation. When will the baby walk? IEEE Spectrum, 19(12):34–39, December 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sumney:1982:VPL

[Sum82b] L. W. Sumney. VHSIC: a promise of leverage. IEEE Spectrum, 19(10):93–94, October 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sugarman:1983:CLS

[SW83] R. Sugarman and P. Wallich. Computers: The limits to simulation: Despite the promise of superfast computers, turbulence and nonlinearity cause problems no foreseeable machine can handle. IEEE Spectrum, 20 (4):36–41, April 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Santo:1988:WSD

[SW88] B. Santo and K. Wollard. The world of silicon: it’s dog eat dog [semiconductor industry]. IEEE Spectrum, 25(9):30–39, September 1988. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Swerling:1982:CCA

[Swe82]  S. Swerling.  Computers:

Computer-aided engineering: Terminals that simulate, test, and debug logic designs promise to change radically the way designers do their jobs. IEEE Spectrum, 19(11):37–41, November 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sykes:1989:RMP

[Syk89] F. E. Sykes. Resonant-mode power supplies: a primer. IEEE Spectrum, 26(5):36–39, May 1989. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Talavage:1983:SMS

[Tal83] J. Talavage. Simulating manufacturing systems: Computer simulations permit run-throughs of systems before they are built and evaluation of proposed changes without shutting existing systems down. IEEE Spectrum, 20(5):53–55, May 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Tannas:1986:ECP

[Tan86] L. E. Tannas. Electroluminescence catches the public eye: Rugged, readable, and compact, electroluminescent technology is breaking the price barrier: Designers are now using it in middle-range products. IEEE Spectrum, 23(10):37–42, October 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Tannas:1989:FPD

[Tan89a] L. E. Tannas. Flat-panel displays displace large, heavy, powerhungry CRTs. IEEE Spectrum, 26(9):34–35, September 1989. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Tannas:1989:HDJ

[Tan89b] L. E. Tannas. HDTV displays in Japan: projection — CRT systems on top. IEEE Spectrum, 26 (10):31–33, October 1989. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Taylor:1989:PNV

[Tay89] B. N. Taylor. Preparing for the new volt and ohm. IEEE Spectrum, 26(7):20–23, July 1989.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Turner:1987:DMB

[TBH87] D. B. Turner, R. D. Burns, and H. Hecht. Designing micro-based systems for fail-safe travel: For reliable control of railroads, aircraft, and space vehicles, designers are harnessing the power of the microprocessor. IEEE Spectrum, 24(2):58–63, February 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Truscello:1984:NEP

[TD84] V. C. Truscello and H. S. Davis.

Nuclear-electric power in space: Nuclear-reactor power systems could produce an abundance of new applications in space, hut design hurdles abound. IEEE Spectrum, 21(12):58–65, December 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Teller:1982:EPN

[Tel82] Edward Teller. Electromagnetic pulses from nuclear explosions. IEEE Spectrum, 19(10):65, October 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Tennant:1982:SSC

[Ten82] S. M. Tennant. Space systems: Can they withstand attack? IEEE Spectrum, 19 (10):99, October 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Tennant:1987:F

[Ten87] D. B. Tennant. Forum. IEEE Spectrum, 24(9):12–13, September 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Thackeray:1983:WTB

[Tha83] D. P. C. Thackeray. When tubes beat crystals: early radio detectors. IEEE Spectrum, 20(3):64–69, March 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Torrero:1980:SRI

[Tor80] E. A. Torrero. A special report: And if hasn’t been, at least not yet. A study of 11 wanted projects requiring technical, economic, or institutional breakthroughs before fruition is possible. IEEE Spectrum, 17(9): 30–32, September 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Torrero:1983:TC

[Tor83] E. A. Torrero. Tomorrow’s computers. IEEE Spectrum, 20(11): 34–35, November 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Torrero:1984:ECC

[Tor84] E. A. Torrero. Education: The challenges are classic. IEEE Spectrum, 21(11):36–37, November 1984. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Torerro:1985:HAF

[Tor85] E. A. Torerro. Hello again the future of telecommunications. IEEE Spectrum, 22(11): 44–45, November 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Trimberger:1983:RMT

[TP83] S. Trimberger and R. F. W.

Pease. Reaching for the milliontransistor chip. IEEE Spectrum, 20(11):100–105, November 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Tribus:1980:SEP

[Tri80] M. Tribus. Sociotechnology: Engineers and politics: Seven survival commandments: a catechism to help engineers champion technology through the morass of political processes in the U.S. IEEE Spectrum, 17 (5):49–51, May 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Trimberger:1982:CAD

[Tri82] S. Trimberger. Computer-aided design: Automating chip layout: New computer-based layout systems are faster than their human counterparts and produce designs that almost match those created manually. IEEE Spectrum, 19(6):38–45, June 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Truxal:1982:CEA

[Tru82] C. Truxal. Consumer electronics: Americanizing videotex: Even as U.S. companies prepare to enter the videotex marketplace, what the consumer wants is not yet clear. IEEE Spectrum, 19(11): 52–57, November 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Truxal:1983:CRE

[Tru83a] C. Truxal. Coping with recession: Engineering schools grapple with hard times by introducing new programs, funding arrangements, and curricula. IEEE Spectrum, 20(1): 79–80, January 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Truxal:1983:MEW

[Tru83b] C. Truxal. Medical electronics: Watching the brain at work: Evoked potentials are a tool for neurologists and psychologists studying the brain’s electrical activity. IEEE Spectrum, 20(3):52–57, March 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Truxal:1983:MWE

[Tru83c] C. Truxal. Minorities: The woman engineer: Women in engineering are no longer the rare phenomenon of yore, yet they continue to face many challenges at the workplace. IEEE Spectrum, 20(4):58–62, April 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Truxal:1983:PES

[Tru83d] C. Truxal. The professionals: Experts are in short supply: Who will design and maintain future factories? A new breed of engineer is needed. Academia and industry seek new avenues for the necessary training. IEEE Spectrum, 20 (5):93–94, May 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Truxal:1983:VHT

[Tru83e] C. Truxal. The very hightech home: Consumers move one step closer to the electronic household as video, audio, and computer-game technologies merge. IEEE Spectrum, 20(1):64–67, January 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Truxal:1984:BST

[Tru84a] C. Truxal. Buying, selling, and trading technology: What technologies can engineers export? What information can they exchange? Here’s an update on U.S. technology-transfer controls. IEEE Spectrum, 21 (2):58–66, February 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Truxal:1984:ECN

[Tru84b] C. Truxal. EE careers: New directions but old issues: To manage or not to manage, how to avoid obsolescence, how to maintain the joy of creating — the old debates haven’t slackened as EES enter their second century. IEEE Spectrum, 21(6):53–58, June 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Tulloss:1983:IAB

[Tul83] R. E. Tulloss. Instrumentation: Automated board testing: Coping with complex circuits: As printed circuit boards become harder to test due to VLSI components, automated systems and new design techniques are coming to the rescue. IEEE Spectrum, 20(7):38–43, July 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Turino:1987:SMC

[Tur87] J. L. Turino. The surfacemounting challenge: Access to circuits. IEEE Spectrum, 24 (5):48–50, May 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Tustin:1986:RRS

[Tus86] W. Tustin. Recipe for reliability: Shake and bake: Environmental stress screening (ESS), a military convention, is moving into other markets as an inexpensive way to improve product reliability and identify design flaws. IEEE Spectrum, 23 (12):37–42, December 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Tan:1989:DAT

[TV89] E. Tan and B. Vermuelen. Digital audio tape for data storage. IEEE Spectrum, 26(10): 34–38, October 1989. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Vaelcker:1986:HCC

[Vae86] J. Vaelcker. Helping computers communicate: The Open Systems Interconnection model promises compatibility for a variety of computer systems, although not all its functions are yet defined. IEEE Spectrum, 23(3):61–70, March 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

VanValkenburg:1984:TTT

[Van84] M. E. Van Valkenburg. Technology as a tool for teaching EEs: As electrical engineering departments become even more computerized, colleges will admit more students and employ innovative techniques for teaching them. IEEE Spectrum, 21 (6):123–127, June 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Voelcker:1986:IPM

[VB86] J. Voelcker and J. S. Birnbaum. Instrumentation: Powerful microprocessors and standard buses help advance modular card instrumentation; meanwhile bench test equipment gets new features. IEEE Spectrum, 23(1):57–60, January 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Voelcker:1987:FFb

[VDT87] J. Voelcker, J. R. Devaney, and R. W. Thomas. Faults failures. IEEE Spectrum, 24(8):17, August 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Vera:1986:BLH

[Ver86] J. G. Vera. Back on line in 72 hours: To reconnect Mexico City’s complex electric grid after last September’s earthquake, engineers zeroed in on the most critical repairs, aided by remote operation of substations. IEEE Spectrum, 23 (6):41–46, June 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Voelcker:1986:MNA

[VH86] J. Voelcker and H. A. Hoeschen. Microprocessors: New architectures increase speed, and RISC processors are moving from university laboratories into commercial products. IEEE Spectrum, 23(1):46–48, January 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Voelcker:1986:GPDa

[Voe86a] J. Voelcker. ‘Gold-plated’ design: Redundancy and extremely high safety factors allow aircraft to withstand forces far exceeding those encountered during routine flights. IEEE Spectrum, 23(11): 56–64, November 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Voelcker:1986:GPDb

[Voe86b] J. Voelcker. ‘Gold-plated’ design: Redundancy and extremely high safety factors allow aircraft to withstand forces far exceeding those encountered during routine flights. IEEE Spectrum, 23(11): 65–66, November 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Voelcker:1986:MYP

[Voe86c] J. Voelcker. Making your PC behave like another. IEEE Spectrum, 23(10):61–67, October 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Voelcker:1986:PSD

[Voe86d] J. Voelcker. Profitability and safety: Deregulation encourages airlines to reduce operating costs to stay competitive, while imperiling traditional safety practices. IEEE Spectrum, 23 (11):77, November 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Voelcker:1987:APP

[Voe87a] J. Voelcker. Ada: From promise to practice? Compilers and other tools put the DOD programming language within the designer’s reach, but embedded applications, like those in avionics, have been slow to arrive. IEEE Spectrum, 24(4):44–49, April 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Voelcker:1987:CSL

[Voe87b] J. Voelcker. Case study: Luxury sedan: Modeled by computer. IEEE Spectrum, 24 (5):60–62, May 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Voelcker:1987:CSP

[Voe87c] J. Voelcker. Case study: PC upgrade: Better graphics and sound. IEEE Spectrum, 24 (5):68–70, May 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Voelcker:1987:CNW

[Voe87d] J. Voelcker. Communications and navigation: From walkietalkies and FM vehicular radios to navigational aids, electronics extended military command and control to unprecedented distances. IEEE Spectrum, 24 (9):42–46, September 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Voelcker:1987:ESP

[Voe87e] J. Voelcker. The experts speak out: The philosophy of competitive automaking has changed, say industry leaders; whatever methods are used, the details must be precisely right. IEEE Spectrum, 24(10):57–60, October 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Voelcker:1987:FFa

[Voe87f] J. Voelcker. Faults failures.

IEEE Spectrum, 24(6):17, June 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Voelcker:1987:IAU

[Voe87g] J. Voelcker. Instrumentation:

Advanced user interfaces appear in oscilloscopes low-cost remote instrument control provided by PCs instrumentation for fiber optics moves ahead. IEEE Spectrum, 24(1):48–49, January 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Voelcker:1987:PVR

[Voe87h] J. Voelcker. Part VI. redesigning:

Why and how: Learning from earlier systems: When redesigning to correct flaws, add capabilities, or update aging features, figuring out what to keep and why is as important as knowing what to change. IEEE Spectrum, 24(5):67–68, May 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Voelcker:1987:PCP

[Voe87i] J. Voelcker. Personal computers: 1986 produced hordes of clones 80386-based PCs approached the power of workstations new laptops look for their market copy protection vanished, desktop publishing flourished. IEEE Spectrum, 24(1): 32–35, January 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Voelcker:1987:WDR

[Voe87j] J. Voelcker. Winchester disks reach for a gigabyte: Poised for a 200-fold increase in capacity over initial models of six years ago, 5 1/4-inch hard disks are emerging as major elements for smallsystem designs. IEEE Spectrum, 24(2):64–67, February 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Voelcker:1988:PFP

[Voe88a] J. Voelcker. 1965 the PDP-8: The first personal computer for engineers and scientists ushered in the minicomputer era. IEEE Spectrum, 25(11):86–92, November 1988. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Voelcker:1988:ASP

[Voe88b] J. Voelcker. Automating software: proceed with caution. IEEE Spectrum, 25(7):25–27, July 1988. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Voelcker:1988:EPF

[Voe88c] J. Voelcker. Electronics puts its foot on the gas. IEEE Spectrum, 25(5):53–57, May 1988. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Voelcker:1988:FSL

[Voe88d] J. Voelcker. Flex in specs: a license to innovate? IEEE Spectrum, 25(12):55–60, November 1988. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Voelcker:1988:IOE

[Voe88e] J. Voelcker. The iffy ‘Orient Express’ [hypersonic transport]. IEEE Spectrum, 25 (8):31–33, August 1988. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Voelcker:1988:TPC

[Voe88f] J. Voelcker. Technology ’88:

personal computers. IEEE Spectrum, 25(1):32–34, January 1988. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Voelcker:1988:TS

[Voe88g] J. Voelcker. Technology ’88:

software. IEEE Spectrum, 25 (1):38–40, January 1988. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Voelcker:1989:GE

[Voe89a] J. Voelcker. The Gunn effect.

IEEE Spectrum, 26(7):24–24–, July 1989. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Voelcker:1989:RAL

[Voe89b] J. Voelcker. Reemployment assistance limited. IEEE Spectrum, 26(11):52–55, November 1989.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Voelcker:1989:TPC

[Voe89c] J. Voelcker. Technology ’89: personal computers. IEEE Spectrum, 26(1):31–34, January 1989. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Voelcker:1989:TS

[Voe89d] J. Voelcker. Technology ’89: software. IEEE Spectrum, 26 (1):37–39, January 1989. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Voelcker:1986:HDP

[VW86] J. Voelcker and P. Wallich. How disks are ‘padlocked’. IEEE Spectrum, 23(6):32–40, June 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Voelcker:1986:LLC

[VWZ86a] J. Voelcker, P. Wallich, and Glenn Zorpette. Lessons learned: Challenges abound as designers tackle integrated software, artificial intelligence, and userfriendliness. IEEE Spectrum, 23(5):62–64, May 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Voelcker:1986:PAD

[VWZ86b] J. Voelcker, P. Wallich, and Glenn Zorpette. PC applications: Directions: In the offing is more intuitive, more useful software than today’s versions, thanks to more powerful microprocessors. IEEE Spectrum, 23(5):64–68, May 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Voelcker:1986:PAE

[VWZ86c] J. Voelcker, P. Wallich, and Glenn Zorpette. PC applications: Engineering: Only recently endowed with the power to do engineering chores, personal computers am already saving designers time and money. IEEE Spectrum, 23(5):68–75, May 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Voelcker:1986:PDA

[VWZ86d] J. Voelcker, P. Wallich, and Glenn Zorpette. PC design: Architecture: Trading off technical innovation against the dependability of established standards is a principal consideration for PC designers. IEEE Spectrum, 23(5):50–56, May 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Voelcker:1986:PDM

[VWZ86e] J. Voelcker, P. Wallich, and Glenn Zorpette. PC design: Manufacturing: Although relatively simple to assemble, PCs contain highly complex parts; manufacturing them is more akin to making consumer goods than computer products. IEEE Spectrum, 23(5):56–61, May 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Voelcker:1986:PCL

[VWZ86f] J. Voelcker, P. Wallich, and Glenn Zorpette. Personal computers: Lessons learned. IEEE Spectrum, 23(5):44–45, May 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wallich:1982:ICI

[Wal82a] P. Wallich. Information control II: Technology transfer at issue: The industry viewpoint: Economics makes industry a cautious ally in academia’s fight against U.S. government attempts to stem the free flow of ideas. IEEE Spectrum, 19 (5):69–73, May 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wallich:1982:ICD

[Wal82b] P. Wallich. Information control: The dilemma of technology transfer: Industry–government cooperation is sought in developing priorities for the protection of technical information developed in the U.S. IEEE Spectrum, 19 (9):66–70, September 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wallich:1982:TAF

[Wal82c] P. Wallich. Today’s air force: The chips are flying. IEEE Spectrum, 19(10):48–52, October 1982. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Wallich:1982:WTA

[Wal82d] P. Wallich. War 2: Today’s arsenals: The modern army, navy, and air force turn to the technologies that can ensure a battlefield edge: Today’s armies: Plodding toward automation. IEEE Spectrum, 19(10):38–40, October 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wallich:1983:DNG

[Wal83a] P. Wallich. Designing the next generation: Machines that are much faster, much smarter, and far more accessible than current ones require advances on a broad technological front. IEEE Spectrum, 20(11):73–74, November 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wallich:1983:DMI

[Wal83b] P. Wallich. Design/manufacturing:

Integrating CAD with CAM is the major challenge for automating a complete design-to-manufacturing process. IEEE Spectrum, 20(1):55–58, January 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wallich:1983:ESA

[Wal83c] P. Wallich. Electronics in space:

Applications of the Space Shuttle, major advances in remote sensing, and an energy-efficient aircraft are highlighted. IEEE Spectrum, 20(1):71–74, January 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wallich:1983:OAM

[Wal83d] P. Wallich. Office automation: Most of the pieces of the automated business puzzle now exist, but they have yet to be fitted together. IEEE Spectrum, 20(1):41–42, January 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wallich:1983:PIT

[Wal83e] P. Wallich. Putting it together: an electronic panel discussion among engineers and scientists elicits a call for a national computer network that spans disciplines. IEEE Spectrum, 20(11): 105–109, November 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wallich:1984:DCD

[Wal84a] P. Wallich. Design: Chip designers are allocated 14 working days to complete their part of the team effort; they must rely on scare, expensive software tools. IEEE Spectrum, 21(9): 30–34, September 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wallich:1984:ESC

[Wal84b] P. Wallich. EE specialties: Computers are pervasive: Solid-state and computer technologies underlie the shifting emphasis in technical disciplines. IEEE Spectrum, 21(6):44–48, June 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wallich:1984:EWF

[Wal84c] P. Wallich. Engineering workstations: a field report: Interviews with engineers having access to CAE workstations yield mixed responses; amidst the praise are complaints of insufficient power and excess complexity. IEEE Spectrum, 21(12):52–57, December 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wallich:1984:EJI

[Wal84d] P. Wallich. The engineer’s job: It moves toward abstraction: Automated modeling and analysis minimize the practitioner’s contact with real hardware. IEEE Spectrum, 21 (6):32–37, June 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wallich:1984:HFC

[Wal84e] P. Wallich. On the horizon: Fast chips quickly: It’s done using silicon compilers. Though they’ve built some significant circuits, their mare ambitious roles are still in the future. IEEE Spectrum, 21(3):28–34, March 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wallich:1984:REW

[Wal84f] P. Wallich. A review of engineering workstations: These powerful computers are to the engineer what word processors are to the secretary: No panacea, hut an important aid. IEEE Spectrum, 21(10):48–53, October 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wallich:1984:SPR

[Wal84g] P. Wallich. Software: Progress is reported in expert systems and software to build such systems and in object-oriented programming. IEEE Spectrum, 21 (1):47–49, January 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wallich:1985:LSS

[Wal85a] P. Wallich. Lighting sets the scene. IEEE Spectrum, 22 (5):48–50, May 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wallich:1985:MMS

[Wal85b] P. Wallich. Minis and mainframes: Superminicomputers push mainframe performance, mainframes operate at supercomputer speeds, and supercomputers reach 400 million operations per second. IEEE Spectrum, 22(1):42–44, January 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wallich:1985:NDN

[Wal85c] P. Wallich. New directions for a ‘national resource’ [telecommunications]. IEEE Spectrum, 22 (11):90–96, November 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wallich:1985:SAI

[Wal85d] P. Wallich. Software: Artificialintelligence software becomes available for minis and mainframes, and the U.S. Department of Defense converts to Ada and pushes to produce secure systems. IEEE Spectrum, 22 (1):50–52, January 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wallich:1985:TSF

[Wal85e] P. Wallich. Toward simpler, faster computers: By omitting unnecessary functions, designers of reduced-instruction-set computers increase system speed and hold down equipment costs. IEEE Spectrum, 22 (8):38–45, August 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wallich:1986:KTF

[Wal86a] P. Wallich. Keeping them flying: airlines sometimes ignore maintenance problems to honor flight schedules — rather than canceling flights to make aircraft worthy by government standards. IEEE Spectrum, 23 (11):69–75, November 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wallich:1986:SDP

[Wal86b] P. Wallich. Software ‘doctor’ prescribes remedies. IEEE Spectrum, 23(10):43–48, October 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wallich:1986:USI

[Wal86c] P. Wallich. U.S. semiconductor industry getting it together: Despite their history of intense competition, U.S. manufacturers are cooperating in research to gain an edge in the international marketplace. IEEE Spectrum, 23(4):75–78, April 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wallich:1987:BCL

[Wal87a] P. Wallich. Better cars for less: Automation is only one of several roads to auto plants with improved quality and production efficiency. IEEE Spectrum, 24 (10):43–46, October 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wallich:1987:CSP

[Wal87b] P. Wallich. Case study: Phone chip: Shared circuits save space. IEEE Spectrum, 24(5):54, May 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wallich:1987:MIU

[Wal87c] P. Wallich. Making it under duress. IEEE Spectrum, 24(11):

47–51, November 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wallich:1987:PIO

[Wal87d] P. Wallich. Part II. Optimizing the design: How and when to make tradeoffs: Aircraft must be lighter, consumer products more functional implantable devices more reliable but at what cost? IEEE Spectrum, 24 (5):33–35, May 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wallich:1987:PNSa

[Wal87e] P. Wallich. Program notes software topics. IEEE Spectrum, 24(6):25–26, June 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wallich:1987:PSR

[Wal87f] P. Wallich. Putting speech recognizers to work: While advances in signal processing and algorithms would extend their usefulness, limited models are already meeting many inspection and inventory applications. IEEE Spectrum, 24(4):55–57, April 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wallich:1987:SAI

[Wal87g] P. Wallich. Software: Artificialintelligence applications appear at last large-scale software projects gain tools strategic defense initiative presses on. IEEE Spectrum, 24(1):36–38, January 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wallich:1987:VAW

[Wal87h] P. Wallich. Volts and amps are not what they used to be: Nor are several other common units of measurement, now that technology can link them to natural constants and provide portable standards of unprecedented accuracy. IEEE Spectrum, 24 (3):44–49, March 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

White:1985:TEM

[WAO85] D. R. J. White, K. Atkinson, and J. D. M. Osburn. Taming EMI in microprocessor systems: The authors propose a systematic approach to limiting the effects of electromagnetic interference. IEEE Spectrum, 22 (12):30–37, December 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Warner:1982:RDF

[War82] V. F. Warner. The rapiddeployment force: How it should work. IEEE Spectrum, 19(10): 87–89, October 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Watson:1987:WHD

[Wat87] G. F. Watson. Whatever happened to direct-broadcast satellites? IEEE Spectrum, 24 (7):16–16L, July 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wallich:1986:SSD

[WB86] P. Wallich and L. A. Belady.

Software: The strategic defense initiative brings into focus issues of software design, software development productivity, and complexity. IEEE Spectrum, 23 (1):43–45, January 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wallich:1987:PNSb

[WBK87] P. Wallich, L. Belady, and J. Kellum. Program notes software news. IEEE Spectrum, 24 (8):22–23, August 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wolff:1987:MTd

[WC87] M. F. Wolff and R. F. Cotellessa.

Managing technology. IEEE Spectrum, 24(11):20–22, November 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Weinschel:1980:PTP

[Wei80] B. O. Weinschel. Politics of technology: Proposal: a national engineering foundation: Would it not help to halt the slide in U.S. productivity and its loss in share of world markets? IEEE Spectrum, 17(2):58–60, February 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Weinstein:1984:SCC

[Wei84] S. B. Weinstein. Smart credit cards: The answer to cashless shopping: Pioneered in France, microprocessor-based credit and debit cards are being used in large-scale tests in three cities there. IEEE Spectrum, 21(2): 43–49, February 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Weinstein:1987:TCD

[Wei87]  S. B. Weinstein.  Telecom-

munications in the coming decades. IEEE Spectrum, 24 (11):62–67, November 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wells:1984:MUN

[Wel84] P. N. T. Wells. Medical ultrasonics: As new techniques are perfected, ultrasound extends its usefulness from diagnosis to research, therapy, and even surgery. IEEE Spectrum, 21 (12):44–51, December 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Weldon:1985:PPP

[Wel85] W. F. Weldon. Pulsed power packs a punch: Lasers, electromagnetic launchers, and fusion reactors require enormous, instantaneous pulses of power not available from the utilities’ mains. IEEE Spectrum, 22 (3):59–67, March 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Werner:1987:DAP

[Wer87a] K. I. Werner. Design automation: Personal computers became important platforms for computer-aided engineering integrated software tools with a common graphics interface proliferated workstations became faster and more cost-effective. IEEE Spectrum, 24(1):39–41, January 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Werner:1987:CSD

[Wer87b] K. L. Werner. Case study:

Digital scopes: Assembly in record tim. IEEE Spectrum, 24(5):52–53, May 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wetzler:1982:CPA

[Wet82a] F. U. Wetzler. Communications:

AT t plans ahead: The telecommunications giant will enter the commercial marketplace, drawing on its traditional technical strengths. IEEE Spectrum, 19(7):49–54, July 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wetzler:1982:PEC

[Wet82b] F. U. Wetzler. Power/energy:

Connecting renewable power sources into the system: Windmills, fuel cells, and photovoltaics may serve existing grids, but first many technical, legal, and economic issues must be addressed. IEEE Spectrum, 19(11):

42–45, November 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

White:1983:CMD

[Whi83] R. M. White. Computers: Magnetic disks: Storage densities on the rise: New materials and new techniques, such as vertical and magnetooptical recording, are driving down costs per stored bit. IEEE Spectrum, 20 (8):32–38, August 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Whiting:1985:WN

[Whi85]  C. Whiting.  What’s in a

name? IEEE Spectrum, 22 (11):62–63, November 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Whitney:1987:PID

[Whi87] D. E. Whitney. Part IV. designing for producibility: Manufacturing and design: a symbiosis: If robots and other advanced assembly methods are to be used efficiently, products must be scrutinized in terms of manufacturability early in the design cycle. IEEE Spectrum, 24(5):47, May 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Willenbrock:1980:PTU

[Wil80] F. K. Willenbrock. Politics of technology: U.S. technological progress: Alternative approaches: Technical developments might best be served by modifying existing agencies or by legislation, rather than by creating a new agency. IEEE Spectrum, 17(10):51–52, October 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Williams:1981:LND

[Wil81] W. C. Williams. Lessons from NASA: Design reviews, failure analysis, and redundancy are favored over predictions and life testing for spacecraft success. IEEE Spectrum, 18(10): 79–84, October 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wiley:1987:PAC

[Wil87] P. Wiley. A parallel architecture comes of age at last! Hypercube machines reach the market on the promise of executing billions of operations a second to solve difficult simulation problems. IEEE Spectrum, 24(6):46–50, June 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Winner:1984:LMH

[Win84] L. Winner. Lifestyle: Mythinformation in the high-tech era: The romanticization of the personal computer as a social panacea threatens to blind society to the fact that without guiding wisdom even the best tool can be misused. IEEE Spectrum, 21(6):90–96, June 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wise:1982:ISW

[Wis82] G. Wise. Innovation: Swan’s way: a study in style: With the birth of the electric light bulb, two modern styles of inventing emerged: Thomas Edison’s and Joseph Swan’s. IEEE Spectrum, 19(4):66–71, April 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Withrow:1984:EBL

[Wit84] F. B. Withrow. Education: Bringing learning home: Instructional TV: New experiments with television as a teaching medium are promising a democratization of learning beyond even that accomplished by the public schools. IEEE Spectrum, 21(6):119, June 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Watkins:1984:SCG

[WM84] H. S. Watkins and J. S. Moore. A survey of color graphics printing: User needs for permanent copies of charts, graphs, and other illustrations are growing rapidly, and the technology is catching up to meet the demand. IEEE Spectrum, 21(7):26–37, July 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wolf:1980:PIF

[Wol80] M. Wolf. Photovoltaics II: Flat panels: Fixed-base silicon arrays are expected to exploit efficient, lightly doped cells. IEEE Spectrum, 17(2):32–33, February 1980. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wolff:1983:AGM

[Wol83] M. F. Wolff. Among the giants:

Mervin J. Kelly: Manager and motivator: The director of Bell Laboratories during an exciting 23-year period is credited with asking the right questions and finding the best people to answer them. IEEE Spectrum, 20 (12):71–75, December 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wolff:1984:RBW

[Wol84a] M. F. Wolff. Riding the biggest wave: Entrepreneur Andrew Kay is betting that his brand of humanistic management will ensure Kaypro’s survival in the turbulent PC waters. IEEE Spectrum, 21(12):66–71, December 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wolff:1984:WDC

[Wol84b] M. F. Wolff. William D.

Coolidge: Shirt-sleeves manager: Famous for two wideranging inventions, this engineer proved himself an innovative manager as he guided the renowned GE Research Laboratory through some of its most turbulent years. IEEE Spectrum, 21(5):81–85, May 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wolff:1985:SSM

[Wol85] M. F. Wolff. The secret sixmonth project: Why Texas Instruments decided to put the first transistor radio on the market by Christmas 1954 and how it was accomplished. IEEE Spectrum, 22(12):64–69, December 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wolff:1987:MTa

[Wol87a] M. F. Wolff. Managing technology. IEEE Spectrum, 24 (1):20–24, January 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wolff:1987:MTb

[Wol87b] M. F. Wolff. Managing technology. IEEE Spectrum, 24(5):24, May 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wolff:1987:MTc

[Wol87c] M. F. Wolff. Managing technology. IEEE Spectrum, 24 (7):18–20, July 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wollard:1988:GP

[Wol88a] K. Wollard. Garbage in, power out. IEEE Spectrum, 25(6): 42–44, June 1988. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wollard:1988:TE

[Wol88b] K. Wollard. Technology ’88: education. IEEE Spectrum, 25 (1):68–69, January 1988. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wollard:1988:TSS

[Wol88c] K. Wollard. Technology ’88:

Solid state. IEEE Spectrum, 25 (1):44–46, January 1988. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wollard:1988:LIA

[Wol88d] K. Wollard. Uno lamm: inventor and activist. IEEE Spectrum, 25(3):42–45, March 1988. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wolff:1989:CDP

[Wol89] M. F. Wolff. Carlson’s dry printer. IEEE Spectrum, 26 (12):44, December 1989. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Waxman:1989:VLD

[WSC89] R. Waxman, L. Saunders, and H. Carter. VHDL links design, test, and maintenance. IEEE Spectrum, 26(5):40–44, May 1989. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wallich:1985:IRA

[WZ85] P. Wallich and Glenn Zorpette. The innovation revolution awaits: Expectations of a technological bonanza might have been inflated, but new openings are seen for the entrepreneurial opportunist. IEEE Spectrum, 22(11):104–110, November 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wallich:1987:MMV

[WZ87] P. Wallich and Glenn Zorpette. Minis and mainframes: VAX family grows at a dizzying pace minicomputer makers merge to survive minisupers emulate VAXes and Cray’s Connection Machine is an AI supercomputer. IEEE Spectrum, 24 (1):28–31, January 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wollard:1988:TPE

[WZ88] K. Wollard and Glenn Zorpette. Technology ’88: power and energy. IEEE Spectrum, 25(1): 53–55, January 1988. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wollard:1989:TPE

[WZ89] K. Wollard and Glenn Zorpette. Technology ’89: power and energy. IEEE Spectrum, 26(1): 56–58, January 1989. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wallich:1986:MMM

[WZB86] P. Wallich, Glenn Zorpette, and C. G. Bell. Minis and mainframes: Microcomputers as we know them may disappear as sophisticated microprocessors reach toward mainframe capability; more powerful supercomputers are introduced. IEEE Spectrum, 23(1):36–39, January 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Xi:1989:ECF

[XM89] X.-I. Xi and G. Morgan. Energizing China: first itself, next the world. IEEE Spectrum, 26 (3):59–63, March 1989. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Yeh:1983:SE

[Yeh83] R. Yeh. Software engineering. IEEE Spectrum, 20(11): 91–94, November 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Yould:1984:APB

[You84] E. P. Yould. Alaska plans a big switch: Most of its electricity is oil-generated, but the construction of new hydroelectric stations is expected to increase electrification throughout the state. IEEE Spectrum, 21 (5):71–76, May 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zadeh:1984:MCT

[Zad84] L. A. Zadeh. Making computers think like people [fuzzy set theory]. IEEE Spectrum, 21 (8):26–32, August 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zdeblick:1983:PDC

[Zde83] W. J. Zdeblick. Process design by computer: In the ideal case the computer, rather than shop personnel, will select machine stations and operation sequences. IEEE Spectrum, 20 (5):56–58, May 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1986:PCF

[ZJ86] Glenn Zorpette and E. Juliussen. Personal computers: Fewer developments hut nevertheless important ones mark a year characterized by less-than-expected growth. IEEE Spectrum, 23 (1):40–42, January 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1984:CIF

[Zor84a] Glenn Zorpette. The current issues: Formal education: EE programs. IEEE Spectrum, 21 (11):44–50, November 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1984:GPH

[Zor84b] Glenn Zorpette. Gs/e programs: How to design and apply computers is the objective of a growing class of computer-based curricula that are called by various names. IEEE Spectrum, 21 (11):50–54, November 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1984:HTB

[Zor84c] Glenn Zorpette. High-tech batteries for power utilities: Load leveling and peak shaving are attractive reasons for power companies and customers alike to encourage development of advanced batteries. IEEE Spectrum, 21(10):40–47, October 1984. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1985:BBN

[Zor85a] Glenn Zorpette. The beauty of 32 bits: This near-optimum bit width has unprecedented potential for the well-informed designer of microprocessor-based systems. IEEE Spectrum, 22(9): 65–71, September 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1985:CNF

[Zor85b] Glenn Zorpette. Computers that are ‘never’ down [fault-tolerant computers]. IEEE Spectrum, 22(4):46–54, April 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1985:HTH

[Zor85c] Glenn Zorpette. The high-tech hobbyhorse. IEEE Spectrum, 22(5):96–101, May 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1985:HWL

[Zor85d] Glenn Zorpette. HVDC: Wheeling lots of power: Moving power from where it is plentiful to where it is needed may be more efficient than building oil-fired or nuclear plants locally. IEEE Spectrum, 22 (6):30–36, June 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1985:MNB

[Zor85e] Glenn Zorpette. Microprocessors: New 32-bit microprocessors in CMOS perform virtualmemory functions on the chip and show promise in applications ranging from multiprocessing to fault-tolerant computing. IEEE Spectrum, 22(1): 53–55, January 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1985:TBS

[Zor85f] Glenn Zorpette. The telecommunications bazaar: Suppliers of telecommunications equipment and services find opportunities but crowded markets in everything from telephones to toll switches. IEEE Spectrum, 22 (11):56–61, November 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1986:MMW

[Zor86a] Glenn Zorpette. The menacing microburst: Wind shear, the single greatest cause of air crashes in the United States, may soon begin to yield to technological solutions. IEEE Spectrum, 23 (11):50–56, November 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1986:MTM

[Zor86b] Glenn Zorpette. Monitoring the tests: Missile flights and test explosions are watched closely to determine whether an adversary’s new weapons might upset the strategic balance. IEEE Spectrum, 23(7):57–66, July 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1986:RFP

[Zor86c] Glenn Zorpette. Robots for fun and profit: a new breed of experimenter is more artist than scientist, yet this work sometimes pays off handsomely. IEEE Spectrum, 23(3):71–75, March 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1986:SUE

[Zor86d] Glenn Zorpette. The shrinking U.S. electric equipment market: What is eating away at U.S. manufacturers’ share of the world market? An alleged European–Japanese cartel may be compounding the slowdown in U.S. growth. IEEE Spectrum, 23 (9):63–65, September 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1987:BEC

[Zor87a] Glenn Zorpette. Breaking the enemy’s code: British intelligence deciphered Germany’s top-secret military communications with Colossus, an early vacuum-tube computer. IEEE Spectrum, 24 (9):47–51, September 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1987:EPL

[Zor87b] Glenn Zorpette. Evacuation planning for Lilco’s Shoreham plant: Lack of an approved emergency evacuation strategy may prevent full operation of a $5 billion nuclear power plant on Long Island’s north shore. IEEE Spectrum, 24(7):22–24, July 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1987:PP

[Zor87c] Glenn Zorpette. Power politics. IEEE Spectrum, 24(11): 34–37, November 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1987:SS

[Zor87d] Glenn Zorpette. The Shoreham saga. IEEE Spectrum, 24 (11):24, November 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1987:WOA

[Zor87e] Glenn Zorpette. Weather outlook: Acid rain or shine? Engineers, scientists, and government officials together seem able to do little to stop the ravages of airborne industrial pollution. IEEE Spectrum, 24 (3):24–29, March 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1987:WWW

[Zor87f] Glenn Zorpette. What went wrong? A fault, a failure, and a debate: Even the best manufacturers may find problems in unexpected places. IEEE Spectrum, 24(10):61–63, October 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1988:PIS

[Zor88a] Glenn Zorpette. 1970 the Pacific intertie: Swedish technology helped harness hydropower to energize booming Southern California. IEEE Spectrum, 25 (11):92–97, November 1988. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1988:FJA

[Zor88b]  Glenn Zorpette.  Fractals:

Not just another pretty picture. IEEE Spectrum, 25(10): 29–31, October 1988. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1988:GOF

[Zor88c] Glenn Zorpette. Government oversight: fighting waste and fraud. IEEE Spectrum, 25(12): 44–47, November 1988. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1988:MTP

[Zor88d] Glenn Zorpette. Military testing:

pass or fail? IEEE Spectrum, 25 (12):64–67, November 1988. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1988:TMM

[Zor88e] Glenn Zorpette. Technology ’88: minis and mainframes. IEEE Spectrum, 25(1):29–31, January 1988. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1988:UVA

[Zor88f] Glenn Zorpette. An underwater visit to Aquarius. IEEE Spectrum, 25(3):24–29, March 1988. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1989:CEP

[Zor89a] Glenn Zorpette. At the cutting edge of power engineering. IEEE Spectrum, 26(12):39–43, December 1989. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1989:KPL

[Zor89b] Glenn Zorpette. Keeping the phone lines open. IEEE Spectrum, 26(6):32–36, June 1989. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1989:KNR

[Zor89c] Glenn Zorpette. Kwajalein’s new role: radars for SDI. IEEE Spectrum, 26(3):64–69, March 1989. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1989:MPT

[Zor89d] Glenn Zorpette. Moving power through the northeast corridor. IEEE Spectrum, 26(8): 46–47, August 1989. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1989:PGD

[Zor89e] Glenn Zorpette. Parkinson’s gun director. IEEE Spectrum, 26(4):43, April 1989. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1989:PTG

[Zor89f] Glenn Zorpette. Photovoltaics:

technical gains and an uncertain market. IEEE Spectrum, 26(7):42–43, July 1989. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1989:RWS

[Zor89g] Glenn Zorpette. Rethinking weapons and strategy. IEEE Spectrum, 26(11):36–42, November 1989. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1989:TMM

[Zor89h] Glenn Zorpette. Technology ’89:

minis and mainframes. IEEE Spectrum, 26(1):29–31, January 1989. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1989:TME

[Zor89i] Glenn Zorpette. Technology ’89:

the main event. IEEE Spectrum, 26(1):28, January 1989. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1989:UNI

[Zor89j] Glenn Zorpette. The US nuclear industry: a changing complexion. IEEE Spectrum, 26 (10):50–51, October 1989. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).